A Complete Bibliography of Proceedings of the ACM on Human-Computer Interaction (PACMHCI)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 November 2024 Version 1.04

Title word cross-reference

#BlackLivesMatter [ASS18]. #pray4victims [BL18].

1 [JFS^{+}23]. 2 [JTB22, OSM^{+}20, Zha23]. 3 [CPB18, DYF^{+}23, ERH^{+}22, GGL^{+}22, KIK20, KSG^{+}24, LDM^{+}22, OKM^{+}23, OSM^{+}20, SEC^{+}20]. 360 [HKL^{+}22]. 6 [ZLB^{+}22]. 8 [GN23]. μV [MRV22].

-degree [HKL^{+}22]. -DoF [ZLB^{+}22]. -year [GN23].

100m [CWR22]. 19 [CEJC23, CBV22, HSVR21, JKC23, KJTC22, KUD^{+}23, KBSP^{+}22, PKLK22, PAC22, SLB23, WBS^{+}23, XKRW23, YGHR21, dSdSdSF^{+}23].

2 [CBWD24]. 2.0 [KE18]. 2020 [CM20, KLS20, KLS21]. 2021 [BKL^{+}21, BLS21, FJIH21, GMM21]. 2022 [ACI22, BLS^{+}22, BMdS^{+}21, BMdS^{+}22, FBNM22, GKKW22, KRW^{+}22]. 2023 [CDG^{+}23, CDGdC23, DKBG23, LMS23, LS23]. 2024 [CDG^{+}24, DKKL24, KvBW24, MBTW24, SD24].

360Anywhere [SCY^{+}18]. 360theater [SLN21]. 3DPFIX [KSG^{+}24].

A/B [VZV19]. Aadhaar [SJ21]. AARP [MM23]. AB4Web [VZV19]. Abilities [HBJR21]. Absolute [DWMV21]. Abstracting [AV19]. Abstraction [KE18]. Abstractions [HJCM22]. Abuse [Bel24, RYD21]. Abusive [Bel24]. Academia [KG21]. Academic [ARH^{+}23, ZD19]. Accelerating [XRS22]. Acceptance [AASK23, BBW23, SAE^{+}21]. Access [FF22, GN23, LZAD21, PKC^{+}21, PNS^{+}24b]. Accessibility [BP17b, BMPP20, SCP^{+}21, UAAR20, WDD^{+}22]. Accessible [BLB^{+}24, LCG24, ZMS^{+}22]. Account [HLV21, WFL^{+}22]. Accountability [SSH^{+}21]. Accounting [JA21]. Accounts [HV22]. Accuracy [BKG^{+}21, CYGW21, MPEC24, UYSM22, YU20]. Accurate [VMD^{+}20]. Achievement [OS24]. Acknowledgment [GTC22]. Acoustic [ZWZ^{+}21]. Acquisition [MdLGG23, PDPS22, UYM21]. Across [BGM^{+}23, BBNT21, DNBB^{+}22, FD20, HV22, KBSP^{+}22, CHS^{+}24, HWF^{+}24, LHT^{+}20, WLZL23]. Acting [ASS18, FA22]. Action [HMK24, HWS21, JKZ^{+}22, JWK22, JPKM24, LO17, NHH20, SRZ24, SRP^{+}24]. Action-Adventure [SRZ24]. Action-State [JPKM24]. Action2Score [JWK22]. Actionable [GYD17]. Active [CMSA22, GLZ^{+}21, LMB^{+}22]. Activism [AS23]. Activists [BWDP20, GYD17]. Activities [BDB^{+}23, BF17, FCE^{+}18, FAMS22, DRDS24, HFBL19, JKZ^{+}22, LMT^{+}22, MTA^{+}17]. Activity [BZC^{+}17, BRM21, CTWM23, EPW^{+}23, GRG20, KBKH21, MFSK22, MMM^{+}21]. Actuated [SLT^{+}21]. Ad [FX22, KKK22a, SCY^{+}18, TRN^{+}22]. Ad-hoc [FX22, KKK22a, SCY^{+}18]. Adaptation [ACDC23, PNS24a, TRO^{+}21]. Adaptations [AMVK23, YHR^{+}19]. Adapting [BS21, CTWM23, NMPDJ24, RZP^{+}22]. Adaptive [BWE^{+}22, BCCV17, LCA^{+}24]. AdaptiX [PGK^{+}24]. Address [AL18, ESCR23]. Addressing [AAHM22, RYD21]. Adherence [RAZ^{+}20]. Administrators [GM23]. Adolescents [BUSA^{+}21, MBK^{+}22]. Adoption [ETS^{+}23, KBSP^{+}22, SAE^{+}21]. ADQDA [LE21]. Adult [CNH^{+}23, VVBK23, ZRL21]. Adulthood [BBK22]. Adults [CMRH22, KSX21, WGM^{+}24, XKRW23]. Advancing [SRZ24]. Advantages [KDH^{+}21]. Adventure [SRZ24]. Adversarial [VHZ22]. Advertise [AOT^{+}18]. Advertising [HPY^{+}22]. Advice [BBK22, HKM^{+}20, SSB^{+}22]. Advisory [BDE^{+}24]. Advocacy [DDV23, DCM^{+}23]. Advocating [SLMB24]. Aesthetic [HHLC23]. Aesthetic-Based [HHLC23]. Aesthetics [ZBV23]. Affect [MFM^{+}20]. Affected [IOES20]. Affective [HC24, IB24, SVS^{+}24, SCSD24]. Affects [DBS^{+}21]. Affinity [LE21, SPT^{+}20]. Affordable [SS22]. Affordances [JNLtB22, MCM23]. Afforded [Ann20]. afraid [CCWH21]. After [LHPH20, SSH^{+}21, CH20]. Aftermath [PNS24a]. Again [AGM^{+}23]. Against [WWSS22]. Age [HKD23, LTP23]. Age-Integrated [LTP23]. Agency [BSS18, GPC^{+}23, GW24, HMB^{+}21, LLAK23, RYD21]. Agenda [BHH^{+}22]. Agent [HLV21, JFL23, JRP^{+}23, KSKH24, LCA^{+}24, LYD^{+}23, SFM^{+}22]. Agent-based [KSKH24]. Agents [BDE^{+}24, JV22, SSB^{+}22, ZMR21]. Aggregated [KSP^{+}23]. Aging [BP17b, SAE^{+}21]. Agreements [KM22]. Agriculture [NPT19]. Ah [YYT^{+}24]. Ah-Aloud [YYT^{+}24]. Ahead [ITKB23]. AI [BYRL23, PF24, VHZ22, ALD^{+}20, BP23, CBSQ24, DBL22, ESCR23, EHG22, GLZ^{+}21, GMB23, GHHC21, GHHS21, KG20, KSG^{+}24, LBM21, MES^{+}22, MBT^{+}20, MHN^{+}21, MWK^{+}22, MSHP23, MAD^{+}23, PF24, PVC^{+}23, RYCC21, RGR^{+}21, SSR24, SHJ^{+}22, VHZ22]. AI-Enabled [GHHC21]. AI-Led [KG20]. AI-Mediated [MHN^{+}21]. AI-On-Skin [BP23]. Aids [YPS^{+}22]. AIMEE [PVC^{+}23].

Aiming [GN23, HK24b]. Air [HGG20, LRP^{+}22, YMG^{+}20, BLC^{+}21, DRL^{+}22, GTO^{+}23, NM22a, VMHO^{+}22, VWHW24]. Aircraft [AARJ23]. Airflow [MXYS23]. Airtime [STM^{+}19]. Alert [ZSK^{+}23]. Alerts [CSL^{+}24, MAS22, SRKK22]. Alexa [CMSA22]. Algorithm [GC19, SHS22]. Algorithm-in-the-Loop [GC19]. Algorithmic [KDEA21, KB22, NM21, SB20b, SFA^{+}23]. Algorithms [BKG^{+}21]. Alienated [HTTH22]. Aligning [HBM23]. Alignment [CS18a, CT18, LRP^{+}22, WDD^{+}22]. Allocation [HBJR21]. Alone [AKVHZ23]. along [WAD^{+}23]. Aloud [MCI20, YYT^{+}24]. Alternative [Gil23, HRR^{+}22, YK23]. Alternatives [HZ22, VZV19]. Altruism [Vya19]. Alzheimer [HSC^{+}23]. Ambient [FRH^{+}22, HCM^{+}23, MSK22]. Ambiguity [BBK22]. Ambiguous [CKP^{+}18]. Ambivalent [PVP22]. Americans [MHP22]. Among [BKKD21, LPZV23, MAR^{+}19, SKV^{+}21, DB22, JV22, KSKH24, SELS24]. Amplification [GPC^{+}23, LHPH20]. Analogies [CCH^{+}18]. Analyses [GPC^{+}23]. Analysing [MPB^{+}19, WJH^{+}24]. Analysis [BHM^{+}23, FOH^{+}22, FYS^{+}24, GW22, HRRS20, HK24b, KML^{+}21, KMR20, LTP23, LE21, MB22, MRRS19, NAS^{+}22, SS23, SCP^{+}21, SB17, TAS^{+}19, TD19, WGGS20, XRS22, YHR^{+}19, dSdSdSF^{+}23]. Analytics [GW24, LMZ^{+}24, SRZ24, SB17, ZZ23]. Analyze [SLMB24]. Analyzing [BHSP22, CAATL22, GH21, HSG19, IB24, JSJ22, LL24, ZMZL21]. Anchors [BBW23]. Android [MP19]. Anganwadi [PVP22]. AngleCAD [LDM^{+}22]. Animations [WZFA22]. Annotation [DBPA22, MSY20, OKM^{+}23]. Annotations [CSS^{+}17, WPH^{+}22b]. Anomaly [PCMP18]. Anonymity [MCYZ19]. Answer [MWM^{+}19]. Answering [DMC^{+}20]. answers [MK23]. Antecedents [XZL^{+}20].

Anticipation [ECA23]. Anxiety [DBS^{+}21]. Any [SCY^{+}18]. Apart [FAMS22]. APEX [LMB^{+}22]. API [BGM^{+}23, RH23]. API-based [RH23]. App [ARH^{+}23, CEM17, LLC^{+}24, MP23, PLY23, SCSD24]. AppealMod [AIRH24]. Appeals [AIRH24]. Appearance [ERH^{+}24]. Appearance-based [ERH^{+}24]. Applicability [YTF^{+}23]. Application [BB20, CMP^{+}24, Sal17, SB17, SLD21, SHY^{+}18, TD19, TPNL22]. Applications [ASO^{+}22, CMN^{+}21, HMB^{+}21, JV22, KVM23, KOL^{+}24, LAR21, PGK^{+}24, SELS24, VMD^{+}20, ZMTH22]. Applied [TFI^{+}22]. Applying [XMS24]. Approach [ARH^{+}23, BCCV17, BPB22, CFG^{+}17, dACZ19, CXL17, DDFK19, FSC21, FSAN23, GGF^{+}20, HBLN24, JWK22, KBKH21, KZZW22, LDP18, MPKEN24, MFSK22, MRK^{+}22, NM21, SWT20, TVS17, ZBV23, ZZK^{+}22]. Approaches [AGM^{+}23]. Appropriateness [HC24]. Approximation [HPL22]. Apps [AZ23, ATH24, BWE^{+}22, EMS^{+}22, GN23, GMM19, KUD^{+}23, MBS22, SCSD24, ZMTH22]. April [BKL^{+}21, BLS^{+}22, CDG^{+}23, CDG^{+}24]. AR-Based [CSH^{+}23]. AR4CAD [MCDR22]. Arabian [ADAY21]. Arc [MPEC24]. Archetypes [PSTK24, RIC^{+}24]. Architectural [KSP^{+}23]. Architecture [SVW19, ZZK^{+}22]. Architectures [BHM^{+}23]. Archive [CLKEF21]. Areas [WvECM22]. Arm [CDIH23, GGL^{+}22]. Array [BHM^{+}23]. ARST [MRV22]. Art [BVM21, GMB23]. Artefacts [WPH^{+}22b]. Article [HJC^{+}22]. Articulation [LWTW23, MRV22, dCRM^{+}23]. Artifact [CAATL22]. Artifacts [KSP^{+}23, WQCZ24]. Artificial [BYRL23, WNT^{+}22]. Arts [LKS21]. Ask [BSA23, MWM^{+}19]. AskHistorians [Gil20]. Asking [MCYZ19]. Aspects [AS17, ACDC23, SB20b, dCRM^{+}23]. Ass [PKC^{+}21]. Assemblages [SSB^{+}24a].

Assembling [HFL19]. Assembly [CSH^{+}23, HBJR21, KFSC18, MAN^{+}20, YOFC19]. Assertions [BR17]. Assesment [RSL24]. Assessing [CBSQ24, KMM^{+}22, KKM24, MES^{+}22, SC23, SFM^{+}22]. Assessment [ASM23, BahRKM18, BZSM20, JF24, ND24, ZZK^{+}22]. Assessments [WHW^{+}22]. Assets [HTS24]. Assigned [CSL^{+}24]. Assignment [EHH^{+}18, HvBKG20]. Assist [CXL17, DDFK19, YPS^{+}22]. Assistance [ECA23, KFSC18, KKV^{+}22]. Assistant [CLL^{+}21, SSR24]. Assistants [CR20]. Assisted [CRM^{+}24, DRDS24, JFS^{+}23, JCD19, SWQ^{+}23]. Assisting [BBW21, KKK^{+}24, LPGG22]. Assistive [HOHB22, PGK^{+}24]. Association [CHCG24]. Assume [TPNL22]. Assurance [CCS^{+}18]. Asthma [SHJ^{+}22]. Asylum [NM22b]. Asymmetric [HRP23, OG21, WACdA^{+}21]. Asymmetries [NSJ21]. Asynchronous [AR22, CCNY19, CMY^{+}17, WQCZ24]. Attach [BR20]. attached [CST^{+}23]. Attack [ASJK23]. Attacks [WVG^{+}24]. Attendance [CSL^{+}24]. Attention [BSM23, BBD^{+}24, CI22, DWSF24, DRP^{+}23, KCC20, LAHW19, RKFK22, SSD23]. Attention-based [DRP^{+}23, RKFK22]. Attitudes [KUD^{+}23, RA21b, SSD23]. Attribute [CLP^{+}21]. Attribution [BB18]. Audio [CSS^{+}17, GKHD23, ND24, TRO^{+}21, XMS24, ZRP^{+}22]. AudioMove [XMS24]. Audit [HJM20]. Auditing [SMSM22]. Auditive [AARJ23]. Auditory [MRGN20, SAR^{+}24]. Augmentation [BRM21, DMPK22]. Augmentative [YK23]. Augmented [ALG^{+}21, BHH^{+}22, CWC^{+}24, CKK22, DGR^{+}22, DMPK22, DR24, DRL^{+}22, FSSK22, GGE23, GGL^{+}22, HCG^{+}22, KMM^{+}22, LZY^{+}21, LL24, LGG^{+}23, LRP^{+}22, MAN^{+}20, MP23, MGRM22, MdLGG23, MBK^{+}22, MCDR22, NAO23, Poh24, SAR^{+}23, SWQ^{+}23, SRKK22, SCY^{+}18, SHY^{+}18, SLN21, STM24, TRO^{+}21, TG18, WO20, WXL^{+}22, ZKL22]. Augmented-Walking [CKK22]. Augmenting [FZRJ22, WDP^{+}20]. Authentication [BLB^{+}24, WVG^{+}24]. Author [EHH^{+}18]. Authoring [KVM23, WZFA22]. Authority [HFL19]. Autism [RSL24, WXL^{+}22]. Autistic [LLAK23, VVBK23]. Auto [JFL23]. Auto-Response [JFL23]. Autogrip [BR20]. Automated [CWR22, CERR23, HCR22, ITKB23, KSW22, MCDR22, RSL24, SWT20, SCR22, TD19, ZMR21, NAS^{+}22]. Automatic [BMPP20, DTE^{+}22, GB18, HNS^{+}24, LYD^{+}23, RA21b, WvECM22]. Automatically [HSG19, LWS24]. Automation [CFR22, CVL22, SFA^{+}23]. Automations [MP23]. Autonomous [WWH^{+}23]. Autonomy [HFL19]. Avatar [DBS^{+}21, KRMM21, SFLB23]. Avoid [KKV^{+}22]. Aware [HCG^{+}22, JSJ22, MFSK22]. Awareness [CEM17, CHS^{+}24, DB22, EPW^{+}23, KTF^{+}21, LPW^{+}23, MRGN20, SRAG22, STM^{+}19, SPT^{+}20, WDP^{+}20, ZSK^{+}23]. Away [HFBL19]. Awe [CCW22]. Aww [CCW22]. Axis [SLT^{+}21, SLT^{+}21].

B [VZV19]. B2B [BPB22]. Baby

[ATH24, MMMM22]. Back

[AGM^{+}23, CKSP22, FC20, TFI^{+}22]. Back-of-Device [FC20]. Back-print [CKSP22]. Backchanneling [CMSA22]. Backgrounds [HWYW21]. Bad [CLL^{+}21, HV22]. Badge [BAAP20]. Balancing [KTF^{+}21, MPEC24, TBL^{+}24]. Ban [CPS^{+}17]. Bangladesh [AHCD17, RAA22, SAR21]. Barrier [UAAR20]. Barriers [BR24]. Based [ALG^{+}21, ABQ17, ALSK23, BJ21, BahRKM18, BAAP20, BLB^{+}24, CMN^{+}21, CLwH^{+}20, CSH^{+}23, CTWM23, DHH18, DR24, GSLH24, GB18, GKHD23, HTS24, HMK24, HBLN24, HHLC23, JKZ^{+}22, KMM21, LYL^{+}23, LDM^{+}22, LMB^{+}22, LNM21, MHP22, MSK22, NAO23, PA22, Poh24, RLGG21, SSB^{+}22, See20, SWT20, SSGB21, SM22, TFI^{+}22, WGS^{+}24, YTF^{+}23, YHR^{+}19, ZKL22, ZMS^{+}22, AV19, AMVK23, AGM^{+}23, Bel24, BKM20, BCCV17, BDE^{+}24, CMP^{+}24, dACZ19, CHC^{+}17, CHS^{+}24, DBS^{+}21, DRP^{+}23, ERH^{+}24, FWY^{+}22, FSAN23, FYS^{+}24, GH21, HSC^{+}23, HBJR21, HvBKG20, JNLtB22, KSKH24, KMV^{+}24, KNT22, LG22, LO17, LPGG22, MTA^{+}17, MNP^{+}22, RH23, RKFK22, SWQ^{+}23, SDO21, SBSM24, SORV22, TBR20, VN19, Vya19, YGHR21, YML^{+}23, ZBV23, ZGL^{+}22, ZWZ^{+}21, KKE^{+}21]. Basis [CSH^{+}23]. Batches [CLP^{+}21]. BCI [HRC^{+}21]. Be [DFJ^{+}21, PVP22]. Because [LC24]. Becoming [BEHB^{+}23]. BeeAR [MdLGG23]. Beeline [MdLGG23]. been [KBSP^{+}22]. before [ETS^{+}23]. begets [MIS^{+}20]. Beginner [WDP^{+}20]. Beginnings [GKHD23]. Behavior [CSA^{+}21, CFR22, CVL22, HZ22, HSG19, IB24, JJJ^{+}23, KVM24, KYB21, MMdSP21, MBT^{+}20, PCMP18, SWT20, VMP21]. Behavior-Based [SWT20]. Behavioral [CGS^{+}21, MYS^{+}21, RCM21, TD19]. Behaviors [HWF^{+}24, RJVS24, WRM^{+}24]. Behaviour [SRP^{+}24, WJH^{+}24]. Behind [KUWM22]. Beidian [CCX^{+}20]. Being [BYRL23, DFJ^{+}21, HKL^{+}22, SCSD24, CCWH21, KKP^{+}22]. Belief [SHM24]. Beliefs [KJTC22, SHM24]. Believe [ATH24]. Belonging [MBK^{+}22]. Beneath [FSRC24]. Benefit [TBL^{+}24, ZZ23]. Benefit-Driven [ZZ23]. Benefits [CBSQ24, FM22, HPH^{+}23]. Better [CRM^{+}24, KUWM22, SCSD24, TDC17, WB21]. Between [BLH^{+}17, CR20, DWMV21, FCE^{+}18, MSY20, SCSD24, XHH17, BZC^{+}17, BBD^{+}24, CCH^{+}18, LAM22, LZAD21, MPKEN24, PBC20, PDPS22, SWT20, SSNC20, WACdA^{+}21]. Beyond [BSA23, CYW18, JS22, LFK24, LGG^{+}23, MB23, RGW^{+}24, UHH23]. Bezel [RZP^{+}22]. Bezel-to-Bezel [RZP^{+}22]. Bézier [GVNK20]. Bias [JSJ22]. Biases [AOT^{+}18]. Bicycles [MMMM22]. Bike [MMMM22]. Bikers [BBW21]. Biking [GBZ^{+}24]. Bimanual [ZKL22]. Binary [SS21]. Binaural [ZRP^{+}22]. Bindings [KKS19]. Biological [MDM^{+}19]. BioNets [MDM^{+}19]. Biophilia [RVBC^{+}21]. Bipolar [XYDV23]. Bird [SBAK21]. Birds [VMP21]. Bit [BVM21]. Bitcoin [KPW19]. Black [ERT21, KKM^{+}22, KF23, MHP22, RI20]. ble [GKHD23]. Blending [AGM^{+}23]. BlendMR [HCM^{+}23]. Blind [BLB^{+}24, KKV^{+}22, LC24, MRGN20, PNS^{+}24b, SPL^{+}23, WS23]. Blogging [BP17b]. ’Blue [RAZ^{+}20, KKM^{+}22]. Board [PPT^{+}22, SBAK21, TRO^{+}21]. Boardgame [FHF24]. Bodies [EPW^{+}23]. Bodily [GBN21]. Body [EPW^{+}23, GB18, HOSK21, IOES20, Poh24, RJVS24, SFLB23, BP23]. Body-Based [Poh24]. Body-Focused [RJVS24]. Body-Worn [EPW^{+}23]. boilerplate [MK23]. Bolster [STM^{+}19]. Book [KKM^{+}22]. Books [KKM^{+}22]. Boolean [BSA23]. Boost [AOT^{+}18]. Borders [AWM21]. Bot [PK23]. Both [MKRM23, SAR^{+}24]. Bots [WWSG21, BKJ^{+}20]. bought [IRA^{+}19]. Boundaries [ZPMK24]. Boundary [CS18b, CBV22]. Bounded [YK23]. Box [WAD^{+}23]. Boxes [WZT^{+}24]. BPCoach [LMZ^{+}24]. Brain [HMML^{+}22, KMV^{+}24, NAO23]. Brain-Computer [NAO23]. BrainEx [HMML^{+}22]. Brainstorming [HSG19, LW19]. Break [HSH22, ZRL21]. Breakdowns [GMM19]. Breaking [BR24, UAAR20]. Breakout [MKRM23]. Breathe [ERT21]. breeds [BTM^{+}21]. Brevity [GAW19]. BrickStARt [SHC^{+}23]. Bridge [WQCZ24]. Bridges [WQCZ24]. Bridging [KG21, MHP22]. Bringing [BBW23, LAR21]. Browser [GGF^{+}20].

Browsing [CWC^{+}24]. Build [HSVR21, LC24]. Building [CSN19, DPF22, GB18, LLAK23, VAE^{+}22, WWT^{+}21, ZGL^{+}22]. Built [FKK19, Gil23, ZRL21]. Built-In [ZRL21]. Burden [TBL^{+}24]. Burnout [NHL^{+}24]. Business [LRK^{+}24b]. Businesses [KBFK21]. Button [BLC^{+}21, DNBB^{+}22]. Buy [EMS^{+}22]. Bystander [AFKL20, KTF^{+}21, LPW^{+}23]. Bystander-Awareness [LPW^{+}23]. Bystanders [WHW^{+}22, WM22].

CADTrack [BWFG23]. Calibrated [KA23]. Call [HDPL21]. called [BTM^{+}21]. Camera [JF24, NSV22, ZLB^{+}22]. Cameras [CHS^{+}24, Wil20]. Campaign [DDFK19, Kee19]. Campaigns [STM^{+}19]. Can [AGK^{+}22, ASJK23, AOT^{+}18, BYRL23, CMSA22, HHVB22, MNS^{+}21, PKC^{+}21, TKB^{+}22, TCH20, ETE^{+}19, MMMM22]. Canada [MB23, SAM23]. Cancer [SWF^{+}20]. Candy [BBS19]. Cane [SRKK22]. Can’t [BRM22, CPS^{+}17, ERT21]. Canvases [KDH^{+}21]. Capacities [KJTC22]. Capacitive [GBZ^{+}24, KIK20]. CAPath [KIK20]. Captioning [LLL^{+}22]. Capture [LAHW19, LLKC22, RLL22]. Capturing [DRAQ18, JKW^{+}21, CEJC23]. Car [YTF^{+}23]. Card [LKG21]. Cards [LKG21]. Care [BKJ22, CRM^{+}24, COKL20, FCR^{+}17, Gil23, KJTC22, KWdCL^{+}21, PVP22, PLL^{+}23, RCM21, SHJ^{+}22, SWF^{+}20, SMLH23, Vya19]. Care-mongering [KJTC22]. Career [BKKD21]. Caregivers [CRM^{+}24]. Caret [LZY^{+}21]. Caring [MDZ19]. CaringBridge [SHL23]. Cars [BDB^{+}23, YMG^{+}20]. Case [ADD21, AKIS21, BB18, CFG^{+}17, CEM17, CSH^{+}23, CHC^{+}17, FB19, Gar21, Gil20, HFL19, KOL^{+}24, KFSC18, KG20, LLDH21, LTP23, PLY23, XZL^{+}20, YTF^{+}23, vBGH^{+}19]. CASS [WWT^{+}21]. Caste [VMP21]. Catalogs [AWM21]. Catalysts [SKV^{+}21]. Catch [ASJK23]. Categorization [CT18]. Causal [BB18, DCM24, GAW19, LLAK23, SHM24]. causing [RKNES20]. CAVE [GB18]. CCBL [TDC17]. Centered [KRK^{+}21, TDC17]. Centralized [XDC^{+}24]. Centres [WXL^{+}22]. Centric [AFKL20, KKE^{+}21, LAHW19]. cesspool [Gil20]. Chain [CMP^{+}24, WAD^{+}23]. Chains [MHG^{+}21]. Chairs [KFMH17]. Challenge [CWR22, RAZ^{+}20]. Challenges [AHCD17, BK24, CNM^{+}23, CCWH21, CCNY19, DRA^{+}17, FSC21, GSC21, GF22, HK24a, KOOL19, KBFK21, KWdCL^{+}21, LKS21, LLL^{+}22, MMdSP21, MES^{+}22, SGF20, SHY^{+}18, SWF^{+}20, WS22, WWSG21, XYDV23, ZMGK24, dCRM^{+}23, dSdSdSF^{+}23]. Challenging [GRG20, PK23]. Change [AS23, FSRC24, FKK19, GH21, GSLH24, GYD17, KJH19, SRP^{+}24, SSB24b]. Changes [HGG20, ZBL19]. Changing [GSLH24, MKRM23, PBC20, PCL^{+}21]. Channel [ABQ17, MNP^{+}20, NSJ21, RLGG21]. Channeling [LFM22]. Channels [dSdSdSF^{+}23]. Character [WZFA22]. Characteristics [DMC^{+}20, MCM23]. Characterization [DDZ18]. Characterizing [AGO^{+}24, EHZN22, FZM22, HWF^{+}24, HRR^{+}22, MMH21, PM21, SC24, SKV^{+}21]. Charged [CNM^{+}23]. Charity [SS19]. Charter [KM22]. Charting [ESCR23, SFA^{+}23]. Chat [BKJ^{+}20, LLC^{+}24, LYL^{+}23]. Chat-Based [LYL^{+}23]. Chat-bots [BKJ^{+}20]. Chatbot [BKM20, JRP^{+}23, LYH20, WWT^{+}21]. Chatbot-based [BKM20]. Chats [HWYW21]. Chatting [SELS24]. CheckMyFit [YPS^{+}22]. Checks [KKM^{+}22]. CHI [APW^{+}21, FBNM22, GMM21, MDFG24].

Child [MFM21]. Children [RCM21, WXL^{+}22, YWL^{+}23, YGHR21, RSL24]. China [LX20, LAW19]. ChitChatGuide [KKK^{+}24]. Choice [AOT^{+}18, SSGB21]. Choices [Gar21, HWYW21]. Choke [BKM21, JPKM24]. Choreobot [vDBL22]. Chromatin [MMI^{+}22]. Chronic [BLH^{+}17, DB22]. Cine [EHG22]. Cine-AI [EHG22]. Circles [PLL^{+}23]. Circular [RZP^{+}22]. Citizen [HFL19, JØC^{+}20, MM19, MHG^{+}21]. Citizens [SJ21]. City [SBAK21, VAE^{+}22, WNT^{+}22]. City-Building [VAE^{+}22]. Civic [FF22, JKC23, MB23]. Civil [DRA^{+}17]. Clarifying [UYS23]. Class [SCSD24, VN19]. Classes [BPO21, GF22]. Classification [BDSL17, BMR^{+}23, CGM^{+}18, HSC^{+}23]. Classroom [SS23]. Cleaner [BVM21]. Click [LCA^{+}24]. Clickbaits [CSMG17]. Client [KG21]. Climate [GH21, GSLH24]. Clinical [CTH24, KMV^{+}24, TBL^{+}24, JF24]. Clipboard [RRA23]. Close [AGM^{+}23, FB19, ICB^{+}22, KBSP^{+}22, LPW^{+}23]. Close-Knit [FB19]. Closer [LAR21]. Cloud [MV19, OGRV22]. Clouds [AV19]. Clues [AMF18]. Clustering [SPT^{+}20]. Clutch [BKM21]. Co [BUSA^{+}21, BPO21, CSN19, DGR^{+}22, JFL23, KDEA21, KSKH24, LBM21, MFM21, NM22a, OG21, PF24, SAM23, SSNC20, WNT^{+}22, WPH22a, WGM^{+}24, ZMGK24, ZMS^{+}22]. Co-Designing [JFL23, LBM21, WGM^{+}24, BUSA^{+}21, SAM23, ZMS^{+}22]. Co-Editing [CSN19]. Co-Editors [SSNC20]. Co-Ideation [ZMGK24]. Co-learner [BPO21]. Co-Located [DGR^{+}22, WPH22a, KSKH24, NM22a, OG21]. Co-play [MFM21]. Co-Produce [KDEA21]. Co-production [WNT^{+}22]. Co-Worker [PF24]. Coach [CKS18]. Coaching [CKS18]. Coarse [JFS^{+}23]. Cockpit [AARJ23]. Cocode [BPO21]. Code [BSM23, BKJ^{+}20, GCSB23, JK24, KVM24, LPFD21]. Code-mix [BKJ^{+}20]. Codenames [SSR24]. CodeTree [JK24]. Coerciveness [BBV19]. Cognitive [ASM23, CGS^{+}21, CERR23, DWSF24, GSLH24, HvBKG20, IYS23, KFSC18, WGGS20]. Cohabitation [XHH17]. Cohesion [TRN^{+}22]. Coin [KPW19]. Collaborate [ZMW20]. Collaboration [AR22, ALG^{+}21, ALD^{+}20, CCNY19, FZRJ22, GMB23, HCG^{+}22, HHVB22, KSG^{+}24, LFM22, MRS22, OR19, SRAG22, SBPP20, SSR24, SCY^{+}18, WMBO19, WQCZ24, WPH22a]. Collaborations [KG21, MWM^{+}19]. Collaborative [ACA22, BP17a, BDB^{+}23, BLL^{+}21, BRM22, CXL17, DDZ18, FCR^{+}17, FYS^{+}24, FAMS22, KDH^{+}21, KTBL^{+}20, KY24, LAM22, LMB^{+}22, LLAK23, LGG^{+}23, MQX^{+}23, MMdSP21, MAR^{+}19, MRGN20, MDZ19, NAS^{+}22, OGRV22, SWF^{+}20, TBDB21, YDW^{+}22]. Collaboratively [BHNA18, CNH^{+}17]. Collaborators [ZD19]. collecting [CJ23]. Collective [SS21, MNS^{+}21]. Collectives [ERVR^{+}23]. College [GSC21]. Collocated [NSJ21]. Color [WS23]. Color-Blind [WS23]. Coloring [DRAQ18]. Combating [Wil20]. Combining [Hig20, LRP^{+}22, MFSK22]. Comes [HOHB22]. Comfort [AS23]. Command [BDE^{+}24, SB20a]. Command-based [BDE^{+}24]. Commands [PKB^{+}19]. Comment [BGRK17]. Commentator [LUFW20]. Comments [LHPH20]. commerce [CCX^{+}20]. Commercial [ATH24, JSJ22]. Common [HCR22]. Communal [BHM19]. Communicate [ADD21]. Communicating [FMBW24, WHW^{+}22]. Communication [BB20, BLH^{+}17, CYW18, CFR22, ERVR^{+}23, FLC17, FZM22, GMM19, HCR22, HWS21, JV22, KKK22a, LOSD20, LLW^{+}23, LTP23, MWBB23, MM19, MHN^{+}21, NSJ21, OM20, SHL23, SAR21, TRN^{+}22, TSD23, YK23, dSdSdSF^{+}23]. Communications [MM23]. Communities [CILM18, CWB^{+}18, DPA17, ERVR^{+}23, FB19, FD20, PLY23, PSM21, SDF^{+}21, SK20, TKB^{+}22]. Community [Bel24, CXL17, CCS^{+}18, DAQ17, DMC^{+}20, FZM22, FKF^{+}22, GTC22, HMB^{+}21, HTS24, HV23, JS22, JVP^{+}19, JKW^{+}21, KJH19, KKK22b, KJTC22, KG20, LTP23, MHP22, MB22, NPT19, NCM^{+}22, PSTK24, See20, SK23, TKB^{+}22, ULG^{+}22, VQ23, WWT^{+}21, WNT^{+}22, XKRW23, ZD19, WBS^{+}23]. Community-AI [KG20]. Community-Based [HTS24, MHP22, See20, Bel24]. Community-in-the-loop [WNT^{+}22]. Community-led [HV23]. CommunityBots [JRP^{+}23]. CommunityClick [JKW^{+}21]. Companies [SSH^{+}21]. Companion [JJJ^{+}23, SVS^{+}24]. Company [ADD21]. Comparative [RA21a]. compare [CNH^{+}23]. Comparing [KKM^{+}22, MGRM22, MPC^{+}23, VZV19, SAR^{+}24]. Comparison [CWK^{+}23, DR24, GPC^{+}23, HKL^{+}22, KSP^{+}23, ÖBK24, ZWG^{+}23]. Compass [LYL^{+}23]. Compensation [ZLB^{+}22]. Competence [KMV^{+}24, SSB^{+}22]. Competencies [GCSB23]. Competing [DDV23]. Competitive [BKM21, FKF^{+}22, MMdSP21]. Complex [AAHM22, CKP^{+}18, CMY^{+}17, DR24, DWZ^{+}20, ETS^{+}23, HRRS20, KDH^{+}21, MRRS19, ZSK^{+}23]. Complexity [CSS^{+}17, CTWM23, KMR20, SSNC20]. Component [RH19]. Composite [CYGW21, OG21]. Composition [AR22]. Comprehend [YYT^{+}24]. Comprehension [ZRP^{+}22]. Compromise [RJZ^{+}21]. Computation [ERH^{+}24]. Computational [CJ23, GW22, GPC^{+}23, GMB23, HCM^{+}23, KHY^{+}22, WMBO19]. Computer [AM21, Hao21, KTBL^{+}20, LW19, MSY20, NAO23, PLL^{+}23, RSL24, SHD21, TD19, TBL^{+}24].

Computer-Mediated [LW19, PLL^{+}23]. Computers [Ame18]. Computing [AMJ^{+}19, BWV20, DWW^{+}17, FLP^{+}21, HTS24, JVP^{+}19, LPQW22, LS23, OGRV22, SD24, XOEK22]. Concept [BSS18, DFJ^{+}21, FYS^{+}24, LLGH24]. Concepts [HCR22]. Conceptual [MCC^{+}18, RS19]. Conceptualizations [RA21b]. Conceptualize [MHG^{+}21]. Concerned [WFL^{+}22]. Concerns [FZM22, GP22, KBSP^{+}22, LKS21, TSD23, WBS^{+}23, WM22, WSM24, XDC^{+}24]. Condition [CBWD24]. Conditions [BLH^{+}17, JF24, LWC^{+}23]. Conduct [LPFD21]. Conducting [BUSA^{+}21, BV21]. Conductive [KIK20, SM22]. Conductor [ZKL22]. Conferencing [LMT^{+}22]. Configurable [ACF^{+}22, ZMR21]. Configuration [LWB22, TG18]. Configurations [WPH22a]. Configurators [LAD^{+}22]. Conflicts [CT18]. Conformity [FTP20]. confrontation [RKNES20]. Congress [MM19]. Connect [ZD19]. Connected [MCDR22]. Connectedness [XKRW23]. Connection [MBK^{+}22, HFBL19]. Connections [GEEJ23]. Conscious [LPZV23]. Consent [ETS^{+}23]. Consequences [BDSL17, EMS^{+}22, XZL^{+}20]. Considerations [CCNY19, Fie19, FZP^{+}24, VMD^{+}20]. Considering [CAATL22, MWBB23, WGG20]. Consistencies [BL18]. Consistency [SWT20, SSNC20]. Conspiracy [EHZN22, HKD23, PSM21]. Constitutes [NM22b]. Constrained [NMBS21, YTF^{+}23]. Constraints [RIK21]. Construct [BS21]. Constructed [BHNA18]. Construction [FYS^{+}24]. Constructionist [Ame18]. Constructive [KKK22a]. Consultations [YBE^{+}23]. Consumer [GP22]. Consumption [CSMG17]. Contact [KBSP^{+}22].

Contagion [LNM21]. Content [CWB^{+}18, CCW22, DBPA22, ESA^{+}24, FTP20, KKM^{+}22, LLL^{+}22, MMI^{+}22, MAN^{+}20, RGW^{+}24, RJZ^{+}21, RS19, See20, VXHK21, ZGL^{+}22]. Contestability [VXHK21]. Context [GGE23, JJJ^{+}23, KVM24, KMV^{+}24, MH19, MYS^{+}21, MFSK22, SAR21, SMLH23, TDC17, TCH20, TVS17, YHR^{+}19]. Context- [YHR^{+}19]. Context-Aware [MFSK22]. Contexts [DCM24, DHH18, FWY^{+}22, LM20, PBC20, ZWG^{+}23, dCRM^{+}23]. Contextual [LWC^{+}24]. ContextWing [ZWG^{+}23]. Contingent [ALSK23, KKM24]. Continued [KLS21]. Continuum [DFJ^{+}21]. Contribute [DMC^{+}20, GPH21]. Contributor [BZC^{+}17]. Control [AGK^{+}22, AMVK23, BWE^{+}22, CWK^{+}23, HRC^{+}21, HGG20, LZAD21, PGK^{+}24, RMB^{+}22, VP24b]. Controlled [COKL20, GSLH24, YMG^{+}20]. Controller [CDIH23, SORV22]. Controller-Free [CDIH23]. Controllers [KMM21]. Controlling [MP23]. Controls [GSK^{+}17, HPY^{+}22, ZRA22]. Conversation [CMP^{+}21, CBPC23, LKS21, MK23]. Conversational [ACA22, BDE^{+}24, CR20, KKK^{+}24, QGB20, SSB^{+}22, SPT^{+}20]. Conversations [BGRK17, CA22, LPFD21, Poh24, SKV^{+}21, WWSS22]. Converting [LWS24]. ConvEx [CBPC23]. Convolutional [MFSK22]. Cool [BJ21]. Cooperation [Ame18, FR24, JP22, dCRM^{+}23]. Cooperative [ALD^{+}20, BF17, Pri19, SSR24]. Coordinates [RRF20]. Coordinating [ZPMK24]. Coordination [MAS22, Pri19, RSL24, RCM21]. Coping [dQPMdHRdAN21, TSHK20, XYDV23]. Coproducing [JKC23]. COPSE [MTA^{+}17]. Copy [LZY^{+}21]. Copy-and-Paste [LZY^{+}21]. Copyright [FB19]. Cord [BHNA18]. Core [HK24b]. Coronavirus [Wil20]. Corporate [LRSS20]. Correctness [SSNC20]. Corridor [KKV^{+}22]. Corridor-Walker [KKV^{+}22]. Cost [LZW23, MCYZ19, SRKK22, JF24, ZYY23]. Costs [HWS21, LRSS20, MRS22]. COTS [ASZW23]. Couch [SS19]. Could [CBWD24]. Counseling [KG21]. Counted [PVP22]. Counting [PVP22]. Countries [KBSP^{+}22]. Couples [KSKH24]. Courses [XOEK22]. cousin [IRA^{+}19]. COVID [CEJC23, CBV22, HSVR21, JKC23, KJTC22, KUD^{+}23, KBSP^{+}22, PAC22, SLB23, WBS^{+}23, XKRW23, YGHR21, dSdSdSF^{+}23, PKLK22]. COVID-19 [CEJC23, CBV22, HSVR21, JKC23, KJTC22, KUD^{+}23, KBSP^{+}22, PAC22, SLB23, WBS^{+}23, XKRW23, YGHR21, dSdSdSF^{+}23, PKLK22]. Craft [CJ23, Vya19]. Craft-based [Vya19]. Crafting [HC24]. Craftwork [JSKA22]. CRDT [SSNC20]. Create [HCM^{+}23]. Created [CNH^{+}17]. Creating [AL21, JRP^{+}23, KDH^{+}21, KKS19, MP23, WNT^{+}22]. Creation [MCDR22]. Creative [KBFK21, SSB^{+}24a]. Creativity [CILM18, FB19, LFM22, LW19, VAE^{+}22, YWL^{+}23]. Creator [MK23]. Creators [LLL^{+}22, RGW^{+}24, ZMGK24]. Credibility [BZSM20, NM22b]. Creepy [SFLB23, SCSD24]. Crime [WWSS22]. Crises [SHL23]. Crisis [CH20, ESA^{+}24]. Criteria [BZSM20]. Critical [Ame18, BBS19, BKJ22, GMB23, RI20]. Critical-Reflective [GMB23]. cro [WGN^{+}21]. Cropping [HHLC23]. Cross [ASZW23, EMHW23, HKL^{+}22, LE21, PCL^{+}21, RDS18, SY24, SHY^{+}18, VN19]. Cross-Device [EMHW23, RDS18, SHY^{+}18, VN19, LE21]. Cross-disciplinary [PCL^{+}21]. Cross-Domain [ASZW23]. Cross-platform [HKL^{+}22]. Cross-Reality [SY24]. Crossing [FLC20]. Crossmodal [ASMB22]. Crowd [BB18, CILM18, CKS18, CMP^{+}21, GPH21, HBM23, MBT^{+}20, OTH20, TSS21, ZZK^{+}22]. Crowd-Feedback [HBM23]. Crowd-guided [ZZK^{+}22]. CrowdCog [HvBKG20]. CrowdFolio [FKDG21]. Crowdfunding [KKK22b]. CrowdGraph [KZZW22]. CrowdNAS [ZZK^{+}22]. Crowds [FM21]. Crowdsourced [AS17, BZSM20, CSS^{+}17, FGG17, UYS23]. Crowdsourcing [HvBKG20, KZZW22, MMdSP21, PLY23, QGB20, WYY^{+}21, vBGH^{+}19]. CrowdUI [OTH20]. Crowdwork [WMM^{+}19]. Crowdworker [HBM23, SGF20]. Crowdworkers [DHH18, WMM^{+}19]. CSCW [BL17, CPB18, ERT21, WGG20]. CSCW1 [BKL^{+}21, BLS^{+}22, CDG^{+}23, CDG^{+}24]. CSCW2 [KLS20, BLS21, CDGdC23, KRW^{+}22]. CSCW3 [KLS21]. Cues [AMR21, BBW23, BLC^{+}21, KUWM22, LG22, LPW^{+}23]. Cultivating [NCM^{+}22]. Cultural [KSP^{+}23]. Culture [DRA^{+}17, IRA^{+}19, LAM22, MHP22, MKRM23, SCSD24, VKI21, Vya19]. Cultures [LHT^{+}20, MPC^{+}23]. Curation [MRK^{+}22]. Curious [CEM17]. Current [ZMS^{+}22]. Curveball [MMM^{+}21]. Curved [YTF^{+}23]. Curves [GVNK20]. Customizable [MPB^{+}19]. Customization [dACZ19, DBS^{+}21, SPL^{+}23]. Customizations [GMM19]. Customize [dQPMdHRdAN21]. Customized [WvECM22]. Cutscenes [EHG22]. cutting [Hig20]. Cyber [ECP^{+}24, TPNL22]. Cyber-Physical [ECP^{+}24]. Cybersafety [MM23]. Cybersecurity [NHL^{+}24, TSD23]. Cyclist [KMM^{+}22].

D [CPB18, DYF^{+}23, ERH^{+}22, GGL^{+}22, JFS^{+}23, KIK20, KSG^{+}24, LDM^{+}22, OKM^{+}23, OSM^{+}20, SEC^{+}20, Zha23].

D-Printed [GGL^{+}22, KIK20]. D-Touch [JFS^{+}23]. DAIC [HHLC23]. Daily [FM22]. Damage [ZZK^{+}22]. Danger [DNBB^{+}22, LL24]. DANTE [SPT^{+}20]. Dash [JTB22]. Dashboard [SS23, vDBL22]. Dashboards [JTB22, WHW^{+}22]. Data [AL18, BWE^{+}22, CFD^{+}21, CNH^{+}23, CCS^{+}18, DDV23, DCM^{+}23, DB22, EJB^{+}20, FSN^{+}20, Fie19, FCE^{+}18, FSC21, FMBW24, GYD17, GSK^{+}17, HSC^{+}23, HOZ^{+}21, HRRS20, JSKA22, KKE^{+}21, KR24, KG21, LPZV23, LLDH21, LE21, LLKC22, MWM^{+}19, MFM^{+}20, MDZ19, MSY20, MNS^{+}21, MRK^{+}22, MRRS19, NM22b, RWB22, RA21b, SB17, SLT^{+}21, SLMB24, SM22, SHJ^{+}22, SMLH23, TBDB21, TBL^{+}24, VP24b, WMBO19, WS22, YHR^{+}19, ZZ23, ZMW20, ZWG^{+}23, FSC21]. Data-Centric [KKE^{+}21]. Data-Driven [CFD^{+}21, CCS^{+}18, SMLH23, YHR^{+}19]. Datafication [FF22]. Dataset [SHD21]. Datasets [MMI^{+}22, SHD21]. Dating [AZ23, KOL^{+}24, ZMTH22]. Daughters [RYD21]. Death [KMV^{+}24]. Debugging [BYL^{+}24]. Decals [NMBS21]. Deceive [BYRL23]. December [KLS21]. Decentralization [LSLD22]. Decentralized [LSLD22, WS23, XDC^{+}24]. Deceptive [FH24, HSZKN24]. Decision [FMBW24, GSC21, GPH19, GC19, KOOL19, MAS22, MSHP23, MAD^{+}23, NM22b]. Decision-Making [KOOL19, MSHP23, NM22b]. Decisions [AMR21, EJGL24, HJCM22, LMB^{+}22]. Deck [LKG21]. Declarative [DS20]. Decomposed [FKDG21]. Decreasing [HRC^{+}21]. Decrypting [SELS24]. Deep [AH18, CXL17, HSC^{+}23, LYH20, SM22, SPT^{+}20]. Deep-learning [HSC^{+}23]. Deeper [KML^{+}21, VH21]. Defaulting [MK23]. Defender [WS22]. Deficits [HTS24]. Defined [CLwH^{+}20, MLSH21, NM22a, SAY^{+}24, VNSVL22]. Defining [ACF^{+}22]. Definitions [FOH^{+}22].

Deformable [NMBS21]. Degree [LWB22, HKL^{+}22]. Delay [ZSK^{+}23]. Delayed [HKF19]. Delaying [ASJK23]. Delays [HKF19]. Delivery [KB22, SLB23]. Delusio [DMRA24]. Demand [JOFM24, HPL22]. Dementia [KRK^{+}21]. Democracy [MB23, MDZ19]. Democratic [PK23]. Demographically [XOEK22]. Demographically-Labeled [XOEK22]. Demography [SS21]. Demonstrating [MPB^{+}19]. Demystifying [KM22]. Deniability [DMRA24]. Departing [BSS18]. Dependable [CMP^{+}24]. Deployable [AIHS22]. Deployment [RH23]. DePO [LSLD22]. Depression [BRN^{+}19, BRM22, IB24, SWF^{+}20, ZBL19]. Depth [CHS^{+}24, Wil20]. Description [HHLC23, MNP^{+}20, SGF20]. Design [AZ23, AHJV22, BKM20, BAAP20, BWDP20, BDSL17, BDW21, BS21, BP17b, BKJ22, BV21, CLKEF21, CMRH22, CSH^{+}23, CMP^{+}21, CCNY19, DBL22, DMC^{+}20, DPF22, EMT^{+}24, EJGL24, EJB^{+}20, FSRC24, FSSK22, FGG17, GN20, DRDS24, HSZKN24, Hao21, HBM23, HV23, JCD19, KRK^{+}21, KWdCL^{+}21, KSG^{+}24, LLGH24, LRK^{+}24b, LCG24, LDP18, LKG21, LMB^{+}22, MK23, MRGN20, MPEC24, NPT19, OS24, OTH20, PA22, RI20, RJVS24, RIC^{+}24, RDF^{+}22, SHC^{+}23, VZV19, WPH^{+}22b, ZZ23, ZSK^{+}23, ZMGK24, ZP24, vDBL22, LRK^{+}24b]. Designers [SLN21]. Designing [AKL^{+}23, BDB^{+}23, BCCV17, BLV23, CRN21, CWC^{+}24, FKK19, FLP^{+}21, GEEJ23, GHHC21, ICB^{+}22, JFL23, JJJ^{+}23, KKK22b, KKP^{+}22, LYH20, LWB22, LBM21, LGG^{+}23, MAS22, MRK^{+}22, MRRS19, NKV21, PF24, SSB24b, VN19, WQCZ24, WGM^{+}24, WRF^{+}22, YMG^{+}20, ZD19, BUSA^{+}21, SAM23, ZMS^{+}22]. Designs [AFKL20, BLC^{+}21, CYGW21, LAM22]. Desk [LAHW19]. Desk-Centric [LAHW19]. Desktop [DNBB^{+}22]. Detect [ZMTH22]. Detecting [CSA^{+}21]. Detection [DNBB^{+}22, IB24, KSW22, KZZW22, PCMP18, RKFK22, TRO^{+}21, ZBL19]. Developer [ASM23, HWF^{+}24, LLDH21, PF24]. Developers [GM23, LLDH21, MHG^{+}21, PKLK22, PVC^{+}23]. Developing [CSH^{+}23, GKHD23, HBJR21, HBLN24, MPEC24, SS23]. Development [BR24, DRDS24, HHJ^{+}22, HWF^{+}24, HTS24, HZ22, LFM22, PF24, PKLK22, PGK^{+}24, SRZ24, SHD21, SY24, SK20, SHY^{+}18, VP24a, VHZ22, dRSV22, dSdSdSF^{+}23]. Device [EMHW23, FC20, HDPL21, HRP23, HOSK21, LAHW19, RDS18, SLT^{+}21, SHY^{+}18, VN19, WPH22a, LE21]. Devices [AASK23, BR20, CMP^{+}24, DBPA22, JP22, LL24, LWC^{+}23, SCMS18, SAHE23]. DG3 [DS20]. Diagnosis [GHHC21, MT21]. Diagram [VN19]. Diagramming [LE21]. Diary [ABC21]. Did [BZC^{+}17]. didn’t [MK23]. Difference [HZ22]. Differences [BZSM20, DMC^{+}20, SSNC20]. Different [KUD^{+}23]. Differential [FMBW24]. Differentiating [CYC^{+}21]. Difficulties [KR24]. Difficulty [Hig20, LWS24, RGR^{+}21]. Digital [AHCD17, AMR21, AFT^{+}18, BEHB^{+}23, CR20, DBS^{+}21, ECA23, ERVR^{+}23, EMT^{+}24, ECP^{+}24, FF22, FKK19, GN23, Hao21, HFBL19, HBLN24, HGG20, KF23, LO17, LFK24, MFM21, PVP22, RAA22, SRP^{+}24, SS21, TAS^{+}19, WHW^{+}22, ZSOJ20]. Digitalisation [dCRM^{+}23]. Digitization [MAR^{+}19]. Dimension [SLT^{+}21]. Dimensional [UYSM22, YRIV20, UYM21]. Dimensions [MK23]. Direct [GP22, JFS^{+}23]. Direct-to-Consumer [GP22]. Direction [Zha23]. Directional [XMS24]. Directors [EHG22]. Directs [CI22]. Dis [HFBL19]. Disabilities [HCR22, HBJR21, LCG24, Tan21]. Disability [HOHB22]. Disagree [JSJ22].

Disambiguation [BWFG23]. Disaster [BL18, JKC23, MWK^{+}22, PNS24a, ZZK^{+}22]. Disciplinary [SHD21, PCL^{+}21]. Disclose [ZMTH22]. Disclosure [CMSA22, DRA^{+}17, LYH20, MPC^{+}23, MDN^{+}24]. Disclosures [AMF18, ERB^{+}17]. Discomfort [SCSD24]. Disconnects [TBDB21]. Discord [CBPC23]. Discourse [ASS18, GKS21, Hao21, MPKEN24, TAS^{+}19]. Discovering [SHM24]. Discovery [BP17a, HMML^{+}22]. Discriminatory [RGW^{+}24]. Discuss [BSA23]. Discussing [WGS^{+}24, YBE^{+}23]. Discussion [CMY^{+}17, SPL^{+}23]. Discussions [SC24, ULG^{+}22, XZL^{+}20]. Disengagement [AGO^{+}24]. Disinfect [KSF^{+}23]. Disinfection [KSF^{+}23]. Disjunctures [LWTW23]. Disorder [NCM^{+}22, XYDV23]. Disorders [FTP20, TBDB21]. Display [CKSP22, LPED22, WACdA^{+}21, ZMZL21]. Displays [AIHS22, BHM^{+}23, CKP22, CHS^{+}24, ITKB23, PDPS22, RRF20, WLZL23]. Disregard [LL24]. Distance [HHVB22, ICB^{+}22]. Distributed [BKM20, HCG^{+}22, LFM22, MV19, SBPP20]. Distribution [GSK^{+}17, YU20]. Disturb [WWSG21]. Disturbance [TCH20]. Dive [KML^{+}21]. Diverse [KTBL^{+}20, RVH24]. Diversifies [CYW18]. Diversity [dQPMdHRdAN21]. Divides [DDV23]. DIY [HOHB22, SRP^{+}24]. Djnn [MCC^{+}18]. Djnn/Smala [MCC^{+}18]. Do [AZS^{+}22, BKJ^{+}20, GPH19, GP22, RJZ^{+}21, SHD21, SK23, TKB^{+}22, WS23, CNH^{+}23, MHG^{+}21, RH22, ZMW20]. Document [LLGH24, VP24b]. Documents [KM22]. Does [DR24, FKK19, KVK^{+}22, RWL23, SSB^{+}22]. Doesn’t [EMS^{+}22]. DoF [ZLB^{+}22]. Dog [HDPL21]. Dog-to-Human [HDPL21]. Doing [CI22, LRK^{+}24b, VQ23]. Domain [ASZW23, DWZ^{+}20, JSKA22, KBK^{+}22, MWM^{+}19, Rib19]. Domain-independent [DWZ^{+}20]. Domestic [RYD21]. Domestication [SHS22]. Donation [FMBW24, SLMB24]. Donations [SS19]. Don’t [AMR21, CEM17, HSH22, KSF^{+}23, LC24, TSHK20, WWSG21, MM23, EPW^{+}23]. Dota [CBWD24]. Down [IOES20]. Doxing [KSW22]. DR [KM22]. Drafting [LMZ^{+}24]. Drawing [GEEJ23]. Dread [MKRM23]. Drinking [JJJ^{+}23]. Drive [KPW19]. Driven [CFD^{+}21, CCS^{+}18, GBZ^{+}24, JTB22, SMLH23, ZZ23, MMM^{+}21, MRK^{+}22, YHR^{+}19]. Driver [DC17, KMR20, RKFK22, DCM24]. Drivers [CFR22, MH19]. Driving [BDB^{+}23, ITKB23, MMMM22, MCDR22, YTF^{+}23, YML^{+}23, BLV23]. DrivingVibe [YML^{+}23]. Drone [HMB^{+}21]. Dropout [MMH21]. Drowsiness [KMR20]. Dual [DRP^{+}23, NSJ21, YU20]. Dual-channel [NSJ21]. Dual-Sequence [DRP^{+}23]. Dumping [CA22]. Duress [FF22]. During [BBD^{+}24, GSC21, Hig20, Kee19, KJTC22, LWB22, LHPH20, Poh24, SLB23, ULG^{+}22, WBS^{+}23, WSM24, XKRW23, YYT^{+}24, YBE^{+}23, YGHR21, ACA22, DRP^{+}23, ESA^{+}24, HSVR21, JP22, JKC23, PKLK22, WVG^{+}24, dSdSdSF^{+}23]. Duty [CNM^{+}23]. Dwell [DRL^{+}22, IYS23, NMPDJ24]. Dwell-time [IYS23, NMPDJ24]. Dyadic [ÇH24]. Dynamic [BZC^{+}17, CKP22, DCM24, GBN21, HBJR21, LWC^{+}23, NMBS21]. Dynamically [SC23]. DynamicRead [LWC^{+}23]. Dynamics [DWSF24, Kee19, MSY20].

E-commerce [CCX^{+}20]. e-waste [WAD^{+}23]. each [BBK22]. Ear [CLwH^{+}20, JF24, SAY^{+}24, YPS^{+}22]. Ear-Based [CLwH^{+}20]. Ear-Level [SAY^{+}24]. Earbuds [RLL22]. Early [KYB21, MBK^{+}22, SS22, SLD21]. EarMonitor [JF24]. easier [RKNES20]. Eating [FTP20, NCM^{+}22, PBC20, TBDB21]. Echoes [ND24]. Ecological [BahRKM18, KBKH21]. Ecology [CAATL22]. Economy [DWW^{+}17, FR24, NKV21]. Ecosystem [DWB18]. Ecosystems [GMM19]. Edge [OGRV22, UYS23]. EdgeXAR [ZLB^{+}22]. Edit [MD20, PVC^{+}23]. Edit-a-thons [MD20]. Editing [CXL17, CSN19, DDZ18, RRA23]. Editor [KFMH17]. Editorial [ACI22, BKL^{+}21, BLS21, BLS^{+}22, BMdS^{+}21, BMdS^{+}22, CNC^{+}17, CM20, CDG^{+}23, CDGdC23, CDG^{+}24, DKBG23, DKKL24, FJIH21, FBNM22, GKKW22, GMM21, KLS20, KLS21, KvBW24, KRW^{+}22, LMS23, LPQW22, LS23, MDFG24, MBTW24, PP23, SD24]. Editors [CT18, SSNC20, WEM24, EVC^{+}19, KMHLF18, LGS19]. Education [AMJ^{+}19, CCWH21, LKS21, SS22, SS23, TSD23, ZMS^{+}22]. Educational [FYS^{+}24, HKL^{+}22, KRMM21]. Educators [AL21, PM21]. EEG [NAO23]. EEG-Based [NAO23]. Effect [AMR21, ERLW17, GRG20, MRGN20, OM20, UYS23, WPH22a, dCRM^{+}23]. Effectiveness [BHSP22, CSL^{+}24, CCW22, YUSM22]. Effects [ALG^{+}21, ASP20, AOT^{+}18, ACA22, BBNT21, BMR^{+}23, CSS^{+}17, CFR22, CERR23, FWY^{+}22, GAW19, HSH22, HWYW21, KRMM21, KYB21, LG22, LWB22, LPED22, MCYZ19, MHN^{+}21, STSS22, UHH23, VAE^{+}22, Zha23, ZRP^{+}22, AZS^{+}22]. Efficacy [CPS^{+}17, MAD^{+}23]. Efficient [BB18, SBPP20, SBSM24, VP24b]. Effort [ASM23, GBN21]. Egyptian [SELS24]. EICS [JVP^{+}19]. Elderly [PCMP18, SVS^{+}24]. Electrodermal [CTWM23]. Electroencephalography [KFSC18]. Electromyography [KTP^{+}22]. Electronics [FSSK22]. ElectronicsAR [FSSK22]. Elements [GBN21, NMBS21, SAR^{+}24]. Eleven [HHJ^{+}22]. Eleven-Item [HHJ^{+}22]. Elicit [CMSA22, SLMB24]. Elicitation [CWK^{+}23, JRP^{+}23, MV19, WO20, XJBH22]. Eliciting [ITKB23, NM22a, SB20a]. elite [GN20]. EllSeg [KBK^{+}22]. EllSeg-Gen [KBK^{+}22]. Embedded [BYL^{+}24]. Embedding [AKL^{+}23, JWK22]. Embeddings [ASMB22, SBB24]. Embodied [FA22, LMB^{+}22, WRF^{+}22, ZES^{+}23]. Embodiment [FAMS22, SRAG22]. Embody [SLT^{+}21, KKE^{+}21]. Emergent [HK24a]. Emerging [HMB^{+}21, VP24a, WBS^{+}23, dCRM^{+}23]. EMG [KKE^{+}21]. EMG-Based [KKE^{+}21]. emoji [SHL23]. Emoticontrol [AMVK23]. Emotion [ASMB22, BKM20, DTE^{+}22, GKS21, HC24, KA23, RA21b, SVS^{+}24, YYT^{+}24]. Emotional [CCW22, GBN21, LHPH20, WWH^{+}23]. Emotions [AMVK23, ACDC23, FCE^{+}18, GBN21, HPH^{+}23]. Emotions-based [AMVK23]. Empathosphere [KKK22a]. Empathy [WB21, ZAS^{+}21]. Empirical [ALG^{+}21, CWK^{+}23, CSJ^{+}18, JCD19, KML^{+}21, LAD^{+}22, PSTK24, ZP24]. Employee [AKL^{+}23]. Employees [LRSS20]. Empower [MNS^{+}21]. Empowerment [LLC^{+}24]. Emulation [CLKEF21]. Enable [HKM^{+}20]. Enabled [RA21b, SWF^{+}20, GHHC21]. Enablers [RYCC21]. Enabling [AIHS22, BP23, BR20, FLC20, HCG^{+}22, SAR^{+}23, SHC^{+}23, SPL^{+}23, TFI^{+}22, WACdA^{+}21, YWL^{+}23]. Enacting [DCM^{+}23]. Encumbered [RZP^{+}22]. End [ACF^{+}22, BYRL23, BR20, BahRKM18, DRDS24, LFM22, dQPMdHRdAN21, VP24a, dRSV22]. End-User [BR20, BahRKM18, DRDS24, LFM22, VP24a, dRSV22]. End-Users [BYRL23, dQPMdHRdAN21]. Ended [BP17a]. Endoscope [JF24]. Endoscopic [MON^{+}21]. Enforcement [FB19]. Enforcing [WWSS22]. Engage [EHZN22, KDEA21, MK23]. Engagement [AKIS21, DAQ17, DC17, FF22, FM22, GH21, GBN21, JØC^{+}20, JJJ^{+}23, LMB^{+}22, MMH21, PSM21, RGR^{+}21, SS23]. Engaging [FCE^{+}18, FAMS22, GTC22]. Engine [ERLW17, VH21]. Engineer [CMP^{+}24]. Engineering [ASJK23, AGM^{+}23, BHH^{+}22, BWV20, BMR^{+}23, CMN^{+}21, FLP^{+}21, JVP^{+}19, LLGH24, LPQW22, LS23, MdOKT23, MNP^{+}22, OGRV22, SD24, SVV21, SORV22, WPH^{+}22b]. Engineers [JSJ22]. English [LHT^{+}20]. Enhanced [FKF^{+}22, VWHW24, ASMB22]. Enhancers [FKF^{+}22]. Enhancing [KY24, LLC^{+}24, LAR21, SRAG22, YML^{+}23, ZSOJ20]. Enjoy [JJJ^{+}23, CCWH21]. Enough [BYRL23, CNH^{+}17]. Enriching [SLMB24]. Enrichment [GKHD23, SHJ^{+}22]. Ensuring [PNS^{+}24b, SWT20]. Entering [GSK^{+}17]. Enterprise [CWB^{+}18]. Entertainment [GH21]. Entity [RH19]. Entity-Component-System [RH19]. Entrepreneurs [DDFK19]. Entry [ASP20, LPGG22]. Environment [ECP^{+}24, HRP23, SCY^{+}18, YOFC19]. Environmental [RH23]. Environments [GB18, LAHW19, LPED22, NFTK22, SVW19, SBPP20, TD19, VWHW24, VAE^{+}22, WM22]. Envision [LCG24]. Envisioned [SHJ^{+}22]. Envisioning [PF24]. Ephemeral [EJB^{+}20]. Ephemerality [MCYZ19]. Epilepsy [MT21]. Episodic [KSX21]. Epistemic [AS17, LLGH24]. Equals [AGK^{+}22]. Equity [EHH^{+}18, XOEK22]. Era [CSMG17, FF22]. Ergonomics [AFT^{+}18]. Error [YK23]. Errors [ASP20, CLL^{+}21, PLZ^{+}22]. Escalator [AKIS21]. Escaping [STSS22]. Esports [CBWD24, FOH^{+}22, FKF^{+}22, KHY^{+}22, LUFW20, XRS22]. Essential [LAD^{+}22]. Estimating [QGB20].

Estimation

[DYF^{+}23, ERH^{+}24, KCC20, KNT22].

Ethical

[BUSA^{+}21, BS21, Fie19, FZP^{+}24, ZMGK24]. Ethics [LRK^{+}24b]. Ethnographic [BEHB^{+}23, KB22, SLB23]. Ethnographically [MRK^{+}22]. Ethnographically-driven [MRK^{+}22]. ETRA [DKBG23, DKKL24, GKKW22]. EU [dACZ19]. Evaluate [LHI20]. Evaluating [BBSV23, DHH18, JRP^{+}23, KMM21, MSHP23, RAZ^{+}20, SSD23, STSS22, WGGS20, YTF^{+}23, YOFC19, SSB^{+}22]. Evaluation [AFT^{+}18, ALSK23, BLC^{+}21, CHCG24, CWR22, DRL^{+}22, FSSK22, GN23, HCR22, Hig20, KVM24, PA22, PGK^{+}24, RA21a, SRZ24, SY24, WvECM22, ZSK^{+}23]. Evaluations [DRK18]. Event [CH20, CHC^{+}17]. Event-based [CHC^{+}17]. Events [LHPH20, PKB^{+}19]. Ever [MNS^{+}21]. every [BLH^{+}17]. Everyday [BB20, FAMS22, XHH17]. Everyone [SM22]. Everything [KHY^{+}22, TKB^{+}22, BBK22]. Everywhere [PKLK22]. Evidence [LMB^{+}22, LHPH20, RJZ^{+}21]. Evidence-Based [LMB^{+}22]. evil [BTM^{+}21]. Evolution [CNH^{+}17, CAATL22, SFA^{+}23]. Evolving [MT21, ZWG^{+}23]. Examination [KMV^{+}24, VWHW24]. Examined [CPS^{+}17]. Examining [ABC21, ASS18, BBD^{+}24, HC24, SFA^{+}23]. Example [FSAN23, KMR20]. Example-based [FSAN23]. Examples [JK24]. Exchange [RDS18, WB21]. Execution [WYY^{+}21]. Exercise [GF22, XMS24]. Exercises [AGM^{+}23]. Exercising [HFL19]. Exergame [BRM21, KRK^{+}21]. Exergames [ASZW23, CTH24]. Exertion [HPH^{+}23]. Exhibits [LBM21, LMB^{+}22]. Existing [LFK24, MHP22, TRO^{+}21, MKRM23]. Exiting [AGO^{+}24]. EXpanded [LMB^{+}22, CSMV21]. Expanding [YRIV20]. Expectations [LAD^{+}22, SSH^{+}21]. Experience [ALG^{+}21, AGO^{+}24, AZS^{+}22, BVM21, CWC^{+}24, CMSA22, CBV22, GBN21, GM23, HHJ^{+}22, JNLtB22, JOFM24, KVM24, LAR21, MBS22, NSV22, ND24, POG^{+}21, SFLB23, SC23, Tan21, UHH23, YML^{+}23, ZRP^{+}22]. Experiences [ATH24, AMF18, CSN19, DMPK22, HBJR21, JS22, JSP^{+}21, PKLK22, PSTK24, WRF^{+}22, XJBH22, YGHR21, KVK^{+}22]. Experiencing [LAW19]. Experiment [GSLH24, JCD19, OR19, LRK^{+}24a]. Experimental [SSD23]. Experimentation [KKE^{+}21]. Experiments [AOT^{+}18, BV21, MMI^{+}22, UYS23]. Expert [BZSM20, FGG17, NHH20]. Expertise [ASM23]. Experts [JSKA22, MWM^{+}19]. Explain [PVC^{+}23]. Explainable [ESCR23, GLZ^{+}21, KZZW22]. Explaining [RWL23]. Explains [JPKM24]. Explanation [ETS^{+}23, KG20, MSHP23]. Explanations [GLZ^{+}21, JFL23]. Exploiting [DS20]. Exploration [BB18, CBWD24, CBPC23, GGE23, HMML^{+}22, LG22, LLL^{+}22, MCDR22, RVBC^{+}21, RDS18, SHC^{+}23, TBHM23, WAD^{+}23, WGM21, ZSOJ20]. Exploratory [CCX^{+}20, PVC^{+}23]. Explore [GGF^{+}20, KKK^{+}24]. Exploring [BKM21, BLC^{+}21, BAAP20, BMR^{+}23, CLKEF21, CSL^{+}24, CLwH^{+}20, CTH24, CBV22, DWSF24, ERVR^{+}23, EJB^{+}20, GMB23, HTTH22, HMK24, HJC^{+}22, IYS23, JZB^{+}21, JNLtB22, JOFM24, JSP^{+}21, KJH19, KOL^{+}24, LLC^{+}24, LG22, LLW^{+}23, LWC^{+}23, LMZ^{+}24, LX20, LPGG22, MCM23, MDM^{+}19, MMI^{+}22, MMMM22, NSJ21, NHL^{+}24, NSV22, OS24, RDF^{+}22, SAY^{+}24, SWQ^{+}23, SAHE23, UHH23, WS22, WVG^{+}24, WSM24, WGG20, XZL^{+}20, XJBH22, YWL^{+}23, SCSD24]. Exposure [DBS^{+}21]. Expression [GMM19]. Expressions [GBN21, SB20a]. Extend [KIK20]. Extended [BF17, ECP^{+}24, VMHO^{+}22]. Extending [dACZ19, WEM24]. Extensible [BHSP22, SS22]. External [CFR22, FRH^{+}22, HCR22, ITKB23]. Extract [NM21]. Extraction [HMB^{+}21]. Extrafamilial [LTP23]. Extreme [JKZ^{+}22]. Extremist [PM21]. Eye [ASM23, AASK23, BSM23, BK24, FH24, HK24a, HNS^{+}24, HSG^{+}22, JV22, JDX24, KCC20, KR24, LRP^{+}22, MSK22, MFSK22, ÖBK24, PA22, RSL24, SC23, VP24b, WGN^{+}21, WGN^{+}21]. Eye-Gaze [HSG^{+}22, PA22, WGN^{+}21]. Eye-hand [RSL24]. Eye-Head [HNS^{+}24]. Eye-Tracking [JV22, MSK22]. Eyeblinks [ÇH24]. Eyes [AH18, WWH^{+}23]. Eyetracking [KBK^{+}22]. Eyettention [DRP^{+}23].

Fabric [VMD^{+}20]. Fabricated [HRC^{+}21]. Fabrication [Ann20, SHC^{+}23]. Face [CNM^{+}23, DMRA24, KKP^{+}22, SFA^{+}23]. Facebook [BHM19, DWB18, HPY^{+}22, SSGB21]. FacePsy [IB24]. faceted [KOL^{+}24]. Facial [GBN21, IB24]. Facilitated [CYW18]. Facilitating [AR22, CMY^{+}17, HHVB22]. Facilitation [YGHR21]. Facilities [CRM^{+}24]. Factor [BLB^{+}24]. Factors [AFT^{+}18, LWC^{+}24, NHL^{+}24, PSM21]. Factory [ECP^{+}24]. Faculty [LKS21]. Failure [HTTH22]. Fair [vBGH^{+}19]. Fairness [BKG^{+}21, BDB^{+}23, MES^{+}22]. Fake [RWL23]. Fakey [MAMP21]. Fall [MM23]. Familie [Gar21, TPNL22]. Family [AGK^{+}22, IRA^{+}19, JA21, MFM21, PBC20]. fancy [ZPMK24]. Fandom [FD20]. Farm [HMB^{+}21]. Fast [BP23, EPW^{+}23]. Fast-Paced [EPW^{+}23]. Fatherhood [JAB19]. Fatigue [KMR20]. Fauxtography [KZZW22]. Fear [DBS^{+}21]. Feasibility [HSG^{+}22]. Feature [BDL21, BMR^{+}23]. Features [ASP20, BF17, MDM^{+}19].

Federated [ERH^{+}24]. Feedback [AARJ23, BR20, BBW21, CMP^{+}21, DC17, DRL^{+}22, FLC17, FGG17, FKDG21, HBM23, HOSK21, HTS^{+}22, JKW^{+}21, KBFK21, KYB21, MSK22, Poh24, RLGG21, RWB22, TRO^{+}21, WB21, XOEK22, YHR^{+}19, YOFC19, YML^{+}23]. Feedforward [ALRS24, CLV^{+}19]. Feeling [RMB^{+}22]. Feminism [DCM^{+}23]. Feminist [RI20]. Feminization [HPL22]. Fertility [FCR^{+}17, FCE^{+}18, FSC21]. Fertilization [KA23]. Few [AH18, CST^{+}23]. Fiction [LCG24]. Fidelity [FSSK22, SLN21]. Fiducial [SM22]. Field [KKM24, KVK^{+}22, MBS22, VP24a, LRK^{+}24a]. Fighting [WWSS22]. File [WS22]. Final [CWR22]. Financial [PKC^{+}21]. Find [BKJ^{+}20, MWM^{+}19]. Finding [CCH^{+}18]. Findings [WGGS20]. Fine [DWZ^{+}20, NB20]. Fine-grained [DWZ^{+}20, NB20]. Finger [DYF^{+}23, KMH23, RLGG21, YRIV20]. Fingerprint [DYF^{+}23]. Fingers [BR20]. Finland [PKLK22]. Finsta [HV22]. Fire [NHL^{+}24]. First [BMdS^{+}21, GGE23]. Fitness [GF22]. Fixed [JCD19]. Fixed-Gaze [JCD19]. Fixes [GGF^{+}20]. Flamers [PM21]. Flecto [KVDG20]. Flex [LHI20]. Flex-ER [LHI20]. Flexible [BGM^{+}23, Blo24, BMPP20, KVM23, SCMS18]. Flicks [YRIV20]. Flight [HGG20]. Flipping [JOFM24]. Floating [RRA23]. Flood [SLD21]. Floor [EPW^{+}23]. Flow [GW24, HSH22]. Fluid [AKVHZ23, LE21]. Focus [DWSF24]. Focused [RJVS24]. Foldable [KVDG20, LDM^{+}22]. Folk [KDEA21, WZT^{+}24]. Fomites [Wil20]. Food [KB22, PBC20, SLB23, SAHE23]. FoodChoices [RDF^{+}22]. Force [BR20, HTS^{+}22, RA21a, TFI^{+}22]. Forces [HOSK21]. Foregrounding [AZ23]. Forget [KSF^{+}23]. Form [UHH23]. Formal [CMN^{+}21, PB22]. Formation [FB19].

Formats [WS22]. Formatting [RRA23]. Former [FZM22]. Forming [HDPL21]. Forms [UAAR20, JØC^{+}20]. FortClash [CVL22]. FORTE [CST^{+}23]. Fortunettes [CLV^{+}19]. Forum [LLDH21]. Forums [BTM^{+}21, SPL^{+}23]. Foster [GYD17, FZM22]. Fostering [EPW^{+}23, FZM22, ITKB23]. Foundations [WGM^{+}24]. Four [DRL^{+}22, JSJ22, KBSP^{+}22, RA21a]. FPS [SRZ24]. Fragmentation [WMM^{+}19]. Frame [CMP^{+}24]. Frame-based [CMP^{+}24]. Frames [KJH19, OKM^{+}23]. Framework [ACDC23, CGM^{+}18, ESCR23, HWS21, HHLC23, HRRS20, HK24b, JV22, JTB22, KE18, MCC^{+}18, OG21, PGK^{+}24, PSTK24, RH23, RI20, RDS18, RS19, SCMS18, SORV22, TRO^{+}21, WPH^{+}22b, ZLB^{+}22, vDBL22]. Framing [AOT^{+}18, DRK18, KG21, WB21]. Franca [CYW18]. Frappé [RH23]. Free [CDIH23, GBZ^{+}24, HSZKN24, HNS^{+}24]. Free-to-Play [HSZKN24]. Freedom [HPL22]. Freeform [NMBS21]. Freehand [LPGG22]. Freelancers [BKKD21, FR24, LRSS20]. Freelancing [MRS22]. Freemium [EMS^{+}22]. Freezing [JDX24]. Friction [AIRH24, SB20b]. Frictions [BKJ22]. friendly [ZLB^{+}22]. Friends [CEM17, SK20]. Friendscope [NSV22]. Front [TFI^{+}22]. Frontline [SFA^{+}23, SMLH23]. Frustration [HRC^{+}21, SC23]. Full [GB18]. Full-Body [GB18]. Fun [BZC^{+}17, KTP^{+}22]. Function [LW19]. Functional [KKM24]. Fungible [XDC^{+}24]. Future [BDB^{+}23, BWV20, BDW21, CLV^{+}19, GSC21, JOFM24, LCG24, SHJ^{+}22, BL17]. Futures [KF23].

G [CS18a, HOSK21, HHLC23]. G-DAIC [HHLC23]. G-Forces [HOSK21]. G-Gene [CS18a]. Gain [BYRL23]. Gains [WBSM24]. Game [AL21, ALD^{+}20, BP17a, BR24, CGS^{+}21, DBS^{+}21, EHG22, FcLDM19, FRH^{+}22, GH21, GSLH24, GBN21, HSZKN24, HTTH22, Hao21, HMK24, HBLN24, HZ22, KML^{+}21, LOSD20, LFM22, MAMP21, MBK^{+}22, MFM21, MPEC24, PF24, PKLK22, PPT^{+}22, RI20, RVH24, RDF^{+}22, RGR^{+}21, SBAK21, SRZ24, SAR^{+}24, SSB24b, WGS^{+}24, XRS22, ZMGK24, RMB^{+}22]. Game-Based [GSLH24, HMK24, HBLN24, WGS^{+}24, DBS^{+}21, GH21]. GameBus [GN23]. Gameful [LNM21]. Gamemasters [AL21]. Gameplay [CSA^{+}21, HK24b, LL24, YYT^{+}24]. Gamers [TRO^{+}21, ZRL21]. Games [AGO^{+}24, BKM21, CKK22, CI22, DFJ^{+}21, FM22, GH21, GHHS21, HSH22, HTTH22, Hao21, HRC^{+}21, ICB^{+}22, KRMM21, KMM21, KBKH21, KML^{+}21, LG22, MHG^{+}21, ND24, RVBC^{+}21, SSR24, STSS22, TRO^{+}21, VKI21, WZT^{+}24, WRM^{+}24, WXL^{+}22, ZMGK24]. Gamification [AZS^{+}22, LRK^{+}24a, LNM21]. Gamified [SAR^{+}24]. Gaming [BJ21, FKF^{+}22, LCG24, MFM21, NM22a, XJBH22, ZRL21]. Gap [ESCR23]. Gaps [BDW21, DMC^{+}20, HK24a]. Garments [GBZ^{+}24]. Gastroenterological [MON^{+}21]. Gather [SS19]. Gaussian [YU20]. Gay [DWB18]. Gaze [ASMB22, ALSK23, BBD^{+}24, ERH^{+}24, HSC^{+}23, HHLC23, HSG^{+}22, JCD19, KCC20, KNT22, KKM24, LG22, LKYK24, LWC^{+}23, LPGG22, LRP^{+}22, MON^{+}21, NFTK22, PA22, PLZ^{+}22, RSL24, SWQ^{+}23, SBSM24, SBB24, WJH^{+}24, WvECM22, WGN^{+}21]. Gaze-assisted [SWQ^{+}23]. Gaze-based [LG22, SBSM24]. Gaze-Contingent [KKM24]. Gaze-enhanced [ASMB22]. Gaze-Hand [LRP^{+}22]. Gaze-sensitive [RSL24]. GazeCues [LG22]. GazeIntent [NMPDJ24]. GazeSwitch [HNS^{+}24]. Gboard [SR22]. Gelicit [MV19]. Gen [KBK^{+}22]. Gender [BKG^{+}21, BSA23, DMC^{+}20, EHH^{+}18, MWH19, MGG^{+}21, SS21]. Gender-Imbalanced [MWH19]. Gene [CS18a, CS18a]. General [CAATL22, SHM24]. Generalization [KBK^{+}22]. Generated [LL24, ZGL^{+}22, ZMGK24]. Generating [BR17, DCM24, EHG22, SSGB20]. Generation [CFG^{+}17, DS20, JTB22, KVM24, RS19, SDO21, WvECM22]. Generative [PF24]. Generic [CPB18, WPH^{+}22b]. Genetic [GP22, Sal17, SDO21]. Genre [AWM21]. Genres [SRZ24]. genuine [MK23]. Geo [WYY^{+}21]. Geo-Social [WYY^{+}21]. Geodetic [MPEC24]. Geometric [VP24b]. German [HBLN24]. GesFabri [JNLtB22]. GeShort [RRA23]. Gestatten [KCC20]. Gesto [PKB^{+}19]. Gestural [RRA23, SORV22]. Gesture [ASZW23, AFT^{+}18, CWK^{+}23, DS20, DWZ^{+}20, EMT^{+}24, GBZ^{+}24, GVNK20, HG19, IB24, JNLtB22, JFS^{+}23, KCC20, MV19, MRV22, RLL22, WO20, YMG^{+}20]. Gesture-based [JNLtB22]. Gesture-Driven [GBZ^{+}24]. GestureCards [HG19]. Gestures [CS18a, CLwH^{+}20, CST^{+}23, FLC20, LW19, MLSH21, NM22a, OSM^{+}20, PKB^{+}19, SAY^{+}24, VNSVL22]. Get [TSHK20]. gets [HV22]. ggViz [XRS22]. Ghana [WAD^{+}23]. Ghost [SELS24]. Gig [FR24]. Giraffe [GKHD23]. Girls [KBKH21]. Github [LPFD21]. Giving [SSB^{+}22]. Glasses [DBPA22, MSK22, MFSK22, NSV22]. Glimpse [BWV20]. Gluing [WPH^{+}22b]. Go [CYC^{+}21, GPH19, ITKB23, IRA^{+}19, LRK^{+}24a]. Goal [AAHM22, KYB21, OS24]. Goals [DRDS24]. Going [LAW19, XKRW23]. Golden [HKD23]. Golfers [WDP^{+}20]. Good [CBSQ24, LRK^{+}24b, RMB^{+}22, BLH^{+}17]. Gooey [PPT^{+}22]. Got [BVM21]. Governance [SDF^{+}21, WS23, ZMGK24].

Grace [ZBV23]. Gradual [LSLD22, ZRL21]. grained [DWZ^{+}20, NB20]. Grandchildren [LAM22]. Grandparents [LAM22]. Graph [KZZW22]. Graphic [DMC^{+}20, ZMS^{+}22]. Graphic-Based [ZMS^{+}22]. Graphical [KVDG20, NMBS21, PB22, SDO21, SVV21]. Graphics [ZSOJ20]. Grassroots [AL18, NPT19]. Gratifications [LX20]. Great [GF22]. Green [KKM^{+}22]. Grid [KIK20, SDO21]. Grid-based [SDO21]. Grief [BHM19]. Grips [ZP24]. Ground [KMR20]. Grounding [CR20, HWS21]. Group [AS23, AHJV22, DPA17, DRK18, GF22, HJCM22, HBLN24, KWdCL^{+}21, LYL^{+}23, LW19, PP23, RIC^{+}24, VKI21, ZBL19, ZPMK24, dCRM^{+}23, BMdS^{+}21, BMdS^{+}22, MBTW24]. Group-Based [HBLN24]. Grouping [SBSM24]. Groups [HSVR21, LLC^{+}24, LYD^{+}23, MWH19, MAR^{+}19, YGHR21]. Groupware [MRRS19]. Growth [CKS18]. Guess [TPNL22]. GUI [BBSV23, BV21, CLV^{+}19, YTF^{+}23]. GUI-Based [YTF^{+}23]. Guidance [JP22, XMS24]. Guide [JVP^{+}19]. Guided [HRRS20, WGM21, WRF^{+}22, ZZK^{+}22]. Guidelines [CSH^{+}23, RAZ^{+}20]. Guiding [PSTK24, WGM21]. Gut [PPT^{+}22].

Habits [SELS24]. Hackathons [NHH20]. Hackers [Ame18]. Halls [JKW^{+}21]. Hand [CSMV21, CST^{+}23, DNBB^{+}22, FC20, KSF^{+}23, LL24, LO17, LW19, LRP^{+}22, SWQ^{+}23, WVG^{+}24, ZRA22, ZSOJ20, ZP24, RSL24]. Hand-based [LO17, SWQ^{+}23]. Hand-held [LL24]. Handed [FC20, RRA23]. Handheld [AASK23, HRP23, LWC^{+}23, LGG^{+}23]. Handling [AIRH24, Blo24, GVNK20]. Hands [GBZ^{+}24, HNS^{+}24, LMT^{+}22]. Hands-Free [GBZ^{+}24, HNS^{+}24]. Hands-On [LMT^{+}22]. Handwritten [ZWZ^{+}21]. HandyGaze [NFTK22]. Happening [KUWM22]. Happiness [EMS^{+}22]. Haptic [DRL^{+}22, HOSK21, YMG^{+}20]. Haptic2FA [BLB^{+}24]. Haptics [BLB^{+}24, GGL^{+}22, VMHO^{+}22, VWHW24]. Haptics-Based [BLB^{+}24]. HAR [MFSK22]. Harassers [SSD23]. Harassment [BDSL17, SSH^{+}21, SFLB23]. Hardware [BYL^{+}24]. Harm [HHVB22]. Harmful [MBT^{+}20, ZMGK24]. Hate [CPS^{+}17, MIS^{+}20]. Having [LWS24]. HCEye [DWSF24]. HCI [CM20, ERT21, NKV21, PCL^{+}21, WGG20]. Head [GBZ^{+}24, GGE23, HOSK21, HNS^{+}24, IB24, JP22, JCD19, KBK^{+}22, MFSK22, YML^{+}23]. Head-Mounted [GGE23, KBK^{+}22]. Head-Worn [HOSK21]. Health [AAHM22, ASO^{+}22, BLL^{+}21, BKJ22, CRM^{+}24, DB22, EMS^{+}22, ETS^{+}23, FCE^{+}18, GPC^{+}23, GN23, JF24, LYH20, MYS^{+}21, MHP22, MBK^{+}22, OS24, RCM21, SAM23, SHL23, ULG^{+}22, WWT^{+}21, XYDV23]. Healthcare [BLH^{+}17, CMP^{+}24, FSC21, SLMB24]. Hearables [SAY^{+}24]. Hearing [LLL^{+}22, YPS^{+}22]. Heart [MHP22, SHL23]. HeartMob [BDSL17]. Heavy [ZRL21]. Heck [LC24]. held [LL24]. Helmet [HOSK21]. Help [BRM22, ETE^{+}19]. Helpful [SGF20]. Here [CPS^{+}17]. Heritage [KSP^{+}23, MGRM22]. Hero [LMZ^{+}24]. Heterogeneous [CSN19, HvBKG20]. Hexa [BLC^{+}21]. Hexa-metric [BLC^{+}21]. Hexad [AZS^{+}22]. Hexagon [JPKM24]. Hexed [JPKM24]. Hidden [CSJ^{+}18, KUWM22]. Hide [HWYW21]. Hiding [UHH23]. Hierarchical [VP24b]. Hierarchies [JK24, VMP21]. High [CYC^{+}21, FSSK22, MMI^{+}22]. High-Content [MMI^{+}22]. High-Fidelity [FSSK22]. Higher [SS23]. Higher-Education [SS23]. Highlighting [DWSF24]. Highly [CWR22, CERR23, FCR^{+}17, MSK22, SCR22]. HIIT [EPW^{+}23]. Hiking [HPH^{+}23]. hint [LAM22]. Hiring [LRSS20]. Historical [BDW21, GMB23, MGRM22, MPEC24, Gil20]. History [Ame18, MHP22]. HistoryViewer [SB17]. HMD [HTS^{+}22, WACdA^{+}21]. HMDs [LWB22]. Hoaxes [AH18]. Hoc [TRN^{+}22, FX22, KKK22a, SCY^{+}18]. Holding [DNBB^{+}22]. Hole [TBHM23]. Holistic [FKDG21, LE21]. HoloMentor [SAR^{+}23]. Home [CNH^{+}23, CWK^{+}23, CBV22, CVL22, LLC^{+}24, LPZV23, PLL^{+}23, TDC17, TPNL22, WHW^{+}22]. Homepreneurship [Ann20]. Homes [MP23, SAE^{+}21]. Honor [CSA^{+}21]. Hope [SGF20]. Horizontal [CKP22, PDPS22]. Hospital [CAATL22, HKM^{+}20]. Hospitality [SFA^{+}23]. Housed [GKHD23]. Households [WGG20]. Hum [GKHD23]. Hum-ble [GKHD23]. Human [AFT^{+}18, ALD^{+}20, BSM23, BBV19, BLV23, CGM^{+}18, CLP^{+}21, DRP^{+}23, EHG22, FZP^{+}24, GYD17, GMB23, HDPL21, HOSK21, IYS23, KRK^{+}21, KTBL^{+}20, KSG^{+}24, MFSK22, MMM^{+}21, MSHP23, MAD^{+}23, POG^{+}21, PPT^{+}22, RSL24, SFM^{+}22, SBB24, TD19, VP24a, YK23, GHHS21, SSR24, SHJ^{+}22]. Human-Agent [SFM^{+}22]. Human-AI [ALD^{+}20, GMB23, KSG^{+}24, MSHP23, MAD^{+}23, GHHS21, SSR24, SHJ^{+}22]. Human-Centered [KRK^{+}21]. Human-Computer [KTBL^{+}20, TD19]. Human-Machine [CGM^{+}18]. Human-Microbial [PPT^{+}22]. Human-Robot [VP24a]. Human-side [CLP^{+}21]. Humanitarian [DC17, ECA23]. Humanities [LKS21]. Hurry [HKF19]. Hybrid [HG19, NSJ21, ZRP^{+}22]. HybridPointing [DWMV21]. Hyper [DRA^{+}17]. Hyper-Selective [DRA^{+}17]. Hypotheses [AV17].

ICT [ETE^{+}19, TAS^{+}19]. ICTs [DRA^{+}17]. IDE [SB20a]. Ideation [KTBL^{+}20, ZMGK24]. Identification [SJ21, ZRP^{+}22]. Identifying [DPA17, HPY^{+}22, KFSC18]. Identities [SS21]. Identity [DWB18, FA22, KDEA21]. Ideological [KPW19]. ideologies [GN20]. if [ASJK23]. Ill [PF24]. Ill-Informed [PF24]. Illegitimate [JDX24]. illness [DB22]. Illumination [CKSP22]. Illusions [SHM24]. I’m [PF24, WFL^{+}22]. Image [AZS^{+}22, BKG^{+}21, BMR^{+}23, CLP^{+}21, GMB23, HHLC23, LSLD22, MGG^{+}21, MB22, SAR^{+}24, SGF20]. Image-Tagging [SAR^{+}24]. Images [DYF^{+}23]. ImageSense [KTBL^{+}20]. Imaginaries [DDV23, HMB^{+}21]. Imagination [LCG24]. Imagine [FHF24]. Imbalanced [MWH19]. Immersion [KTF^{+}21]. Immersive [BBW23, GEEJ23, GW24, HOZ^{+}21, KRK^{+}21, LLW^{+}23, LPED22, LHI20, NAO23, SEC^{+}20, YDW^{+}22, ZES^{+}23]. Immigrant [LAM22]. Immigrants [LM20]. Impact [CBV22, CLP^{+}21, EJGL24, FC20, MH19, MCI20, MGG^{+}21, MSHP23, MP19]. Impacts [DRK18, LWC^{+}24, RGW^{+}24]. Impairments [JSP^{+}21, KKK^{+}24, LLL^{+}22, TRO^{+}21, ZMS^{+}22]. Imperfect [GYD17, YK23]. Implementing [RH22]. Implications [Bel24, GN20, Hao21, JCD19, MWBB23, MAS22, NPT19, YK23, YBE^{+}23]. Implicit [CHCG24, SCR22]. Implicit-Association [CHCG24]. Importance [VH21]. Imposition [MSY20]. Imprecise [KTP^{+}22]. Improve [JKW^{+}21, MCM23, MAMP21, WB21]. Improved [DMPK22, OS24]. Improvement [Hig20]. Improves [CTWM23, CLL^{+}21]. Improving [CSL^{+}24, KSG^{+}24, LYD^{+}23, MWH19, SPT^{+}20, WWH^{+}23]. IMU [LZW23]. IMUs [SAY^{+}24]. In-Context [KVM24]. In-game [RMB^{+}22]. In-Place [MCM23]. In-Situ [KFSC18, SHC^{+}23]. In-the-Moment [NSV22]. In-the-Wild [KMH23]. In-Vitro [KA23]. Inaccurate [ABC21]. Incels [RJZ^{+}21]. Incentives [STM^{+}19]. Incident [NHL^{+}24]. Inclusion [IRA^{+}19]. Inclusive [HBJR21, RI20, RVH24, TRO^{+}21]. Inclusivity [JKW^{+}21]. Income [MAR^{+}19]. Incongruent [FZRJ22]. Increase [BBW23, CI22, SSB^{+}22, WMM^{+}19]. Independence [BHNA18, LC24]. independent [DWZ^{+}20]. Index [YRIV20]. India [SLB23]. Indian [VMP21]. Indicate [TRN^{+}22]. Indicating [ERB^{+}17]. Indications [MMMM22]. Indicator [PLZ^{+}22]. Indicators [FRH^{+}22, GPH21]. Indie [LFM22, PF24]. Indirect [AMF18]. Individual [FWY^{+}22, HBJR21, LLL^{+}22, MDN^{+}24, ZPMK24]. Individuals [BRN^{+}19]. Indoor [KKV^{+}22]. Inducing [AIRH24]. Industrial [HMB^{+}21, TG18]. Industry [KG21]. Inertia [YML^{+}23]. Inertia-based [YML^{+}23]. Infected [KBSP^{+}22]. Infer [TG18]. Inference [BP23]. Inferring [NCM^{+}22]. INFEX [RDS18]. Influence [BJ21, BWE^{+}22, BDL21, FKDG21, GW24, KVK^{+}22, NCM^{+}22]. Influencers [LNM21]. Informal [LM20, LBM21]. Informatics [EJB^{+}20, MHP22]. Information [ASS18, Kee19, KOOL19, KA23, MGRM22, MM19, MPC^{+}23, NM21, NPT19, PAC22, RDS18, TSD23, VN19, WQCZ24]. Informational [SB20b]. Informed [ETS^{+}23, FMBW24, PF24]. Informing [RJVS24]. Infrastructure [BBV19, BBS19, PNS24a, SJ21, WBS^{+}23]. Infrastructures [JAB19, PVP22]. Infrastructuring [LC24]. Ingroup [WB21]. Initial [RGW^{+}24, YYT^{+}24]. Initialized [HHLC23]. Initiative [ACA22, CCH^{+}18]. Injury [BHNA18]. Inline [BYL^{+}24]. Inmates [TAS^{+}19]. Innovate [GBZ^{+}24]. Innovation [KJH19, LCG24]. Input [BAAP20, CMP^{+}24, GKHD23, GSK^{+}17, HRC^{+}21, JRP^{+}23, LLKC22, MXYS23, PLZ^{+}22, SAY^{+}24, SCMS18, SSR22, TFI^{+}22, YRIV20, ZMZL21]. Inputs [KIK20, MCM23]. Insertion [YPS^{+}22]. Insights [BDSL17, FH24, FR24, MHP22, XYDV23]. Inspection [MCI20]. Inspired [YWL^{+}23]. Instagram [HV22, RGW^{+}24]. Instant [CKP^{+}18, CA22, COKL20, HKF19, YHR^{+}19]. Instantiating [MTA^{+}17]. Institutions [FKK19, PK23]. Instructing [DR24]. Instruction [SAR^{+}23]. Instructional [LWS24]. Instructions [BWFG23, CSH^{+}23, KMM21]. Instrumenting [SB17]. Integer [LZS^{+}21]. Integrated [LTP23, MSK22]. Integrating [MQX^{+}23, SCMS18, JV22]. Integration [CMP^{+}24, CI22, ECP^{+}24, OR19]. Intellectual [HCR22]. Intelligence [BYRL23, WNT^{+}22]. Intelligent [CFG^{+}17, CLL^{+}21, KTBL^{+}20]. Intelligible [vDBL22]. Intent [KUWM22, NMPDJ24]. Intentions [CBSQ24, FH24, SB20a]. Inter [DPA17, HHVB22]. Inter-Group [DPA17]. Inter-Team [HHVB22]. Interactants [SPT^{+}20]. Interacting [CKP22, FHF24, LRP^{+}22, WWSG21]. Interaction [AHJV22, AV17, ALRS24, AARJ23, AM21, BLC^{+}21, BLV23, ÇH24, CYC^{+}21, CSMV21, CERR23, CLL^{+}21, GGF^{+}20, GBZ^{+}24, GB18, JNLtB22, KKK^{+}24, KTP^{+}22, KG20, LAR21, LWC^{+}23, LHI20, MCC^{+}18, MAN^{+}20, NMPDJ24, NAS^{+}22, RH22, RSL24, RLL22, RRF20, RIC^{+}24, SVW19, SVS^{+}24, SCR22, TD19, TBR20, UAAR20, UHH23, VP24a, WACdA^{+}21, WXL^{+}22, ZLB^{+}22]. Interaction-Oriented [MCC^{+}18]. Interactionism [PBC20]. Interactions [Blo24, BDE^{+}24, CLwH^{+}20, DGR^{+}22, FSAN23, GTO^{+}23, JDX24, KMH23, MDM^{+}19, MMI^{+}22, NAO23, PPT^{+}22, RH19, RZP^{+}22, WO20, WSM24, WLZL23, ZBL19, GHHS21]. Interactive [AS17, AIHS22, BWV20, BDW21, BT23, CMN^{+}21, DDFK19, FBG17, FLP^{+}21, Gar21, HMML^{+}22, JV22, JVP^{+}19, KE18, KKS19, LKG21, LPQW22, LS23, MdOKT23, MXYS23, OKM^{+}23, OGRV22, PB22, SD24, SVS^{+}24, SB17, SSB^{+}24a, TBR20, WGGS20, WGM21, WPH^{+}22b, WHW^{+}22, ZMR21]. Interconnected [WSM24]. Intercultural [AMJ^{+}19]. Interdisciplinary [SFA^{+}23]. Interest [WvECM22, ZMTH22]. Interface [ACDC23, CMY^{+}17, CHCG24, KKE^{+}21, MXYS23, MNP^{+}20, Sal17, TFI^{+}22, VZV19, YHR^{+}19, ZBV23, ZSOJ20, dRSV22]. Interfaces [AMVK23, AFT^{+}18, BP23, CLKEF21, CPB18, CWC^{+}24, FMBW24, GLZ^{+}21, GKHD23, HCM^{+}23, JNLtB22, KIK20, KVDG20, MTA^{+}17, MNP^{+}20, NMBS21, dQPMdHRdAN21, PCL^{+}21, RH19, SY24, SVV21, SORV22, SLN21, VN19, XHH17, XJBH22, YMG^{+}20]. Interferometry [MSK22]. Intergenerational [LOSD20, LTP23, PBC20]. Intermediaries [CCX^{+}20]. Intermediate [CRN21]. International [ADD21, PKLK22, RVH24, SAM23]. Internet [BT23, CSJ^{+}18, HKD23, HDPL21]. Interpersonal [ÇH24, LLW^{+}23, LFK24, MWBB23, MHN^{+}21, MPC^{+}23, TSHK20]. Interplay [AS17, BZC^{+}17, FCE^{+}18]. Interpretations [HKF19]. Interruption [SBPP20, YK23]. Interruptions [WGN^{+}21]. Intersection [ZKL22]. Intersection-Based [ZKL22]. Intersectional [Gil23, PLL^{+}23, WGM^{+}24]. Intersections [CFR22, KKV^{+}22]. Intersubjectivity [LWTW23]. Intervention [MHP22, MAMP21, RJVS24, WRM^{+}24, ZRL21]. Interventions [JZB^{+}21, RA21b, SRP^{+}24]. Intimate [BTM^{+}21]. Intra [HHVB22]. Intra-Team [HHVB22]. Intrinsic [CI22, LW19]. Introducing [SLMB24]. Introduction [BMdS^{+}21, BMdS^{+}22, LPQW22, LS23, PP23, SD24]. Invariant [MRV22]. Inventions [CBSQ24]. Inventory [HHJ^{+}22]. Investigating [AMR21, AASK23, BZSM20, CILM18, CSS^{+}17, CFR22, DBPA22, ERH^{+}22, IOES20, JP22, LWC^{+}24, LLKC22, MWH19, POG^{+}21, RVBC^{+}21]. Investigation [BWDP20, ESA^{+}24, HWGV24, VH21, YYT^{+}24, ZMGK24]. Invisible [MRS22, TSS21]. Involving [Fie19, GSLH24]. IoT [ETS^{+}23]. IRL [DGR^{+}22, SK20]. Islamic [IRA^{+}19, RAA22]. ism [GHHS21]. ISS [CM20, FJIH21]. Issue [SC24]. Issues [CSN19, XOEK22]. Item [HHJ^{+}22]. Items [PNS^{+}24b]. Iterative [BPB22, KVM24, LMB^{+}22, LLAK23, ZSK^{+}23]. IVR [BP17b].

Jam [FcLDM19]. Jams [FcLDM19].

January [MBTW24]. Japanese [YYT^{+}24]. Jarrah [ESA^{+}24]. JavaFX [FBG17]. Join [PSM21]. Joint [AGK^{+}22, HWS21]. Journaling [SAHE23]. Journey [SAE^{+}21]. Journeys [CAATL22, SWF^{+}20]. Judgments [XDC^{+}24]. June [LS23, SD24]. Jupyter [WEM24]. Just [BRM22, JS22, LFK24, TSHK20, TPNL22, VQ23]. Justice [HV23, SSH^{+}21]. Justifications [BTM^{+}21]. JustSpeak [ZMR21].

Kart [SS22]. Kart-ON [SS22]. KAVE [GB18]. Keeping [CNM^{+}23, XKRW23]. Kenya [HOHB22]. Kenyan [WGG20]. Key [BLC^{+}21, BBV19]. Keyboard [BB20, CSMV21, LZW23]. Keyboards [ASP20]. Kick [PKC^{+}21]. Kinds [CCW22]. Kinect [GB18]. Kit [dRSV22]. Knit [FB19]. Knitted [VMD^{+}20]. Knob [SSR22]. Know [LC24, BLH^{+}17]. Knowing [SRKK22]. Knowledge [ABC21, HOHB22, KDEA21, KZZW22, MdLGG23, OR19, TPNL22]. Korea [KOL^{+}24]. Korero [CMY^{+}17].

Labeled [XOEK22]. Labeling [CLP^{+}21]. Labels [ZYY23]. Labor [HPL22, TSS21]. Landing [Sal17]. Lands [FD20]. Language [BB20, BBSV23, CYW18, HSG19, KKK^{+}24, KVM24, MCC^{+}18, MHN^{+}21, MNP^{+}20, PAC22, TDC17, WBSM24, ZAS^{+}21]. LanguageLogger [BB20]. Laptop [CSMV21]. Large [BB18, BBSV23, DWMV21, FLC20, KKK^{+}24, KVM24, KNT22, LYL^{+}23, RRF20, WBSM24, WACdA^{+}21, XRS22, XOEK22]. Large-Scale [XRS22]. largest [Gil20]. Laser [MSK22]. Lasso [YUSM22]. Last [BZC^{+}17]. Lasting [RGW^{+}24]. Latency [FX22, HSH22, ZLB^{+}22]. Later [KUD^{+}23]. Laws [YTF^{+}23]. Layer [SDF^{+}21]. layered [TVS17]. Layout [KIK20, LPED22, SDO21]. Layouts [BBSV23, LZS^{+}21, SEC^{+}20]. Lead [AOT^{+}18]. Leader [LRK^{+}24b]. League [RWB22, TRN^{+}22]. Leakage [CHC^{+}17]. Leap [SORV22]. Learn [HKL^{+}22]. Learn2Earn [STM^{+}19]. Learned [GHHC21, KRK^{+}21, Rib19]. learner [BPO21]. Learners [DHH18, FWY^{+}22]. Learnersourcing [JK24]. Learning [Ame18, AMJ^{+}19, BBD^{+}24, CMRH22, CXL17, CI22, DHH18, DMPK22, ETS^{+}23, FcLDM19, FWY^{+}22, GSLH24, GLZ^{+}21, HRP23, HZ22, JØC^{+}20, JA21, KVM24, KHY^{+}22, LCA^{+}24, LM20, LBM21, LX20, MON^{+}21, RIK21, STSS22, SM22, SBB24, WBS^{+}23, WGS^{+}24, YGHR21, vBGH^{+}19, HSC^{+}23]. Lecture [BBD^{+}24]. Led [KG20, HV23]. Legal [DBL22]. Legends [RWB22, TRN^{+}22]. Lenient [KB22]. Lens [HOHB22, LGG^{+}23, NM22b, SB20b, WGM^{+}24, NKV21]. Lenses [GKS21]. Lessons [AMJ^{+}19, BLV23, FF22, GHHC21, KRK^{+}21]. Let [EPW^{+}23, SFM^{+}22, BKJ^{+}20]. Letter [SN18]. Level [SAY^{+}24]. Levels [Hig20]. Leveraging [LFM22, LM20, RLL22]. LGBTQ [DWB18, SHS22]. LibraryThing [AWM21]. Life [AV19, GSC21, KVK^{+}22, LOSD20, SBAK21, WMM^{+}19]. Life-Tags [AV19]. Lifecourse [WGM^{+}24]. Light [HFBL19, MSK22, NHL^{+}24]. Lighting [FRH^{+}22]. LightMeUp [CKSP22].

LightPlay [FRH^{+}22]. Lightweight [DAQ17, JZB^{+}21, RH23]. Like [SJ21, VQ23]. Likert [RDF^{+}22]. Likert-style [RDF^{+}22]. Limb [LCG24, XMS24]. Limits [GC19]. LINA [MBK^{+}22]. Line [CYGW21, HBJR21, VZV19]. Lingua [CYW18]. Linguistic [ERB^{+}17, HWF^{+}24]. Links [DPA17, DRAQ18, GEEJ23, VH21]. Listen [FLC17]. Listener [CMSA22]. Literacy [HBLN24, LBM21, MAMP21]. Literature [APW^{+}21, DWW^{+}17, DFJ^{+}21, ND24, SFA^{+}23, SCR22, STM24, WRM^{+}24]. Little [AM21, BVM21]. Live [CCWH21, GF22, LUFW20, LFM22, LHPH20, RWB22, SS19]. Live-Streaming [LHPH20]. Lives [KKM^{+}22]. Livestreaming [LAW19, SK20]. Living [CRM^{+}24, CWB^{+}18, SCSD24, CS18b]. LLM [SSR24]. Load [CERR23, DWSF24]. Loads [NB20]. Local [ERVR^{+}23, ZD19]. Located [DGR^{+}22, WPH22a, KSKH24, NM22a, OG21]. Location [BJ21, LNM21]. Location-Based [BJ21, LNM21]. locations [FZRJ22]. Locomotion [CDIH23, LKYK24]. Log [BYL^{+}24]. Logo [Ame18]. Long [CWB^{+}18, HFBL19, SMSM22, SMLH23, DB22]. Long-Term [CWB^{+}18, SMSM22, SMLH23, HFBL19, DB22]. Longitudinal [AAHM22, BKKD21, HSG^{+}22, KVK^{+}22]. Look [KSP^{+}23]. Looking [WAD^{+}23, Pri19]. Loop [AGM^{+}23, GC19, KG21, WNT^{+}22]. Loot [WZT^{+}24]. Low [BLB^{+}24, CYC^{+}21, JF24, LAR21, LZW23, MAR^{+}19, SJ21, ZYY23]. Low-Cost [LZW23, JF24, ZYY23]. Low-Income [MAR^{+}19]. Low-resolution [SJ21]. Low-Vision [LAR21]. Lower [CSL^{+}24]. Lying [IOES20].

M [WGN^{+}21]. Machine [CGM^{+}18, ETS^{+}23, GLZ^{+}21, ZGL^{+}22, vBGH^{+}19]. Machinery [MBS22]. Macro [CSJ^{+}18]. Macrotask [MRS22]. MADE [SLT^{+}21].

MADE-Axis [SLT^{+}21]. Magic [KG21, LAM22]. Magnetic [TFI^{+}22, ZYY23]. Magnetic-Based [TFI^{+}22]. Maintain [RKNES20]. Maintenance [SSNC20]. Major [JSJ22]. Make [SLB23, ZPMK24]. Maker [HV23, MWH19, Vya19]. Makers [CJ23, MWH19]. Makerspaces [LZAD21]. Makes [PSM21]. Making [BKKD21, DDFK19, GYD17, GPH19, GC19, KJTC22, KOOL19, MAS22, MSHP23, MAD^{+}23, VQ23, NM22b]. Mall [AKIS21, KKK^{+}24]. Manage [AGK^{+}22, SFLB23]. Management [BKM20, BL17, BRM22, CNH^{+}23, CRN21, GB18, KB22, LWTW23, NSJ21, SBPP20, SFA^{+}23]. Managing [BRN^{+}19]. Mandarin [LHT^{+}20]. Maneuver [MMMM22]. Mani [LO17]. Mani-Pull-Action [LO17]. Manifest [VMP21]. Manipulating [HRC^{+}21]. Manipulation [BYL^{+}24, CFG^{+}17, ERLW17, MMM^{+}21, WO20]. Manipulative [HSZKN24]. Manual [CFR22, DR24, GPC^{+}23, KFSC18]. Manuals [LWS24]. Manufacturing [MBS22]. Many [CT18]. Map [DCM24, SSR22]. Mapping [DC17, GSLH24, MdOKT23, PKB^{+}19]. Mappings [CDIH23, ZMZL21]. Maps [FYS^{+}24, GEEJ23, NM21, SEC^{+}20]. MAR [ZLB^{+}22]. Marginalized [RGW^{+}24]. Marital [ADAY21]. Markers [ERB^{+}17, KNT22]. Market [XDC^{+}24]. MARLUI [LCA^{+}24]. Mask [WVG^{+}24]. Matching [AZ23, ZD19]. Matchmaking [ADAY21]. Material [VWHW24]. Materialities [MM19]. Materials [CJ23, CMY^{+}17, PAK^{+}22]. Math [KR24]. Matter [AZS^{+}22, DR24, ECA23]. Matters [HHVB22]. Mature [LM20]. Maximization [WYY^{+}21]. may [KBSP^{+}22, DKBG23, DKKL24, GKKW22]. MDE [LDP18]. Me [ASJK23, SEC^{+}20, WWSG21, IRA^{+}19].

Meaningful [SLMB24]. Means [LLDH21]. Measure [ASM23, HHJ^{+}22]. Measurement [APW^{+}21]. Measurements [CYGW21]. Measures [GBN21]. Measuring [ALD^{+}20, HJM20, ZBV23]. MechanicalHeart [CGM^{+}18]. Mechanics [HTTH22, HMK24, KML^{+}21]. Mechanisms [MAS22, ZSK^{+}23]. Media [AMF18, BSS18, BBK22, BWDP20, CSMG17, CH20, DWB18, DR24, EJB^{+}20, ERB^{+}17, GAW19, GSC21, GPH19, GPH21, GN20, HKD23, JZB^{+}21, JSJ22, KBFK21, MAMP21, RKNES20, RA21b, RAZ^{+}20, SSH^{+}21, SSD23, VVBK23, YBE^{+}23, ZWG^{+}23]. Mediated [LW19, MHN^{+}21, PLL^{+}23, XYDV23]. Mediating [CVL22, KG20]. Mediation [BEHB^{+}23, KSKH24]. Mediator [LYH20, VKI21]. Medical [KKK22b]. Meditation [VQ23]. Medium [DR24, SLN21]. Medium-Fidelity [SLN21]. Medusa3D [JDX24]. Meeter [HSG19]. Meeting [WQCZ24]. Meetings [KDH^{+}21, WQCZ24, ZRP^{+}22]. meets [RYCC21]. Meetup [CHC^{+}17]. Members [GTC22, HMB^{+}21, PBC20]. Memes [MB22, PAC22]. Memory [CTWM23, JA21, KSX21, LPED22, WGGS20, WGM21]. Men [RYD21, VQ23]. Mental [AAHM22, EMS^{+}22, GSLH24, LYH20, MBK^{+}22, SFM^{+}22, SAM23, ULG^{+}22, VHZ22, WSM24, XYDV23]. Mentoring [NHH20]. Mentorship [LYL^{+}23]. Menu [ZP24]. Menus [BCCV17, FLC20, LRP^{+}22]. Merlynne [CGS^{+}21]. Meso [CSJ^{+}18]. Message [CA22, KMHLF18, LGS19]. Messaging [ABQ17, CKP^{+}18, CA22, COKL20, HKF19, JFL23, LLC^{+}24, MFM^{+}20, RAZ^{+}20]. Meta [OR19]. Meta-Knowledge [OR19]. Metaphoric [LOSD20]. Method [BPB22, CS18a, CBSQ24, GM23, HCM^{+}23, KVM24, PA22, SLMB24, SSR22, TBHM23, WGM21, YYT^{+}24]. Methodologies [PSTK24]. Methodology [SHM24].

Methods [ABC21, BDL21, DRL^{+}22, GB18, HSC^{+}23, LWC^{+}23, SSGB20, TRN^{+}22, WGM^{+}24]. metric [BLC^{+}21]. Mexico [GYD17]. MHCI [ACI22, KvBW24, LMS23]. Micro [CSJ^{+}18]. Microbial [PPT^{+}22]. Microgestures [WGN^{+}21]. Microscopy [MMI^{+}22]. Microtask [QGB20]. Mid [BLC^{+}21, DRL^{+}22, GTO^{+}23, LRP^{+}22, NM22a, VMHO^{+}22, VWHW24, YMG^{+}20]. Mid-Air [LRP^{+}22, YMG^{+}20, BLC^{+}21, DRL^{+}22, GTO^{+}23, NM22a, VMHO^{+}22, VWHW24]. Midair [ZMZL21]. middle [LRSS20]. Migrant [HV23]. Migration [FD20, KWdCL^{+}21]. Migrations [RJZ^{+}21]. Mile [ETE^{+}19]. Military [DRA^{+}17, GGE23]. Millimetre [PAK^{+}22]. Millimetre-Wave [PAK^{+}22]. Mind [AMR21]. Mindfulness [HPH^{+}23]. Miner [SHM24]. Miniaturized [LZY^{+}21]. Minimalistic [ARH^{+}23]. Mining [TD19]. miniPXI [HHJ^{+}22, HWGV24]. Minoritised [WGM^{+}24]. Minority [AS23]. Miracle [WAD^{+}23]. Misaligned [DPA17, TBDB21]. Misfires [TBDB21]. Misinformation [HKD23, HJM20, JZB^{+}21]. Misinterpretations [CKP^{+}18]. Missed [TBDB21]. Mist [MXYS23]. mitigating [GN20]. Mitigation [LPZV23, ZSK^{+}23]. mix [BKJ^{+}20]. Mixed [ABC21, BPB22, CCH^{+}18, CHS^{+}24, HCM^{+}23, PDPS22, Pri19, SRZ24, SHC^{+}23, TRN^{+}22, TBHM23]. Mixed-Method [TBHM23]. Mixed-Methods [ABC21, TRN^{+}22]. Mixed-Presence [CHS^{+}24]. ML [ADD21, SSB^{+}24a]. mmWave [ASZW23]. MOBA [LMZ^{+}24]. Mobile [AHCD17, AGK^{+}22, ASO^{+}22, ASP20, AZ23, AASK23, ABQ17, BB20, BWE^{+}22, BR24, CEM17, CSL^{+}24, COKL20, CBSQ24, DGR^{+}22, DBPA22, DTE^{+}22, DNBB^{+}22, FSSK22, GGL^{+}22, GRG20, IB24, KCC20, KMH23, KOL^{+}24, KFSC18, KBSP^{+}22, KKV^{+}22, LZY^{+}21, LWC^{+}23, LWB22, MP23, MBS22, MPEC24, NM22a, RH23, RRA23, RZP^{+}22, SAR^{+}23, SCY^{+}18, STM^{+}19, TBHM23, TCH20, WYY^{+}21, WPH22a, YRIV20, ZP24, dRSV22, IRA^{+}19]. MoCaDiX [VN19]. Mod [JTB22]. modal [BLC^{+}21, KZZW22]. Mode [HNS^{+}24]. Model [AAHM22, BCCV17, BBSV23, CMN^{+}21, DRP^{+}23, EMS^{+}22, Gil23, HC24, JKZ^{+}22, JTB22, KVM24, KNT22, LAHW19, RH19, RS19, SR22, TBR20, VP24b, VHZ22, YU20, ZBV23]. Model-Based [CMN^{+}21, BCCV17, KNT22, TBR20, ZBV23]. Model-Driven [JTB22]. Modeling [BSM23, BDL21, EJGL24, MPKEN24, MDN^{+}24, NMPDJ24, TG18, UYM21, Wil20, YUSM22, YK23]. Modelling [BDW21, BT23, LDM^{+}22, TVS17]. Models [ADD21, AFT^{+}18, AM21, BR17, BF17, BMR^{+}23, CFG^{+}17, CMN^{+}21, CNH^{+}17, DS20, GSLH24, KKK^{+}24, PCMP18, PVC^{+}23, SFM^{+}22, See20, SWT20, WBSM24, WZT^{+}24, WSM24]. Moderated [CWK^{+}23]. Moderately [JJJ^{+}23]. Moderates [SK23]. Moderating [Gil20]. Moderation [ESA^{+}24, FTP20, Gil23, JJJ^{+}23, KJH19, KG20, MK23, RGW^{+}24, RJZ^{+}21, See20, SK23, VXHK21, WS23]. Moderator [AIRH24, BHSP22]. Moderators [CBPC23, CCW22]. Modern [MFM21]. Modular [PCL^{+}21, SDF^{+}21, SLT^{+}21]. Molecular [PCL^{+}21]. Mom [LLC^{+}24]. Moment [NSV22]. Momentary [BahRKM18]. Moments [LWC^{+}24]. Monetization [HRR^{+}22]. Money [EMS^{+}22, LFK24]. Moneywork [PKC^{+}21]. mongering [KJTC22]. Monitoring [CRM^{+}24, MBS22, WGM21]. MOOCs [KCC20]. Moods [MMH21]. Most [KPW19]. Motherhood [BBS19]. Mothers [BBS19, LLC^{+}24]. Motion [BLV23, GB18, SORV22]. Motivating [AS17, BRM21, CGS^{+}21, HSZKN24, SSB24b]. motivation [OS24]. Motivations [GPH19, KOL^{+}24, SGF20, ULG^{+}22]. Motivators [PM21]. Motives [BZC^{+}17]. Motor [LCG24]. Motorist [KKM^{+}22]. Mountain [BBW21]. Mounted [GGE23, JP22, KBK^{+}22]. Mourning [BHM19]. Mouse [SC23]. Move [STM24]. Movement [IOES20, JCD19, UYM21, Zha23]. Movements [ASM23, BSM23, CDIH23, FH24, MFSK22, PM21, WVG^{+}24]. Moving [FD20, OKM^{+}23]. MR [LWB22]. MTBalance [BBW21]. Multi [ABQ17, BLC^{+}21, BDE^{+}24, CKSP22, ERH^{+}24, GVNK20, HOZ^{+}21, JRP^{+}23, KOL^{+}24, KZZW22, LCA^{+}24, LAHW19, LDP18, MRV22, MDM^{+}19, MMI^{+}22, OKM^{+}23, OSM^{+}20, RLGG21, SCP^{+}21, SBPP20, SAE^{+}21, TD19, TVS17, WEM24, XMS24, ZLB^{+}22]. Multi-Agent [JRP^{+}23, LCA^{+}24]. Multi-Application [TD19]. Multi-Channel [ABQ17, RLGG21]. Multi-Device [LAHW19]. Multi-Directional [XMS24]. Multi-faceted [KOL^{+}24]. Multi-layered [TVS17]. Multi-modal [BLC^{+}21, KZZW22]. Multi-Paradigm [WEM24]. Multi-Party [BDE^{+}24, ERH^{+}24]. Multi-Resident [SAE^{+}21]. Multi-stable [CKSP22]. Multi-Staged [LDP18]. Multi-Stakeholder [SCP^{+}21]. Multi-stroke [GVNK20, MRV22, OSM^{+}20]. Multi-Surface [MDM^{+}19, MMI^{+}22]. Multi-target [ZLB^{+}22]. Multi-tasking [SBPP20]. Multi-Touch [HOZ^{+}21]. Multi-view [OKM^{+}23]. Multidisciplinary [MAS22]. Multilevel [ALRS24]. Multilingual [BKJ^{+}20, CYW18, LYD^{+}23]. Multimodal [ALRS24, BCCV17, CMP^{+}24, CCNY19, KUWM22, SS23, WO20, YOFC19]. Multimodality [SAHE23]. Multiparty [CYW18]. Multiplayer [HRP23]. Multiple [BGM^{+}23, BLH^{+}17, LDP18, MON^{+}21, SAHE23]. Multitasking [WGN^{+}21]. Multitouch [RRF20]. Multiview [SEC^{+}20]. Museum [LMB^{+}22]. Music [CILM18, HJC^{+}22, JCD19]. My [ATH24, BVM21, CYC^{+}21, CLL^{+}21, GSC21, IRA^{+}19, PF24, SFLB23, WWH^{+}23, HSH22, HV22, MMMM22]. Mysteries [JA21].

’n [GBZ^{+}24, LZY^{+}21]. NaCanva

[YWL^{+}23]. Nairobi [AMJ^{+}19]. Named [CH20]. Narrative [BTM^{+}21, HKL^{+}22, NM21, WB21]. Narratives [CBWD24, NM21, YBE^{+}23]. Native [LYD^{+}23]. Natural [AKVHZ23, DTE^{+}22, KMH23]. Naturalistic [IB24]. Nature [MT21, MYS^{+}21, RVBC^{+}21, YWL^{+}23, LRK^{+}24a]. Nature-Inspired [YWL^{+}23]. Navigate [KG21, LKG21]. Navigation [AMR21, LZY^{+}21, MdLGG23, SSR22]. NCAlt [HZ22]. Near [BWFG23, CSMV21]. Near-Keyboard [CSMV21]. Near-Symmetrical [BWFG23]. Nearby [ZMTH22]. Need [CEM17, HJC^{+}22, LPZV23, RH22]. Needs [BDW21, ERVR^{+}23, HPY^{+}22, MES^{+}22]. Negative [EMS^{+}22]. Negatively [DRK18]. Negotiating [CS18b]. Negro [KKM^{+}22]. NeighboAR [SBSM24]. Neighborhood [HSVR21]. Neoliberalism [HTS24]. Net [TD19]. Network [ALG^{+}21, HSVR21, SS21, SPT^{+}20]. Networked [BHM19]. Networking [Ann20]. Networks [BB18, CHC^{+}17, FYS^{+}24, KR24, LSLD22, MDM^{+}19, NCM^{+}22, WYY^{+}21]. Neural [BSM23, KR24, ZZK^{+}22]. never [IRA^{+}19]. Newcomer [JVP^{+}19]. Newcomers [NHH20]. News [BZSM20, HKD23, LHPH20, MAMP21, RAZ^{+}20]. Next [RS19]. NEXTGEN [SLD21]. NFTs [XDC^{+}24, EMT^{+}24]. Nicer [WB21]. Nile [SELS24]. NLP [JFS^{+}23]. NLP-Assisted [JFS^{+}23]. No [TKB^{+}22]. Nobody [SS21]. Non [AZS^{+}22, BSS18, BDB^{+}23, EMT^{+}24, FLC17, GPH19, JF24, LYD^{+}23, MQX^{+}23, RVBC^{+}21, SSGB20, SSGB21, SSR22, UAAR20, XDC^{+}24, ZSOJ20, ZYY23, ZMTH22, Gar21]. Non- [AZS^{+}22, ZMTH22, Gar21]. Non-/ [Gar21]. Non-clinical [JF24]. Non-Driving [BDB^{+}23]. Non-Fungible [XDC^{+}24]. Non-Native [LYD^{+}23]. Non-Occluding [SSR22]. Non-Real-Time [MQX^{+}23]. Non-Transferable [EMT^{+}24]. Non-use [GPH19, SSGB21]. Non-Verbal [FLC17]. Non-Violent [RVBC^{+}21]. Non-Visual [UAAR20, ZSOJ20, ZYY23]. Non/use [BSS18, SSGB20]. Noncompliance [BKJ22]. Noncontact [MMM^{+}21]. Nondominant [YGHR21]. Norm [CSJ^{+}18, FB19, NSJ21]. Norms [MPC^{+}23, RKNES20]. Notation [HG19]. Notations [MPB^{+}19]. Note [FWY^{+}22, KFMH17]. Note-taking [FWY^{+}22]. Note/Chairs [KFMH17]. Notebooks [WMBO19]. NoteWordy [LLKC22]. NotiBike [KMM^{+}22]. Notification [CSL^{+}24]. Notifications [CSL^{+}24, COKL20, KMM^{+}22]. Novel [CKK22, RH22, SCSD24, WS22]. Novelty [CTH24]. November [CM20, FJIH21, KRW^{+}22]. Novice [BBW21, DDFK19, FGG17, KSG^{+}24, WGM21]. Novices [DR24]. Nowhere [PKLK22]. Nudge [BKJ^{+}20]. Nuisance [LL24]. Number [RA21a].

Object [LLGH24, MMM^{+}21, RKFK22, SBSM24, VP24b, ZMZL21]. Objects [BWFG23, CWK^{+}23, CS18b, OKM^{+}23, WGG20, XHH17]. Obligations [BR17]. Observation [YYT^{+}24]. Observer [WvECM22]. Observing [DNBB^{+}22]. Obstacle [SRKK22]. Obstacles [KKV^{+}22]. OCC [HC24]. Occluding [SSR22]. Occupations [MGG^{+}21]. Oct [KLS20].

October [BLS21, CDGdC23, FBNM22, MDFG24]. Odin [HOSK21]. Off [TCH20, BKG^{+}21, FMBW24]. Office [LAR21]. Offline [LFK24]. Offshored [LWTW23]. Older [CNH^{+}23, CMRH22, KSX21, WGM^{+}24, XKRW23]. Omni [MNP^{+}20]. Omni-Channel [MNP^{+}20]. On-body [BP23]. On-demand [HPL22]. On-Hand [ZSOJ20]. On-Line [VZV19]. On-screen-start [YU20]. On-Skin [BP23]. One [FC20, PNS^{+}24b, RRA23, UYM21, UYSM22, ZRA22]. One-Dimensional [UYSM22, UYM21]. One-Hand [ZRA22]. One-Handed [FC20, RRA23]. Online [AGK^{+}22, ADAY21, AS23, AOT^{+}18, BTM^{+}21, BEHB^{+}23, BDSL17, BKKD21, BBNT21, BPO21, CNM^{+}23, CILM18, CS18a, CMRH22, CCX^{+}20, CCWH21, CSH^{+}23, CWB^{+}18, CCW22, DHH18, DS20, FcLDM19, FD20, FKDG21, FZM22, HV23, JØC^{+}20, JS22, JAB19, LMT^{+}22, LTP23, LSLD22, LX20, OM20, OR19, PM21, PNS^{+}24b, SDF^{+}21, SSH^{+}21, SSD23, See20, TSD23, TKB^{+}22, WWT^{+}21, WZFA22, WRM^{+}24, XZL^{+}20, vDBL22]. Open [BP17a, GN23, HWF^{+}24, IB24, LPFD21, SC24, WWSG21]. Open-Ended [BP17a]. Open-Source [IB24]. OpenUIDL [MNP^{+}20]. Operating [HBJR21, ZRA22]. Operations [ASS18, LLGH24, LZY^{+}21, MNP^{+}22]. Operations-based [MNP^{+}22]. Operators [SDO21]. Opportune [LWC^{+}24]. Opportunities [Ann20, BWDP20, KOOL19, MWH19, OS24, SAHE23, SHY^{+}18, SWF^{+}20, SMLH23, WXL^{+}22, BL17]. Optimised [HNS^{+}24]. Optimization [Sal17, ZP24]. Optional [BRM21]. Order [EHH^{+}18]. Organizational [PNS24a, RYCC21]. Organizations [DDV23, XKRW23]. organized [ERVR^{+}23]. Organizing [BSS18]. Orientation [BWFG23, JPKM24]. Oriented [MCC^{+}18, SSB24b, ZGL^{+}22].

Orienting [BHM19, KG21]. Origins [MB22]. OsciHead [HTS^{+}22]. Oscillation [HTS^{+}22]. OT [SSNC20]. other [CNH^{+}23]. Ourselves [CA22]. Outages [ALG^{+}21]. Outcomes [ALD^{+}20, ERB^{+}17, HBM23]. Outdoor [HFBL19, LAW19, MGRM22]. Outer [KG21]. Outlets [HKD23]. Outpace [CKK22]. Output [SCMS18]. outreach [Gil20]. Overcome [ETE^{+}19]. Overlays [LSLD22]. Oversight [AGK^{+}22]. Owned [LX20]. Owner [ZZ23]. Owners [WHW^{+}22]. Ownership [EMT^{+}24].

P2P [LFK24]. Paced [EPW^{+}23]. PACM [CM20]. PACMHCI [ACI22, BKL^{+}21, BLS21, BLS^{+}22, CDG^{+}23, CDGdC23, CDG^{+}24, DKBG23, DKKL24, FJIH21, FBNM22, GKKW22, GMM21, KLS20, KLS21, KvBW24, KRW^{+}22, LMS23, LS23, MDFG24, MBTW24, SD24]. Page [NB20]. Pages [Sal17]. Paint [AIHS22]. Painter [GGF^{+}20]. Pair [ZWG^{+}23]. Pair-wise [ZWG^{+}23]. Pandemic [FF22, HSVR21, PKLK22, RIK21, SLB23, XKRW23, YGHR21, dSdSdSF^{+}23]. Pandemic-Era [FF22]. Paper [CKP22, CKSP22, Hig20, HGG20, SS22]. Paper-cutting [Hig20]. Papers [CCH^{+}18]. Paradigm [KKM24, PK23, WEM24]. Paradox [HHVB22, STSS22]. Parallax [GB18]. Parallel [CAATL22, RRF20, SWF^{+}20]. Parent [ADAY21, MFM21]. Parent-Child [MFM21]. Parental [AGK^{+}22, YGHR21]. Parents [AGK^{+}22]. Part [ASS18, MH19]. Part-Time [MH19]. Partial [CS18a]. Participant [KVM24]. Participants [GM23]. Participate [TKB^{+}22]. Participation [AS17, CYW18, EHZN22, FF22, HOHB22, Hao21, KWdCL^{+}21, LYD^{+}23, SLMB24, TAS^{+}19]. Participatory [AZ23, BKM20, CMRH22, DBL22, FKK19, DRDS24, LCG24, TBL^{+}24, WGM^{+}24]. Partisan [DDV23]. Partner [BTM^{+}21, Bel24]. Partnerships [KTBL^{+}20]. Party [BDE^{+}24, ERH^{+}24, GN20, PLY23, PKC^{+}21]. Pass [GP22]. Passenger [LWB22]. Passing [VKI21]. Passive [ZBL19]. Password [SELS24]. passwords [BBK22]. Past [BWV20, WAD^{+}23]. Paste [LZY^{+}21, LZY^{+}21]. Patent [SLD21]. Path [YUSM22]. Path-Segmentation [YUSM22]. Pathologists [GHHC21]. Paths [YUSM22, YTF^{+}23]. Patient [BKJ22, FSC21, HKM^{+}20, YBE^{+}23]. Patients [BLH^{+}17, KA23, SWF^{+}20]. Patterns [AV17, ZWG^{+}23]. PAWS [CDIH23]. Payment [CEM17]. Payments [LFK24]. PC [JOFM24]. PCG [VAE^{+}22]. PDF [UAAR20]. Pedagogies [CJ23]. Pedagogy [CLKEF21, MD20]. Peer [BZC^{+}17, CGS^{+}21, CKS18, DC17, DRK18, HKM^{+}20, Kee19, LM20, PLL^{+}23, XOEK22]. Peer-Produced [Kee19]. Peer-Ranked [XOEK22]. Peer-to-Peer [CGS^{+}21]. People [AM21, BLB^{+}24, CR20, HCR22, KKK^{+}24, KRK^{+}21, KKV^{+}22, LLL^{+}22, LC24, MRGN20, PSM21, SKV^{+}21, SVS^{+}24, SSB^{+}24a, Tan21, TKB^{+}22, WB21, XYDV23, ZZ23, ZMTH22, BRN^{+}19, CNH^{+}23, DB22]. People-Nearby [ZMTH22]. Perceived [DRAQ18, GPH21, HRC^{+}21, KML^{+}21, KYB21, LPZV23, SSB^{+}22, WWH^{+}23, ZBL19, DB22, LWC^{+}24, GP22]. Perception [ADAY21, ALD^{+}20, ALSK23, FSRC24, HWYW21, MCM23, VWHW24]. Perception-Based [ALSK23]. Perceptions [AASK23, CH20, vBGH^{+}19]. Perforated [CKP22]. Perform [BRM21, DR24]. Performance [ARH^{+}23, AZS^{+}22, BDB^{+}23, BRM21, CTWM23, ERH^{+}22, FC20, FKF^{+}22, GW24, LUFW20, SFM^{+}22, WXL^{+}22, YK23]. Peripheral [LKYK24, SRAG22, SVW19]. Perpetual [WMM^{+}19]. Person [KBSP^{+}22]. Personal [AAHM22, Ann20, BLL^{+}21, CNH^{+}23, DWB18, DB22, EJB^{+}20, FZRJ22, HFL19, LLDH21, LLKC22, MHP22, SHC^{+}23]. Personalised [HLV21, SBB24]. Personalization [KVM23, SR22]. Personalized [AZS^{+}22, CDIH23, FCR^{+}17, LZS^{+}21, ZGL^{+}22]. Personas [BPB22]. Perspective [KKK22a, KHY^{+}22, MWK^{+}22, VKI21]. Perspective-taking [KKK22a]. Perspectives [Bel24, BHM19, BK24, CNH^{+}23, DTE^{+}22, FKF^{+}22, HBM23, LRK^{+}24b, MWBB23, RYCC21, RVH24, WFL^{+}22]. Persuasive [DDFK19]. Pervasive [NMBS21, ÖBK24]. Pests [FM22]. Petri [TD19]. Phase [ASJK23, KHY^{+}22]. Phish [DNBB^{+}22]. Phone [AHCD17, TBHM23, IRA^{+}19]. Phones [YRIV20]. Phonocardiograms [CGM^{+}18]. Photochromic [AIHS22]. Photography [AS23, WGG20]. Photos [RWL23]. Physical [CTH24, ECP^{+}24, EPW^{+}23, ERH^{+}22, GRG20, DRDS24, HPH^{+}23, JOFM24, MLSH21, WHW^{+}22, XHH17]. Physicalizations [HOZ^{+}21]. Piano [DMPK22]. Pick [PKC^{+}21]. picked [WBS^{+}23]. Pickers [RA21a]. Picture [ATH24]. pictures [HV22]. Piecework [SLB23]. Pillowfort [DPF22]. Pin [BHM^{+}23, WVG^{+}24]. Pinball [JOFM24]. Pinch [DRL^{+}22]. Pinhole [CKP22]. Pipeline [CHS^{+}24]. Pitfalls [FM22]. Place [FKK19, MCM23, PNS^{+}24b, SAE^{+}21]. Placemaking [CBV22]. Placements [MGRM22]. Plan [ABC21, KYB21]. Planar [GTO^{+}23]. Planning [ABC21, LLAK23]. Platform [DPF22, FD20, GN23, JRP^{+}23, KB22, LHI20, LX20, MH19, MV19, RSL24, RJZ^{+}21, HKL^{+}22]. Platformed [WWSS22]. Platforms [BBNT21, ETS^{+}23, HJM20, NKV21]. Plausible [DMRA24]. Play [AGO^{+}24, AMJ^{+}19, GHHS21, HSZKN24, JS22, LKG21, PPT^{+}22, MFM21, APW^{+}21, FBNM22, GMM21, MDFG24]. Playability [MPEC24]. Player [ASZW23, HHJ^{+}22, JWK22, JOFM24, ND24, VHZ22, WRM^{+}24, XJBH22]. Players [BRM21, FHF24, LCG24, MHG^{+}21, RGR^{+}21, WXL^{+}22]. Playful [AKIS21, DGR^{+}22, KTP^{+}22, VAE^{+}22]. Playing [AL21, BJ21, SSR24]. Please [ITKB23, KWdCL^{+}21]. Poetics [RAA22]. Point [LCA^{+}24]. Point-and-Click [LCA^{+}24]. Pointing [DWMV21, ERH^{+}22, HNS^{+}24, HK24b, KMH23, LRP^{+}22, PA22, UYS23, YU20, ZMZL21, ZKL22, Zha23]. Points [KIK20]. Pokémon [LRK^{+}24a]. polarisation [GN20]. Policies [STSS22]. Policy [CCS^{+}18, DPF22, KM22]. Policymaking [AS17]. Politeness [LHT^{+}20]. Political [DPF22, Kee19, SCP^{+}21, GN20]. Politicians [VMP21]. Politics [GN20, RAA22, SHD21, SDF^{+}21, SSB^{+}24a]. Polymodal [BCCV17]. Polyphony [RH19]. Pool [MMH21]. Populations [SHM24]. Porous [CS18b]. Portfolios [FKDG21]. Portrait [WvECM22]. Portuguese [GGE23]. Positive [CCW22, HPH^{+}23, JS22, KF23]. Possessions [EMT^{+}24]. Post [ASJK23, CCS^{+}18, GSC21]. Post-Attack [ASJK23]. Post-College [GSC21]. Posting [JZB^{+}21]. Posture [IOES20, WDP^{+}20]. Postures [CSMV21]. Potential [CLKEF21, GH21]. Power [Gil23, LPZV23, LNM21, MSY20]. powerful [BRN^{+}19]. Practical [ALSK23]. Practicality [HC24]. Practice [CTH24, HV23]. Practices [AL18, BL17, CEJC23, CFD^{+}21, CAATL22, FSC21, Gar21, GF22, HGG20, JA21, JSKA22, LUFW20, LLL^{+}22, PBC20, RYD21, RYCC21, SK23, SLMB24, SAR21, SMLH23, ULG^{+}22, WAD^{+}23, WMM^{+}19, WLZL23, ZMS^{+}22]. Practitioner [MES^{+}22, MWK^{+}22, RYCC21]. Prayers [SHL23]. Pre [HBLN24, MKRM23]. Pre-existing [MKRM23]. Pre-Readers [HBLN24]. Predict [ARH^{+}23, GHHS21]. Predicting [CVL22, DRP^{+}23, RGR^{+}21, UYSM22, WJH^{+}24, YU20]. Prediction [ALSK23, DWSF24, KR24, MMH21, SMSM22, SBB24]. Predictive [ASP20]. Predictors [SSGB21, vBGH^{+}19]. Prefer [BKJ^{+}20]. Preference [ZPMK24]. Preferences [dACZ19, HJCM22, MDN^{+}24, ZP24]. Pregnancy [ATH24]. Presence [BPO21, CHS^{+}24, GW24, SSB^{+}22]. Present [BWV20, CWB^{+}18]. Presentation [HV22, LUFW20, RS19]. Presenting [DWB18]. Preserving [KSX21, ÖBK24]. Presidential [Kee19]. Press [LZY^{+}21]. Press-n-Paste [LZY^{+}21]. Pressure [BKM21, JPKM24, LZY^{+}21, TFI^{+}22]. Pressure-sensitive [LZY^{+}21]. Prevent [MBT^{+}20]. Prevention [Bel24, RAZ^{+}20]. Primary [BBV19]. Principal [HLV21]. Principal-Agent [HLV21]. Principles [EJB^{+}20, GC19]. print [CKSP22]. Printed [GGL^{+}22, KIK20]. Printing [KSG^{+}24, ZYY23]. Priority [HJC^{+}22]. Prison [ETE^{+}19, TAS^{+}19]. Privacy [AFKL20, AHCD17, AGK^{+}22, AASK23, CHC^{+}17, DB22, FMBW24, GP22, KSX21, KM22, KY24, KBSP^{+}22, LPZV23, LLDH21, LSLD22, ÖBK24, PLY23, RKNES20, SB20b, STSS22, TSD23, WHW^{+}22, WM22, WSM24, YBE^{+}23, BTM^{+}21, WBS^{+}23]. Privacy-Conscious [LPZV23]. Privacy-Preserving [KSX21, ÖBK24]. Privacy-Utility [FMBW24]. PrivacyPrimer [KSX21]. Private [COKL20, DC17]. PrivatEyes [ERH^{+}24]. Proactive [CCS^{+}18, JFL23]. Probability [GSK^{+}17]. Probe [HKM^{+}20]. Problem [BP17a, FCR^{+}17, MTA^{+}17, SCP^{+}21]. Problematic [WFL^{+}22]. Problems [AAHM22, BBV19, SMLH23]. Procedures [MAN^{+}20]. Process [AGO^{+}24, CMN^{+}21, CMP^{+}24, CMP^{+}21, CNH^{+}17, PAC22, WNT^{+}22]. Processes [IYS23, MES^{+}22, SSB^{+}24a]. Produce [KDEA21]. Produced [Kee19]. Product [BHM^{+}23]. Production [BZC^{+}17, CSMG17, Hig20, LWTW23, dCRM^{+}23, WNT^{+}22]. Productivity [LAR21, LWB22, WBSM24]. Products [SSB^{+}24a]. Professional [LYH20, LMZ^{+}24]. Profiles [SSD23]. Program [XKRW23]. Programmers [BSA23]. Programming [BahRKM18, BPO21, FSAN23, LZS^{+}21, MQX^{+}23, MCC^{+}18, RH19, SS22, Sal17, SRP^{+}24, TDC17, WBSM24]. Programs [BYL^{+}24]. Progression [JKZ^{+}22]. Project [BL17, Gil20, AMJ^{+}19, DGR^{+}22, SJ21]. Projection [GB18, MSK22]. Projections [KFSC18]. Projects [LPFD21, SC24, WWSG21]. Prolonged [LMB^{+}22]. Promote [BGRK17, CMSA22, ETE^{+}19, LOSD20]. Promoting [dACZ19, KKK22a, LYH20]. Prompts [DAQ17]. Propensity [JPKM24]. Properties [CPB18, PB22]. Proprioceptive [BBW21]. Props [MLSH21]. Prosocial [ITKB23]. Prospects [CCWH21]. Protagonist [WGS^{+}24]. Protecting [DB22]. Protection [BUSA^{+}21]. ProtoChat [CMP^{+}21]. Protocol [KE18]. Protocols [BahRKM18]. Prototype [GN23, ZSK^{+}23]. Prototypes [DMPK22, MQX^{+}23]. Prototyping [HDPL21, KKE^{+}21, KVDG20, RH23, SVS^{+}24, SLN21]. Provenance [AR22, ZES^{+}23]. Provider [KYB21]. Providers [BLH^{+}17, CNM^{+}23, FSC21]. Providing [BPO21]. Proxemic [LPW^{+}23]. Proxemics [LGG^{+}23]. Proximity [GKHD23, SBSM24]. Proximity- [SBSM24]. Proximity-Input-Based [GKHD23]. Proxy [DHH18, RDF^{+}22]. Psycho [CBV22]. Psycho-Social [CBV22]. Psychophysiological [BDL21]. Psychosocial [ETE^{+}19]. Psychotherapy [YBE^{+}23]. Public [AL18, AKIS21, Fie19, GPC^{+}23, Gil20, JRP^{+}23, MD20, RAA22, STM^{+}19, TAS^{+}19, WS22]. Publications [KKS19]. Pull [LO17]. Pupil [FH24]. Pupils [ZMS^{+}22]. Puppetry [LO17]. Purposes [KUD^{+}23]. Push [DRL^{+}22]. Puts [SS21]. PWA [dACZ19]. PWA-EU [dACZ19].

Q [RDF^{+}22]. Q&A [CXL17, CCS^{+}18, CT18]. Q&As [DMC^{+}20]. QAnon [EHZN22]. Qualitative [BWDP20, LE21]. Quality [ADD21, CCS^{+}18, WYY^{+}21, ZBV23]. Quantify [WS23]. Quantifying [TSS21, SK23]. QuantumLeap [SORV22]. Quaranteens [PAC22]. Queer [FA22, HTTH22]. Queerness [HTTH22]. Querying [HOZ^{+}21]. Question [MCYZ19, MWM^{+}19, MMH21]. Questionnaires [RDF^{+}22]. Questions [DMC^{+}20]. Quests [FM22]. Quilters [MWH19]. quo [GN23]. Quotidian [AL18].

r [RJZ^{+}21, Gil20]. r/AskHistorians [Gil20]. r/Incels [RJZ^{+}21]. r/The Donald [RJZ^{+}21]. Rabbit [TBHM23]. Racial [MGG^{+}21]. Racism [TSHK20, WWSS22]. Racist [WS23]. Radar [ASZW23, CST^{+}23, PAK^{+}22]. Ramadan [CEJC23]. Ranked [XOEK22]. Rapid [KVDG20, KMH23, RH23]. Rapidly [MTA^{+}17]. Raters [BZSM20]. Rates [UYM21]. Rationality [YK23]. Raw [KR24]. Re [AWM21, CEJC23, CJ23, FZRJ22, LFK24, WWSS22]. Re-Enforcing [WWSS22]. Re-locations [FZRJ22]. Re-Working [AWM21]. Reachability [YRIV20]. React [ZRL21]. Reactions [SHL23]. Readers [HBLN24, SELS24]. Reading [DRP^{+}23, JCD19, KKP^{+}22, LWC^{+}23]. Real [BBW21, BYL^{+}24, CWC^{+}24, DDZ18, FLC17, GHHS21, KR24, LYH20, MQX^{+}23, MBK^{+}22, MRRS19, NMPDJ24, PCMP18, RWB22, SBPP20, SSNC20, WMBO19]. Real-Time [BBW21, BYL^{+}24, DDZ18, MQX^{+}23, MRRS19, PCMP18, RWB22, SBPP20, WMBO19, FLC17, KR24, NMPDJ24]. Real-world [CWC^{+}24, MBK^{+}22]. Reality [ALG^{+}21, ACF^{+}22, AARJ23, BBW23, BLC^{+}21, BRM21, BHH^{+}22, CWC^{+}24, CTWM23, CDIH23, CKK22, DGR^{+}22, DR24, ECP^{+}24, FSN^{+}20, FA22, FAMS22, GGE23, GGL^{+}22, HPH^{+}23, HCM^{+}23, HCG^{+}22, HRP23, HOSK21, JDX24, KMM21, KVM23, KMM^{+}22, LG22, LZY^{+}21, LL24, LWB22, LPW^{+}23, LGG^{+}23, LRP^{+}22, MAN^{+}20, MP23, MGRM22, MdLGG23, MBK^{+}22, MLSH21, MCDR22, NAO23, OG21, PDPS22, Poh24, Pri19, RYCC21, SRZ24, SY24, SFLB23, SAR^{+}23, SCMS18, SWQ^{+}23, SCY^{+}18, SHY^{+}18, SLN21, STM24, SHC^{+}23, TRO^{+}21, TG18, VMHO^{+}22, WZFA22, WVG^{+}24, WO20, WLZL23, ZKL22]. Really [VHZ22]. Reappropriation [CA22]. Recallability [WJH^{+}24]. Receipt [PLY23]. Receipts [CEM17]. Recidivism [vBGH^{+}19]. Reciprocal [JKC23]. Recognition [ASMB22, ASZW23, BKG^{+}21, CS18a, DS20, DWZ^{+}20, DMRA24, DTE^{+}22, GRG20, MFSK22, PLZ^{+}22, RA21b, SVS^{+}24]. Recognition-Enabled [RA21b]. Recognize [KKV^{+}22]. Recognizer [MRV22]. Recognizing [CST^{+}23, MH19, OSM^{+}20, SAY^{+}24]. Recommendation [AHJV22, HvBKG20]. Recommendations [BBNT21, TD19]. Recommender [MDN^{+}24]. Recommending [CRN21, MP23]. Reconsidering [HFBL19, See20]. Recording [AV19]. Recovery [HMK24]. Recruitment [AOT^{+}18, KOOL19]. Rectangular [UYSM22]. Reddit [CPS^{+}17, CSJ^{+}18, EHZN22, FZP^{+}24, GKS21, Gil20, LLDH21, MPKEN24, PSTK24, WWSS22, XZL^{+}20]. Redditor [SS21]. Redirection [WVG^{+}24]. Reduce [AIRH24, JZB^{+}21, RVBC^{+}21, SRKK22]. Reducing [FX22]. Reduction [CCW22]. Reference [vDBL22]. Referencing [CMY^{+}17, CWB^{+}18]. Referential [MHN^{+}21]. Reflect [GBN21, PSTK24]. Reflection [BLL^{+}21, ICB^{+}22, LLAK23]. Reflections [ERT21]. Reflective [GMB23, HTTH22]. Reflects [JAB19]. Reframing [BKJ22]. Refugee [ECA23, HV23]. Region [SWQ^{+}23]. Regulate [KA23]. Regulated [KHY^{+}22]. Rehabilitation [AGM^{+}23, CMP^{+}24, HRC^{+}21]. Reinforcement [LCA^{+}24]. Related [BDB^{+}23, CSL^{+}24]. Relating [SBAK21]. Relation [ARH^{+}23]. Relational [MT21, WGG20]. Relationship [BBD^{+}24, HKD23, SK20]. Relationships [ABQ17, HLV21, LFK24, MFM21]. Relative [DWMV21]. Release [SHL23]. Relevance [ICB^{+}22]. Reliability [HWGV24]. Reliable [KMR20]. Relief [JKC23]. Religion [MPKEN24]. Religious [CEJC23]. Remapping [VP24b]. Remember [FZP^{+}24, LAM22]. Remote [ALG^{+}21, FZRJ22, JP22, KSG^{+}24, MMM^{+}21, PBC20, PKLK22, RIK21, SRAG22, SAR^{+}23, WBS^{+}23, ZMS^{+}22]. RemoteCoDe [SRAG22]. Removal [RGW^{+}24]. Rendering [ALSK23, HTS^{+}22]. Repair [SSH^{+}21]. Repairing [CLL^{+}21]. Repetitive [RJVS24]. Replay [MON^{+}21]. Report [AL18]. reported [BBD^{+}24]. Reporting [APW^{+}21, JKW^{+}21]. Reports [RAZ^{+}20]. Represent [NM21]. Representation [MON^{+}21, MGG^{+}21]. Representations [BLC^{+}21, SAR^{+}24]. Reproducibility [FSN^{+}20]. Reproducing [BV21]. Requirements [SWT20]. Research [BUSA^{+}21, BHH^{+}22, BL17, CLKEF21, CCH^{+}18, FSN^{+}20, Fie19, FZP^{+}24, HTS24, MRK^{+}22, NHH20, PK23, PSTK24, See20]. Researchers [GPC^{+}23, RH22, ZD19]. Reshapes [SHL23]. Resident [SAE^{+}21]. Residents [CRM^{+}24]. Resilience [AKVHZ23]. Resistance [KDEA21]. Resocialization [ETE^{+}19]. resolution [SJ21]. Resource [RIK21]. Resources [CH20, PNS24a]. Respond [CNM^{+}23]. Responders [NHL^{+}24]. Response [BL18, HKF19, JFL23]. Responsible [MNS^{+}21, RYCC21]. Responsive [LZS^{+}21, XHH17]. Results [MGG^{+}21, UYSM22, VP24a, VH21]. Retention [BGRK17]. Retest [HWGV24]. Rethinking [BP17b, Hao21, YU20]. Retinal [MSK22]. Retrieval [GMB23, SBSM24]. Returning [BSS18]. Retweet [VMP21]. Reusable [JV22]. Reuse [MRK^{+}22]. Reveal [ÇH24]. Revealing [UHH23]. Reveals [GSC21]. Reversion [GPH19]. Review [APW^{+}21, ASO^{+}22, AHJV22, DWW^{+}17, DFJ^{+}21, FZP^{+}24, FOH^{+}22, GCSB23, GM23, LOSD20, ND24, NAO23, SFA^{+}23, SCR22, STM24, WRM^{+}24, dRSV22]. Reviews [DAQ17]. Revise [KYB21]. Revisiting [CA22, SB17]. Reviving [CJ23]. Rewards [FM22]. RFID [DWZ^{+}20, MdOKT23]. Rhetoric [DDV23]. Ride [MH19, MMMM22]. Ride-Share [MH19]. Ride-Sharing [MH19]. Right [MWM^{+}19]. Rights [GYD17]. Ring [KVK^{+}22, LZW23]. Ring-Shaped [LZW23]. RingVKB [LZW23]. Rinsta [HV22]. Rise [LX20]. Risk [MWK^{+}22, SMSM22]. Risks [CBSQ24, LPZV23, SFLB23]. Risky [BUSA^{+}21, CBSQ24]. Road [DCM24, STM24]. Roblox [ZMGK24]. Robot [DRDS24, JJJ^{+}23, KUWM22, SVS^{+}24, VP24a, ZRA22]. Robot-Assisted [DRDS24]. Robotic [BLV23, GTO^{+}23, SRAG22, WWH^{+}23, vDBL22]. Robotics [HC24, PGK^{+}24]. Robust [LWC^{+}23, MSK22]. Role [AL21, CCX^{+}20, CR20, FSN^{+}20, HMK24, IRA^{+}19, KYB21, LLC^{+}24, OR19, PSM21, See20, VKI21, WGS^{+}24]. Roles [LLGH24, LDP18, PBC20, PM21, ZMW20]. Romantic [ABQ17]. Room [GTO^{+}23, HRP23, NFTK22]. Room-Scale [GTO^{+}23, HRP23, NFTK22]. Rooms [MKRM23]. Rotary [SSR22]. Rotation [DYF^{+}23, LKYK24, MRV22]. Routes [LRK^{+}24a]. Routines [ZBL19]. Routing [MCYZ19]. RPG [LKYK24]. Rule [FSAN23]. Rule-based [FSAN23]. Rules [CSJ^{+}18, JS22, PVC^{+}23]. run [Gil20]. Running [SSB24b]. Runtime [MNP^{+}20]. Rural [SAR21, WGG20]. Rwandan [HMB^{+}21].

SaaS [BPB22]. Safe [CNM^{+}23, RAZ^{+}20]. SaferHome [WHW^{+}22]. Safety [AGK^{+}22, AZ23, AL18, DCM24, TSD23]. Saliency [DWSF24, SBB24]. same [BBK22]. Sample [AOT^{+}18, DS20]. Samples [CST^{+}23]. Sampling [CLP^{+}21, GM23, KVM24]. Sanctioning [RKNES20]. SAPIENS [SVW19]. Sarcasm [OM20]. Satisfaction [GW24, YHR^{+}19]. Saudi [ADAY21, RYD21]. Savings [MAR^{+}19]. say [CNH^{+}17]. Scaffolding [CMRH22, RIC^{+}24]. Scalable [BP23, ZBL19]. Scale [GTO^{+}23, HRP23, KMV^{+}24, LKS21, MB22, NFTK22, POG^{+}21, XRS22]. Scales [CSJ^{+}18]. Scaling [MRV22]. Scandinavia [JAB19]. Scanpath [BMR^{+}23, EJGL24, ÖBK24]. Scanpaths [DRP^{+}23, KSP^{+}23]. Scarcity [BBK22]. Scenarios [CWC^{+}24, RZP^{+}22, SSR22]. Scene [MON^{+}21]. Schizophrenia [ERB^{+}17]. Scholarship [Gil20]. Science [JØC^{+}20, JSKA22, MHG^{+}21, ZMW20]. Sciences [LKS21]. Scientific [AR22, BP17a, CRN21, FSN^{+}20, HRRS20, KKS19, MWM^{+}19, MRRS19, NHH20]. Scientists [KG21, MWM^{+}19, WMBO19].

Score [DRK18, JWK22, JCD19]. Score-Group [DRK18]. Screen [BPO21, GB18, YU20]. Screening [HBLN24]. Screens [BCCV17, DWMV21, KNT22]. Scripting [HRRS20]. Sculptures [Hao21]. Search [ACA22, ERLW17, HJM20, KKM24, MGG^{+}21, VH21]. Searching [BBSV23, ZZK^{+}22]. Seat [RI20]. SeatmateVR [LPW^{+}23]. Second [BMdS^{+}22]. Secure [ERH^{+}24]. Security [BBK22, GP22, KY24, TBL^{+}24, TPNL22, WBS^{+}23]. See [BKG^{+}21, VHZ22]. Seeing [CSS^{+}17, MXYS23, SJ21, SAR21]. Seek [HWYW21]. Seeking [KBFK21, KYB21]. Segmentation [YUSM22]. Selection [BDL21, DRL^{+}22, IYS23, JFS^{+}23, KMM^{+}22, LPGG22, NMPDJ24, PA22, SWQ^{+}23]. Selection-based [LPGG22]. Selective [DRA^{+}17]. Self [BR20, BLV23, BBD^{+}24, BRM22, CNH^{+}23, CMSA22, ERVR^{+}23, FCR^{+}17, FCE^{+}18, HV22, KRMM21, KHY^{+}22, LYH20, LUFW20, MMMM22, OS24, PAC22, See20]. Self-Attach [BR20]. Self-Disclosure [CMSA22, LYH20]. Self-Driving [MMMM22, BLV23]. Self-Management [BRM22, CNH^{+}23]. Self-Moderation [See20]. Self-motivation [OS24]. Self-organized [ERVR^{+}23]. Self-Presentation [HV22, LUFW20]. Self-Quaranteens [PAC22]. Self-reported [BBD^{+}24]. Self-Similar [KRMM21]. Self-Tracking [FCR^{+}17, FCE^{+}18, OS24]. Selfie [YPS^{+}22]. SEME [ERLW17]. Semi [JFS^{+}23, NAS^{+}22]. Semi-automated [NAS^{+}22]. Semi-Direct [JFS^{+}23]. Sender [COKL20]. Sender-Controlled [COKL20]. Sending [AMF18]. Sense [BSS18, GW24, KKK22b, MBK^{+}22, PSTK24]. Sensemaking [FGG17, MNS^{+}21, YDW^{+}22, ZES^{+}23]. Sensing [FLC17, GBZ^{+}24, IB24, PAK^{+}22, VMD^{+}20, ZBL19]. Sensitive [HKF19, LZY^{+}21, RSL24]. Sensitivity [BS21, CDIH23]. Sensor [AFKL20, MSK22, MNS^{+}21]. Sensors [EPW^{+}23]. Sentiment [JSJ22]. September [ACI22, GMM21, KvBW24, LMS23]. Sequence [ASZW23, DRP^{+}23, HMML^{+}22]. Sequences [TRN^{+}22, TBR20]. Sequential [ZWG^{+}23]. Serendipity [HTTH22]. Series [YYT^{+}24]. Serious [AL21, CGS^{+}21, GH21, KML^{+}21, RVH24, RDF^{+}22, STSS22]. Sermons [RAA22]. Server [JS22]. Service [CNM^{+}23, SFA^{+}23, WNT^{+}22]. Services [BBV19, BDE^{+}24, GN20, HLV21]. Session [SB17]. Sessions [AGO^{+}24]. Sets [FLC20]. Setting [AAHM22, ALD^{+}20, DTE^{+}22, KYB21]. Settings [IB24, JSKA22]. Sexual [ZMTH22]. Shape [FSRC24, PCL^{+}21, LFK24]. Shape-changing [PCL^{+}21]. Shaped [LZW23]. Shaping [JS22]. Share [COKL20, LPZV23, MH19, MDN^{+}24]. Shared [AHCD17, FBG17, FZRJ22, MAS22, NSV22, PGK^{+}24, RCM21, SFM^{+}22, WWH^{+}23]. Sharing [BWE^{+}22, BBNT21, BPO21, COKL20, DWW^{+}17, EMHW23, EJB^{+}20, HLV21, JZB^{+}21, LAM22, MH19, MRK^{+}22, NSV22, NKV21, PBC20, PLL^{+}23, SELS24, VKI21, WFL^{+}22]. Shear [TFI^{+}22]. Shed [VQ23]. Sheikh [ESA^{+}24]. Shielding [ESA^{+}24]. Shift [RIK21]. Shifting [DWSF24, JØC^{+}20, RYCC21, WVG^{+}24]. Shooting [HSH22]. Shopping [KKK^{+}24, PNS^{+}24b]. shops [IRA^{+}19]. Short [Pri19]. Short-term [Pri19]. Shortcuts [RRA23]. Should [FLC17]. Shoulder [GBZ^{+}24, WVG^{+}24]. Shoulder-Surfing [WVG^{+}24]. Shouldn’t [LPZV23]. Side [FLC20, YRIV20, CLP^{+}21]. Side-Crossing [FLC20]. Sides [MKRM23]. Sighted [LC24]. Sightseeing [MGRM22]. Signals [HMML^{+}22, MFM^{+}20, ZWZ^{+}21]. Signature [ZWZ^{+}21]. Significant [WBSM24]. Signifier [MCM23]. Silencing [ESA^{+}24]. Silent [PA22]. Similar [KRMM21, TKB^{+}22]. Similarities [HGG20]. Similarity [HMML^{+}22]. Simplified [LAHW19]. simply [Pri19]. Simulate [HOSK21]. Simulated [AM21]. Simulating [HTS^{+}22, KVM24]. Simulation [SM22]. Simulation-Based [SM22]. Simulator [MMMM22]. Single [NFTK22]. Site [DMC^{+}20, Gil20]. Sites [CXL17, CCS^{+}18, MD20, MGRM22]. Situ [KFSC18, SHC^{+}23]. Situated [BHM19, CMRH22, GEEJ23, KG20, MT21]. Situating [JSKA22, RAA22]. Situational [ZAS^{+}21]. Situations [HLV21]. six [BBK22]. Size [FH24, KKM24]. Sized [CHS^{+}24]. Sizes [UYM21]. SketchADoodle [GVNK20]. Sketched [SB20a]. Skewed [WM22]. Skill [CKS18, DR24, HvBKG20, Hig20, MON^{+}21, MHG^{+}21, RSL24]. Skill-Based [DR24]. Skills [KMV^{+}24]. Skin [BP23, BP23]. Sleep [KVK^{+}22]. Slidable [SVV21]. Sliders [GSK^{+}17]. Slime [SVV21]. Slogans [AOT^{+}18]. SMAC [LAHW19]. Smala [MCC^{+}18]. Small [BCCV17, FLC20, KBFK21]. Smart [CNH^{+}23, CWK^{+}23, Gar21, LPZV23, MP23, MFSK22, SVW19, TDC17, TPNL22, VHZ22, WHW^{+}22, WM22, ZMZL21]. Smartest [GSC21]. Smartglasses [AV19]. Smartglasses-based [AV19]. Smartphone [CST^{+}23, IOES20, KSKH24, NFTK22, TFI^{+}22, ZRA22, ZWZ^{+}21]. Smartphone-attached [CST^{+}23]. Smartphone-based [ZWZ^{+}21]. Smartphones [JF24, LLKC22]. Smartwatch [FSRC24, RLL22]. Smartwatches [RA21a, RZP^{+}22, SSR22]. Smile [BRM22]. SnappView [dRSV22]. So-called [BTM^{+}21]. Sochiatrist [MFM^{+}20]. Sociability [WWH^{+}23]. Social [AMR21, ASJK23, AZ23, AS23, AMF18, Ann20, AWM21, ALD^{+}20, AM21, BSS18, BBK22, BWDP20, BBNT21, BPO21, CNM^{+}23, CEM17, CSMG17, CH20, CCX^{+}20, CBV22, CHC^{+}17, DWB18, DBS^{+}21, DRAQ18, EJB^{+}20, ERB^{+}17, ETE^{+}19, FWY^{+}22, FB19, FA22, FAMS22, GYD17, GAW19, GSC21, GPH19, GPH21, GN20, HLV21, HWF^{+}24, Hao21, HFL19, JZB^{+}21, JSJ22, JJJ^{+}23, KKP^{+}22, KBFK21, LKS21, LSLD22, LNM21, MCYZ19, MAMP21, MBK^{+}22, PSM21, PK23, RKNES20, RA21b, SKV^{+}21, SSB^{+}22, SSH^{+}21, SSD23, SFLB23, SPT^{+}20, TSHK20, TCH20, VVBK23, VKI21, WWT^{+}21, WYY^{+}21, WXL^{+}22, XKRW23, YBE^{+}23, ZWG^{+}23, ZD19]. Social-Spiritual [KKP^{+}22]. Social-Support [WWT^{+}21]. Sociality [BRN^{+}19]. Socially [AMF18]. Societal [JAB19]. Society [MGG^{+}21]. Socio [AKVHZ23, KBKH21, SCP^{+}21]. Socio-Ecological [KBKH21]. Socio-Political [SCP^{+}21]. Socio-Technical-Natural [AKVHZ23]. Sociocultural [OM20]. Sociotechnical [ESCR23, LLGH24, RYD21, WBS^{+}23]. Soft [CWK^{+}23, FSRC24]. Software [BHH^{+}22, CMN^{+}21, CLKEF21, CAATL22, LPFD21, MMdSP21, MdOKT23, OGRV22, SC24, SVW19, TG18, WWSG21, dRSV22, dSdSdSF^{+}23]. Soli [PAK^{+}22]. Solicitors [PM21]. Solidarity [NKV21]. Solids [PAK^{+}22]. Solitary [JJJ^{+}23]. Solo [PF24]. Solutions [CWR22, RJVS24, RS19]. SOLVENT [CCH^{+}18]. Solving [BP17a, MTA^{+}17]. Some [LC24]. Someone [PKC^{+}21]. Something [WS23]. Sometimes [BRM22]. Soul [BVM21]. Sound [CSS^{+}17]. Source [BSM23, HWF^{+}24, IB24, LPFD21, SC24, WWSG21]. South [KOL^{+}24]. Space [BAAP20, HV23, KDH^{+}21, LKG21, NSJ21, PCL^{+}21, WLZL23]. Spaces [AKIS21, FZRJ22, KKK22a, KF23, LBM21, SEC^{+}20, WWH^{+}23, WDD^{+}22, dCRM^{+}23]. Spacetime [DDZ18]. Spark [SKV^{+}21]. Sparse [OKM^{+}23]. Spatial [LWB22, LPED22, MdLGG23, POG^{+}21, SR22, TRO^{+}21, TCH20, XMS24]. Spatially [HCG^{+}22]. Spatially-Aware [HCG^{+}22]. Speak [WWH^{+}23]. Speaker [LYD^{+}23, ZRP^{+}22]. Speakers [Gar21, YYT^{+}24]. Speaking [RAA22]. Special [ZMS^{+}22]. Specifying [BF17]. Spectator [WACdA^{+}21]. Spectators [WACdA^{+}21]. Speculation [CSL^{+}24, KF23]. Speculative [Fie19]. Speculatively [KM22]. Speech [CPS^{+}17, JKZ^{+}22, LLKC22, MIS^{+}20, PA22, WO20, ZMR21]. Speech-Based [JKZ^{+}22, PA22]. Speed [ASP20, RLGG21]. Speedrunning [SS19]. Speeds [SRKK22]. Spillover [NSJ21]. Spinal [BHNA18]. Spirit [CEJC23]. Spiritual [KKP^{+}22]. Spit [GP22]. Spoil [CT18]. Spontaneous [LUFW20]. Sport [SSB24b]. Spread [Wil20]. SQUAMATA [CMN^{+}21]. Squeeze [VNSVL22]. SSIO [PK23]. stable [CKSP22]. Stage [SLN21]. Staged [LDP18, CS18b]. Stairlifts [SAE^{+}21]. Stakeholder [SCP^{+}21]. Stakeholders [XDC^{+}24]. start [YU20]. State [CLV^{+}19, JPKM24, LX20]. State-Owned [LX20]. Statement [CNC^{+}17]. States [CRN21]. Stations [KSF^{+}23]. Status [BL17, COKL20, GN23]. Statuses [HWF^{+}24]. Stay [CPS^{+}17, GPH19, LLC^{+}24]. Stay-At-Home [LLC^{+}24]. Steering [HK24b, YTF^{+}23]. Step [AMR21, CRM^{+}24]. Stepper [RA21a]. Stereotype [BKG^{+}21]. Stewardship [TBL^{+}24]. Sticker [CKP^{+}18]. Sticky [BGRK17]. Stigmatized [AMF18]. Stimulation [CCW22]. Stoop [KF23]. Stories [JA21]. Story [MRK^{+}22, MKRM23]. Storyboards [HWS21]. Storytelling [LAM22, MRK^{+}22]. stranger [BLH^{+}17]. Strangers [SK20]. Strategic [BWDP20]. Strategies [AHJV22, DNBB^{+}22, HK24a, HFL19, HRR^{+}22, KOL^{+}24, LPZV23, MSHP23, MAD^{+}23, RKNES20, XYDV23]. Strategy [SS22]. Stream [GF22]. Streamed [SS19]. streamer [CCWH21]. Streamers [JSP^{+}21]. Streaming [CCWH21, LFM22, LHPH20, WLZL23]. Streams [CEM17, ZWG^{+}23]. Strenuous [BRM21]. Stress [CCW22, GSC21, HMK24, RVBC^{+}21, ZBL19]. StressMon [ZBL19]. Strip [FRH^{+}22, HGG20]. Strips [HGG20]. Stroke [CFD^{+}21, CS18a, HRC^{+}21, GVNK20, MRV22, OSM^{+}20]. Stronger [HSVR21]. Structural [MDM^{+}19]. Structure [TSD23]. Structured [FWY^{+}22]. Structures [VP24b]. Structuring [PCL^{+}21]. Struggles [FCE^{+}18]. Struve [MPEC24]. Stuck [DFJ^{+}21, LRSS20]. Student [ARH^{+}23, FYS^{+}24, KR24, MMH21, MKRM23, TSD23, XOEK22]. Students [LKS21, LLAK23, RVH24, SAM23, WGS^{+}24]. Studies [CWK^{+}23, MV19]. Studious [LUFW20]. Study [ABC21, ALG^{+}21, AZ23, ADD21, AKIS21, BKM20, BEHB^{+}23, BKKD21, BS21, BHNA18, CSJ^{+}18, CCX^{+}20, CSH^{+}23, CNH^{+}17, CHC^{+}17, FB19, Gil20, GM23, DRDS24, HKM^{+}20, HJM20, KOL^{+}24, KFSC18, KVK^{+}22, KB22, LLDH21, LAW19, LC24, MAN^{+}20, MIS^{+}20, MB23, MBS22, NHH20, PDPS22, PVC^{+}23, RWB22, SSD23, SLB23, SSR24, TRN^{+}22, WPH22a, WGGS20, WDD^{+}22, WLZL23, XZL^{+}20, XJBH22, ZP24, vBGH^{+}19, DB22]. Studying [ASM23, BB20, LHT^{+}20, SB20b]. Style [EHG22, GAW19, HKL^{+}22, RDF^{+}22]. Styles [QGB20]. Subgoal [JK24]. Subject [BL17, RA21b]. Subjective [AASK23, CLP^{+}21, GBN21]. Subjectivity [MSY20]. Subreddit [SS21]. Subtletee [WDP^{+}20]. Success [GAW19, UYM21]. Successful [OR19]. Suicide [RAZ^{+}20]. Summarization [BSM23]. Super [JPKM24]. Supervised [BMR^{+}23]. Support [BWFG23, BMPP20, CPB18, FCR^{+}17, FZRJ22, GKS21, GB18, HKM^{+}20, JKC23, KSX21, KKP^{+}22, KHY^{+}22, KTBL^{+}20, LLC^{+}24, LW19, LM20, LDP18, MES^{+}22, MMM^{+}21, MRK^{+}22, MRRS19, NHH20, PF24, PCMP18, PVC^{+}23, PLL^{+}23, RLL22, RMB^{+}22, SVW19, SB17, TSD23, WWT^{+}21, XKRW23]. Supported [KSF^{+}23, MPB^{+}19, TG18, AGM^{+}23]. Supporting [BLL^{+}21, BYL^{+}24, BWDP20, BBS19, BHNA18, CFD^{+}21, CR20, CMP^{+}21, CHS^{+}24, CAATL22, ETS^{+}23, FMBW24, GTO^{+}23, HRRS20, LYL^{+}23, LMB^{+}22, MBK^{+}22, OTH20, PP23, See20, TDC17, XHH17, dRSV22, vDBL22]. Supportive [SHL23]. Suppressing [ERLW17]. Surface [EMHW23, GTO^{+}23, GRG20, LZY^{+}21, LDM^{+}22, MDM^{+}19, GVNK20, MMI^{+}22]. Surface-Based [LDM^{+}22]. SurfaceCast [EMHW23]. Surfaces [BGM^{+}23, CWK^{+}23, CYGW21, ERH^{+}22, FX22, FBG17, FLC20, KE18, XHH17]. Surfacing [XOEK22]. Surfing [WVG^{+}24]. Surgery [MON^{+}21]. Surgical [MON^{+}21]. Surrogate [YK23]. Surround [GB18]. Surround-Screen [GB18]. Surrounding [MKRM23]. Surveillance [BTM^{+}21, WS22]. Survey [AM21, BDW21, DMPK22, LAD^{+}22, RDF^{+}22, SSB^{+}22, SSGB21]. Survey-Based [SSGB21]. Sustainable [CMRH22, JKC23, NPT19]. Swipe [RA21a]. Switching [CYW18, DWMV21, HSH22, HNS^{+}24]. Symbiosis [CR20]. Symbolic [PBC20]. Symmetrical [BWFG23]. Sympathizers [PM21]. Sympathy [ITKB23]. Sympathy-Eliciting [ITKB23]. Sync [BBD^{+}24]. Synchronization [ÇH24]. Synchronized [MON^{+}21]. Synchronized-Video [MON^{+}21]. Synchronous [NSJ21, Pri19, RIC^{+}24, WQCZ24]. Synchrony [BBD^{+}24]. System [AKVHZ23, AV19, ACA22, Blo24, BDW21, BT23, BP17b, BGRK17, CRN21, CCH^{+}18, CYW18, CBPC23, DTE^{+}22, HKM^{+}20, HBM23, HvBKG20, IB24, JK24, LNM21, MCYZ19, MON^{+}21, dQPMdHRdAN21, RH19, SBPP20, SBSM24, SLD21, VN19, ZSK^{+}23, ZRL21, ZD19]. System-Facilitated [CYW18]. Systematic [APW^{+}21, ASO^{+}22, AHJV22, CMP^{+}24, CWR22, DWW^{+}17, DAQ17, DFJ^{+}21, FZP^{+}24, FOH^{+}22, MdOKT23, NAO23, SCR22, STM24, WRM^{+}24]. Systems [AHJV22, BWV20, CPB18, CRM^{+}24, CSN19, FWY^{+}22, FLP^{+}21, HGG20, JVP^{+}19, LLGH24, LDP18, LPQW22, LS23, MES^{+}22, MdOKT23, MRRS19, MDN^{+}24, NPT19, PB22, SD24, TG18, TBL^{+}24, TBR20, WRM^{+}24, WPH^{+}22b, YK23, vDBL22].

Table [MMM^{+}21, RI20]. Tabletop [MTA^{+}17, WXL^{+}22]. Tabletops [MRGN20, PDPS22]. Tabloids [CSMG17]. Tactile [BHM^{+}23, CYGW21, GRG20, RLGG21]. Tag [AV19, CT18]. Tagging [AZS^{+}22, BKG^{+}21, SAR^{+}24]. Tags [AV19, AWM21]. Tailored [LDP18]. Take [ZRL21]. taking [FWY^{+}22, KKK22a]. Tale [MPC^{+}23, PLY23]. Talk [LLDH21]. Talking [BDE^{+}24, LC24]. Tame [SHS22, WRM^{+}24]. Tandem [MMMM22]. Tangible [AFKL20, ERH^{+}22, FSSK22, KE18, MDM^{+}19, MMI^{+}22, MTA^{+}17, SHC^{+}23]. Tangibles [SM22]. Tap [DRL^{+}22]. Taps [YRIV20]. Target [KMM^{+}22, ZLB^{+}22]. Targets [ERH^{+}22, RMB^{+}22, UYM21, UYSM22, UYS23]. TARPS [KY24]. Task [AZS^{+}22, BDW21, BF17, BBD^{+}24, CFG^{+}17, DBS^{+}21, DBL22, HBJR21, HvBKG20, HK24b, JOFM24, MPB^{+}19, MHN^{+}21, PCMP18, PDPS22, RSL24, SAR^{+}24, SWT20, SGF20, UYSM22, WYY^{+}21]. tasking [SBPP20]. Tasks [AGM^{+}23, BZSM20, DR24, SRAG22, WO20, WRF^{+}22, YU20, YOFC19, Zha23].

Tasks-based [AGM^{+}23]. Taste [BEHB^{+}23]. Taster [BEHB^{+}23]. Taxing [MYS^{+}21]. Taxonomy [AM21, MCDR22]. Teachers [GLZ^{+}21, ZMS^{+}22]. Teaching [BK24, KMV^{+}24, LMT^{+}22, ZMS^{+}22]. Team [BDB^{+}23, CYC^{+}21, HHVB22, MAD^{+}23, TRN^{+}22]. Teams [ADD21, BKM20, BS21, CYC^{+}21, KKK22a, KJH19, MAD^{+}23, SFM^{+}22, TRN^{+}22, dSdSdSF^{+}23]. Teamwork [JKZ^{+}22, MKRM23, ZSK^{+}23]. Tech [MWBB23, MB23]. Technical [AKVHZ23, VMD^{+}20]. Technique [ASJK23, CKK22, LKYK24, NFTK22, SSB24b]. Technique-Oriented [SSB24b]. Techniques [CJ23, KMM^{+}22, LDM^{+}22, LHI20, MQX^{+}23, RH22, WVG^{+}24]. Techno [CEJC23]. Techno-Religious [CEJC23]. Technological [HTS24, IRA^{+}19, KJH19]. Technologies [ADAY21, Ann20, BBW23, BHM^{+}23, HFBL19, MWBB23, SAM23, TSD23, TSHK20]. Technology [ADD21, Bel24, BDL21, BS21, BKJ22, CFD^{+}21, FHF24, HKM^{+}20, HOHB22, KSF^{+}23, KOOL19, SCSD24, SWF^{+}20, XKRW23, XYDV23, YGHR21]. Technology-based [YGHR21]. Technology-Enabled [SWF^{+}20]. Technology-Mediated [XYDV23]. Technology-Supported [KSF^{+}23]. Tedious [MYS^{+}21]. Teens [AGK^{+}22, BUSA^{+}21]. Teleconsultation [CAATL22]. Telehealth [FLC17]. Telephony [RIC^{+}24]. Telepresence [ZRA22]. Telerehabilitation [CFD^{+}21]. Telework [Tan21]. Temperature [VWHW24]. TEMPEST [BahRKM18]. Templates [ACF^{+}22]. Temporal [ACDC23, MIS^{+}20]. Temporalities [ECA23]. Temporality [HTTH22]. Tennis [MMM^{+}21]. Tensions [DPF22, KPW19, LZAD21, SCSD24]. Term [CWB^{+}18, SMSM22, SMLH23, DB22, HFBL19, Pri19]. Terms [KM22]. Test [CFG^{+}17, GN23, HWGV24]. Test-Retest [HWGV24]. Tester [VZV19]. Testing [AMF18, AV17, CMN^{+}21, GP22, MP19, TBR20]. Tests [BR17, CHCG24]. TetraForce [TFI^{+}22]. Text [ASP20, BB20, JFS^{+}23, LPGG22, RRA23]. Textile [JNLtB22]. Texting [CA22]. Textual [ZGL^{+}22]. Texture [GRG20]. The Donald [RJZ^{+}21]. Their [BYRL23, DTE^{+}22, HWF^{+}24, HWYW21, ACDC23, CS18b, dSdSdSF^{+}23]. Thematic [FOH^{+}22, GW22, GPC^{+}23]. Theoretically [VNSVL22]. Theoretically-Defined [VNSVL22]. Theorie [HKD23]. Theories [EHZN22, KDEA21]. Theory [KKK22b]. Therapeutic [ERB^{+}17]. Therapists [CTH24]. Therapy [CGS^{+}21, HPL22]. There [NAS^{+}22, TPNL22, HKL^{+}22]. Thing [LAM22]. Things [BT23, LAR21]. Think [SFM^{+}22, BRN^{+}19]. Thinking [MCI20]. Thinking-Aloud [MCI20]. Third [PLY23, PKC^{+}21]. thons [MD20]. Thought [RI20]. Thoughts [SHL23]. Threaded [BGRK17]. Threading [BGRK17]. Throughout [DWB18]. Thumb [YRIV20]. Ties [DRAQ18]. TikTok [KDEA21, SHS22]. Tilt [RA21a, ZRA22]. TiltWalker [ZRA22]. Time [BBW21, BYL^{+}24, CEJC23, DDZ18, ECA23, HKF19, JZB^{+}21, KDH^{+}21, KHY^{+}22, KUD^{+}23, MH19, MQX^{+}23, MRRS19, MFM21, PCMP18, RWB22, SBPP20, WMBO19, YYT^{+}24, ZRL21, BLH^{+}17, FLC17, IYS23, KR24, NMPDJ24]. Time-Sensitive [HKF19]. Time-Series [YYT^{+}24]. Times [UYM21, YUSM22]. Tinder [KOL^{+}24]. TL [KM22]. Today [FcLDM19]. Together [AKVHZ23, FAMS22, GF22, HSVR21, JKC23, MWM^{+}19, PK23, SFM^{+}22, SAE^{+}21]. Token [EMT^{+}24]. Tokens [XDC^{+}24]. Toker [LC24]. Tomorrow [FcLDM19]. Tone [CSL^{+}24]. Too [CT18, ICB^{+}22, DWB18]. Tool [AGM^{+}23, CMN^{+}21, GHHC21, KTBL^{+}20, LE21, MPB^{+}19, MP19, NB20, SRZ24, WGGS20, ZMS^{+}22, ZPMK24]. Tool-Supported [MPB^{+}19, AGM^{+}23]. Toolbox [CPB18, KY24]. Tooling [WMM^{+}19]. Toolkit [FSSK22, FBG17, GW22, HCG^{+}22, KKE^{+}21, LMB^{+}22, SM22, TVS17]. Tools [GMB23, GM23, JSJ22, LMT^{+}22, MHP22, RMB^{+}22, WAD^{+}23, WGM21, ZMW20]. Topic [ABQ17, KWdCL^{+}21, MPKEN24]. Topic-Based [ABQ17]. Topophilia [CBV22]. Touch [CYGW21, DWMV21, FX22, GTO^{+}23, GKHD23, GVNK20, HOZ^{+}21, KIK20, LLKC22, MCM23, NM22a, TRO^{+}21, UYSM22, UYS23, YU20, YRIV20, Zha23, JFS^{+}23]. Touch- [GKHD23]. Touch-surface [GVNK20]. Touchable [UYM21]. Touchscreens [LDM^{+}22]. Tournaments [LMZ^{+}24]. Town [JKW^{+}21]. Toxic [CSA^{+}21, WRM^{+}24]. Toxicity [CSA^{+}21, RMB^{+}22, XZL^{+}20]. Toys [Gar21]. Traces [KDH^{+}21]. Tracing [CYGW21, KBSP^{+}22, RSL24]. Trackball [SSR22]. TrackballWatch [SSR22]. Tracking [AASK23, BK24, CHS^{+}24, FX22, FCR^{+}17, FCE^{+}18, HSG^{+}22, JKZ^{+}22, JV22, KCC20, KR24, KVK^{+}22, MSK22, MNS^{+}21, NFTK22, ÖBK24, SC23, OS24]. Trade [BKG^{+}21, FMBW24]. Trade-off [BKG^{+}21, FMBW24]. Tradeoffs [HTS24]. Traditional [SAR21]. Traffic [BLV23, HGG20]. Trail [PPT^{+}22]. Training [DRDS24, KVM23, KMV^{+}24, MNP^{+}22, RWB22, SSB24b]. Trajectories [BKKD21, OSM^{+}20]. Transaction [LRSS20]. Transactional [MB23]. Transactions [LFK24]. Transcription [PLY23]. Transferable [EMT^{+}24]. Transit [BBW23]. Transition [BBS19, GSC21, HSZKN24, LLAK23, OGRV22]. Transitional [PGK^{+}24, SY24]. Transitioning [DRA^{+}17]. Transitions [BSS18, PDPS22]. Translation [MRV22]. Translator [ZGL^{+}22]. Transparency [APW^{+}21, BWE^{+}22, MK23]. Transparent [KNT22]. Travel [WDD^{+}22]. Treatment [BUSA^{+}21, JP22, TBDB21]. Trees [HZ22]. Trends [BDW21, VP24a]. Triage [DAQ17]. Trial [WS22]. Trigger [KMH23, SRP^{+}24]. Trigger-Action [SRP^{+}24]. Troi [DTE^{+}22]. Troubleshooting [KSG^{+}24]. Trust [BYRL23, BDL21, KPW19, KBSP^{+}22, SFM^{+}22, SSH^{+}21]. Trustworthiness [CH20, GPH21]. Trustworthy [BYRL23, KPW19]. Truth [KMR20]. trying [BBK22]. TUIO [KE18]. TUIOFX [FBG17]. Turkey [ECA23]. Turn [HTS24, TCH20]. Tutorial [KMM21]. Tutorial-Based [KMM21]. Tutoring [ZMR21]. Twins [ECP^{+}24]. Twitch [SK23, SK20, ULG^{+}22, WLZL23]. Twitter [BL18, CSMG17, KSW22, KKM^{+}22, LHT^{+}20, MPC^{+}23, NCM^{+}22]. Two [ASZW23, AGM^{+}23, BLB^{+}24, CCW22, CAATL22, MPC^{+}23, PLY23, YRIV20]. Two-Dimensional [YRIV20]. Two-Factor [BLB^{+}24]. Two-Player [ASZW23]. Type [Blo24]. Typealike [CSMV21]. Types [AZS^{+}22, HTS^{+}22]. Typing [HK24a]. Typologies [SSGB20].

U [MFSK22]. U-HAR [MFSK22]. U.S. [Kee19, MM19]. UbiChromics [AIHS22]. Ubiquitous [EMHW23, MBS22]. Ubiquitously [AIHS22]. UbiSurface [GTO^{+}23]. UI [BR17, LYL^{+}23, PKB^{+}19, RH23]. UiLab [BV21]. UIs [LCA^{+}24]. UK [TPNL22]. Ultra [RH23]. Ultrasonic [DRL^{+}22]. Ultrasound [MMM^{+}21, VMHO^{+}22, VWHW24]. Ultrasound-driven [MMM^{+}21]. UMI [CPB18]. Uncertain [GSK^{+}17, UYM21]. Uncertainty [KA23]. Uncommon [DBL22]. Unconscious [SC23]. Unconstrained [WO20]. Undermine [PK23]. Understand [ARH^{+}23, GKS21, PPT^{+}22]. Understanding [Ann20, BSS18, CKP^{+}18, CSMG17, CCX^{+}20, CCWH21, CRM^{+}24, CLP^{+}21, CWB^{+}18, DTE^{+}22, EHH^{+}18, FWY^{+}22, FKDG21, Gar21, GF22, HOHB22, KUWM22, KSF^{+}23, KOOL19, KBSP^{+}22, LTP23, LW19, LWS24, MPKEN24, MDN^{+}24, NM22b, PAC22, RZP^{+}22, RCM21, SGF20, Tan21, TPNL22, VVBK23, WO20, YK23, ZSK^{+}23, Zha23, ZMS^{+}22]. Unemployment [GKS21, SMSM22]. Unified [KSP^{+}23, TVS17]. Unifying [RDS18]. Unintended [CVL22]. Unique [KBFK21]. Unity [CPB18]. University [WFL^{+}22, SAM23]. Unknown [FSC21, SHL23]. Unmoderated [CWK^{+}23]. Untold [JA21]. Untrustworthy [BYRL23]. Unveiling [FH24]. Upper [LCG24]. Upwork [FR24]. Urban [SCP^{+}21, WDD^{+}22]. Urgency [CERR23]. Usability [MCI20, SC24, SY24, ZP24]. Usable [PNS^{+}24b]. Usage [ARH^{+}23, KSKH24, LW19]. Use [AHCD17, BB20, BWDP20, BSA23, DBPA22, GYD17, Gar21, GGE23, GBN21, HSZKN24, HFBL19, JP22, KA23, MH19, MHN^{+}21, MNS^{+}21, RVBC^{+}21, SSGB21, TAS^{+}19, VVBK23, WMBO19, XKRW23, BSS18, BBK22, GPH19, SSGB20, ZPMK24]. Used [AOT^{+}18, FC20, ZMTH22]. User [AFKL20, ALG^{+}21, AMVK23, ACDC23, ATH24, AZS^{+}22, AIRH24, BR20, BahRKM18, Blo24, BPB22, BGRK17, dACZ19, CPB18, CSL^{+}24, CLwH^{+}20, CMSA22, CHCG24, DWSF24, DNBB^{+}22, DRDS24, HPY^{+}22, HV22, KCC20, KVDG20, KJH19, LL24, LWC^{+}24, LLDH21, LFM22, LWS24, MAN^{+}20, MNP^{+}20, MLSH21, MBS22, NM22a, PSTK24, dQPMdHRdAN21, PCL^{+}21, RLGG21, Sal17, SAY^{+}24, SWT20, SHS22, SVV21, SORV22, SC23, SBB24, ULG^{+}22, VP24a, VZV19, VN19, VNSVL22, WGGS20, WSM24, WRF^{+}22, XJBH22, YPS^{+}22, YHR^{+}19, YMG^{+}20, ZBV23, ZMR21, ZGL^{+}22, ZLB^{+}22, ZMGK24, ZRP^{+}22, ZP24, dRSV22].

User-Assigned [CSL^{+}24]. User-Centric [AFKL20]. User-Configurable [ZMR21]. User-Defined [MLSH21, NM22a, SAY^{+}24, VNSVL22]. User-friendly [ZLB^{+}22]. User-Generated [LL24, ZGL^{+}22, ZMGK24]. User-Interfaces [AMVK23]. User-oriented [ZGL^{+}22]. User-perceived [LWC^{+}24]. Users [ACF^{+}22, BYRL23, BKJ^{+}20, DTE^{+}22, EHZN22, JFL23, JDX24, KDEA21, KOL^{+}24, LAD^{+}22, LPZV23, PLY23, PNS^{+}24b, SRKK22, STM24, SPL^{+}23, WGM21, WDD^{+}22, YK23, ZMTH22, dQPMdHRdAN21]. Uses [CBSQ24, LX20]. Using [AR22, BahRKM18, BRM21, CCWH21, CSH^{+}23, CHS^{+}24, EMS^{+}22, ERH^{+}24, EPW^{+}23, FSC21, FRH^{+}22, GKS21, GHHS21, JF24, JCD19, KKK^{+}24, KTP^{+}22, KOL^{+}24, LOSD20, LZW23, LCG24, LKG21, MdOKT23, MBT^{+}20, NMBS21, RGR^{+}21, STM^{+}19, UYSM22, WO20, ZBL19, ZYY23, ASZW23, FSAN23, HSC^{+}23, KR24, KKS19, LKYK24, LHT^{+}20, LSLD22, MAN^{+}20, MGRM22, MMMM22, NFTK22, RSL24, Sal17, SBSM24, SCY^{+}18, SHC^{+}23, SM22, SC23, TRO^{+}21, TBR20, VP24b, WPH22a, YRIV20, YML^{+}23, ZWZ^{+}21]. UsyBus [JV22]. Utility [FMBW24, GPH21]. UX [GGF^{+}20]. UX-Painter [GGF^{+}20].

V4 [CM20, KLS20, KLS21]. V5 [BKL^{+}21, BLS21, FJIH21, GMM21]. V6 [ACI22, BLS^{+}22, FBNM22, GKKW22, KRW^{+}22]. V7 [CDG^{+}23, CDGdC23, DKBG23, LMS23]. V8 [CDG^{+}24, DKKL24, MDFG24, MBTW24]. Validation [BMPP20, HHJ^{+}22]. Value [CRM^{+}24, KML^{+}21, WAD^{+}23, XDC^{+}24]. Values [BLH^{+}17, BLL^{+}21, CS18b, GN20, HJC^{+}22, NPT19, SHD21, SLMB24]. Variables [OM20]. Various [HTS^{+}22]. Varying [YUSM22]. Vehicle [ITKB23].

Vehicles [CWR22, CERR23, HCR22, SCR22, WWH^{+}23]. Venmo [CEM17]. Verbal [FLC17]. Verification [PB22, ZWZ^{+}21]. Versatile [HTS^{+}22]. Versus [CCW22, MYS^{+}21]. Vertical [PDPS22, RSL24, RRF20]. VI [KvBW24]. Via [LAM22, HTTH22, JFS^{+}23, LLW^{+}23, LMZ^{+}24, MMM^{+}21, NSV22]. Viability [CYC^{+}21]. VibHand [ZSOJ20]. Vibrotactile [AARJ23, YML^{+}23, ZSOJ20]. Vicariously [LAW19]. Video [BKM21, BBD^{+}24, DHH18, EHG22, FWY^{+}22, FLC17, FRH^{+}22, HSH22, HDPL21, HJM20, HWYW21, LMT^{+}22, MON^{+}21, SSB^{+}22, SCY^{+}18, WZT^{+}24, WRM^{+}24]. Video-Based [DHH18, SSB^{+}22, FWY^{+}22]. Video-Conferencing [LMT^{+}22]. Videogames [BVM21, MPKEN24]. VideoPoseVR [WZFA22]. Videos [CSH^{+}23, DDFK19, HKL^{+}22, LL24, LWS24, OKM^{+}23, WZFA22, WvECM22]. VidLyz [DDFK19]. View [KKM24, OKM^{+}23, WACdA^{+}21]. Views [SAR^{+}23]. Violations [CSJ^{+}18, GYD17, NSJ21]. Violence [Bel24]. Violent [RVBC^{+}21]. ViRgilites [ALRS24]. Virtual [ACF^{+}22, AARJ23, BHSP22, BLC^{+}21, BRM21, CTWM23, CDIH23, CKK22, FA22, FAMS22, GB18, HSVR21, HPH^{+}23, HRP23, HOSK21, HWYW21, JDX24, KMM21, KVM23, KKK22a, KMV^{+}24, LG22, LZW23, LPW^{+}23, MAN^{+}20, MLSH21, NAO23, OG21, PDPS22, PSTK24, SFLB23, SCMS18, SK20, SLN21, VMHO^{+}22, VWHW24, WZFA22, WVG^{+}24, WLZL23, XHH17, YOFC19, ZKL22]. Visibility [FA22, UHH23]. Visible [KJTC22]. Vision [BLB^{+}24, LAR21, MSY20, SHD21]. VisRecall [WJH^{+}24]. Visual [ASM23, AARJ23, BGM^{+}23, BV21, CTWM23, CBPC23, CMY^{+}17, DWSF24, HMML^{+}22, JSP^{+}21, KKK^{+}24, LG22, LMZ^{+}24, MCM23, MSHP23, PDPS22, SRZ24, SAR^{+}24, SB17, SAR21, TRO^{+}21, UAAR20, ZSOJ20, ZMS^{+}22, ZWG^{+}23, vDBL22, ZYY23]. Visualisation [WJH^{+}24]. Visualization [BYL^{+}24, LDP18, MCDR22, NB20, PAC22]. Visualizations [CSS^{+}17, HZ22, LHI20, RWB22]. Visualizing [HJCM22, HWS21]. Visually [HRRS20]. Visuals [CKSP22]. Vitro [KA23]. VizSciFlow [HRRS20]. Vocabulary [RIC^{+}24]. Voice [AKL^{+}23, CLL^{+}21, KRMM21, PKB^{+}19, SVS^{+}24]. Volitionality [SSGB21]. Volunteer [GTC22, JØC^{+}20]. VR [ALRS24, CTH24, CCNY19, GTO^{+}23, HPH^{+}23, JOFM24, KMV^{+}24, KTF^{+}21, LKYK24, NMPDJ24, VAE^{+}22, WRF^{+}22, YTF^{+}23, YML^{+}23]. VR-based [KMV^{+}24]. VR-Hiking [HPH^{+}23]. vs [BDE^{+}24, RVBC^{+}21, VNSVL22]. Vulnerable [STM24].

Wake [PNS24a]. Walker [KKV^{+}22]. Walking [CWC^{+}24, CKK22, KKV^{+}22, SRKK22, LRK^{+}24a]. Walkthrough [WGGS20]. Wall [CHS^{+}24, KUWM22]. Wall-Sized [CHS^{+}24]. WAMS [BGM^{+}23]. Want [CSL^{+}24, VHZ22]. Warning [SLD21]. Was [FKK19, Rib19, BZC^{+}17, CCWH21, Pri19]. waste [WAD^{+}23]. watched [Pri19]. Watchful [JDX24]. Waters [AMF18]. Wave [BMdS^{+}21, BMdS^{+}22, PAK^{+}22]. Wearable [BP23, CBSQ24, GGL^{+}22, KVK^{+}22, MBS22, RLL22, XJBH22]. Wearables [LCG24]. Wearers [WACdA^{+}21]. Weather [BJ21]. Web [AV17, BahRKM18, BMPP20, LZS^{+}21, LAD^{+}22, MB22, NB20, OTH20, dQPMdHRdAN21, SCMS18]. Web-Based [BahRKM18]. Web-Scale [MB22]. Web5VR [SCMS18]. Webcam [HSC^{+}23]. Webcam-based [HSC^{+}23]. website [Gil20, KKK22b]. Websites [MMH21]. Weeks [KUD^{+}23]. Weibo [LHT^{+}20].

Welcome [EVC^{+}19, KFMH17, SN18]. Welfare [HFL19]. Well [KKP^{+}22, SCSD24]. Well-Being [SCSD24, KKP^{+}22]. Wellbeing [RA21b, Vya19]. Wellness [ASO^{+}22]. Western [HOHB22]. WevQuery [AV17]. Whale [RAZ^{+}20]. Wheelchair [WDD^{+}22]. Where [RYCC21, RKFK22, Pri19]. which [MH19]. White [SRKK22]. Who [EHZN22, HRC^{+}21, PKC^{+}21, SKV^{+}21, SK23, SSGB21]. Wider [HKD23]. Widgets [CLV^{+}19]. Width [YUSM22]. Width-Varying [YUSM22]. Wikipedia [HJC^{+}22, MD20, VH21]. Wild [BB20, KMH23]. Will [CYC^{+}21, CT18, LAM22]. Willed [HRC^{+}21]. Willingness [BWE^{+}22, GPH21]. WIMP [CMN^{+}21]. Win [CBWD24]. Wind [MXYS23]. Wine [BEHB^{+}23]. Wings [SBAK21]. Winning [MMdSP21]. Winter [BJ21]. wise [ZWG^{+}23]. Within [ASS18, NCM^{+}22]. Without [HBJR21, LAW19, TRO^{+}21]. Witness [UHH23]. Wolt [KB22]. Woman [HJC^{+}22]. Women [AZ23, ERT21, HV23, IRA^{+}19, SELS24, SFLB23, RYD21]. Work [ABC21, BKKD21, BBV19, CSL^{+}24, CBV22, DCM^{+}23, ECA23, FSN^{+}20, HGG20, JSKA22, LWTW23, LC24, MWM^{+}19, MYS^{+}21, MDZ19, MNS^{+}21, MRS22, PKLK22, PP23, RWL23, SLB23, SHJ^{+}22, SMLH23, TSS21, Vya19, WMM^{+}19, WDD^{+}22, ZBL19, SFA^{+}23]. Work-From-Home [CBV22]. Work-Related [CSL^{+}24]. Workbench [BV21]. Worker [CKS18, MWBB23, MBS22, PF24]. Workers [HMB^{+}21, HBJR21, PVP22, PLL^{+}23, SMLH23, ZMW20]. Workflow [AR22, CRN21, MQX^{+}23]. Workflows [FKDG21, ZMW20]. Working [AWM21, CTWM23, FAMS22, KB22, PK23, WS22]. Workload [ASM23, AIRH24]. Workplace [HSG^{+}22, MNS^{+}21, MDN^{+}24]. Workspace [MRGN20]. Workspaces [BGM^{+}23, FZRJ22, LWB22]. World [GHHS21, CWC^{+}24, Gil20, MBK^{+}22]. WorldPoint [KMH23]. Worn [EPW^{+}23, HOSK21]. Wounds [JP22]. WProfX [NB20]. Wrappers [BBS19]. Wrist [CDIH23]. Write [FM21, LM20].

XAI [ESCR23]. XAL [GLZ^{+}21]. XD [SHY^{+}18]. XD-AR [SHY^{+}18]. xPress [BP17b]. XR [FSAN23, PGK^{+}24]. XRSpotlight [FSAN23]. XSpace [HCG^{+}22].

year [GN23]. Years [HK24a]. Young [YGHR21]. You’re [LC24, BLH^{+}17]. Yours [SLN21]. Youth [BWDP20, CNM^{+}23, FZM22, JS22, SSH^{+}21]. YouTube [BBNT21, HRR^{+}22, HJM20, LLL^{+}22].

Zoo [GKHD23]. Zoo-Housed [GKHD23]. Zoom [GF22]. Zoomorphic [SVS^{+}24].

References

Agapie:2022:LGS

[AAHM22] Elena Agapie, Patricia A.

Areán, Gary Hsieh, and Sean A. Munson. A longitudinal goal setting model for addressing complex personal problems in mental health. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):270:1–270:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555160.

Auer:2023:ACI

[AARJ23] Stefan Auer, Christoph Anthes, Harald Reiterer, and Hans-Christian Jetter. Aircraft cockpit interaction in virtual reality with visual, auditive, and vibrotactile feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):445:1–445:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626481.

Alsakar:2023:IPP

[AASK23] Noora Alsakar, Yasmeen Abdrabou, Simone Stumpf, and Mohamed Khamis. Investigating privacy perceptions and subjective acceptance of eye tracking on handheld mobile devices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):164:1–164:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591133.

Ahmetoglu:2021:PPM

[ABC21] Yoana Ahmetoglu, Duncan P. Brumby, and Anna L. Cox. To plan or not to plan? a mixedmethods diary study examining when, how and why knowledge work planning is inaccurate. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):222:1–222:20, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432921.

Andalibi:2017:MCT

[ABQ17] Nazanin Andalibi, Frank Bentley, and Katie Quehl. Multi-channel topic-based mobile messaging in romantic relationships. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):20:1–20:18, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134655.

Avula:2022:ESI

[ACA22] Sandeep Avula, Bogeum Choi, and Jaime Arguello. The effects of system initiative during conversational collaborative search. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):66:1–66:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512913.

Alipour:2023:FUI

[ACDC23] Mina Alipour, Éric Céret, and Sophie Dupuy-Chessa. A framework for user interface adaptation to emotions and their temporal aspects. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 186:1–186:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593238.

Artizzu:2022:DCV

[ACF^{+}22] Valentino Artizzu, Gianmarco Cherchi, Davide Fara, Vittoria Frau, Riccardo Macis, Luca Pitzalis, Alessandro Tola, Ivan Blecic, and Lucio Davide Spano. Defining configurable virtual reality templates for end users. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 163:1–163:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3534517.

Alexander:2022:PVM

[ACI22] Jason Alexander, Robin Bing-Yu Chen, and Petra Isenberg. PACMHCI V6, MHCI, September 2022 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 175:1–175:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546710.

Al-Dawood:2021:SAP

[ADAY21] Adel Al-Dawood, Serene Alhajhussein, and Svetlana Yarosh. Saudi Arabian parents’ perception of online marital matchmaking technologies. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW3): 211:1–211:32, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432910.

Almahmoud:2021:HTC

[ADD21] Jumana Almahmoud, Robert DeLine, and Steven M. Drucker. How teams communicate about the quality of ML models: a case study at an international technology company. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):222:1–222:24, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3463934.

Ahmad:2020:TPT

[AFKL20] Imtiaz Ahmad, Rosta Farzan, Apu Kapadia, and Adam J. Lee. Tangible privacy: Towards user-centric sensor designs for bystander privacy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):116:1–116:28, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415187.

Aromaa:2018:DHM

[AFT^{+}18] Susanna Aromaa, Nikos Frangakis, Domenico Tedone, Juhani Viitaniemi, and Iina Aaltonen. Digital human models in human factors and ergonomics evaluation of gesture interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):6:1–6:14, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3229088.

Akter:2022:PCJ

[AGK^{+}22] Mamtaj Akter, Amy J.

Godfrey, Jess Kropczynski, Heather R. Lipford, and Pamela J. Wisniewski. From parental control to joint family oversight: Can parents and teens manage mobile online safety and privacy as equals? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):57:1–57:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512904.

Axel:2023:ERB

[AGM^{+}23] Carayon Axel, Juan E.

Garrido, Célia Martinie, Philippe Palanque, Eric Barboni, María Dolores Lozano, and Víctor M. R. Penichet. Engineering rehabilitation: Blending two tool-supported approaches to close the loop from tasks-based rehabilitation to exercises and back again. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 177:1–177:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593229.

Alexandrovsky:2024:DGC

[AGO^{+}24] Dmitry Alexandrovsky, Kathrin Gerling, Merlin Steven Opp, Christopher Benjamin Hahn, Max V. Birk, and Meshaiel Alsheail. Disengagement from games: Characterizing the experience and process of exiting play sessions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):301:1– 301:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677066.

Asthana:2018:FEA

[AH18] Sumit Asthana and Aaron Halfaker. With few eyes, all hoaxes are deep. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 21:1–21:18, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274290.

Ahmed:2017:DPC

[AHCD17] Syed Ishtiaque Ahmed, Md. Romael Haque, Jay Chen, and Nicola Dell. Digital privacy challenges with shared mobile phone use in Bangladesh. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):17:1–17:20, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134652.

Alvarado:2022:SRI

[AHJV22] Oscar Alvarado, Nyi Nyi Htun, Yucheng Jin, and Katrien Verbert. A systematic review of interaction design strategies for group recommendation systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):271:1–271:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555161.

Alkayyali:2022:UEU

[AIHS22] Amani Alkayyali, Yasha Iravantchi, Jaylin Herskovitz, and Alanson P. Sample. UbiChromics: Enabling ubiquitously deployable interactive displays with photochromic paint. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ISS): 561:1–561:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3567714.

Atreja:2024:AIF

[AIRH24] Shubham Atreja, Jane Im, Paul Resnick, and Libby Hemphill. AppealMod: Inducing friction to reduce moderator workload of handling user appeals. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):19:1–19:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637296.

Araake:2021:PEP

[AKIS21] Koichi Araake, Michinari Kono, Eiji Iwata, and Norio Sasaki. Playful engagement for public spaces: a case study on a mall escalator. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 498:1–498:19, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488543.

Abdulgalimov:2023:DEE

[AKL^{+}23] Dinislam Abdulgalimov, Reuben Kirkham, Stephen Lindsay, James Nicholson, Vasilis Vlachokyriakos, Emily Dao, Daniel Kos, Daniel Jitnah, Pam Briggs, and Patrick Olivier. Designing for the embedding of employee voice. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):45:1–45:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579478.

Abad:2023:ATR

[AKVHZ23] Beatriz Palacios Abad, Michael Koohang, Morgan Vigil-Hayes, and Ellen Zegura. Alone and together: Resilience in a fluid socio-technical-natural system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):24:1–24:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579457.

AlvaradoGarcia:2018:QRG

[AL18] Adriana Alvarado Garcia and Christopher A. Le Dantec. Quotidian report: Grassroots data practices to address public safety. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 17:1–17:18, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274286.

Ahmad:2021:EGC

[AL21] Azhan Ahmad and Effie Lai-Chong Law. Educators as gamemasters: Creating serious role playing game with. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):230:1– 230:29, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474657.

Ashktorab:2020:HAC

[ALD^{+}20] Zahra Ashktorab, Q. Vera Liao, Casey Dugan, James Johnson, Qian Pan, Wei Zhang, Sadhana Kumaravel, and Murray Campbell. Human-AI collaboration in a cooperative game setting: Measuring social perception and outcomes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):96:1–96:20, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415167.

Ahsen:2021:ENO

[ALG^{+}21] Tooba Ahsen, Zi Yi Lim, Aaron L. Gardony, Holly A. Taylor, Jan P. de Ruiter, and Fahad Dogar. The effects of network outages on user experience in augmented reality based remote collaboration — an empirical study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):313:1–313:27, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476054.

Artizzu:2024:VMF

[ALRS24] Valentino Artizzu, Kris Luyten, Gustavo Rovelo Ruiz, and Lucio Davide Spano. ViRgilites: Multilevel feedforward for multimodal interaction in VR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 247:1–247:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3658645.

Aziz:2023:PPB

[ALSK23] Samantha Aziz, Dillon J. Lohr, Razvan Stefanescu, and Oleg Komogortsev. Practical perception-based evaluation of gaze prediction for gaze contingent rendering. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):165:1–165:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591134.

Azad:2021:LCP

[AM21] Sasha Azad and Chris Martens. Little computer people: a survey and taxonomy of simulated models of social interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):245:1– 245:30, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474672.

Ames:2018:HCC

[Ame18] Morgan G. Ames. Hackers, computers, and cooperation: a critical history of logo and constructionist learning. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 18:1–18:19, November 2018. CODEN ????

ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274287.

Andalibi:2018:TWS

[AMF18] Nazanin Andalibi, Margaret E. Morris, and Andrea Forte. Testing waters, sending clues: Indirect disclosures of socially stigmatized experiences on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 19:1–19:23, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274288.

Arawjo:2019:CEI

[AMJ^{+}19] Ian Arawjo, Ariam Mogos, Steven J. Jackson, Tapan Parikh, and Kentaro Toyama. Computing education for intercultural learning: Lessons from the Nairobi Play Project. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 52:1–52:24, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359154.

Albarrak:2021:DMS

[AMR21] Luluah Albarrak, Oussama Metatla, and Anne Roudaut. (don’t) mind the step: Investigating the effect of digital social cues on navigation decisions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 492:1–492:18, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488537.

Alipour:2023:EEB

[AMVK23] Mina Alipour, Mahyar T. Moghaddam, Karthik Vaidhyanathan, and Mikkel Baun Kjærgaard. Emoticontrol: Emotions-based control of user-interfaces adaptations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 175:1–175:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593227.

Annett:2020:UHO

[Ann20] Michelle Annett. Understanding the homepreneurship opportunities afforded by social networking and personal fabrication technologies.

Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):99:1–99:48, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415170.

August:2018:FEC

[AOT^{+}18] Tal August, Nigini Oliveira, Chenhao Tan, Noah Smith, and Katharina Reinecke. Framing effects: Choice of slogans used to advertise online experiments can boost recruitment and lead to sample biases. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 22:1–22:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274291.

Aeschbach:2021:TMR

[APW^{+}21] Lena Fanya Aeschbach, Sebastian A. C. Perrig, Lorena Weder, Klaus Opwis, and Florian Brühlmann. Transparency in measurement reporting: a systematic literature review of CHI PLAY. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):233:1– 233:21, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474660.

AbediniAla:2022:FAC

[AR22] Mostafa AbediniAla and Banani Roy. Facilitating asynchronous collaboration in scientific workflow composition using provenance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 166:1–166:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3534520.

Ahmed:2023:MAP

[ARH^{+}23] Md Sabbir Ahmed, Rahat Jahangir Rony, Mohammad Abdul Hadi, Ekram Hossain, and Nova Ahmed. A minimalistic approach to predict and understand the relation of app usage with students’ academic performance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):193:1–193:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604240.

Aitamurto:2017:MPC

[AS17] Tanja Aitamurto and Jorge Saldivar. Motivating participation in crowdsourced policymaking: The interplay of epistemic and interactive aspects. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):18:1–18:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134653.

Alshehri:2023:CAO

[AS23] Majdah Alshehri and Norman Makoto Su. Comfort activism: Online photography for social change in a minority group. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):35:1–35:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579468.

Alizadeh:2023:CMI

[ASJK23] Fatemeh Alizadeh, Gunnar Stevens, Timo Jakobi, and Jana Krüger. Catch me if you can: “Delaying” as a social engineering technique in the post-attack phase.

Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):32:1–32:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579465.

Aljehane:2023:SDE

[ASM23] Salwa D. Aljehane, Bonita Sharif, and Jonathan I. Maletic. Studying developer eye movements to measure cognitive workload and visual effort for expertise assessment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):166:1–166:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591135.

Abdou:2022:GEC

[ASMB22] Ahmed Abdou, Ekta Sood, Philipp Müller, and Andreas Bulling. Gaze-enhanced crossmodal embeddings for emotion recognition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):138:1–138:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530879.

AlSlaity:2022:MAH

[ASO^{+}22] Alaa AlSlaity, Banuchitra Suruliraj, Oladapo Oyebode, Jonathon Fowles, darren steeves, and Rita Orji. Mobile applications for health and wellness: a systematic review. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 171:1–171:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3534525.

Alharbi:2020:EPF

[ASP20] Ohoud Alharbi, Wolfgang Stuerzlinger, and Felix Putze. The effects of predictive features of mobile keyboards on text entry speed and errors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 183:1–183:16, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427311.

Arif:2018:APE

[ASS18] Ahmer Arif, Leo Graiden Stewart, and Kate Starbird. Acting the part: Examining information operations within #Black-LivesMatter discourse.

Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (CSCW):20:1–20:27, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274289.

Akbar:2023:CDG

[ASZW23] Ahsan Jamal Akbar, Zhiyao Sheng, Qian Zhang, and Dong Wang. Cross-domain gesture sequence recognition for two-player exergames using COTS mmWave radar. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ISS): 441:1–441:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3626477.

Alsebayel:2024:BBP

[ATH24] Ghada Alsebayel, Giovanni Troiano, and Casper Harteveld. “I Believe the Baby in the Picture is My Baby”: User experiences with commercial pregnancy apps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):266:1–266:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676511.

Apaolaza:2017:WTH

[AV17] Aitor Apaolaza and Markel Vigo. Wev-Query: Testing hypotheses about Web interaction patterns. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):4:1–4:17, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095806.

Aiordachioae:2019:LTS

[AV19] Adrian Aiordachioae and Radu-Daniel Vatavu. Life-tags: a smartglassesbased system for recording and abstracting life with tag clouds. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):15:1–15:22, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331157.

Antoniak:2021:TBC

[AWM21] Maria Antoniak, Melanie Walsh, and David Mimno. Tags, borders, and catalogs: Social re-working of genre on LibraryThing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):29:1–29:29, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449103.

Aljasim:2023:FWS

[AZ23] Hanan Khalid Aljasim and Douglas Zytko. Foregrounding women’s safety in mobile social matching and dating apps: a participatory design study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):9:1–9:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567559.

Altmeyer:2022:DHU

[AZS^{+}22] Maximilian Altmeyer, Berina Zenuni, Hanne Spelt, Thierry Jegen, Pascal Lessel, and Antonio Krüger. Do hexad user types matter? Effects of (non-) personalized gamification on task performance and user experience in an image tagging task. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):228:1– 228:??, October 2022.

CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549491.

Bhatia:2020:EDS

[BAAP20] Arpit Bhatia, Yajur Ahuja, Suyash Agarwal, and Aman Parnami. Exploring the design space of badge based input. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 185:1–185:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427313.

Batalas:2018:UTE

[BahRKM18] Nikolaos Batalas, Marije aan het Rot, Vassilis Javed Khan, and Panos Markopoulos. Using TEMPEST: Enduser programming of Web-based ecological momentary assessment protocols. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):3:1–3:24, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3179428.

Berenberg:2018:ECE

[BB18] Daniel Berenberg and James P. Bagrow. Efficient crowd exploration of large networks: The case of causal attribution. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 24:1–24:25, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274293.

Bemmann:2020:LMK

[BB20] Florian Bemmann and Daniel Buschek. LanguageLogger: a mobile keyboard application for studying language use in everyday text communication in the wild. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):84:1–84:24, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397872.

Buhler:2024:TSE

[BBD^{+}24] Babette Bühler, Efe Bozkir, Hannah Deininger, Peter Gerjets, Ulrich Trautwein, and Enkelejda Kasneci. On task and in sync: Examining the relationship between gaze synchrony and self-reported attention during video lecture learning. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):230:1–230:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655604.

Bhagavatula:2022:ATE

[BBK22] Sruti Bhagavatula, Lujo Bauer, and Apu Kapadia. “Adulthood is trying each of the same six passwords that you use for everything”: The scarcity and ambiguity of security advice on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):264:1–264:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555154.

Buntain:2021:YRE

[BBNT21] Cody Buntain, Richard Bonneau, Jonathan Nagler, and Joshua A. Tucker. YouTube recommendations and effects on sharing across online social platforms. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):11:1–11:26, April 2021. CODEN ????

ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449085.

Britton:2019:MCW

[BBS19] Lauren Britton, Louise Barkhuus, and Bryan Semaan. “Mothers as Candy Wrappers”: Critical infrastructure supporting the transition into motherhood. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):232:1–232:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3361113.

Brie:2023:ELL

[BBSV23] Paul Brie, Nicolas Burny, Arthur Sluÿters, and Jean Vanderdonckt. Evaluating a large language model on searching for GUI layouts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 178:1–178:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593230.

Bopp:2019:CPK

[BBV19] Chris Bopp, Lehn M. Benjamin, and Amy Voida. The coerciveness of the primary key: Infrastructure problems in human services work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 51:1–51:26, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359153.

Berentsen:2021:MAN

[BBW21] Mark J. Berentsen, Marit Bentvelzen, and Paweł W. Woźniak. MTBalance: Assisting novice mountain bikers with realtime proprioceptive feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 506:1–506:25, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488551.

Bajorunaite:2023:RAB

[BBW23] Laura Bajorunaite, Stephen Brewster, and Julie R. Williamson. Reality anchors: Bringing cues from reality to increase acceptance of immersive technologies in transit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):219:1–219:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604266.

Bouzit:2017:PMM

[BCCV17] Sara Bouzit, Gaëlle Calvary, Denis Chêne, and Jean Vanderdonckt. Polymodal menus: a model-based approach for designing multimodal adaptive menus for small screens. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):15:1–15:19, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3099585.

Berger:2023:DCN

[BDB^{+}23] Melanie Berger, Debargha Dey, Bahareh Barati, Bastian Pfleging, and Regina Bernhaupt. Designing for collaborative non-driving related activities in future cars: Fairness and team performance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI): 202:1–202:??, September 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3604249.

Bucher:2024:TMP

[BDE^{+}24] Andreas Bucher, Mateusz Dolata, Sven Eckhardt, Dario Staehelin, and Gerhard Schwabe. Talking to multi-party conversational agents in advisory services: Command-based vs. conversational interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):7:1–7:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3633072.

BenAjenaghughrure:2021:PMT

[BDL21] Ighoyota Ben Ajenaghughrure, Sonia Cláudia Da Costa Sousa, and David Lamas. Psychophysiological modeling of trust in technology: Influence of feature selection methods. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 203:1–203:25, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3459745.

Blackwell:2017:CCO

[BDSL17] Lindsay Blackwell, Jill Dimond, Sarita Schoenebeck, and Cliff Lampe. Classification and its consequences for online harassment: Design insights from HeartMob. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):24:1–24:19, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134659.

Bowen:2021:TMI

[BDW21] Judy Bowen, Anke Dittmar, and Benjamin Weyers. Task modelling for interactive system design: a survey of historical trends, gaps and future needs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 214:1–214:22, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461736.

Berkholz:2023:BOW

[BEHB^{+}23] Jenny Berkholz, Margarita Esau-Held, Alexander Boden, Gunnar Stevens, and Peter Tolmie. Becoming an online wine taster: an ethnographic study on the digital mediation of taste. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):26:1–26:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579459.

Bellini:2024:APP

[Bel24] Rosanna Frances Bellini.

Abusive partner perspectives on technology abuse: Implications for community-based violence prevention. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):15:1–15:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637292.

Buchholz:2017:EFT

[BF17] Gregor Buchholz and Peter Forbrig. Extended features of task models for specifying cooperative activities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):7:1–7:21, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095809.

Bateman:2023:WFA

[BGM^{+}23] Scott Bateman, Carl Gutwin, Hamid Mansoor, Miguel Nacenta, Michael van der Kamp, Mykyta Baliesnyi, Kolton Gagnon, and Jesse Rollheiser. WAMS: a flexible API for visual workspaces across multiple surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 184:1–184:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593236.

Budak:2017:TSH

[BGRK17] Ceren Budak, R. Kelly Garrett, Paul Resnick, and Julia Kamin. Threading is sticky: How threaded conversations promote comment system user retention. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):27:1–27:20, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134662.

Borsting:2022:SEA

[BHH^{+}22] Ingo Börsting, Markus Heikamp, Marc Hesenius, Wilhelm Koop, and Volker Gruhn. Software engineering for augmented reality — a research agenda. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (EICS):155:1–155:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3532205.

Brubaker:2019:ONG

[BHM19] Jed R. Brubaker, Gillian R.

Hayes, and Melissa Mazmanian. Orienting to networked grief: Situated perspectives of communal mourning on Facebook. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 27:1–27:19, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359129.

Bhatnagar:2023:APA

[BHM^{+}23] Tigmanshu Bhatnagar, Albert Higgins, Nicolai Marquardt, Mark Miodownik, and Catherine Holloway. Analysis of product architectures of pin array technologies for tactile displays. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):432:1–432:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626468.

Buyuktur:2018:SCC

[BHNA18] Ayse G. Buyuktur, Pei-Yao Hung, Mark W. Newman, and Mark S. Ackerman. Supporting collaboratively constructed independence: a study of spinal cord injury. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 26:1–26:25, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274295.

Bagmar:2022:AEE

[BHSP22] Aadesh Bagmar, Kevin Hogan, Dalia Shalaby, and James Purtilo. Analyzing the effectiveness of an extensible virtual moderator. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):18:1–18:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3492837.

Back:2021:PCW

[BJ21] Jon Back and Karin B. Johansson. Playing cool — winter weather’s influence on locationbased gaming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):242:1– 242:16, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474669.

Burch:2024:TET

[BK24] Michael Burch and Kuno Kurzhals. Teaching eye tracking: Challenges and perspectives. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):237:1–237:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655611.

Barlas:2021:SSI

[BKG^{+}21] Pinar Barlas, Kyriakos Kyriakou, Olivia Guest, Styliani Kleanthous, and Jahna Otterbacher. To “see” is to stereotype: Image tagging algorithms, gender recognition, and the accuracy– fairness trade-off. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):232:1–232:31, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432931.

Bawa:2020:DMU

[BKJ^{+}20] Anshul Bawa, Pranav Khadpe, Pratik Joshi, Kalika Bali, and Monojit Choudhury. Do multilingual users prefer chat-bots that code-mix? Let’s nudge and find out! Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):041:1–041:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392846.

Burgess:2022:CFC

[BKJ22] Eleanor R. Burgess, Elizabeth Kaziunas, and Maia Jacobs. Care frictions: a critical reframing of patient noncompliance in health technology design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):281:1–281:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555172.

Blaising:2021:MIW

[BKKD21] Allie Blaising, Yasmine Kotturi, Chinmay Kulkarni, and Laura Dabbish. Making it work, or not: a longitudinal study of career trajectories among online freelancers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):226:1–226:29, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432925.

Bardzell:2021:PVCa

[BKL^{+}21] Shaowen Bardzell, Juho Kim, Siân Lindley, Aleksandra Sarcevic, and Sarita Schoenebeck. PACMHCI V5 CSCW1 April 2021 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):1:1, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449254.

Benke:2020:CBE

[BKM20] Ivo Benke, Michael Thomas Knierim, and Alexander Maedche. Chatbot-based emotion management for distributed teams: a participatory design study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):118:1–118:30, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415189.

Beres:2021:UPE

[BKM21] Nicole A. Beres, Madison Klarkowski, and Regan L. Mandryk. Under pressure: Exploring choke and clutch in competitive video games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):239:1– 239:22, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474666.

Bossen:2017:PMP

[BL17] Claus Bossen and Timo Leimbach. Project management practices as a subject of research for CSCW: Status and future opportunities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):25:1–25:25, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134660.

Buntain:2018:PCR

[BL18] Cody L. Buntain and Jung Kyu Rhys Lim.

                                        #pray4victims: Con-

sistencies in response to disaster on Twitter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 25:1–25:18, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274294.

Bhole:2024:HHB

[BLB^{+}24] Palavi V. Bhole, Ziming Li, Shivang Bokolia, Tae Oh, Garreth W. Tigwell, and Roshan L. Peiris. Haptic2FA: Haptics-based accessible two-factor authentication for blind and low vision people. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 264:1–264:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676509.

Bermejo:2021:EBD

[BLC^{+}21] Carlos Bermejo, Lik Hang Lee, Paul Chojecki, David Przewozny, and Pan Hui. Exploring button designs for mid-air interaction in virtual reality: a hexa-metric evaluation of key representations and multi-modal cues. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 194:1–194:26, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3457141.

Berry:2017:GKY

[BLH^{+}17] Andrew B. L. Berry, Catherine Y. Lim, Andrea L. Hartzler, Tad Hirsch, Evette Ludman, Edward H. Wagner, and James D. Ralston. “It’s good to know you’re not a stranger every time”: Communication about values between patients with multiple chronic conditions and healthcare providers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):23:1–23:20, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134658.

Berry:2021:SCR

[BLL^{+}21] Andrew B. L. Berry, Catherine Y. Lim, Calvin A. Liang, Andrea L. Hartzler, Tad Hirsch, Dawn M. Ferguson, Zoë A. Bermet, and James D. Ralston. Supporting collaborative reflection on personal values and health. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):299:1–299:39, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476040.

Blouin:2024:TSF

[Blo24]  Arnaud Blouin.  A

type system for flexible user interactions handling. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 246:1–246:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3660248.

Bardzell:2021:PVCb

[BLS21] Shaowen Bardzell, Siân Lindley, and Aleksandra Sarcevic. PACMHCI V5 CSCW2 October 2021 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):291:1, October 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3476022.

Bardzell:2022:PVC

[BLS^{+}22] Shaowen Bardzell, Sian Lindley, Aleksandra Sarcevic, Hideaki Kuzuoka, Katharina Reinecke, Hao-Chuan Wang, and Naomi Yamashita. PACMHCI V6, CSCW1, April 2022 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):46:1–46:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512825.

Brown:2023:DML

[BLV23] Barry Brown, Eric Laurier, and Erik Vinkhuyzen. Designing motion: Lessons for self-driving and robotic motion from human traffic interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):5:1–5:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567555.

Baumer:2021:EIG

[BMdS^{+}21] Eric P. S. Baumer, Naja L. Holten Møller, Cleidson R. B. de Souza, Casey Fiesler, Aparecido Fabiano Pinatti de Carvalho, Juliana Baptista dos Santos França, and Maria Menendez-Blanco. Editorial introduction: GROUP 2022 first wave. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(GROUP):215:1–215:2, July 2021. CODEN ????

ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463927.

Baumer:2022:EIG

[BMdS^{+}22] Eric P. S. Baumer, Naja L. Holten Møller, Cleidson R. B. de Souza, Casey Fiesler, Aparecido Fabiano Pinatti de Carvalho, Juliana Baptista dos Santos França, and Maria Menendez-Blanco. Editorial introduction — GROUP 2022 second wave. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):1:1–1:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3492820.

Broccia:2020:FAS

[BMPP20] Giovanna Broccia, Marco Manca, Fabio Paternò, and Francesca Pulina. Flexible automatic support for Web accessibility validation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):83:1–83:24, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397871.

Byrne:2023:EES

[BMR^{+}23] Sean Anthony Byrne, Virmarie Maquiling, Adam Peter Frederick Reynolds, Luca Polonio, Nora Castner, and Enkelejda Kasneci. Exploring the effects of scanpath feature engineering for supervised image classification models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):161:1–161:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3591130.

Bauer:2017:CPS

[BP17a] Aaron Bauer and Zoran Popović. Collaborative problem solving in an open-ended scientific discovery game. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):22:1–22:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134657.

Brewer:2017:XRD

[BP17b] Robin N. Brewer and Anne Marie Piper. xPress: Rethinking design for aging and accessibility through an IVR blogging system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):26:1–26:17, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3139354.

Balaji:2023:AST

[BP23] Ananta Narayanan Balaji and Li-Shiuan Peh. AI-On-Skin: Towards enabling fast and scalable on-body AI inference for wearable on-skin interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 187:1–187:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593239.

Boyle:2022:IMM

[BPB22] Rachael E. Boyle, Ruslana Pledger, and Hans-Frederick Brown. Iterative mixed method approach to B2B SaaS user personas. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 169:1–169:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3534523.

Byun:2021:CPS

[BPO21] Jeongmin Byun, Jungkook Park, and Alice Oh. Cocode: Providing social presence with co-learner screen sharing in online programming classes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):300:1–300:28, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476041.

Bowen:2017:GOA

[BR17] Judy Bowen and Steve Reeves. Generating obligations, assertions and tests from UI models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):5:1–5:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3095807.

Barnaby:2020:AEF

[BR20] Gareth Barnaby and Anne Roudaut. Autogrip: Enabling force feedback devices to selfattach to end-users’ fingers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 184:1–184:14, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427312.

Brewer:2024:BBM

[BR24] Johanna Brewer and Morgan Romine. Breaking barriers in mobile game development. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):296:1– 296:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677061.

Born:2021:MPP

[BRM21] Felix Born, Adrian Rygula, and Maic Masuch. Motivating players to perform an optional strenuous activity in a virtual reality exergame using virtual performance augmentation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):225:1– 225:21, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474652.

Burgess:2022:JCH

[BRM22] Eleanor R. Burgess, Madhu C. Reddy, and David C. Mohr. “I Just Can’t Help But Smile Sometimes”: Collaborative self-management of depression. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):70:1–70:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512917.

Burgess:2019:TPP

[BRN^{+}19] Eleanor R. Burgess, Kathryn E. Ringland, Jennifer Nicholas, Ashley A. Knapp, Jordan Eschler, David C. Mohr, and Madhu C. Reddy. “I think people are powerful”: The sociality of individuals managing depression. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 41:1–41:29, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359143.

Boyd:2021:AES

[BS21] Karen L. Boyd and Katie Shilton. Adapting ethical sensitivity as a construct to study technology design teams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):217:1–217:29, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463929.

Bouma-Sims:2023:BBH

[BSA23] Elijah Bouma-Sims and Yasemin Acar. Beyond the Boolean: How programmers ask about, use, and discuss gender. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):28:1–28:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579461.

Bansal:2023:TMH

[BSM23] Aakash Bansal, Bonita Sharif, and Collin McMillan. Towards modeling human attention from eye movements for neural source code summarization. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):167:1–167:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591136.

Baumer:2018:DRS

[BSS18] Eric P. S. Baumer, Rui Sun, and Peter Schaedler. Departing and returning: Sense of agency as an organizing concept for understanding social media non/use transitions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 23:1–23:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274292.

Bowen:2023:ISM

[BT23] Judy Bowen and Jessica Turner. Interactive system modelling for the Internet of things. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 181:1–181:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593233.

Bellini:2021:CPB

[BTM^{+}21] Rosanna Bellini, Emily Tseng, Nora McDonald, Rachel Greenstadt, Damon McCoy, Thomas Ristenpart, and Nicola Dell. “So-called privacy breeds evil”: Narrative justifications for intimate partner surveillance in online forums.

Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):210:1–210:27, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432909.

Badillo-Urquiola:2021:CRR

[BUSA^{+}21] Karla Badillo-Urquiola, Zachary Shea, Zainab Agha, Irina Lediaeva, and Pamela Wisniewski. Conducting risky research with teens: Codesigning for the ethical treatment and protection of adolescents. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):231:1–231:46, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432930.

Burny:2021:UWC

[BV21] Nicolas Burny and Jean Vanderdonckt. UiLab, a workbench for conducting and reproducing experiments in GUI visual design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 196:1–196:31, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3457143.

Bopp:2021:MSG

[BVM21] Julia A. Bopp, Jan B. Vornhagen, and Elisa D. Mekler. “My Soul Got a Little Bit Cleaner”: Art experience in videogames. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):237:1– 237:19, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474664.

Bisafar:2020:SYA

[BWDP20] Farnaz Irannejad Bisafar, Brooke Foucault Welles, Catherine D’Ignazio, and Andrea G. Parker. Supporting youth activists? strategic use of social media: a qualitative investigation of design opportunities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):109:1–109:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415180.

Bemmann:2022:ITC

[BWE^{+}22] Florian Bemmann, Maximiliane Windl, Jonas Erbe, Sven Mayer, and Heinrich Hussmann. The influence of transparency and control on the willingness of data sharing in adaptive mobile apps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):189:1–189:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546724.

Belo:2023:CIS

[BWFG23] João Marcelo Evangelista Belo, Jon Wissing, Tiare Feuchtner, and Kaj Grønbæk. CAD-Track: Instructions and support for orientation disambiguation of near-symmetrical objects. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ISS): 426:1–426:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3626462.

Bowen:2020:GPP

[BWV20] Judy Bowen, Marco Winckler, and Jean Vanderdonckt. A glimpse into the past, present, and future of engineering interactive computing systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):71:1–71:32, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3394973.

Bianchi:2024:IVM

[BYL^{+}24] Andrea Bianchi, Zhi Lin Yap, Punn Lertjaturaphat, Austin Z. Henley, Kongpyung Justin Moon, and Yoonji Kim. Inline visualization and manipulation of real-time hardware log for supporting debugging of embedded programs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 248:1–248:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3660250.

Banovic:2023:BTE

[BYRL23] Nikola Banovic, Zhuoran Yang, Aditya Ramesh, and Alice Liu. Being trustworthy is not enough: How untrustworthy artificial intelligence (AI) can deceive the end-users and gain their trust. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):27:1–27:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579460.

Balestra:2017:IWF

[BZC^{+}17] Martina Balestra, Lior Zalmanson, Coye Cheshire, Ofer Arazy, and Oded Nov. It was fun, but did it last?: The dynamic interplay between fun motives and contributors’ activity in peer production. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):21:1–21:13, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134656.

Bhuiyan:2020:IDC

[BZSM20] Md Momen Bhuiyan, Amy X. Zhang, Connie Moon Sehat, and Tanushree Mitra. Investigating differences in crowdsourced news credibility assessment: Raters, tasks, and expert criteria. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):93:1–93:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415164.

Cheng:2022:RIM

[CA22] Claire Wei Cheng and Leila Aflatoony. The reappropriation of instant messaging: Texting ourselves, message dumping, and revisiting conversations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):187:1–187:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3546722.

Cormi:2022:CAE

[CAATL22] Clément Cormi, Khuloud Abou-Amsha, Matthieu Tixier, and Myriam Lewkowicz. Considering the artifact ecology when supporting the evolution of practices — analyzing the parallel journeys of two teleconsultation software in a general hospital. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):2:1–2:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3492821.

Choi:2023:CVC

[CBPC23] Frederick Choi, Tanvi Bajpai, Sowmya Pratipati, and Eshwar Chandrasekharan. ConvEx: a visual conversation exploration system for discord moderators. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):262:1–262:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610053.

Constantinides:2024:GIR

[CBSQ24] Marios Constantinides, Edyta Paulina Bogucka, Sanja Scepanovic, and Daniele Quercia. Good intentions, risky inventions: a method for assessing the risks and benefits of AI in mobile and wearable uses. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):262:1–262:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676507.

Cho:2022:TPB

[CBV22] Janghee Cho, Samuel Beck, and Stephen Voida. Topophilia, placemaking, and boundary work:

Exploring the psychosocial impact of the COVID-19 work-from-home experience. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):24:1–24:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492843.

Chitayat:2024:HCT

[CBWD24] Alan Pedrassoli Chitayat, Florian Block, James A. Walker, and Anders Drachen. How could they win? An exploration of win condition for esports narratives in Dota 2. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):314:1– 314:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677079.

Chan:2018:SMI

[CCH^{+}18] Joel Chan, Joseph Chee Chang, Tom Hope, Dafna Shahaf, and Aniket Kittur. SOLVENT: a mixed initiative system for finding analogies between research papers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 31:1–31:21, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274300.

Chow:2019:CDC

[CCNY19] Kevin Chow, Caitlin Coyiuto, Cuong Nguyen, and Dongwook Yoon. Challenges and design considerations for multimodal asynchronous collaboration in VR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):40:1–40:24, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359142.

Chen:2018:DDP

[CCS^{+}18] Chunyang Chen, Xi Chen, Jiamou Sun, Zhenchang Xing, and Guoqiang Li. Data-driven proactive policy assurance of post quality in community Q&A sites. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 33:1–33:22, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274302.

Cook:2022:AVA

[CCW22] Christine L. Cook, Jie Cai, and Donghee Yvette Wohn. Awe versus aww: The effectiveness of two kinds of positive emotional stimulation on stress reduction for online content moderators. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):277:1–277:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555168.

Chen:2021:WAN

[CCWH21] Xinyue Chen, Si Chen, Xu Wang, and Yun Huang. “I was afraid, but now I enjoy being a streamer!”: Understanding the challenges and prospects of using live streaming for online education. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):237:1–237:32, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432936.

Chen:2020:URI

[CCX^{+}20] Zhilong Chen, Hancheng Cao, Fengli Xu, Mengjie Cheng, Tao Wang, and Yong Li. Understanding the role of intermediaries in online social e-commerce: an exploratory study of Beidian. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):114:1–114:24, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415185.

Choudhury:2023:PVCa

[CDG^{+}23] Munmun De Choudhury, Xianghua Ding, Shion Guha, Aparecido Fabiano Pinatti de Carvalho, Hideaki Kuzuoka, Katharina Reinecke, Hao-Chuan Wang, and Naomi Yamashita. PACMHCI V7, CSCW1, April 2023 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):21:1–21:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579454.

Choudhury:2024:PVC

[CDG^{+}24] Munmun De Choudhury, Xianghua Ding, Shion Guha, Aparecido Fabiano Pinatti de Carvalho, Daniel Cardoso Llach, Maryam Mustafa, Daniele Quercia, and Marisol Wong-Villacres. PACMHCI V8, CSCW1, April 2024 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):10:1–10:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637287.

Choudhury:2023:PVCb

[CDGdC23] Munmun De Choudhury, Xianghua Ding, Shion Guha, and Aparecido Fabiano Pinatti de Carvalho. PACMHCI V7, CSCW2, October 2023 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):235:1–235:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3609976.

Chowdhury:2023:PPA

[CDIH23] Sohan Chowdhury, William Delamare, Pourang Irani, and Khalad Hasan. PAWS: Personalized arm and wrist movements with sensitivity mappings for controller-free locomotion in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):217:1–217:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604264.

Caidi:2023:RCS

[CEJC23] Nadia Caidi, Cansu Ekmekcioglu, Rojin Jamali, and Priyank Chandra. (Re)Capturing the spirit of Ramadan: Technoreligious practices in the time of COVID-19. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):249:1–249:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610040.

Caraway:2017:FDN

[CEM17] Monica Caraway, Daniel A.

Epstein, and Sean A. Munson. Friends don’t need receipts: The curious case of social awareness streams in the mobile payment app venmo. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):28:1–28:17, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134663.

Colley:2023:EUC

[CERR23] Mark Colley, Cristina Evangelista, Tito Daza Rubiano, and Enrico Rukzio. Effects of urgency and cognitive load on interaction in highly automated vehicles. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):207:1–207:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604254.

Caldeira:2021:TSD

[CFD^{+}21] Clara Caldeira, Mayara Costa Figueiredo, Lucy Dodakian, Cleidson R. B. de Souza, Steven C. Cramer, and Yunan Chen. Towards supporting data-driven practices in stroke telerehabilitation technology. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):25:1–25:33, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449099.

Campos:2017:MIT

[CFG^{+}17] José Creissac Campos, Camille Fayollas, Marcelo Gonçalves, Célia Martinie, David Navarre, Philippe Palanque, and Miguel Pinto. A more intelligent test case generation approach through task models manipulation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS): 9:1–9:20, June 2017. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3095811.

Colley:2022:IEE

[CFR22] Mark Colley, Tim Fabian, and Enrico Rukzio. Investigating the effects of external communication and automation behavior on manual drivers at intersections. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 176:1–176:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546711.

Callaghan:2018:MHM

[CGM^{+}18] William Callaghan, Joslin Goh, Michael Mohareb, Andrew Lim, and Edith Law. MechanicalHeart: a human-machine framework for the classification of phonocardiograms. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 28:1–28:17, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274297.

Chan:2021:MMP

[CGS^{+}21] Tina Chan, Robert P. Gauthier, Ally Suarez, Nicholas F. Sia, and James R. Wallace. Merlynne: Motivating peerto-peer cognitive behavioral therapy with a serious game. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):250:1– 250:23, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474677.

Chauhan:2020:TPS

[CH20] Apoorva Chauhan and Amanda Lee Hughes. Trustworthiness perceptions of social media resources named after a crisis event. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):044:1–044:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392849.

Cakir:2024:WER

[ÇH24] Mehtap Çakir and Anke Huckauf. What eyeblinks reveal: Interpersonal synchronization in dyadic interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):240:1–240:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655614.

Chung:2017:PLE

[CHC^{+}17] Taejoong Chung, Jinyoung Han, Daejin Choi, Ted Taekyoung Kwon, Jong-Youn Rha, and Hyunchul Kim. Privacy leakage in eventbased social networks: a Meetup case study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):35:1–35:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134670.

Cockburn:2024:UIE

[CHCG24] Andy Cockburn, Declan Hills, Zhe Chen, and Carl Gutwin. User interface evaluation through implicit-association tests. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (EICS):262:1–262:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3664636.

Coppens:2024:SMP

[CHS^{+}24] Adrien Coppens, Johannes Hermen, Lou Schwartz, Christian Moll, and Valérie Maquil. Supporting mixed-presence awareness across wallsized displays using a tracking pipeline based on depth cameras. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (EICS):260:1–260:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3664634.

Cutting:2022:LDI

[CI22] Joe Cutting and Ioanna Iacovides. Learning by doing: Intrinsic integration directs attention to increase learning in games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):240:1– 240:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549503.

Calefato:2018:ICC

[CILM18] Fabio Calefato, Giuseppe Iaffaldano, Filippo Lanubile, and Federico Maiorano. Investigating crowd creativity in online music communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 27:1–27:21, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274296.

Cheatle:2023:RCC

[CJ23] Amy Cheatle and Steven Jackson. (Re)collecting craft: Reviving materials, techniques, and pedagogies of craft for computational makers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):250:1–250:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610041.

Cmentowski:2022:ORN

[CKK22] Sebastian Cmentowski, Fabian Kievelitz, and Jens Harald Krueger. Outpace reality: a novel augmented-walking technique for virtual reality games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):246:1– 246:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549509.

Cha:2018:CAU

[CKP^{+}18] Yoonjeong Cha, Jongwon Kim, Sangkeun Park, Mun Yong Yi, and Uichin Lee. Complex and ambiguous: Understanding sticker misinterpretations in instant messaging. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 30:1–30:22, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274299.

Campos:2022:DPP

[CKP22] Cuauhtli Campos, Matjaz Kljun, and Klen Copic Pucihar. Dynamic pinhole paper: Interacting with horizontal displays through perforated paper. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):567:1–567:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567720.

Chiang:2018:CCP

[CKS18] Chun-Wei Chiang, Anna Kasunic, and Saiph Savage. Crowd coach: Peer coaching for crowd workers’ skill growth. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (CSCW):37:1–37:17, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274306.

Campos:2022:LBP

[CKSP22] Cuauhtli Campos, Matjaž Kljun, Jakub Sandak, and Klen Čopič Pucihar. LightMeUp: Backprint illumination paper display with multistable visuals. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):573:1–573:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3570333.

Cardoso-Llach:2021:AIE

[CLKEF21] Daniel Cardoso-Llach, Eric Kaltman, Emek Erdolu, and Zachary Furste. An archive of interfaces: Exploring the potential of emulation for software research, pedagogy, and design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):294:1–294:22, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476035.

Cuadra:2021:MBR

[CLL^{+}21] Andrea Cuadra, Shuran Li, Hansol Lee, Jason Cho, and Wendy Ju. My bad! repairing intelligent voice assistant errors improves interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):27:1–27:24, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449101.

Chung:2021:UHS

[CLP^{+}21] Chaeyeon Chung, Jungsoo Lee, Kyungmin Park, Junsoo Lee, Minjae Kim, Mookyung Song, Yeonwoo Kim, Jaegul Choo, and Sungsoo Ray Hong. Understanding humanside impact of sampling image batches in subjective attribute labeling. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):296:1–296:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476037.

Coppers:2019:FFA

[CLV^{+}19] Sven Coppers, Kris Luyten, Davy Vanacken, David Navarre, Philippe Palanque, and Christine Gris. Fortunettes: Feedforward about the future state of GUI widgets. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):20:1–20:20, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331162.

Chen:2020:EUD

[CLwH^{+}20] Yu-Chun Chen, Chia-Ying Liao, Shuo wen Hsu, Da-Yuan Huang, and Bing-Yu Chen. Exploring user defined gestures for ear-based interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 186:1–186:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427314.

Chevalier:2020:PHV

[CM20] Fanny Chevalier and Nicolai Marquardt. PACM HCI V4 ISS, November 2020 — editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (ISS):182:1, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427310.

Canny:2021:EMB

[CMN^{+}21] Alexandre Canny, Célia Martinie, David Navarre, Philippe Palanque, Eric Barboni, and Christine Gris. Engineering modelbased software testing of WIMP interactive applications: a process based on formal models and the SQUAMATA tool. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 207:1–207:30, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3461729.

Choi:2021:PSC

[CMP^{+}21] Yoonseo Choi, Toni-Jan Keith Palma Monserrat, Jeongeon Park, Hyungyu Shin, Nyoungwoo Lee, and Juho Kim. ProtoChat: Supporting the conversation design process with crowd feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):225:1–225:27, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432924.

Carayon:2024:SPE

[CMP^{+}24] Axel Carayon, Célia Martinie, Philippe Palanque, Eric Barboni, and Sandra Steere. A systematic process to engineer dependable integration of frame-based input devices in a multimodal input chain: Application to rehabilitation in healthcare. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 259:1–259:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3664633.

Cerna:2022:SSS

[CMRH22] Katerina Cerna, Claudia Müller, Dave Randall, and Martin Hunker. Situated scaffolding for sustainable participatory design: Learning online with older adults. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):12:1–12:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492831.

Cho:2022:AAL

[CMSA22] Eugene Cho, Nasim Motalebi, S. Shyam Sundar, and Saeed Abdullah. Alexa as an active listener: How backchanneling can elicit selfdisclosure and promote user experience. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):273:1–273:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555164.

Chua:2017:KFC

[CMY^{+}17] Soon Hau Chua, Toni-Jan Keith Palma Monserrat, Dongwook Yoon, Juho Kim, and Shengdong Zhao. Korero: Facilitating complex referencing of visual materials in asynchronous discussion interface. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):34:1–34:19, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134669.

Calvary:2017:ES

[CNC^{+}17] Gaelle Calvary, Jeffrey Nichols, José Creissac Campos, Nuno Nunes, and Pedro Campos. Editorial statement. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):1:1–1:2, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3106386.

Chounta:2017:WSE

[CNH^{+}17] Irene-Angelica Chounta, Alexander Nolte, Tobias Hecking, Rosta Farzan, and Thomas Herrmann. When to say “Enough is Enough!”: a study on the evolution of collaboratively created process models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):33:1–33:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134668.

Caldeira:2023:HDC

[CNH^{+}23] Clara Caldeira, Novia Nurain, Anna A. Heintzman, Haley Molchan, Kelly Caine, George Demiris, Katie A. Siek, Blaine Reeder, and Kay Connelly. How do I compare to the other people¿‘: Older adults’ perspectives on personal smart home data for self-management”. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):238:1–238:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610029.

Caddle:2023:DRC

[CNM^{+}23] Xavier V. Caddle, Nurun Naher, Zachary P. Miller, Karla Badillo-Urquiola, and Pamela J. Wisniewski. Duty to respond: The challenges social service providers face when charged with keeping youth safe online. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):6:1–6:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567556.

Cho:2020:SYC

[COKL20] Hyunsung Cho, Jinyoung Oh, Juho Kim, and Sung-Ju Lee. I share, you care: Private status sharing and sender-controlled notifications in mobile instant messaging. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):034:1–034:25, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392839.

Casarin:2018:UUT

[CPB18] Julien Casarin, Nicolas Pacqueriaud, and Dominique Bechmann. UMI3D: a Unity3D toolbox to support CSCW systems properties in generic 3D user interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 29:1–29:20, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274298.

Chandrasekharan:2017:YCS

[CPS^{+}17] Eshwar Chandrasekharan, Umashanthi Pavalanathan, Anirudh Srinivasan, Adam Glynn, Jacob Eisenstein, and Eric Gilbert. You can’t stay here: The efficacy of Reddit’s 2015 ban examined through hate speech. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):31:1–31:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134666.

Cho:2020:RCG

[CR20] Janghee Cho and Emilee Rader. The role of conversational grounding in supporting symbiosis between people and digital assistants. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):033:1–033:28, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392838.

Cherian:2024:STB

[CRM^{+}24] Josh Cherian, Samantha Ray, Thomas Mernar, Paul Taele, Helen Mach, Jung In Koh, Ping Ye, and Tracy Hammond. A step toward better care: Understanding what caregivers and residents in assisted living facilities value in health monitoring systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):13:1–13:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637290.

Chakroborti:2021:DRI

[CRN21] Debasish Chakroborti, Banani Roy, and Sristy Sumana Nath. Designing for recommending intermediate states in a scientific workflow management system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 198:1–198:29, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457145.

Carcangiu:2018:GGG

[CS18a] Alessandro Carcangiu and Lucio Davide Spano. G-Gene: a gene alignment method for online partial stroke gestures recognition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):13:1–13:17, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229095.

Cheon:2018:SLN

[CS18b] EunJeong Cheon and Norman Makoto Su. ’Staged for Living’: Negotiating objects and their values over a porous boundary. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 36:1–36:24, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274305.

Canossa:2021:HTD

[CSA^{+}21] Alessandro Canossa, Dmitry Salimov, Ahmad Azadvar, Casper Harteveld, and Georgios Yannakakis. For honor, for toxicity: Detecting toxic behavior through gameplay. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):253:1– 253:29, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474680.

Chen:2023:UOV

[CSH^{+}23] Niu Chen, Frances Jihae Sin, Laura Mariah Herman, Cuong Nguyen, Ivan Song, and Dongwook Yoon. Using online videos as the basis for developing design guidelines: a case study of AR-based assembly instructions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):428:1–428:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626464.

Chandrasekharan:2018:IHR

[CSJ^{+}18] Eshwar Chandrasekharan, Mattia Samory, Shagun Jhaver, Hunter Charvat, Amy Bruckman, Cliff Lampe, Jacob Eisenstein, and Eric Gilbert. The Internet’s hidden rules: an empirical study of Reddit norm violations at micro, meso, and macro scales. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 32:1–32:25, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274301.

Chang:2024:WLT

[CSL^{+}24] Tang-Jie Chang, Li-Ting Su, Yong-Han Lin, Jie Tsai, Zi-Xun Tang, and Yung-Ju Chang. “I Want Lower Tone for Work-Related Notifications”: Exploring the effectiveness of user-assigned notification alerts in improving user speculation of and attendance to mobile notifications. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):267:1–267:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3676512.

Chakraborty:2017:TES

[CSMG17] Abhijnan Chakraborty, Rajdeep Sarkar, Ayushi Mrigen, and Niloy Ganguly. Tabloids in the era of social media?: Understanding the production and consumption of clickbaits in Twitter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):30:1–30:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134665.

Chhibber:2021:TNK

[CSMV21] Nalin Chhibber, Hemant Bhaskar Surale, Fabrice Matulic, and Daniel Vogel. Typealike: Near-keyboard hand postures for expanded laptop interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 486:1–486:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486952.

Cho:2019:IEB

[CSN19] Bryden Cho, Chengzheng Sun, and Agustina Ng. Issues and experiences in building heterogeneous co-editing systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):245:1–245:28, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361126.

Cartwright:2017:SSI

[CSS^{+}17] Mark Cartwright, Ayanna Seals, Justin Salamon, Alex Williams, Stefanie Mikloska, Duncan Mac-Connell, Edith Law, Juan P. Bello, and Oded Nov. Seeing sound: Investigating the effects of visualizations and complexity on crowdsourced audio annotations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):29:1–29:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134664.

Chioccarello:2023:FFS

[CST^{+}23]  Stefano  Chioccarello,

Arthur Sluÿters, Alberto Testolin, Jean Vanderdonckt, and Sébastien Lambot. FORTE: Few samples for recognizing hand gestures with a smartphone-attached radar. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 179:1–179:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593231.

Choi:2018:WTM

[CT18] Joohee Choi and Yla Tausczik. Will too many editors spoil the tag?: Conflicts and alignment in Q&A categorization. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 38:1–38:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274307.

Chiu:2024:NCP

[CTH24] Maria Chiu, Elina Tochilnikova, and Casper Harteveld. From novelty to clinical practice: Exploring VR exergames with physical therapists. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):303:1– 303:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677068.

Chiossi:2023:AVC

[CTWM23] Francesco Chiossi, Yagiz Turgut, Robin Welsch, and Sven Mayer. Adapting visual complexity based on electrodermal activity improves working memory performance in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):196:1–196:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604243.

Coppers:2022:FPM

[CVL22] Sven Coppers, Davy Vanacken, and Kris Luyten. FortClash: Predicting and mediating unintended behavior in home automation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 154:1–154:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3532204.

Compton:2018:LPU

[CWB^{+}18] Ryan J. Compton, Jeff Warshaw, Hernan Badenes, Barton Smith, and Steve Whittaker. Living in the present: Understanding long-term content referencing in enterprise online communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 39:1–39:21, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274308.

Chang:2024:EDA

[CWC^{+}24] Yu-Cheng Chang, Yen-Pu Wang, Chiao-Ju Chang, Wei Tian Mireille Tan, Yu Lun Hsu, Yu Chen, and Mike Y. Chen. Experience from designing augmented reality browsing interfaces for real-world walking scenarios. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 255:1–255:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3676500.

Chamunorwa:2023:ECM

[CWK^{+}23] Michael Chamunorwa, Mikolaj P. Wozniak, Susanna Krämer, Heiko Müller, and Susanne Boll. An empirical comparison of moderated and unmoderated gesture elicitation studies on soft surfaces and objects for smart home control. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):198:1–198:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604245.

Colley:2022:SES

[CWR22] Mark Colley, Bastian Wankmüller, and Enrico Rukzio. A systematic evaluation of solutions for the final 100m challenge of highly automated vehicles. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):178:1–178:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546713.

Chen:2017:CCD

[CXL17] Chunyang Chen, Zhenchang Xing, and Yang Liu. By the community & for the community: a deep learning approach to assist collaborative editing in Q&A sites. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):32:1–32:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134667.

Cao:2021:MTW

[CYC^{+}21] Hancheng Cao, Vivian Yang, Victor Chen, Yu Jin Lee, Lydia Stone, N’godjigui Junior Diarrassouba, Mark E. Whiting, and Michael S. Bernstein. My team will go on: Differentiating high and low viability teams through team interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):230:1–230:27, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432929.

Chau:2021:CLD

[CYGW21] Edwin Chau, Jiakun Yu, Cagatay Goncu, and Anusha Withana. Composite line designs and accuracy measurements for tactile line tracing on touch surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 491:1–491:17, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488536.

Chen:2018:BLF

[CYW18] Mei-Ling Chen, Naomi Yamashita, and Hao-Chuan Wang. Beyond lingua franca: Systemfacilitated language switching diversifies participation in multiparty multilingual communication. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 34:1–34:22, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274303.

Cardieri:2019:PEE

[dACZ19] Giulia de Andrade Cardieri and Luciana A. M. Zaina. PWA-EU: Extending PWA approach for promoting customization based on user preferences. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):10:1–10:28, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331152.

Dittus:2017:CET

[DAQ17] Martin Dittus, Luca Maria Aiello, and Daniele Quercia. Community engagement triage: Lightweight prompts for systematic reviews. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(CSCW): 39:1–39:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3134674.

Duckert:2022:PPH

[DB22] Melanie Duckert and Louise Barkhuus. Protecting personal health data through privacy awareness: a study of perceived data privacy among people with chronic or long-term illness. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP): 11:1–11:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3492830.

Delgado:2022:UTP

[DBL22] Fernando Delgado, Solon Barocas, and Karen Levy. An uncommon task: Participatory design in legal AI. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):51:1–51:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512898.

DiGioia:2022:IUA

[DBPA22] Francesco Riccardo Di Gioia, Eugenie Brasier, Emmanuel Pietriga, and Caroline Appert. Investigating the use of AR glasses for content annotation on mobile devices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):574:1–574:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567727.

Dechant:2021:HAC

[DBS^{+}21] Martin Johannes Dechant, Max V. Birk, Youssef Shiban, Knut Schnell, and Regan L. Mandryk. How avatar customization affects fear in a game-based digital exposure task for social anxiety. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):248:1– 248:27, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474675.

Dittus:2017:PPF

[DC17] Martin Dittus and Licia Capra. Private peer feedback as engagement driver in humanitarian mapping. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(CSCW): 40:1–40:18, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3134675.

Darian:2023:EDF

[DCM^{+}23] Shiva Darian, Aarjav Chauhan, Ricky Marton, Janet Ruppert, Kathleen Anderson, Ryan Clune, Madeline Cupchak, Max Gannett, Joel Holton, Elizabeth Kamas, Jason Kibozi-Yocka, Devin Mauro-Gallegos, Simon Naylor, Meghan O’Malley, Mehul Patel, Jack Sandberg, Troy Siegler, Ryan Tate, Abigil Temtim, Samantha Whaley, and Amy Voida. Enacting data feminism in advocacy data work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):47:1–47:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579480.

Das:2024:DTG

[DCM24] Debasree Das, Sandip Chakraborty, and Bivas Mitra. DriveR: Towards generating a dynamic road safety map with causal contexts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):253:1–253:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676498.

deCarvalho:2023:GEA

[dCRM^{+}23] Aparecido Fabiano Pinatti de Carvalho, Sarah Reichel, Marcel Manuel Sanchez Martin, Eva Sonja Allen, and Marcus Schweitzer. Group effect aspects in digitalisation production contexts: Articulation spaces for emerging cooperation challenges. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):10:1–10:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567560.

Dey:2019:VIA

[DDFK19] Sanorita Dey, Brittany R. L. Duff, Wai-Tat Fu, and Karrie Karahalios. VidLyz: an interactive approach to assist novice entrepreneurs in making persuasive campaign videos. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 43:1–43:26, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359145.

Darian:2023:CIP

[DDV23] Shiva Darian, Brianna Dym, and Amy Voida. Competing imaginaries and partisan divides in the data rhetoric of advocacy organizations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):259:1–259:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3610050.

DAngelo:2018:SCR

[DDZ18] Gabriele D’Angelo, Angelo Di Iorio, and Stefano Zacchiroli. Spacetime characterization of real-time collaborative editing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 41:1–41:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274310.

Drey:2021:SIC

[DFJ^{+}21] Tobias Drey, Fabian Fischbach, Pascal Jansen, Julian Frommel, Michael Rietzler, and Enrico Rukzio. To be or not to be stuck, or is it a continuum?: a systematic literature review on the concept of being stuck in games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):229:1– 229:35, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474656.

Dagan:2022:PIP

[DGR^{+}22] Ella Dagan, Ana María Cárdenas Gasca, Ava Robinson, Anwar Noriega, Yu Jiang Tham, Rajan Vaish, and Andrés Monroy-Hernández. Project IRL: Playful co-located interactions with mobile augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):62:1–62:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512909.

Davis:2018:ECP

[DHH18] Dan Davis, Claudia Hauff, and Geert-Jan Houben. Evaluating crowdworkers as a proxy for online learners in video-based learning contexts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 42:1–42:16, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274311.

Duchowski:2023:PVE

[DKBG23] Andrew T. Duchowski, Krzysztof Krejtz, Zoya Bylinskii, and Hans Gellersen. PACMHCI V7, ETRA, May 2023 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):156:1–156:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3591125.

Duchowski:2024:PVE

[DKKL24] Andrew Duchowski, Peter Kiefer, Krzysztof Krejtz, and Jochen Laubrock. PACMHCI V8, ETRA, May 2024 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):221:1–221:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655595.

Dubois:2020:GDG

[DMC^{+}20] Patrick Marcel Joseph Dubois, Mahya Maftouni, Parmit K. Chilana, Joanna McGrenere, and Andrea Bunt. Gender differences in graphic design Q&As: How community and site characteristics contribute to gender gaps in answering questions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):113:1–113:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415184.

Deja:2022:SAP

[DMPK22] Jordan Aiko Deja, Sven Mayer, Klen Copic Pucihar, and Matjaz Kljun. A survey of augmented piano prototypes: Has augmentation improved learning experiences? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):566:1–566:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567719.

Dietz:2024:DPD

[DMRA24] Felix Dietz, Lukas Mecke, Daniel Riesner, and Florian Alt. Delusio — plausible deniability for face recognition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 249:1–249:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676494.

Dixon:2022:HYH

[DNBB^{+}22] Matt Dixon, James Nicholson, Dawn Branley-Bell, Pam Briggs, and Lynne Coventry. Holding your hand on the danger button: Observing user phish detection strategies across mobile and desktop. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 195:1–195:??, September 2022. CODEN ????

ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546730.

Datta:2017:IMI

[DPA17] Srayan Datta, Chanda Phelan, and Eytan Adar. Identifying misaligned inter-group links and communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):37:1–37:23, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134672.

Dym:2022:BPP

[DPF22] Brianna Dym, Namita Pasupuleti, and Casey Fiesler. Building a pillowfort: Political tensions in platform design and policy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):16:1–16:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492835.

Proenca:2021:CDS

[dQPMdHRdAN21] Mailson de Queiroz Proença, Vivian Genaro Motti, Kamila Rios da Hora Rodrigues, and Vânia Paula de Almeida Neris. Coping with diversity — a system for end-users to customize Web user interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 201:1–201:27, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457151.

Dhiman:2024:DMM

[DR24] Hitesh Dhiman and Carsten Röcker. Does the medium matter? A comparison of augmented reality media in instructing novices to perform complex, skill-based manual tasks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 249:1–249:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3660249.

Dosono:2017:CTC

[DRA^{+}17] Bryan Dosono, Yasmeen Rashidi, Taslima Akter, Bryan Semaan, and Apu Kapadia. Challenges in transitioning from civil to military culture: Hyper-selective disclosure through ICTs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):41:1–41:23, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134676.

Deri:2018:CLC

[DRAQ18] Sebastian Deri, Jeremie Rappaz, Luca Maria Aiello, and Daniele Quercia. Coloring in the links: Capturing social ties as they are perceived. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 43:1–43:18, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274312.

Gutierrez:2024:DGE

[DRDS24] Jose Pablo De La Rosa Gutierrez, Thiago Rocha Silva, Yvonne Dittrich, and Anders Stengaard Sørensen. Design goals for end-user development of robot-assisted physical training activities: a participatory design study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 258:1–258:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3664632.

Durkin:2018:SGF

[DRK18] Celia Durkin, Federico Rossano, and Scott Klemmer. Score-group framing negatively impacts peer evaluations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (CSCW):45:1–45:19, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274314.

Dube:2022:PTD

[DRL^{+}22] Tafadzwa Joseph Dube, Yuan Ren, Hannah Limerick, I. Scott MacKenzie, and Ahmed Sabbir Arif. Push, tap, dwell, and pinch: Evaluation of four mid-air selection methods augmented with ultrasonic haptic feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):565:1–565:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567718.

Deng:2023:EAB

[DRP^{+}23] Shuwen Deng, David R. Reich, Paul Prasse, Patrick Haller, Tobias Scheffer, and Lena A. Jäger. Eyettention: an attention-based dualsequence model for predicting human scanpaths during reading. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):162:1–162:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591131.

deRyckel:2022:SSD

[dRSV22] Xavier de Ryckel, Arthur Sluÿters, and Jean Vanderdonckt. SnappView, a software development kit for supporting enduser mobile interface review. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 173:1–173:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3534527.

Dessi:2020:DEG

[DS20] Stefano Dessì and Lucio Davide Spano. DG3: Exploiting gesture declarative models for sample generation and online recognition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):82:1–82:21, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397870.

deSouza:2023:CCT

[dSdSdSF^{+}23] Adriano Neves de Souza, Sírius Thadeu Ferreira da Silva, Juliana Baptista dos Santos França, Angélica Fonseca da Silva Dias, Jonice Oliveira, and Adriana S. Vivacqua. Communication channels and their challenges: an analysis of software development teams during the COVID-19 pandemic. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):3:1–3:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3567553.

Dissanayake:2022:TTU

[DTE^{+}22] Vipula Dissanayake, Vanessa Tang, Don Samitha Elvitigala, Elliott Wen, Michelle Wu, and Suranga Nanayakkara. Troi: Towards understanding users perspectives to mobile automatic emotion recognition system in their natural setting. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):203:1–203:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546738.

DeVito:2018:TGF

[DWB18] Michael A. DeVito, Ashley Marie Walker, and Jeremy Birnholtz. ’Too Gay for Facebook’: Presenting LGBTQ+ identity throughout the personal social media ecosystem. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 44:1–44:23, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274313.

Dickson:2021:HTS

[DWMV21] Terence Dickson, Rina R.

Wehbe, Fabrice Matulic, and Daniel Vogel. HybridPointing for touch: Switching between absolute and relative pointing on large touch screens. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 495:1–495:22, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488540.

Das:2024:SFH

[DWSF24] Anwesha Das, Zekun Wu, Iza Skrjanec, and Anna Maria Feit. Shifting focus with HCEye: Exploring the dynamics of visual highlighting and cognitive load on user attention and saliency prediction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):236:1–236:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3655610.

Dillahunt:2017:SEC

[DWW^{+}17] Tawanna R. Dillahunt, Xinyi Wang, Earnest Wheeler, Hao Fei Cheng, Brent Hecht, and Haiyi Zhu. The sharing economy in computing: a systematic literature review. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):38:1–38:26, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134673.

Dian:2020:TDI

[DWZ^{+}20] Cao Dian, Dong Wang, Qian Zhang, Run Zhao, and Yinggang Yu. Towards domain-independent complex and fine-grained gesture recognition with RFID. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 187:1–187:22, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427315.

Duan:2023:FRE

[DYF^{+}23] Yongjie Duan, Jinyang Yu, Jianjiang Feng, Ke He, Jiwen Lu, and Jie Zhou. 3D finger rotation estimation from fingerprint images. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):431:1–431:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626467.

Ekmekcioglu:2023:MTA

[ECA23] Cansu Ekmekcioglu, Priyank Chandra, and Syed Ishtiaque Ahmed. A matter of time: Anticipation work and digital temporalities in refugee humanitarian assistance in Turkey. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):22:1–22:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3579455.

Emporio:2024:IER

[ECP^{+}24] Marco Emporio, Ariel Caputo, Deborah Pintani, Dong Seon Cheng, Thomas De Marchi, Gianmaria Forte, Franco Fummi, and Andrea Giachetti. Integration of extended reality with a cyber-physical factory environment and its digital twins. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 250:1–250:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3660246.

Evin:2022:CAG

[EHG22] Inan Evin, Perttu Hämäläinen, and Christian Guckelsberger. Cine-AI: Generating video game cutscenes in the style of human directors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):223:1– 223:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549486.

Early:2018:UGE

[EHH^{+}18] Kirstin Early, Jessica Hammer, Megan Kelly Hofmann, Jennifer A. Rode, Anna Wong, and Jennifer Mankoff. Understanding gender equity in author order assignment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 46:1–46:21, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274315.

Engel:2022:CRP

[EHZN22] Kristen Engel, Yiqing Hua, Taixiang Zeng, and Mor Naaman. Characterizing Reddit participation of users who engage in the QAnon conspiracy theories. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):53:1–53:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512900.

Epstein:2020:EDP

[EJB^{+}20] Daniel A. Epstein, Siyun Ji, Danny Beltran, Griffin D’Haenens, Zhaomin Li, and Tan Zhou. Exploring design principles for sharing of personal informatics data on ephemeral social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):95:1–95:24, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415166.

Emami:2024:IDD

[EJGL24] Parvin Emami, Yue Jiang, Zixin Guo, and Luis A. Leiva. Impact of design decisions in scanpath modeling. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):228:1–228:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655602.

Echtler:2023:SUC

[EMHW23] Florian Echtler, Vitus Maierhöfer, Nicolai Brodersen Hansen, and Raphael Wimmer. SurfaceCast: Ubiquitous, cross-device surface sharing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):439:1–439:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626475.

Eagle:2022:MDB

[EMS^{+}22] Tessa Eagle, Aman Mehrotra, Aayush Sharma, Alex Zuniga, and Steve Whittaker. “Money Doesn’t Buy You Happiness”: Negative consequences of using the freemium model for mental health apps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):265:1–265:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555155.

Elsden:2024:TGN

[EMT^{+}24] Chris Elsden, Evan Morgan, Ella Tallyn, Suzanne R. Black, Martin Disley, Burkhard Schafer, Dave Murray-Rust, and Chris Speed. A token gesture: Nontransferable NFTs, digital possessions and ownership design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):25:1–25:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637302.

Eska:2023:DLB

[EPW^{+}23] Bettina Eska, Marco Philip, Pawel W. Wozniak, Albrecht Schmidt, and Jakob Karolus. [Don’t] let the bodies HIIT the floor: Fostering body awareness in fastpaced physical activity using body-worn sensors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):203:1–203:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604250.

Ernala:2017:LMI

[ERB^{+}17] Sindhu Kiranmai Ernala, Asra F. Rizvi, Michael L. Birnbaum, John M. Kane, and Munmun De Choudhury. Linguistic markers indicating therapeutic outcomes of social media disclosures of schizophrenia. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):43:1–43:27, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134678.

Everitt:2022:IPP

[ERH^{+}22] Aluna Everitt, Anne Roudaut, Kasper Hornbæk, Mike Fraser, and Jason Alexander. Investigating pointing performance for tangible surfaces with physical 3D targets. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):583:1–583:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567736.

Elfares:2024:PAB

[ERH^{+}24] Mayar Elfares, Pascal Reisert, Zhiming Hu, Wenwu Tang, Ralf Küsters, and Andreas Bulling. PrivatEyes: Appearance-based gaze estimation using federated secure multi-party computation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):232:1–232:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655606.

Epstein:2017:SSE

[ERLW17] Robert Epstein, Ronald E.

Robertson, David Lazer, and Christo Wilson. Suppressing the search engine manipulation effect (SEME). Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):42:1–42:22, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134677.

Erete:2021:CBR

[ERT21] Sheena Erete, Yolanda A. Rankin, and Jakita O. Thomas. I can’t breathe: Reflections from black women in CSCW and HCI. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW3): 234:1–234:23, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3432933.

El-Raheb:2023:EDC

[ERVR^{+}23] Katerina El-Raheb, Vasilis Vlachokyriakos, Maria Roussou, Patrick Olivier, Tom Bartindale, and Andrew Garbett. Exploring digital communication needs of local communities and self-organized collectives. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI): 208:1–208:??, September 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3604255.

Elmimouni:2024:SSI

[ESA^{+}24] Houda Elmimouni, Yarden Skop, Norah Abokhodair, Sarah Rüller, Konstantin Aal, Anne Weibert, Adel Al-Dawood, Volker Wulf, and Peter Tolmie. Shielding or silencing?: An investigation into content moderation during the Sheikh Jarrah Crisis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):6:1–6:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3633071.

Ehsan:2023:CSG

[ESCR23] Upol Ehsan, Koustuv Saha, Munmun De Choudhury, and Mark O. Riedl. Charting the sociotechnical gap in explainable AI: a framework to address the gap in XAI. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):34:1–34:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579467.

Ertl:2019:SMH

[ETE^{+}19] Tanja Ertl, Sebastian Taugerbeck, Margarita Esau, Konstantin Aal, Peter Tolmie, and Volker Wulf. The social mile — how (psychosocial) ICT can help to promote resocialization and to overcome prison. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):248:1–248:31, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3370270.

Eardley:2023:EBA

[ETS^{+}23] Rachel Eardley, Emma L. Tonkin, Ewan Soubutts, Amid Ayobi, Gregory J. L. Tourte, Rachael Gooberman-Hill, Ian Craddock, and Aisling Ann O’Kane. Explanation before adoption: Supporting informed consent for complex machine learning and IoT health platforms. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):49:1–49:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579482.

Erickson:2019:EW

[EVC^{+}19] Ingrid Erickson, Adriana S. Vivacqua, Lars Rune Christensen, Naja L. Holten Moller, Eric Baumer, Yvette Wohn, Louise Barkhuus, and Lionel P. Robert. Editors’ welcome. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(GROUP): 229:1–229:2, December 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3361110.

Freeman:2022:AQI

[FA22] Guo Freeman and Dane Acena. “Acting Out” queer identity: The embodied visibility in social virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):263:1–263:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555153.

Freeman:2022:WTA

[FAMS22] Guo Freeman, Dane Acena, Nathan J. Mc-Neese, and Kelsea Schulenberg. Working together apart through embodiment: Engaging in everyday collaborative activities in social virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):17:1–17:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492836.

Fiesler:2019:CCC

[FB19] Casey Fiesler and Amy S. Bruckman. Creativity, copyright, and close-Knit communities: a case study of social norm formation and enforcement. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):241:1–241:24, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361122.

Fetter:2017:TJT

[FBG17] Mirko Fetter, David Bimamisa, and Tom Gross. TUIOFX: a JavaFX toolkit for shared interactive surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):10:1–10:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095812.

Freeman:2022:PVC

[FBNM22] Guo Freeman, Scott Bateman, Lennart Nacke, and Regan Mandryk. PACMHCI V6, CHI PLAY, October 2022 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):217:1– 217:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3548661.

Fan:2020:IHU

[FC20] Zhuzhi Fan and Céline Coutrix. Impact of hand used on one-handed back-of-device performance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 188:1–188:19, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427316.

Figueiredo:2018:EHD

[FCE^{+}18] Mayara Costa Figueiredo, Clara Caldeira, Elizabeth Victoria Eikey, Melissa Mazmanian, and Yunan Chen. Engaging with health data: The interplay between self-tracking activities and emotions in fertility struggles. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(CSCW): 40:1–40:20, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274309.

Faas:2019:JTJ

[FcLDM19] Travis Faas, I ching Liu, Lynn Dombrowski, and Andrew D. Miller. Jam today, jam tomorrow: Learning in online game jams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):240:1–240:27, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361121.

Figueiredo:2017:STF

[FCR^{+}17] Mayara Costa Figueiredo, Clara Caldeira, Tera L. Reynolds, Sean Victory, Kai Zheng, and Yunan Chen. Self-tracking for fertility care: Collaborative support for a highly personalized problem. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):36:1–36:21, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134671.

Fiesler:2020:MAL

[FD20] Casey Fiesler and Brianna Dym. Moving across lands: Online platform migration in fandom communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):042:1–042:25, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392847.

Fan:2022:AUD

[FF22] Bonnie Fan and Sarah E.

Fox. Access under duress: Pandemic-era lessons on digital participation and datafication in civic engagement. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):14:1–14:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492833.

Foong:2017:NES

[FGG17] Eureka Foong, Darren Gergle, and Elizabeth M. Gerber. Novice and expert sensemaking of crowdsourced design feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):45:1–45:18, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3134680.

Foucher:2024:UDI

[FH24] Valentin Foucher and Anke Huckauf. Unveiling deceptive intentions: Insights from eye movements and pupil size. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):238:1–238:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3655612.

Farkas:2024:HBP

[FHF24] Timea Farkas, Nathan Gerard Jayy Hughes, and Rebecca Fiebrink. How boardgame players imagine interacting with technology. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):313:1– 313:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677078.

Fiesler:2019:ECR

[Fie19]  Casey Fiesler.  Ethi-

cal considerations for research involving (speculative) public data. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):249:1–249:13, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3370271.

Fjeld:2021:PVI

[FJIH21] Morten Fjeld, Hans-Christian Jetter, Petra Isenberg, and Mark Hancock. PACMHCI V5, ISS, November 2021 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (ISS):483:1, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486933.

Foong:2021:CUH

[FKDG21] Eureka Foong, Joy O.

Kim, Mira Dontcheva, and Elizabeth M. Gerber. CrowdFolio: Understanding how holistic and decomposed workflows influence feedback on online portfolios. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):22:1–22:31, April 2021. CODEN ????

ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449096.

Friehs:2022:EEC

[FKF^{+}22] Maximilian A. Friehs, Madison Klarkowski, Julian Frommel, Cody Phillips, and Regan L. Mandryk. Enhanced esports: Community perspectives on performance enhancers in competitive gaming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):231:1– 231:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549494.

Frey:2019:PDW

[FKK19] Seth Frey, P. M. Krafft, and Brian C. Keegan. “This Place Does What It Was Built For”: Designing digital institutions for participatory change. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 32:1–32:31, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359134.

Faucett:2017:SLM

[FLC17] Heather A. Faucett, Matthew L. Lee, and Scott Carter. I should listen more: Real-time sensing and feedback of non-verbal communication in video telehealth. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):44:1–44:19, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3134679.

Fruchard:2020:SCM

[FLC20] Bruno Fruchard, Eric Lecolinet, and Olivier Chapuis. Side-crossing menus: Enabling large sets of gestures for small surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 189:1–189:19, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427317.

Funk:2021:DEI

[FLP^{+}21] Mathias Funk, Rong-Hao Liang, Philippe Palanque, Jun Hu, and Panos Markopoulos. Designing and engineering interactive computing systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 193:1–193:4, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3457140.

Feldman:2021:HWW

[FM21] Molly Q. Feldman and Brian James McInnis. How we write with crowds. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW3): 229:1–229:31, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3432928.

Frommel:2022:DQD

[FM22] Julian Frommel and Regan L. Mandryk. Daily quests or daily pests? The benefits and pitfalls of engagement rewards in games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):226:1– 226:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549489.

Franzen:2024:CPU

[FMBW24] Daniel Franzen, Claudia Müller-Birn, and Odette Wegwarth. Communicating the privacyutility trade-off: Supporting informed data donation with privacy decision interfaces for differential privacy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):32:1–32:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637309.

Formosa:2022:DES

[FOH^{+}22] Jessica Formosa, Nicholas O’Donnell, Ella M. Horton, Selen Türkay, Regan L. Mandryk, Michael Hawks, and Daniel Johnson. Definitions of esports: a systematic review and thematic analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):227:1– 227:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549490.

Fulker:2024:CGE

[FR24] Zachary Fulker and Christoph Riedl. Cooperation in the gig economy: Insights from upwork freelancers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):37:1–37:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637314.

Fung:2022:LUE

[FRH^{+}22] Kin Pong Fung, Katja Rogers, Stuart Hallifax, Gabrielle S. Woodside, Daniel Vogel, and Lennart E. Nacke. Light-Play: Using an external ambient lighting strip for video game indicators. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):241:1– 241:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549504.

Frau:2023:XEB

[FSAN23] Vittoria Frau, Lucio Davide Spano, Valentino Artizzu, and Michael Nebeling. XRSpotlight: Example-based programming of XR interactions using a rule-based approach. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 185:1–185:??, June 2023.

CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593237.

Figueiredo:2021:UDA

[FSC21] Mayara Costa Figueiredo, H. Irene Su, and Yunan Chen. Using data to approach the unknown: Patients’ and healthcare providers? Data practices in fertility challenges. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):227:1–227:35, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432926.

Feinberg:2020:NRR

[FSN^{+}20] Melanie Feinberg, Will Sutherland, Sarah Beth Nelson, Mohammad Hossein Jarrahi, and Arcot Rajasekar. The new reality of reproducibility: The role of data work in scientific research. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):035:1–035:22, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392840.

Fan:2024:DPS

[FSRC24] Zhuzhi Fan, Alexis Sanson, Thomas Rames, and Céline Coutrix. Design and perception of a soft shape change beneath a smartwatch. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 250:1–250:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3676495.

Feger:2022:EDE

[FSSK22] Sebastian S. Feger, Lars Semmler, Albrecht Schmidt, and Thomas Kosch. ElectronicsAR: Design and evaluation of a mobile and tangible highfidelity augmented electronics toolkit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):587:1–587:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567740.

Feuston:2020:CED

[FTP20] Jessica L. Feuston, Alex S. Taylor, and Anne Marie Piper. Conformity of eating disorders through content moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):040:1–040:28, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392845.

Fang:2022:UES

[FWY^{+}22] Jingchao Fang, Yanhao Wang, Chi-Lan Yang, Ching Liu, and Hao-Chuan Wang. Understanding the effects of structured note-taking systems for video-based learners in individual and social learning contexts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):21:1–21:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492840.

Fan:2022:RLT

[FX22] Neil Xu Fan and Robert Xiao. Reducing the latency of touch tracking on ad-hoc surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):577:1–577:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567730.

Freedman:2024:CAC

[FYS^{+}24] Hayden Freedman, Neil Young, David Schaefer, Qingyu Song, André van der Hoek, and Bill Tomlinson. Construction and analysis of collaborative educational networks based on student concept maps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):36:1–36:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637313.

Fowler:2022:FCC

[FZM22] John Fowler, Mark Zachry, and David W. McDonald. Fostering communication: Characterizing the concerns of former Foster youth in an online community. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):15:1–15:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3492834.

Fiesler:2024:RHS

[FZP^{+}24] Casey Fiesler, Michael Zimmer, Nicholas Proferes, Sarah Gilbert, and Naiyan Jones. Remember the human: a systematic review of ethical considerations in Reddit research. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):5:1–5:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3633070.

Fink:2022:RLA

[FZRJ22] Daniel Immanuel Fink, Johannes Zagermann, Harald Reiterer, and Hans-Christian Jetter. Re-locations: Augmenting personal and shared workspaces to support remote collaboration in incongruent spaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):556:1–556:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567709.

Garg:2021:UFN

[Gar21] Radhika Garg. Understanding families’ Non-/ use practices and choices: The case of smart speakers and smart interactive toys. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):295:1–295:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476036.

Gligoric:2019:CEB

[GAW19] Kristina Gligorić, Ashton Anderson, and Robert West. Causal effects of brevity on style and success in social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 45:1–45:23, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359147.

Goncalves:2018:KBK

[GB18] Afonso Gonçalves and Sergi Bermúdez. KAVE: Building Kinect based CAVE automatic virtual environments, methods for surround-screen projection management, motion parallax and fullbody interaction support. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS): 10:1–10:15, June 2018. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3229092.

Greipl:2021:FBE

[GBN21] Simon Greipl, Katharina Bernecker, and Manuel Ninaus. Facial and bodily expressions of emotional engagement: How dynamic measures reflect the use of game elements and subjective experience of emotions and effort. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):240:1– 240:25, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474667.

Geissler:2024:HSG

[GBZ^{+}24] Daniel Geißler, Hymalai Bello, Esther Zahn, Emil Woop, Bo Zhou, Paul Lukowicz, and Jakob Karolus. Head ’n shoulder: Gesture-driven biking through capacitive sensing garments to innovate hands-free interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 265:1–265:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676510.

Green:2019:PLA

[GC19] Ben Green and Yiling Chen. The principles and limits of algorithmin-the-loop decision making. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 50:1–50:24, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359152.

Goncalves:2023:CCR

[GCSB23] Pavlína Wurzel Gonçalves, Gül Calikli, Alexander Serebrenik, and Alberto Bacchelli. Competencies for code review. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):38:1–38:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3579471.

Ghaemi:2023:DCD

[GEEJ23] Zeinab Ghaemi, Barrett Ens, Ulrich Engelke, and Bernhard Jenny. Drawing connections: Designing situated links for immersive maps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):205:1–205:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604252.

Guo:2022:GET

[GF22] Jiajing Guo and Susan R.

Fussell. “It’s Great to Exercise Together on Zoom!”: Understanding the practices and challenges of live stream group fitness classes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):71:1–71:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512918.

Gomes:2023:FEU

[GGE23] Carlos Gomes, Tiago Guedes, and Augusto Esteves. A first exploration on the use of headmounted augmented reality in the context of the Portuguese military. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):192:1–192:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604239.

Gardey:2020:UPA

[GGF^{+}20] Juan Cruz Gardey, Alejandra Garrido, Sergio Firmenich, Julián Grigera, and Gustavo Rossi. UX-Painter: an approach to explore interaction fixes in the browser. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):89:1–89:21, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397877.

Gruenefeld:2022:AHP

[GGL^{+}22] Uwe Gruenefeld, Alexander Geilen, Jonathan Liebers, Nick Wittig, Marion Koelle, and Stefan Schneegass. ARm haptics: 3D-printed wearable haptics for mobile augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):193:1–193:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546728.

Galeote:2021:GBC

[GH21] Daniel Fernández Galeote and Juho Hamari. Game-based climate change engagement: Analyzing the potential of entertainment and serious games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):226:1– 226:21, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474653.

Gu:2021:LLD

[GHHC21] Hongyan Gu, Jingbin Huang, Lauren Hung, and Xiang ‘Anthony’ Chen. Lessons learned from designing an AI-Enabled diagnosis tool for pathologists. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):10:1–10:25, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449084.

Guttman:2021:PRI

[GHHS21] Rotem D. Guttman, Jessica Hammer, Erik Harpstead, and Carol J. Smith. Play for real(ism)

— using games to predict Human-AI interactions in the real world. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):228:1– 228:17, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474655.

Gilbert:2020:RWL

[Gil20] Sarah A. Gilbert. “I run the world’s largest historical outreach project and it’s on a cesspool of a website.” moderating a public scholarship site on Reddit: a case study of r/AskHistorians. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):019:1–019:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392822.

Gilbert:2023:TIM

[Gil23] Sarah Gilbert. Towards intersectional moderation: an alternative model of moderation built on care and power. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):256:1–256:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3610047.

Grant:2023:HBB

[GKHD23] Alana Grant, Vilma Kankaanpää, and Ilyena Hirskyj-Douglas. Humble beginnings: Developing touch- and proximity-input-based interfaces for zoo-housed giraffes’ audio enrichment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ISS): 434:1–434:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3626470.

Gellersen:2022:PVE

[GKKW22] Hans Gellersen, Enkelejda Kasneci, Krzysztof Krejtz, and Daniel Weiskopf. PACMHCI V6, ETRA, May 2022 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):137:1–137:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530878.

Garg:2021:ULE

[GKS21] Radhika Garg, Yash Kapadia, and Subhasree Sengupta. Using the lenses of emotion and support to understand unemployment discourse on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):14:1–14:24, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449088.

Ghai:2021:EAL

[GLZ^{+}21] Bhavya Ghai, Q. Vera Liao, Yunfeng Zhang, Rachel Bellamy, and Klaus Mueller. Explainable active learning (XAL): Toward AI explanations as interfaces for machine teachers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):235:1–235:28, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432934.

Gross:2023:ESM

[GM23] Tom Gross and Tony Malzhacker. The experience sampling method and its tools: a review for developers, study administrators, and participants. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 182:1–182:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593234.

Glinka:2023:CRH

[GMB23] Katrin Glinka and Claudia Müller-Birn. Criticalreflective human-AI collaboration: Exploring computational tools for art historical image retrieval. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):263:1–263:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610054.

Griggio:2019:CEB

[GMM19] Carla F. Griggio, Joanna McGrenere, and Wendy E. Mackay. Customizations and expression breakdowns in ecosystems of communication apps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 26:1–26:26, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359128.

Gerling:2021:PVC

[GMM21] Kathrin Gerling, Elisa Mekler, and Regan L. Mandryk. PACMHCI V5, CHI PLAY, September 2021 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):224:1– 224:2, September 2021.

CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474651.

Gron:2020:PPV

[GN20] Kirsikka Grön and Matti Nelimarkka. Party politics, values and the design of social media services: Implications of political elites’ values and ideologies to mitigating of political polarisation through design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):104:1–104:29, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415175.

Gorp:2023:YEG

[GN23] Pieter Van Gorp and Raoul Nuijten. 8-year evaluation of Game-Bus: Status quo in aiming for an open access platform to prototype and test digital health apps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 171:1–171:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593223.

Grandhi:2022:DSD

[GP22] Sukeshini A. Grandhi and Linda Plotnick. Do I spit or do I pass?: Perceived privacy and security concerns of direct-to-consumer genetic testing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):19:1–19:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492838.

Gauthier:2023:AAC

[GPC^{+}23] Robert P. Gauthier, Catherine Pelletier, Laurie-Ann Carrier, Maude Dionne, Ève Dubé, Samantha Meyer, and James R. Wallace. Agency and amplification: a comparison of manual and computational thematic analyses by public health researchers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):2:1–2:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567552.

Grandhi:2019:DSD

[GPH19] Sukeshini A. Grandhi, Linda Plotnick, and Starr Roxanne Hiltz. Do I stay or do I go?: Motivations and decision making in social media non-use and reversion. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):235:1–235:27, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361116.

Grandhi:2021:CCP

[GPH21] Sukeshini Grandhi, Linda Plotnick, and Starr Roxanne Hiltz. By the crowd and for the crowd: Perceived utility and willingness to contribute to trustworthiness indicators on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):218:1–218:24, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463930.

Guettaf:2020:EPC

[GRG20] Adnane Guettaf, Yosra Rekik, and Laurent Grisoni. Effect of physical challenging activity on tactile texture recognition for mobile surface. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS):

190:1–190:12, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427318.

Gong:2021:SDM

[GSC21] Crystal Gong, Koustuv Saha, and Stevie Chancellor. “The Smartest Decision for My Future”: Social media reveals challenges and stress during post-college life transition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):298:1–298:29, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476039.

Greis:2017:ICE

[GSK^{+}17] Miriam Greis, Hendrik Schuff, Marius Kleiner, Niels Henze, and Albrecht Schmidt. Input controls for entering uncertain data: Probability distribution sliders. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):3:1–3:17, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095805.

Galeote:2024:CCC

[GSLH24] Daniel Fernández Galeote, Nevena Sicevic, Nikoletta-Zampeta Legaki, and Juho Hamari. Changing climate change mental models through gamebased learning: a controlled experiment involving cognitive mapping. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY): 297:1–297:??, 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677062.

Gui:2022:CAE

[GTC22] Fanlu Gui, Chun-Hua Tsai, and John M. Carroll. Community acknowledgment: Engaging community members in volunteer acknowledgment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):20:1–20:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3492839.

Gomi:2023:URT

[GTO^{+}23] Ryota Gomi, Kazuki Takashima, Yuki Onishi, Kazuyuki Fujita, and Yoshifumi Kitamura.

UbiSurface: a robotic touch surface for supporting mid-air planar interactions in roomscale VR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):443:1–443:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626479.

Grolaux:2020:STS

[GVNK20] Donatien Grolaux, Jean Vanderdonckt, Thanh-Diane Nguyen, and Iyad Khaddam. SketchADoodle: Touch-surface multistroke gesture handling by Bézier curves. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):87:1–87:30, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397875.

Gauthier:2022:CTA

[GW22] Robert P. Gauthier and James R. Wallace. The computational thematic analysis toolkit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):25:1–25:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3492844.

Grundling:2024:IAI

[GW24] Jan P. Gründling and Benjamin Weyers. Immersive analytics: The influence of flow, sense of agency, and presence on performance and satisfaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 254:1–254:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3661144.

Garcia:2017:MDA

[GYD17] Adriana Alvarado Garcia, Alyson L. Young, and Lynn Dombrowski. On making data actionable: How activists use imperfect data to foster social change for human rights violations in Mexico. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(CSCW): 19:1–19:19, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3134654.

Hao:2021:CGS

[Hao21]  Yu Hao.  Computer

games as social sculptures: Rethinking the discourse of participation and its implications for digital game design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):241:1– 241:15, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474668.

Heinz:2021:DTA

[HBJR21] Mario Heinz, Sebastian Büttner, Sascha Jenderny, and Carsten Röcker. Dynamic task allocation based on individual abilities — experiences from developing and operating an inclusive assembly line for workers with and without disabilities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 206:1–206:19, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461728.

Holz:2024:DGB

[HBLN24] Heiko Holz, Benedikt Beuttler, Denise Löfflad, and Manuel Ninaus. Developing a group-based literacy screening for German pre-readers: a digital, game-based approach. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 251:1–251:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676496.

Haug:2023:ACP

[HBM23] Saskia Haug, Ivo Benke, and Alexander Maedche. Aligning crowdworker perspectives and feedback outcomes in crowdfeedback system design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):23:1–23:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579456.

Herdel:2024:CEA

[HC24] Viviane Herdel and Jessica R. Cauchard. Crafting for emotion appropriateness in affective robotics: Examining the practicality of the OCC model. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 248:1–248:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676493.

Herskovitz:2022:XAR

[HCG^{+}22] Jaylin Herskovitz, Yi Fei Cheng, Anhong Guo, Alanson P. Sample, and Michael Nebeling. XSpace: an augmented reality toolkit for enabling spatially-aware distributed collaboration. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ISS): 568:1–568:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567721.

Han:2023:BCM

[HCM^{+}23] Violet Yinuo Han, Hyunsung Cho, Kiyosu Maeda, Alexandra Ion, and David Lindlbauer. BlendMR: a computational method to create ambient mixed reality interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):436:1–436:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626472.

Haimerl:2022:ECE

[HCR22] Mathias Haimerl, Mark Colley, and Andreas Riener. Evaluation of common external communication concepts of automated vehicles for people with intellectual disabilities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 182:1–182:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546717.

Hirskyj-Douglas:2021:FDI

[HDPL21] Ilyena Hirskyj-Douglas, Roosa Piitulainen, and Andrés Lucero. Forming the dog Internet: Prototyping a dog-to-human video call device. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 494:1–494:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488539.

Helms:2019:DCR

[HFBL19] Karey Helms, Pedro Ferreira, Barry Brown, and Airi Lampinen. Away and (dis)connection: Reconsidering the use of digital technologies in light of long-term outdoor activities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):230:1–230:20, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361111.

HoltenMoller:2019:ACC

[HFL19] Naja L. Holten Møller, Geraldine Fitzpatrick, and Christopher A. Le Dantec. Assembling the case: Citizens’ strategies for exercising authority and personal autonomy in social welfare. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):244:1–244:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361125.

Hesenius:2019:GHG

[HG19] Marc Hesenius and Volker Gruhn. GestureCards: a hybrid gesture notation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):22:1–22:35, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331164.

Huber:2020:PFS

[HGG20] Stephan Huber, Johanna Gramlich, and Tobias Grundgeiger. From paper flight strips to digital strip systems: Changes and similarities in air traffic control work practices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):028:1–028:21, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392833.

Haider:2022:MDV

[HHJ^{+}22] Aqeel Haider, Casper Harteveld, Daniel Johnson, Max V. Birk, Regan L. Mandryk, Magy Seif El-Nasr, Lennart E. Nacke, Kathrin Gerling, and Vero Vanden Abeele. miniPXI: Development and validation of an eleven-item measure of the player experience inventory. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):244:1– 244:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549507.

Horanyi:2023:GDG

[HHLC23] Nora Horanyi, Yuqi Hou, Ales Leonardis, and Hyung Jin Chang. G-DAIC: a gaze initialized framework for description and aestheticbased image cropping. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):163:1–163:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591132.

Hu:2022:DMP

[HHVB22] Xinlan Emily Hu, Rebecca Hinds, Melissa Valentine, and Michael S. Bernstein. A “Distance Matters” paradox: Facilitating intra-team collaboration can harm interteam collaboration. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):48:1–48:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512895.

Higashi:2020:ESI

[Hig20] Takafumi Higashi. Evaluation of skill improvement by combining skill and difficulty levels during paper-cutting production. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS):

191:1–191:17, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427319.

Houtti:2022:WNW

[HJC^{+}22] Mo Houtti, Isaac Johnson, Joel Cepeda, Soumya Khandelwal, Aviral Bhatnagar, and Loren Terveen. “We Need a Woman in Music”: Exploring Wikipedia’s values on article priority. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):266:1–266:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555156.

Hindalong:2022:AVP

[HJCM22] Emily Hindalong, Jordon Johnson, Giuseppe Carenini, and Tamara Munzner. Abstractions for visualizing preferences in group decisions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):49:1–49:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512896.

Hussein:2020:MMV

[HJM20] Eslam Hussein, Prerna Juneja, and Tanushree Mitra. Measuring misinformation in video search platforms: an audit study on YouTube. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):048:1–048:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392854.

Hamid:2024:YET

[HK24a] Aleesha Hamid and Per Ola Kristensson. 40 years of eye typing: Challenges, gaps, and emergent strategies. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):222:1–222:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655596.

Hougaard:2024:APS

[HK24b] Bastian Ilsø Hougaard and Hendrik Knoche. Aiming, pointing, steering: a core task analysis framework for gameplay. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):292:1– 292:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677057.

Hanley:2023:GAC

[HKD23] Hans W. A. Hanley, Deepak Kumar, and Zakir Durumeric. A golden age: Conspiracy theories’ relationship with misinformation outlets, news media, and the wider Internet. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):252:1–252:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610043.

Hwang:2019:WDH

[HKF19] Sun Young Hwang, Negar Khojasteh, and Susan R. Fussell. When delayed in a hurry: Interpretations of response delays in time-sensitive instant messaging. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):234:1–234:20, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361115.

Hwang:2022:BTL

[HKL^{+}22] Angel Hsing-Chi Hwang, Jaryung Kim, Shane Neil Lobo, Yingyi Shu, and Andrea Stevenson Won. Being there to learn: Narrative style and crossplatform comparison for 360-degree educational videos. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW2): 278:1–278:??, November 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3555169.

Haldar:2020:PAS

[HKM^{+}20] Shefali Haldar, Yoojung Kim, Sonali R. Mishra, Andrea L. Hartzler, Ari H. Pollack, and Wanda Pratt. The patient advice system: a technology probe study to enable peer support in the hospital. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):112:1–112:23, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415183.

Haapoja:2021:PSS

[HLV21] Jesse Haapoja, Airi Lampinen, and Kari Mikko Vesala. Personalised services in social situations: Principal-agent relationships in account sharing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):219:1–219:21, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432918.

Hanrahan:2021:AEE

[HMB^{+}21] Benjamin V. Hanrahan, Carleen Maitland, Timothy Brown, Anita Chen, Fraterne Kagame, and Beatrice Birir. Agency and extraction in emerging industrial drone applications: Imaginaries of Rwandan farm workers and community members. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):233:1–233:21, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432932.

Heinze:2024:ERA

[HMK24] Rafael Alves Heinze, Regan L. Mandryk, and Madison Klarkowski. Exploring the role of action mechanics in gamebased stress recovery. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):305:1– 305:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677070.

Howell-Munson:2022:BIV

[HMML^{+}22] Alicia Howell-Munson, Christopher Micek, Ziheng Li, Michael Clements, Andrew C. Nolan, Jackson Powell, Erin T. Solovey, and Rodica Neamtu. BrainEx: Interactive visual exploration and discovery of sequence similarity in brain signals. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 162:1–162:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3534516.

Hou:2024:GAE

[HNS^{+}24] Baosheng James Hou, Joshua Newn, Ludwig Sidenmark, Anam Ahmad Khan, and Hans Gellersen. GazeSwitch: Automatic eye-head mode switching for optimised hands-free pointing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):227:1–227:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655601.

Hamidi:2022:KCT

[HOHB22] Foad Hamidi, Patrick Mbullo Owuor, Michaela Hynie, and Melanie Baljko. “Knowledge Comes Through Participation”: Understanding disability through the lens of DIY assistive technology in Western Kenya. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):72:1–72:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512919.

Hoppe:2021:OHH

[HOSK21] Matthias Hoppe, Daria Oskina, Albrecht Schmidt, and Thomas Kosch. Odin’s helmet: a headworn haptic feedback device to simulate Gforces on the human body in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 212:1–212:15, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461734.

Herman:2021:MTQ

[HOZ^{+}21] Bridger Herman, Maxwell Omdal, Stephanie Zeller, Clara A. Richter, Francesca Samsel, Greg Abram, and Daniel F. Keefe. Multi-touch querying on data physicalizations in immersive AR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (ISS):497:1–497:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488542.

Haliburton:2023:VHP

[HPH^{+}23] Luke Haliburton, Benedikt Pirker, Paolo Holinski, Albrecht Schmidt, Pawel W. Wozniak, and Matthias Hoppe. VR-Hiking: Physical exertion benefits mindfulness and positive emotions in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):216:1–216:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604263.

Huber:2022:AFD

[HPL22] Linda Huber, Casey Pierce, and Silvia Lindtner. An approximation of freedom: On-demand therapy and the feminization of labor. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):275:1–275:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555166.

Habib:2022:IUN

[HPY^{+}22] Hana Habib, Sarah Pearman, Ellie Young, Ishika Saxena, Robert Zhang, and Lorrie FaIth Cranor. Identifying user needs for advertising controls on Facebook. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):59:1–59:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512906.

Hougaard:2021:WWI

[HRC^{+}21] Bastian Ilsø Hougaard, Ingeborg Goll Rossau, Jedrzej Jacek Czapla, Mózes Adorján Mikó, Rasmus Bugge Skammelsen, Hendrik Knoche, and Mads Jochumsen. Who willed it?: Decreasing frustration by manipulating perceived control through fabricated input for stroke rehabilitation BCI games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):235:1– 235:19, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474662.

Holly:2023:AML

[HRP23] Michael Holly, Sebastian Resch, and Johanna Pirker. An asymmetric multiplayer learning environment for roomscale virtual reality and a handheld device. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):197:1–197:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604244.

Hua:2022:CAM

[HRR^{+}22] Yiqing Hua, Manoel Horta Ribeiro, Thomas Ristenpart, Robert West, and Mor Naaman. Characterizing alternative monetization strategies on YouTube. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):283:1–283:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555174.

Hossain:2020:VVG

[HRRS20] Muhammad M. Hossain, Banani Roy, Chanchal K. Roy, and Kevin A. Schneider. VizSciFlow: a visually guided scripting framework for supporting complex scientific data analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):74:1–74:37, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3394976.

Harisinghani:2023:CAU

[HSC^{+}23] Anuj Harisinghani, Harshinee Sriram, Cristina Conati, Giuseppe Carenini, Thalia Field, Hyeju Jang, and Gabriel Murray. Classification of Alzheimer’s using deeplearning methods on webcam-based gaze data. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):157:1–157:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3591126.

Huber:2019:AAB

[HSG19] Bernd Huber, Stuart Shieber, and Krzysztof Z. Gajos. Automatically analyzing brainstorming language behavior with meeter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 30:1–30:17, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359132.

Hutt:2022:FLE

[HSG^{+}22] Stephen Hutt, Angela E. B. Stewart, Julie Gregg, Stephen Mattingly, and Sidney K. D’Mello. Feasibility of longitudinal eyegaze tracking in the workplace. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):148:1–148:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530889.

Halbhuber:2022:DBM

[HSH22] David Halbhuber, Valentin Schwind, and Niels Henze. Don’t break my flow: Effects of switching latency in shooting video games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):229:1– 229:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549492.

Haesler:2021:STH

[HSVR21] Steffen Haesler, Stefka Schmid, Annemike Sophia Vierneisel, and Christian Reuter. Stronger together: How neighborhood groups build up a virtual network during the COVID-19 pandemic. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):304:1–304:31, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476045.

Hadan:2024:MMU

[HSZKN24] Hilda Hadan, Sabrina Alicia Sgandurra, Leah Zhang-Kennedy, and Lennart E. Nacke. From motivating to manipulative: The use of deceptive design in a game’s free-to-play transition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):309:1– 309:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677074.

Hung:2022:OSV

[HTS^{+}22] Ching-Wen Hung, Hsin-Ruey Tsai, Chi-Chun Su, Jui-Cheng Chiu, and Bing-Yu Chen. Osci-Head: Simulating versatile force feedback on an HMD by rendering various types of oscillation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):180:1–180:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546715.

Hardy:2024:TAC

[HTS24] Jean Hardy and Jacob Thebault-Spieker. A turn to assets in community-based computing research: Tradeoffs, deficits, and neoliberalism in technological development. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):14:1–14:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637291.

Hantsbarger:2022:ASR

[HTTH22] Matthew Hantsbarger, Giovanni Maria Troiano, Alexandra To, and Casper Harteveld. Alienated serendipity and reflective failure: Exploring queer game mechanics and queerness in games via queer temporality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):221:1– 221:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549484.

Huang:2022:FGA

[HV22] Xiaoyun Huang and Jessica Vitak. “Finsta gets all my bad pictures”: Instagram users’ selfpresentation across finsta and rinsta accounts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):69:1–69:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512916.

Hedditch:2023:DJP

[HV23] Sonali Hedditch and Dhaval Vyas. Design justice in practice: Community-led design of an online maker space for refugee and migrant women. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):4:1–4:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567554.

Hettiachchi:2020:CCS

[HvBKG20] Danula Hettiachchi, Niels van Berkel, Vassilis Kostakos, and Jorge Goncalves. Crowd-Cog: a cognitive skill based system for heterogeneous task assignment and recommendation in crowdsourcing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):110:1–110:22, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415181.

Han:2024:CDL

[HWF^{+}24] Yisi Han, Zhendong Wang, Yang Feng, Zhihong Zhao, and Yi Wang. Characterizing developers’ linguistic behaviors in open source development across their social statuses. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):29:1–29:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637306.

Haider:2024:ITR

[HWGV24] Aqeel Haider, Günter Wallner, Kathrin Gerling, and Vero Vanden Abeele. An investigation of the test-retest reliability of the miniPXI. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):293:1– 293:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677058.

Homaeian:2021:JAS

[HWS21] Leila Homaeian, James R. Wallace, and Stacey D. Scott. Joint action storyboards: a framework for visualizing communication grounding costs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):28:1–28:27, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449102.

Hwang:2021:HSC

[HWYW21] Angel Hsing-Chi Hwang, Cheng Yao Wang, Yao-Yuan Yang, and Andrea Stevenson Won. Hide and seek: Choices of virtual backgrounds in video chats and their effects on perception. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):303:1–303:30, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476044.

Hossain:2022:NAD

[HZ22] Md. Yousuf Hossain and Loutfouz Zaman. NCAlt: Alternatives and difference visualizations for behavior trees in game development learning. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):245:1– 245:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549508.

Islam:2024:FOS

[IB24] Rahul Islam and Sang Won Bae. FacePsy: an opensource affective mobile sensing system — analyzing facial behavior and head gesture for depression detection in naturalistic settings. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 260:1–260:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3676505.

Iacovides:2022:CTC

[ICB^{+}22] Ioanna Iacovides, Joe Cutting, Jen Beeston, Marta E. Cecchinato, Elisa D. Mekler, and Paul Cairns. Close but not too close: Distance and relevance in designing games for reflection. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):224:1– 224:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549487.

Ikematsu:2020:IHS

[IOES20] Kaori Ikematsu, Haruna Oshima, Rachel Eardley, and Itiro Siio. Investigating how smartphone movement is affected by lying down body posture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 192:1–192:17, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427320.

Ibtasam:2019:MCB

[IRA^{+}19] Samia Ibtasam, Lubna Razaq, Maryam Ayub, Jennifer R. Webster, Syed Ishtiaque Ahmed, and Richard Anderson. “My cousin bought the phone for me. I never go to mobile shops.”: The role of family in women’s technological inclusion in Islamic culture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 46:1–46:33, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359148.

Ippoliti:2023:PGA

[ITKB23] Hatice Sahin Ippoliti, Nina Trilck, Marion Koelle, and Susanne Boll. Please, go ahead! Fostering prosocial driving with sympathyeliciting automated vehicle external displays. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):218:1–218:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604265.

Isomoto:2023:EDT

[IYS23] Toshiya Isomoto, Shota Yamanaka, and Buntarou Shizuki. Exploring dwell-time from human cognitive processes for dwell selection. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):159:1–159:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3591128.

Jones:2021:LFM

[JA21] Jasmine Jones and Mark S. Ackerman. Learning from family mysteries: Accounting for untold stories in family memory practices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):221:1–221:22, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432920.

Jensen:2019:SWO

[JAB19] Jonas Kjeldmand Jensen, Tawfiq Ammari, and Pernille Bjørn. Into Scandinavia: When online fatherhood reflects societal infrastructures. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):231:1–231:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361112.

Ju:2019:AMS

[JCD19] Qinjie Ju, René Chalon, and Stéphane Derrode. Assisted music score reading using fixed-gaze head movement: Empirical experiment and design implications. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):3:1–3:29, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300962.

Jiao:2024:MWE

[JDX24] Aochen Jiao, Di Duan, and Weitao Xu. Medusa3D: The watchful eye freezing illegitimate users in virtual reality interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):270:1–270:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676515.

Jin:2024:ENC

[JF24] Xiaofu Jin and Mingming Fan. EarMonitor: Non-clinical assessment of ear health conditions using a low-cost endoscope camera on smartphones. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 254:1–254:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676499.

Jain:2023:CDU

[JFL23] Pranut Jain, Rosta Farzan, and Adam J. Lee. Co-designing with users the explanations for a proactive auto-response messaging agent. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI): 201:1–201:??, September 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3604248.

Jiang:2023:TNA

[JFS^{+}23] Peiling Jiang, Li Feng, Fuling Sun, Parakrant Sarkar, Haijun Xia, and Can Liu. 1D-Touch: NLP-assisted coarse text selection via a semidirect gesture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):447:1–447:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626483.

Jung:2023:EMD

[JJJ^{+}23] Yugyeong Jung, Gyuwon Jung, Sooyeon Jeong, Chaewon Kim, Woontack Woo, Hwajung Hong, and Uichin Lee. “Enjoy, but Moderately!”: Designing a social companion robot for social engagement and behavior moderation in solitary drinking context. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):237:1–237:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610028.

Jin:2024:CSL

[JK24] Hyoungwook Jin and Juho Kim. CodeTree: a system for learnersourcing subgoal hierarchies in code examples. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):31:1–31:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637308.

Jo:2023:CST

[JKC23] Jeongwon Jo, Tiffany Knearem, and John M. Carroll. Coproducing support together: Sustainable and reciprocal civic disaster relief during COVID-19. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):41:1–41:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3579474.

Jasim:2021:CCR

[JKW^{+}21] Mahmood Jasim, Pooya Khaloo, Somin Wadhwa, Amy X. Zhang, Ali Sarvghad, and Narges Mahyar. CommunityClick: Capturing and reporting community feedback from town halls to improve inclusivity. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):213:1–213:32, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432912.

Jagannath:2022:SBM

[JKZ^{+}22] Swathi Jagannath, Neha Kamireddi, Katherine Ann Zellner, Randall S. Burd, Ivan Marsic, and Aleksandra Sarcevic. A speech-based model for tracking the progression of activities in extreme action teamwork. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):73:1–73:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512920.

Jiang:2022:GEA

[JNLtB22] Mengqi Jiang, Vijayakumar Nanjappan, Hai-Ning Liang, and Martijn ten Bhömer. Ges-Fabri: Exploring affordances and experience of textile interfaces for gesture-based interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 168:1–168:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3534522.

Jackson:2020:SFE

[JØC^{+}20] Corey Brian Jackson, Carsten Østerlund, Kevin Crowston, Mahboobeh Harandi, and Laura Trouille. Shifting forms of engagement: Volunteer learning in online citizen science. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):036:1–036:19, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392841.

Johnson:2024:FFE

[JOFM24] Daniel Johnson, Nicholas O’Donnell, Julian Frommel, and Regan L. Mandryk. Flipping into the future: Exploring player experience and task demand in physical, VR, and PC pinball. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):306:1– 306:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677071.

Janssen:2022:IUH

[JP22] Marc Janßen and Michael Prilla. Investigating the use of head mounted devices for remote cooperation and guidance during the treatment of wounds. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):3:1–3:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3492822.

Johanson:2024:HPH

[JPKM24] Colby Johanson, Susanne Poeller, Madison Klarkowski, and Regan L. Mandryk. Hexed by pressure: How actionstate orientation explains propensity to choke in super hexagon. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):312:1– 312:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677077.

Jiang:2023:CCE

[JRP^{+}23] Zhiqiu Jiang, Mashrur Rashik, Kunjal Panchal, Mahmood Jasim, Ali Sarvghad, Pari Riahi, Erica DeWitt, Fey Thurber, and Narges Mahyar. CommunityBots: Creating and evaluating a multi-agent chatbot platform for public input elicitation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):36:1–36:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579469.

Jagannath:2022:BJR

[JS22] Krithika Jagannath and Katie Salen. Beyond just rules: Server rules for shaping positive experiences in an online play community for youth. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):222:1– 222:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549485.

Jung:2022:EAC

[JSJ22] Soon-Gyo Jung, Joni Salminen, and Bernard J. Jansen. Engineers, aware! Commercial tools disagree on social media sentiment: Analyzing the sentiment bias of four major tools. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 153:1–153:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3532203.

Jung:2022:HDE

[JSKA22] Ju Yeon Jung, Tom Steinberger, John L. King, and Mark S. Ackerman. How domain experts work with data: Situating data science in the practices and settings of craftwork. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):58:1–58:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512905.

Jun:2021:EES

[JSP^{+}21] Joonyoung Jun, Woosuk Seo, Jihyeon Park, Subin Park, and Hyunggu Jung. Exploring the experiences of streamers with visual impairments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):297:1–297:23, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476038.

Jiang:2022:MAF

[JTB22] Liuyue Jiang, Nguyen Khoi Tran, and Muhammad Ali Babar. Mod2Dash: a framework for modeldriven dashboards generation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 172:1–172:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3534526.

Jambon:2022:UCF

[JV22] Francis Jambon and Jean Vanderdonckt. UsyBus: a communication framework among reusable agents integrating eyetracking in interactive applications. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 157:1–157:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3532207.

Jaquero:2019:NGE

[JVP^{+}19] Víctor Manuel López Jaquero, Radu-Daniel Vatavu, Jose Ignacio Panach, Oscar Pastor, and Jean Vanderdonckt. A Newcomer’s guide to EICS, the engineering interactive computing systems community. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):1:1–1:9, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3300960.

Jang:2022:AEA

[JWK22] Junho Jang, Ji Young Woo, and Huy Kang Kim. Action2Score: an embedding approach to score player action. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):220:1– 220:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549483.

Jahanbakhsh:2021:ELI

[JZB^{+}21] Farnaz Jahanbakhsh, Amy X. Zhang, Adam J. Berinsky, Gordon Pennycook, David G. Rand, and David R. Karger. Exploring lightweight interventions at posting time to reduce the sharing of misinformation on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):18:1–18:42, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449092.

Ku:2023:CUH

[KA23] Melody S. Y. Ku and Mark S. Ackerman. Calibrated uncertainty: How in-vitro fertilization patients use information to regulate emotion. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):31:1–31:??, April 2023. CODEN ????

ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3579464.

Kusk:2022:WWE

[KB22] Kalle Kusk and Claus Bossen. Working with wolt: an ethnographic study of lenient algorithmic management on a food delivery platform. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):4:1–4:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3492823.

Kotturi:2021:UCC

[KBFK21] Yasmine Kotturi, Allie Blaising, Sarah E. Fox, and Chinmay Kulkarni. The unique challenges for creative small businesses seeking feedback on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):15:1–15:27, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449089.

Kothari:2022:EGT

[KBK^{+}22] Rakshit S. Kothari, Reynold J. Bailey, Christopher Kanan, Jeff B.

Pelz, and Gabriel J. Diaz. EllSeg-Gen, towards domain generalization for head-mounted eyetracking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):139:1–139:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530880.

Kapuscinska:2021:SEA

[KBKH21] Adela Kapuscinska, Payal M.

Bhujwala, Melissa Kalarchian, and Jessica Hammer. A socio-ecological approach to activity games for girls. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):246:1– 246:28, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474673.

Kulyk:2022:YBC

[KBSP^{+}22] Oksana Kulyk, Lauren Britton-Steele, Elda Paja, Melanie Duckert, and Louise Barkhuus. ‘You have been in close contact with a person infected with COVID-19 and you may have been infected’: Understanding privacy concerns, trust and adoption in mobile COVID-19 tracing across four countries. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 204:1–204:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546739.

Kar:2020:GEU

[KCC20] Pragma Kar, Samiran Chattopadhyay, and Sandip Chakraborty. Gestatten: Estimation of user’s attention in mobile MOOCs from eye gaze and gaze gesture tracking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):72:1–72:32, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394974.

Karizat:2021:AFT

[KDEA21] Nadia Karizat, Dan Delmonaco, Motahhare Eslami, and Nazanin Andalibi. Algorithmic folk theories and identity: How TikTok users coproduce knowledge of identity and engage in algorithmic resistance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):305:1–305:44, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476046.

Kirshenbaum:2021:TTT

[KDH^{+}21] Nurit Kirshenbaum, Kylie Davidson, Jesse Harden, Chris North, Dylan Kobayashi, Ryan Theriot, Roderick S. Tabalba, Michael L. Rogers, Mahdi Belcaid, Andrew T. Burks, Krishna N. Bharadwaj, Luc Renambot, Andrew E. Johnson, Lance Long, and Jason Leigh. Traces of time through space: Advantages of creating complex canvases in collaborative meetings. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 502:1–502:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488552.

Kaltenbrunner:2018:TPA

[KE18] Martin Kaltenbrunner and Florian Echtler. The TUIO 2.0 protocol: an abstraction framework for tangible interactive surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):8:1–8:35, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3229090.

Keegan:2019:DPP

[Kee19] Brian C. Keegan. The dynamics of peer-produced political information during the 2016 U.S. Presidential campaign. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):33:1–33:20, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359135.

Klassen:2023:SSP

[KF23] Shamika Klassen and Casey Fiesler. The stoop: Speculation on positive futures of black digital spaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):17:1–17:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567567.

Karahalios:2017:ENC

[KFMH17] Karrie Karahalios, Geraldine Fitzpatrick, and Andrés Monroy-Hernández. Editor’s note/Chairs’ welcome. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (CSCW):16:1, December 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3134651.

Kosch:2018:ICA

[KFSC18] Thomas Kosch, Markus Funk, Albrecht Schmidt, and Lewis L. Chuang. Identifying cognitive assistance with mobile electroencephalography: a case study with in-situ projections for manual assembly. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):11:1–11:20, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229093.

Kou:2020:MCA

[KG20] Yubo Kou and Xinning Gui. Mediating community-AI interaction through situated explanation: The case of AI-led moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):102:1–102:27, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415173.

Kross:2021:OFB

[KG21] Sean Kross and Philip Guo. Orienting, framing, bridging, magic, and counseling: How data scientists navigate the outer loop of client collaborations in industry and academia. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):311:1–311:28, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476052.

Kleinman:2022:TPE

[KHY^{+}22] Erica Kleinman, Reza Habibi, Yichen Yao, Christian Gayle, and Magy Seif El-Nasr. “A Time and Phase for Everything” — towards a self regulated learning perspective on computational support for esports. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY): 218:1–218:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3549481.

Kato:2020:CPI

[KIK20] Kunihiro Kato, Kaori Ikematsu, and Yoshihiro Kawahara. CA-Path: 3D-printed interfaces with conductive points in grid layout to extend capacitive touch inputs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 193:1–193:17, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427321.

Kiene:2019:TFU

[KJH19] Charles Kiene, Jialun Aaron Jiang, and Benjamin Mako Hill. Technological frames and user innovation: Exploring technological change in community moderation teams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 44:1–44:23, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359146.

Knearem:2022:MCB

[KJTC22] Tiffany Knearem, Jeongwon Jo, Chun-Hua Tsai, and John M. Carroll. Making community beliefs and capacities visible through care-mongering during COVID-19. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):28:1–28:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492847.

Karolus:2021:EDC

[KKE^{+}21] Jakob Karolus, Francisco Kiss, Caroline Eckerth, Nicolas Viot, Felix Bachmann, Albrecht Schmidt, and Pawel W. Wozniak. EMBody: a datacentric toolkit for EMG-Based interface prototyping and experimentation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 195:1–195:29, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457142.

Khadpe:2022:EPC

[KKK22a] Pranav Khadpe, Chinmay Kulkarni, and Geoff Kaufman. Empathosphere: Promoting constructive communication in ad-hoc virtual teams through perspective-taking spaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):55:1–55:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512902.

Kim:2022:DMC

[KKK22b] Jennifer G. Kim, Robert E.

Kraut, and Karrie Karahalios. Designing a medical crowdfunding Website from sense of community theory. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):290:1–290:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555181.

Kaniwa:2024:CCI

[KKK^{+}24] Yuka Kaniwa, Masaki Kuribayashi, Seita Kayukawa, Daisuke Sato, Hironobu Takagi, Chieko Asakawa, and Shigeo Morishima. ChitChatGuide: Conversational interaction using large language models for assisting people with visual impairments to explore a shopping mall. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):247:1–247:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3676492.

Klassen:2022:BLG

[KKM^{+}22] Shamika Klassen, Sara Kingsley, Kalyn McCall, Joy Weinberg, and Casey Fiesler. Black lives, green books, and blue checks: Comparing the content of the Negro Motorist Green Book to the content on black Twitter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):27:1–27:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3492846.

Krasovskaya:2024:ASF

[KKM24] Sofia Krasovskaya, Árni Kristjánsson, and Joseph Macinnes. Assessing the size of the functional field of view in a gaze-contingent search paradigm. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):223:1–223:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655597.

Kim:2022:SSF

[KKP^{+}22] Inyeop Kim, Minsam Ko, Joonyoung Park, Sung Wook Moon, Gyuwon Jung, Youn kyung Lim, and Uichin Lee. Social-spiritual face: Designing social reading support for spiritual well-being. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):262:1–262:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555162.

Konkol:2019:CIS

[KKS19] Markus Konkol, Christian Kray, and Jan Suleiman. Creating interactive scientific publications using bindings. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):16:1–16:18, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331158.

Kuribayashi:2022:CWM

[KKV^{+}22] Masaki Kuribayashi, Seita Kayukawa, Jayakorn Vongkulbhisal, Chieko Asakawa, Daisuke Sato, Hironobu Takagi, and Shigeo Morishima. Corridor-Walker: Mobile indoor walking assistance for blind people to avoid obstacles and recognize intersections. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 179:1–179:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546714.

Kim:2020:PVC

[KLS20] Juho Kim, Siân Lindley, and Sarita Schoenebeck. PACMHCI V4 CSCW2 Oct 2020 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):90:1, October 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3415161.

Kim:2021:PVC

[KLS21] Juho Kim, Sian Lindley, and Sarita Schoenebeck. PACMHCI V4 CSCW3 December 2020 continued editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):208:1, January 2021. CODEN ????

ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432907.

Kotut:2022:TDC

[KM22] Lindah Kotut and D. Scott McCrickard. The TL;DR charter: Speculatively demystifying privacy policy documents and terms agreements. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):23:1–23:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492842.

Kim:2023:WFP

[KMH23] Daehwa Kim, Vimal Mollyn, and Chris Harrison. WorldPoint: Finger pointing as a rapid and natural trigger for in-the-wild mobile interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):442:1–442:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626478.

Karahalios:2018:EM

[KMHLF18] Karrie Karahalios, Andrés Monroy-Hernández, Airi Lampinen, and Geraldine Fitzpatrick. Editors’ message. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (CSCW):16:1, November 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3274285.

Kniestedt:2021:DDE

[KML^{+}21] Isabelle Kniestedt, Marcello A. Gómez Maureira, Iulia Lefter, Stephan Lukosch, and Frances M. Brazier. Dive deeper: Empirical analysis of game mechanics and perceived value in serious games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):236:1– 236:25, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474663.

Kao:2021:ETB

[KMM21] Dominic Kao, Alejandra J. Magana, and Christos Mousas. Evaluating tutorial-based instructions for controllers in virtual reality games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):234:1– 234:28, September 2021.

CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474661.

Kosch:2022:NAT

[KMM^{+}22] Thomas Kosch, Andrii Matviienko, Florian Müller, Jessica Bersch, Christopher Katins, Dominik Schön, and Max Mühlhäuser. NotiBike: Assessing target selection techniques for cyclist notifications in augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 197:1–197:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546732.

Kundinger:2020:TRG

[KMR20] Thomas Kundinger, Celena Mayr, and Andreas Riener. Towards a reliable ground truth for drowsiness: a complexity analysis on the example of driver fatigue. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):78:1–78:18, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394980.

Kockwelp:2024:VBC

[KMV^{+}24] Pascal Kockwelp, Marcel Meyerheim, Dimitar Valkov, Marvin Mergen, Anna Junga, Antonio Krüger, Bernhard Marschall, Markus Holling, and Benjamin Risse. VR-based competence training at scale: Teaching clinical skills in the context of virtual brain death examination. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 261:1–261:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3664635.

Koshikawa:2022:MBG

[KNT22] Koki Koshikawa, Takashi Nagamatsu, and Kentaro Takemura. Modelbased gaze estimation with transparent markers on large screens. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):147:1–147:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3530888.

Kim:2024:EMF

[KOL^{+}24] Hyolim Kim, Seolim Oh, Gaeun Lee, Sumin Jung, Kyuha Jung, and Hajin Lim. Exploring multi-faceted motivations and strategies using mobile dating applications: Case study of tinder users in South Korea. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):8:1–8:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3633073.

Koivunen:2019:UDM

[KOOL19] Sami Koivunen, Thomas Olsson, Ekaterina Olshannikova, and Aki Lindberg. Understanding decision-making in recruitment: Opportunities and challenges for information technology. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):242:1–242:22, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361123.

Knittel:2019:MTC

[KPW19] Megan Knittel, Shelby Pitts, and Rick Wash. “The Most Trustworthy Coin”: How ideological tensions drive trust in Bitcoin. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 36:1–36:23, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359138.

Kennel:2024:RTP

[KR24] Kathrin Kennel and Stefan Ruzika. Real-time prediction of students’ math difficulties using raw data from eye tracking and neural networks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):233:1–233:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3655607.

Karaosmanoglu:2021:LLH

[KRK^{+}21] Sukran Karaosmanoglu, Sebastian Rings, Lucie Kruse, Christian Stein, and Frank Steinicke. Lessons learned from a human-centered design of an immersive exergame for people with dementia. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):252:1– 252:27, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474679.

Kao:2021:ESS

[KRMM21] Dominic Kao, Rabindra Ratan, Christos Mousas, and Alejandra J. Magana. The effects of a self-similar avatar voice in educational games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):238:1– 238:28, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474665.

Kuzuoka:2022:PVC

[KRW^{+}22] Hideaki Kuzuoka, Katharina Reinecke, Hao-Chuan Wang, Naomi Yamashita, Shaowen Bardzell, Siân Lindley, and Aleksandra Sarcevic. PACMHCI V6, CSCW2, November 2022 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):261:1–261:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555227.

Keppel:2023:DFD

[KSF^{+}23] Jonas Keppel, Marvin Strauss, Sarah Faltaous, Jonathan Liebers, Roman Heger, Uwe Gruenefeld, and Stefan Schneegass. Don’t forget to disinfect: Understanding technologysupported hand disinfection stations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI): 204:1–204:??, September 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3604251.

Kwon:2024:PIR

[KSG^{+}24] Nahyun Kwon, Tong Steven Sun, Yuyang Gao, Liang Zhao, Xu Wang, Jeeeun Kim, and Sungsoo Ray Hong. 3DPFIX: Improving remote novices’ 3D printing troubleshooting through human-AI collaboration design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):11:1–11:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637288.

Khatra:2024:ABM

[KSKH24] Karanmeet Khatra, Jaisie Sin, Anastasia Kuzminykh, and Khalad Hasan. Agentbased mediation on smartphone usage among co-located couples. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):242:1–242:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676487.

Krejtz:2023:ULC

[KSP^{+}23] Krzysztof Krejtz, Patryk Szczecinski, Aneta Pawlowska, Daria Rutkowska-Siuda, Katarzyna Wisiecka, Piotr Milczarski, Artur Hlobaz, Andrew T. Duchowski, and Izabela Krejtz. A unified look at cultural heritage: Comparison of aggregated scanpaths over architectural artifacts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):169:1–169:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591138.

Karimi:2022:ADD

[KSW22] Younes Karimi, Anna Squicciarini, and Shomir Wilson. Automated detection of doxing on Twitter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW2): 276:1–276:??, November 2022. CODEN ????

ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3555167.

Kandappu:2021:PTP

[KSX21] Thivya Kandappu, Vigneshwaran Subbaraju, and Qianli Xu. PrivacyPrimer: Towards privacy-preserving episodic memory support for older adults. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):306:1–306:32, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476047.

Koch:2020:IIC

[KTBL^{+}20] Janin Koch, Nicolas Taffin, Michel Beaudouin-Lafon, Markku Laine, Andrés Lucero, and Wendy E. Mackay. ImageSense: an intelligent collaborative ideation tool to support diverse human-computer partnerships. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):045:1–045:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392850.

Kudo:2021:TBV

[KTF^{+}21] Yoshiki Kudo, Anthony Tang, Kazuyuki Fujita, Isamu Endo, Kazuki Takashima, and Yoshifumi Kitamura. Towards balancing VR immersion and bystander awareness. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 484:1–484:22, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486950.

Karolus:2022:IFP

[KTP^{+}22] Jakob Karolus, Simon Thanheiser, David Peterson, Nicolas Viot, Thomas Kosch, Albrecht Schmidt, and Paweł W. Wozniak. Imprecise but fun: Playful interaction using electromyography. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):190:1–190:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546725.

Kowalewski:2023:WLA

[KUD^{+}23] Marvin Kowalewski, Christine Utz, Martin Degeling, Theodor Schnitzler, Franziska Herbert, Leonie Schaewitz, Florian M. Farke, Steffen Becker, and Markus Dürmuth. 52 weeks later: Attitudes towards COVID-19 apps for different purposes over time. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):251:1–251:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610042.

Kassem:2022:WHB

[KUWM22] Khaled Kassem, Tobias Ungerböck, Philipp Wintersberger, and Florian Michahelles. What is happening behind the wall?: Towards a better understanding of a hidden robot’s intent by multimodal cues. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):196:1–196:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546731.

Koelle:2024:PVM

[KvBW24] Marion Koelle, Niels van Berkel, and Jenny Waycott. PACMHCI, VI, MHCI, September 2024 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 241:1–241:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676486.

Khaddam:2020:TRP

[KVDG20] Iyad Khaddam, Jean Vanderdonckt, Salah Dowaji, and Donatien Grolaux. Towards rapid prototyping of foldable graphical user interfaces with Flecto. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 194:1–194:33, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427322.

Kuosmanen:2022:HDS

[KVK^{+}22] Elina Kuosmanen, Aku Visuri, Saba Kheirinejad, Niels van Berkel, Heli Koskimäki, Denzil Ferreira, and Simo Hosio. How does sleep tracking influence your life?: Experiences from a longitudinal field study with a wearable ring. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):185:1–185:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3546720.

Ketoma:2023:TFA

[KVM23] Vix Kemanji Ketoma, Jean Vanderdonckt, and Gerrit Meixner. Towards flexible authoring and personalization of virtual reality applications for training. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 189:1–189:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593241.

Khanshan:2024:ECG

[KVM24] Alireza Khanshan, Pieter Van Gorp, and Panos Markopoulos. Evaluation of code generation for simulating participant behavior in experience sampling method by iterative in-context learning of a large language model. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 255:1–255:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3661143.

Kruger:2021:WTG

[KWdCL^{+}21] Max Krüger, Anne Weibert, Debora de Castro Leal, Dave Randall, and Volker Wulf. “What is the Topic of the Group, Please?” on migration, care and the challenges of participation in design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):309:1–309:29, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476050.

Krings:2024:TTE

[KY24] Sarah Claudia Krings and Enes Yigitbas. TARPS: a toolbox for enhancing privacy and security for collaborative AR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 251:1–251:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3660251.

Kumaran:2021:PER

[KYB21] Sneha R. Krishna Kumaran, Yue Yin, and Brian P. Bailey. Plan early, revise more: Effects of goal setting and perceived role of the feedback provider on feedback seeking behavior. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):24:1–24:22, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449098.

Kou:2022:CCM

[KZZW22] Ziyi Kou, Yang Zhang, Daniel Zhang, and Dong Wang. CrowdGraph: a crowdsourcing multimodal knowledge graph approach to explainable fauxtography detection. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):287:1–287:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555178.

Leclercq:2022:EEU

[LAD^{+}22] Tony Leclercq, Ebrahim Khalil Abbasi, Bruno Dumas, Marie-Ange Remiche, and Patrick Heymans. Essential expectations of users of Web configurators: an empirical survey. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 165:1–165:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3534519.

Li:2019:SSM

[LAHW19] Zhen Li, Michelle Annett, Ken Hinckley, and Daniel Wigdor. SMAC: a simplified model of attention and capture in multi-device desk-centric environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):2:1–2:47, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300961.

Liaqat:2022:HSW

[LAM22] Amna Liaqat, Benett Axtell, and Cosmin Munteanu. “With a hint she will remember”: Collaborative storytelling and culture sharing between immigrant grandparents and grandchildren via magic thing designs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):268:1–268:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555158.

Lee:2021:BTC

[LAR21] Hae-Na Lee, Vikas Ashok, and I. V. Ramakrishnan. Bringing things closer:

Enhancing low-vision interaction experience with Office productivity applications. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 197:1–197:18, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457144.

Lu:2019:VEI

[LAW19] Zhicong Lu, Michelle Annett, and Daniel Wigdor. Vicariously experiencing it all without going outside: a study of outdoor livestreaming in China. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 25:1–25:28, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359127.

Long:2021:CDA

[LBM21] Duri Long, Takeria Blunt, and Brian Magerko. Co-designing AI literacy exhibits for informal learning spaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):293:1–293:35, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476034.

Lyu:2024:BSS

[LC24] Yao Lyu and John M.

Carroll. “Because Some Sighted People, They Don’t Know What the Heck You’re Talking About:” a study of blind tokers’ infrastructuring work to build independence. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):20:1–20:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637297.

Langerak:2024:MMA

[LCA^{+}24] Thomas Langerak, Sammy Christen, Mert Albaba, Christoph Gebhardt, Christian Holz, and Otmar Hilliges. MAR-LUI: Multi-agent reinforcement learning for adaptive point-and-click UIs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 253:1–253:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3661147.

Loewen:2024:IIU

[LCG24] Georgia Loewen, Karen Anne Cochrane, and Audrey Girouard. From imagination to innovation: Using participatory design fiction to envision the future of accessible gaming wearables for players with upper limb motor disabilities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):308:1– 308:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677073.

Liu:2022:ASB

[LDM^{+}22] Can Liu, Chenyue Dai, Qingzhou Ma, Brinda Mehra, and Alvaro Cassinelli. AngleCAD: Surface-based 3D modelling techniques on foldable touchscreens. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):582:1–582:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567735.

Logre:2018:MSV

[LDP18] Ivan Logre and Anne-Marie Déry-Pinna. MDE in support of visualization systems design: a multi-staged approach tailored for multiple roles. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):14:1–14:17, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229096.

Liu:2021:ACD

[LE21] Jiali Liu and James Eagan. ADQDA: a crossdevice affinity diagramming tool for fluid and holistic qualitative data analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 489:1–489:19, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488534.

Li:2024:BJM

[LFK24] Lingyuan Li, Guo Freeman, and Bart Knijnenburg. Beyond just money transactions: How digital P2P payments (re)shape existing offline interpersonal relationships. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):24:1–24:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637301.

Li:2022:CEU

[LFM22] Lingyuan Li, Guo Freeman, and Nathan J. Mc-Neese. Channeling enduser creativity: Leveraging live streaming for distributed collaboration in indie game development. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):282:1–282:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555173.

Lankes:2022:GEE

[LG22] Michael Lankes and Argenis Ramirez Gomez. GazeCues: Exploring the effects of gaze-based visual cues in virtual reality exploration games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):237:1– 237:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549500.

Lunding:2023:RBP

[LGG^{+}23] Mille Skovhus Lunding, Jens Emil Sloth Grønbæk, Nicolai Grymer, Thomas Wells, Steven Houben, and Marianne Graves Petersen. Reality and beyond:

Proxemics as a lens for designing handheld collaborative augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ISS): 427:1–427:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3626463.

Lampinen:2019:EM

[LGS19] Airi Lampinen, Darren Gergle, and David A. Shamma. Editors’ message. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):23:1, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359125.

Lobo:2020:FPE

[LHI20] María Jesús Lobo, Christophe Hurter, and Pourang Irani. Flex-ER: a platform to evaluate interaction techniques for immersive visualizations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 195:1–195:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427323.

Luo:2020:EAD

[LHPH20] Mufan Luo, Tiffany W. Hsu, Joon Sung Park, and Jeffrey T. Hancock. Emotional amplification during livestreaming: Evidence from comments during and after news events. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):047:1–047:19, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392853.

Li:2020:SPA

[LHT^{+}20] Mingyang Li, Louis Hickman, Louis Tay, Lyle Ungar, and Sharath Chandra Guntuku. Studying politeness across cultures using English Twitter and Mandarin Weibo. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):119:1–119:15, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415190.

Lomas:2021:DSC

[LKG21] James Derek Lomas, Mihovil Karac, and Mathieu Gielen. Design space cards: Using a card deck to navigate the design space of interactive play. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):227:1– 227:21, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474654.

Li:2021:AAC

[LKS21] Tiffany Wenting Li, Karrie Karahalios, and Hari Sundaram. “It’s all about conversation”: Challenges and concerns of faculty and students in the arts, humanities, and the social sciences about education at scale. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):216:1–216:37, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432915.

Lee:2024:RRT

[LKYK24] Jaeyoon Lee, Hanseob Kim, Yechan Yang, and Gerard Jounghyun Kim. RPG: Rotation technique in VR locomotion using peripheral gaze. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):235:1–235:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3655609.

Lee:2024:DND

[LL24] Lik-Hang Lee and Zijun Lin. Danger, nuisance, disregard: Analyzing user-generated videos for augmented reality gameplay on handheld devices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):298:1– 298:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677063.

Lowy:2023:BCA

[LLAK23] Rachel Lowy, Chung Eun Lee, Gregory D. Abowd, and Jennifer G. Kim. Building causal agency in autistic students through iterative reflection in collaborative transition planning. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):246:1–246:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3610037.

Lai:2024:ERM

[LLC^{+}24] Yu-Ju Lai, Yi-Chieh Lee, Chia-Chi Chang, Wan-Ting Dai, and Ying-Yu Chen. Exploring the role of mom’s chat groups in the messaging app: Enhancing support and empowerment for stayat-home mothers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):3:1–3:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3633068.

Li:2021:HDT

[LLDH21] Tianshi Li, Elizabeth Louie, Laura Dabbish, and Jason I. Hong. How developers talk about personal data and what it means for user privacy: a case study of a developer forum on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):220:1–220:28, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432919.

LeDoux:2024:COE

[LLGH24] Ridley Jones LeDoux, Charlotte P. Lee, Sucheta Ghoshal, and Mark Haselkorn. Concept of operations as epistemic object: The sociotechnical design roles of a systems engineering document. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):34:1–34:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637311.

Luo:2022:NIT

[LLKC22] Yuhan Luo, Bongshin Lee, Young-Ho Kim, and Eun Kyoung Choe. Note-Wordy: Investigating touch and speech input on smartphones for personal data capture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):581:1–581:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567734.

Li:2022:ECP

[LLL^{+}22] Franklin Mingzhe Li, Cheng Lu, Zhicong Lu, Patrick Carrington, and Khai N. Truong. An exploration of captioning practices and challenges of individual content creators on YouTube for people with hearing impairments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):75:1–75:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512922.

Lee:2023:EII

[LLW^{+}23] Kyungjun Lee, Hong Li, Muhammad Rizky Wellyanto, Yu Jiang Tham, Andrés Monroy-Hernández, Fannie Liu, Brian A. Smith, and Rajan Vaish. Exploring immersive interpersonal communication via AR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):50:1–50:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579483.

Liaqat:2020:LPS

[LM20] Amna Liaqat and Cosmin Munteanu. Leveraging peer support for mature immigrants learning to write in informal contexts. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):106:1–106:24, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415177.

Long:2022:APE

[LMB^{+}22] Duri Long, Tom McKlin, Nylah Akua Adjei Boone, Dezarae Dean, Mirina Garoufalidis, and Brian Magerko. Active Prolonged Engagement EXpanded (APEX): a toolkit for supporting evidence-based iterative design decisions for collaborative, embodied museum exhibits. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):50:1–50:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512897.

Leiva:2023:PVM

[LMS23] Luis A. Leiva, Maristella Matera, and Johannes Schöning. PACMHCI V7, MHCI, September 2023 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):191:1–191:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604238.

Labrie:2022:TVC

[LMT^{+}22] Audrey Labrie, Terrance Mok, Anthony Tang, Michelle Lui, Lora Oehlberg, and Lev Poretski. Toward videoconferencing tools for hands-on activities in online teaching. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):10:1–10:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492829.

Liu:2024:BEH

[LMZ^{+}24] Shiyi Liu, Ruofei Ma, Chuyi Zhao, Zhenbang Li, Jianpeng Xiao, and Quan Li. BPCoach: Exploring hero drafting in professional MOBA tournaments via visual analytics. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):26:1–26:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637303.

Loria:2021:SCG

[LNM21] Enrica Loria, Lennart E.

Nacke, and Annapaola Marconi. On social contagion in gamification: The power of influencers in a locationbased gameful system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):243:1– 243:20, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474670.

Leite:2017:MPA

[LO17] Luis Leite and Veronica Orvalho. Mani-pull-action: Hand-based digital puppetry. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):2:1–2:16, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095804.

Lee:2020:LRU

[LOSD20] Seyeon Lee, Hyunyoung Oh, Chung-Kon Shi, and Young Yim Doh. Life review using a life metaphoric game to promote intergenerational communication. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):98:1–98:21, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415169.

Liu:2022:EDL

[LPED22] Jiazhou Liu, Arnaud Prouzeau, Barrett Ens, and Tim Dwyer. Effects of display layout on spatial memory for immersive environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):576:1–576:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567729.

Li:2021:CCC

[LPFD21] Renee Li, Pavitthra Pandurangan, Hana Frluckaj, and Laura Dabbish. Code of conduct conversations in open source software projects on Github. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):19:1–19:31, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449093.

Lystbaek:2022:EGA

[LPGG22] Mathias N. Lystbæk, Ken Pfeuffer, Jens Emil Sloth Grønbæk, and Hans Gellersen. Exploring gaze for assisting freehand selection-based text entry in AR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):141:1–141:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530882.

Luyten:2022:EIC

[LPQW22] Kris Luyten, Philippe Palanque, Aaron John Quigley, and Marco Winckler. Engineering interactive computing systems 2022: Editorial introduction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 149:1–149:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3532089.

Li:2023:SPC

[LPW^{+}23] Jingyi Li, Hyerim Park, Robin Welsch, Sven Mayer, and Andreas Butz. SeatmateVR: Proxemic cues for close bystander-awareness in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):438:1–438:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626474.

Lenhart:2023:YSN

[LPZV23] Anna Lenhart, Sunyup Park, Michael Zimmer, and Jessica Vitak. “You Shouldn’t Need to Share Your Data”: Perceived privacy risks and mitigation strategies among privacy-conscious smart home power users. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):247:1–247:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610038.

Laato:2024:GWN

[LRK^{+}24a] Samuli Laato, Sampsa Rauti, Bastian Kordyaka, Konstantinos Papangelis, Sangwon Jung, Timo Nummenmaa, and Juho Hamari. Gamification of walking in nature: a field experiment with Pokémon GO routes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):310:1– 310:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677075.

Lindberg:2024:DGB

[LRK^{+}24b] Sharon Lindberg, Chiara Rossitto, Ola Knutsson, Petter Karlström, and Sirkku Männikkö Barbutiu. Doing good business? Design leaders’ perspectives on ethics in design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):2:1–2:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3633067.

Lystbaek:2022:GHA

[LRP^{+}22] Mathias N. Lystbæk, Peter Rosenberg, Ken Pfeuffer, Jens Emil Grønbæk, and Hans Gellersen. Gaze-hand alignment: Combining eye gaze and mid-air pointing for interacting with menus in augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):145:1–145:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3530886.

Lustig:2020:SMY

[LRSS20] Caitlin Lustig, Sean Rintel, Liane Scult, and Siddharth Suri. Stuck in the middle with you: The transaction costs of corporate employees hiring freelancers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):037:1–037:28, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392842.

Luyten:2023:PEI

[LS23] Kris Luyten and Carmen Santoro. PACMHCI — engineering interactive computing systems, June 2023: Editorial introduction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 170:1–170:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593222.

Logas:2022:IDT

[LSLD22] Jacob Logas, Ari Schlesinger, Zhouyu Li, and Sauvik Das. Image DePO: Towards gradual decentralization of online social networks using decentralized privacy overlays. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):60:1–60:??, April 2022. CODEN ????

ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512907.

Li:2023:UEI

[LTP23] Lin Li, Xinru Tang, and Anne Marie Piper. Understanding extrafamilial intergenerational communication: a case analysis of an age-integrated online community. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):261:1–261:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3610052.

Li:2020:SYS

[LUFW20] Lingyuan Li, Jirassaya Uttarapong, Guo Freeman, and Donghee Yvette Wohn. Spontaneous, yet studious: Esports commentators’ live performance and self-presentation practices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):103:1–103:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415174.

Liao:2019:GIC

[LW19] Jingxian Liao and Hao-Chuan Wang. Gestures as intrinsic creativity support: Understanding the usage and function of hand gestures in computermediated group brainstorming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):243:1–243:16, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361124.

Li:2022:DMM

[LWB22] Jingyi Li, Luca Woik, and Andreas Butz. Designing mobile MR workspaces: Effects of reality degree and spatial configuration during passenger productivity in HMDs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):181:1–181:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546716.

Lei:2023:DER

[LWC^{+}23] Yaxiong Lei, Yuheng Wang, Tyler Caslin, Alexander Wisowaty, Xu Zhu, Mohamed Khamis, and Juan Ye. DynamicRead: Exploring robust gaze interaction methods for reading on handheld mobile devices under dynamic conditions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):158:1–158:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3591127.

Lee:2024:IUP

[LWC^{+}24] Yu-Jen Lee, Meng-Hsin Wu, Chung-Chiao Chang, Xi-Jing Chang, and Yung-Ju Chang. Investigating user-perceived impacts of contextual factors on opportune moments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 269:1–269:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3676514.

Liu:2024:HDU

[LWS24] Songsong Liu, Shu Wang, and Kun Sun. Having difficulty understanding manuals? Automatically converting user manuals into instructional videos. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 243:1–243:??, June 2024.

CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3660245.

Liu:2023:AWM

[LWTW23] Jingjing Liu, Xun Wang, Peter Tolmie, and Volker Wulf. Articulation work and the management of intersubjectivity disjunctures in offshored production. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):25:1–25:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579458.

Lu:2020:LRC

[LX20] Alex Jiahong Lu and Xuecong Xu. “Learning for the Rise of China”: Exploring uses and gratifications of state-owned online platform. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):030:1–030:25, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392835.

Li:2023:INN

[LYD^{+}23] Xiaoyan Li, Naomi Yamashita, Wen Duan, Yoshinari Shirai, and Susan R. Fussell. Improving non-native speakers’ participation with an automatic agent in multilingual groups. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):12:1–12:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567562.

Lee:2020:DCM

[LYH20] Yi-Chieh Lee, Naomi Yamashita, and Yun Huang. Designing a chatbot as a mediator for promoting deep selfdisclosure to a real mental health professional. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):031:1–031:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392836.

Lei:2023:CSL

[LYL^{+}23] Kehua Lei, Mingrui Yu, Marissa Lewellen, Venus Ku, and David T. Lee. Compass: Supporting large group mentorship in a chat-based UI. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):37:1–37:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579470.

Logas:2021:TBA

[LZAD21] Jacob Logas, Ruican Zhong, Stephanie Almeida, and Sauvik Das. Tensions between access and control in Makerspaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):215:1–215:33, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432914.

Laine:2021:RPW

[LZS^{+}21] Markku Laine, Yu Zhang, Simo Santala, Jussi P. P. Jokinen, and Antti Oulasvirta. Responsive and personalized Web layouts with integer programming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 213:1–213:23, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461735.

Li:2023:RRS

[LZW23] Zhenjiang Li, Xinglin Zhang, and Chenshu Wu.

RingVKB: a ring-shaped virtual keyboard using low-cost IMU. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):220:1–220:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604267.

Lee:2021:PPC

[LZY^{+}21] Lik Hang Lee, Yiming Zhu, Yui-Pan Yau, Pan Hui, and Susanna Pirttikangas. Press-n-Paste: Copy-and-paste operations with pressuresensitive caret navigation for miniaturized surface in mobile augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 199:1–199:29, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457146.

Munyaka:2023:DMS

[MAD^{+}23] Imani Munyaka, Zahra Ashktorab, Casey Dugan, J. Johnson, and Qian Pan. Decision making strategies and team efficacy in human-AI teams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):43:1–43:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579476.

Micallef:2021:FGI

[MAMP21] Nicholas Micallef, Mihai Avram, Filippo Menczer, and Sameer Patil. Fakey: a game intervention to improve news literacy on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):6:1–6:27, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449080.

Marques:2020:IVC

[MAN^{+}20] Bernardo Marques, João Alves, Miguel Neves, Inês Justo, André Santos, Raquel Rainho, Rafael Maio, Dany Costa, Carlos Ferreira, Paulo Dias, and Beatriz Sousa Santos. Interaction with virtual content using augmented reality: a user study in assembly procedures. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 196:1–196:17, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427324.

Mehmood:2019:TDC

[MAR^{+}19] Hamid Mehmood, Tallal Ahmad, Lubna Razaq, Shrirang Mare, Maryem Zafar Usmani, Richard Anderson, and Agha Ali Raza. Towards digitization of collaborative savings among low-income groups. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 35:1–35:30, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3274304.

Mastrianni:2022:ACM

[MAS22] Angela Mastrianni, Lynn Almengor, and Aleksandra Sarcevic. Alerts as coordination mechanisms: Implications for designing alerts for multidisciplinary and shared decision making. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):9:1–9:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3492828.

Morina:2022:WSA

[MB22] Durim Morina and Michael S.

Bernstein. A Web-scale analysis of the community origins of image memes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):74:1–74:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512921.

McCord:2023:BTD

[MB23] Curtis W. McCord and Christoph Becker. Beyond transactional democracy: a study of civic tech in Canada. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):29:1–29:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579462.

Mittmann:2022:LSA

[MBK^{+}22] Gloria Mittmann, Adam Barnard, Ina Krammer, Diogo Martins, and João Dias. LINA — a social augmented reality game around mental health, supporting realworld connection and sense of belonging for early adolescents. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):242:1– 242:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549505.

Muller:2022:UMM

[MBS22] Sebastian Müller, Matthias Baldauf, and Arne Seeliger. Ubiquitous machinery monitoring — a field study on manufacturing workers’ user experience of mobile and wearable monitoring apps. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):198:1–198:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3546733.

Mandel:2020:UCP

[MBT^{+}20] Travis Mandel, Jahnu Best, Randall H. Tanaka, Hiram Temple, Chansen Haili, Sebastian J. Carter, Kayla Schlechtinger, and Roy Szeto. Using the crowd to prevent harmful AI behavior. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):97:1–97:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415168.

Menendez-Blanco:2024:PVG

[MBTW24] Maria Menendez-Blanco, Matthieu Tixier, and Donghee Yvette Wohn. PACMHCI V8, GROUP, January 2024 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):1:1–1:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3633066.

Magnaudet:2018:DSC

[MCC^{+}18]  Mathieu  Magnaudet,

Stéphane Chatty, Stéphane Conversy, Sébastien Leriche, Celia Picard, and Daniel Prun. Djnn/Smala: a conceptual framework and a language for interaction-oriented programming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):12:1–12:27, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229094.

Muller:2022:ACE

[MCDR22] Tobias Müller, Mark Colley, Gülsemin Dogru, and Enrico Rukzio. AR4CAD: Creation and exploration of a taxonomy of augmented reality visualization for connected automated driving. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 177:1–177:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546712.

McDonald:2020:ITA

[MCI20] Sharon McDonald, Gilbert Cockton, and Alastair Irons. The impact of thinking-aloud on usability inspection. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):88:1–88:22, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397876.

Mackamul:2023:EVS

[MCM23] Eva Mackamul, Géry Casiez, and Sylvain Malacria. Exploring visual signifier characteristics to improve the perception of affordances of in-place touch inputs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):210:1–210:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604257.

Ma:2019:EAE

[MCYZ19] Haiwei Ma, Hao-Fei Cheng, Bowen Yu, and Haiyi Zhu. Effects of anonymity, ephemerality, and system routing on cost in social question asking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):238:1–238:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361119.

March:2020:WET

[MD20] Laura March and Sayamindu Dasgupta. Wikipedia edit-a-thons as sites of public pedagogy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):100:1–100:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415171.

Mandryk:2024:PVC

[MDFG24] Regan L. Mandryk, Alena Denisova, Julian Frommel, and Kathrin Gerling. PACMHCI V8, CHI PLAY, October 2024 editorial. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):288:1– 288:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677053.

Mazurkiewicz:2023:BAR

[MdLGG23] Bartosz Mazurkiewicz, Marcelo de Lima Galvão, and Ioannis Giannopoulos. BeeAR: Augmented reality beeline navigation for spatial knowledge acquisition. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI): 199:1–199:??, September 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3604246.

Manshaei:2019:TBM

[MDM^{+}19] Roozbeh Manshaei, Sean DeLong, Uzair Mayat, Dhrumil Patal, Matthew Kyan, and Ali Mazalek. Tangible BioNets: Multisurface and tangible interactions for exploring structural features of biological networks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):14:1–14:22, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331156.

Musick:2024:SSU

[MDN^{+}24] Geoff Musick, Wen Duan, Shabnam Najafian, Subhasree Sengupta, Christopher Flathmann, Bart Knijnenburg, and Nathan McNeese. To share or not to share: Understanding and modeling individual disclosure preferences in recommender systems for the workplace. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):9:1–9:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3633074.

Maia:2023:URE

[MdOKT23] Vitor Carneiro Maia, Kathia Marçal de Oliveira, Christophe Kolski, and Guilherme Horta Travassos. Using RFID in the engineering of interactive software systems: a systematic mapping. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 183:1–183:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593235.

Meng:2019:CDW

[MDZ19] Amanda Meng, Carl DiSalvo, and Ellen Zegura. Collaborative data work towards a caring democracy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 42:1–42:23, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359144.

Madaio:2022:AFA

[MES^{+}22] Michael Madaio, Lisa Egede, Hariharan Subramonyam, Jennifer Wortman Vaughan, and Hanna Wallach. Assessing the fairness of AI systems: AI practitioners’ processes, challenges, and needs for support. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):52:1–52:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512899.

Massachi:2020:SSA

[MFM^{+}20] Talie Massachi, Grant Fong, Varun Mathur, Sachin R. Pendse, Gabriela Hoefer, Jessica J. Fu, Chong Wang, Nikita Ramoji, Nicole R. Nugent, Megan L. Ranney, Daniel P. Dickstein, Michael F. Armey, Ellie Pavlick, and Jeff Huang. Sochiatrist: Signals of affect in messaging data. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):111:1–111:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415182.

Musick:2021:GFT

[MFM21] Geoff Musick, Guo Freeman, and Nathan J. Mc-Neese. Gaming as family time: Digital game coplay in modern parentchild relationships. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):251:1– 251:25, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474678.

Meyer:2022:UHC

[MFSK22] Johannes Meyer, Adrian Frank, Thomas Schlebusch, and Enkelejda Kasneci. U-HAR: a convolutional approach to human activity recognition combining head and eye movements for context-aware smart glasses. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA): 143:1–143:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3530884.

Metaxa:2021:ISG

[MGG^{+}21] Danaë Metaxa, Michelle A. Gan, Su Goh, Jeff Hancock, and James A. Landay. An image of society: Gender and racial representation and impact in image search results for occupations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):26:1–26:23, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449100.

Matviienko:2022:ASC

[MGRM22] Andrii Matviienko, Sebastian Günther, Sebastian Ritzenhofen, and Max Mühlhäuser. AR sightseeing: Comparing information placements at outdoor historical heritage sites using augmented reality.

Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):194:1–194:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546729.

Ma:2019:PTR

[MH19] Ning F. Ma and Benjamin V. Hanrahan. Part-time ride-sharing: Recognizing the context in which drivers rideshare and its impact on platform use. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):247:1–247:17, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361128.

Miller:2021:HDP

[MHG^{+}21] Josh Aaron Miller, Britton Horn, Matthew Guthrie, Jonathan Romano, Guy Geva, Celia David, Amy Robinson Sterling, and Seth Cooper. How do players and developers of citizen science games conceptualize skill chains? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):244:1– 244:29, September 2021.

CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3474671.

Mieczkowski:2021:AMC

[MHN^{+}21] Hannah Mieczkowski, Jeffrey T. Hancock, Mor Naaman, Malte Jung, and Jess Hohenstein. AI-mediated communication: Language use and interpersonal effects in a referential communication task. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):17:1–17:14, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3449091.

Martin-Hammond:2022:BCH

[MHP22] Aqueasha Martin-Hammond and Tanjala S. Purnell. Bridging community, history, and culture in personal informatics tools: Insights from an existing community-based heart health intervention for black Americans. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):29:1–29:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492848.

Mathew:2020:HBH

[MIS^{+}20] Binny Mathew, Anurag Illendula, Punyajoy Saha, Soumya Sarkar, Pawan Goyal, and Animesh Mukherjee. Hate begets hate: a temporal study of hate speech. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):92:1–92:24, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415163.

Ma:2023:DBA

[MK23] Renkai Ma and Yubo Kou. “Defaulting to boilerplate answers, they didn’t engage in a genuine conversation”: Dimensions of transparency design in creator moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):44:1–44:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579477.

Moster:2023:BSS

[MKRM23] Makayla Moster, Ella Kokinda, Paige Rodeghero, and Nathan McNeese.

Both sides of the story: Changing the “Preexisting Culture of Dread” surrounding student teamwork in breakout rooms. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):30:1–30:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579463.

Moran-Ledesma:2021:UDG

[MLSH21] Marco Moran-Ledesma, Oliver Schneider, and Mark Hancock. Userdefined gestures with physical props in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 488:1–488:23, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486954.

McDonald:2019:IMC

[MM19]  Samantha  McDonald

and Melissa Mazmanian. Information materialities of citizen communication in the U.S. Congress. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 47:1–47:20, November 2019. CODEN ????

ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359149.

McDonald:2023:DFC

[MM23] Nora McDonald and Helena M. Mentis. ’Don’t fall for this’: Communications about cybersafety from the AARP. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):248:1–248:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610039.

Machado:2021:CBW

[MMdSP21] Leticia S. Machado, Ricardo Rodrigo M. Melo, Cleidson R. B. de Souza, and Rafael Prikladnicki. Collaborative behavior and winning challenges in competitive software crowdsourcing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):220:1–220:25, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463932.

Mogavi:2021:CSE

[MMH21] Reza Hadi Mogavi, Xiaojuan Ma, and Pan Hui. Characterizing student engagement moods for dropout prediction in question pool Websites. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):12:1–12:22, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449086.

Manshaei:2022:TCT

[MMI^{+}22] Roozbeh Manshaei, Uzair Mayat, Syeda Aniqa Imtiaz, Veronica Andric, Kazeera Aliar, Nour Abu Hantash, Kashaf Masood, Gabby Resch, Alexander Bakogeorge, Sarah Sabatinos, and Ali Mazalek. Tangible chromatin: Tangible and multi-surface interactions for exploring datasets from highcontent microscopy experiments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):558:1–558:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567711.

Morisaki:2021:UDC

[MMM^{+}21] Tao Morisaki, Ryoma Mori, Ryosuke Mori, Kohki Serizawa, Yasutoshi Makino, Yuta Itoh, Yuji Yamakawa, and Hiroyuki Shinoda. Ultrasound-driven curveball in table tennis: Human activity support via noncontact remote object manipulation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 503:1–503:20, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488548.

Matviienko:2022:BYC

[MMMM22] Andrii Matviienko, Damir Mehmedovic, Florian Müller, and Max Mühlhäuser. “Baby, You can Ride my Bike”: Exploring maneuver indications of self-driving bicycles using a tandem simulator. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):188:1–188:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546723.

Moldovan:2020:OUI

[MNP^{+}20] Alex Moldovan, Vlad Nicula, Ionut Pasca, Mihai Popa, Jaya Krishna Namburu, Anamaria Oros, and Paul Brie. OpenUIDL, a user interface description language for runtime omnichannel user interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):86:1–86:52, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397874.

Martinie:2022:EOB

[MNP^{+}22] Célia Martinie, David Navarre, Philippe Palanque, Eric Barboni, and Sandra Steere. Engineering operations-based training. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 164:1–164:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3534518.

Moller:2021:CWT

[MNS^{+}21] Naja Holten Møller, Gina Neff, Jakob Grue Simonsen, Jonas Christoffer Villumsen, and Pernille Bjørn. Can workplace tracking ever empower? Collective sensemaking for the responsible use of sensor data at work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):219:1–219:21, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463931.

Matsuda:2021:SSR

[MON^{+}21] Akira Matsuda, Toru Okuzono, Hiromi Nakamura, Hideaki Kuzuoka, and Jun Rekimoto. A surgical scene replay system for learning gastroenterological endoscopic surgery skill by multiple synchronizedvideo and gaze representation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 204:1–204:22, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461726.

Morgado:2019:ITA

[MP19] Inês Coimbra Morgado and Ana C. R. Paiva. The iMPAcT tool for Android testing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):4:1–4:23, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300963.

Mattioli:2023:MAR

[MP23] Andrea Mattioli and Fabio Paternò. A mobile augmented reality app for creating, controlling, recommending automations in smart homes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):195:1–195:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604242.

Martinie:2019:ADT

[MPB^{+}19] Célia Martinie, Philippe Palanque, Elodie Bouzekri, Andy Cockburn, Alexandre Canny, and Eric Barboni. Analysing and demonstrating toolsupported customizable task notations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):12:1–12:26, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331154.

Mondal:2023:TTC

[MPC^{+}23] Mainack Mondal, Anju Punuru, Tyng-Wen Scott Cheng, Kenneth Vargas, Chaz Gundry, Nathan S. Driggs, Noah Schill, Nathaniel Carlson, Josh Bedwell, Jaden Q. Lorenc, Isha Ghosh, Yao Li, Nancy Fulda, and Xinru Page. A tale of two cultures: Comparing interpersonal information disclosure norms on Twitter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):254:1–254:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610045.

Mustonen:2024:BPH

[MPEC24] Sanni Mustonen, Siiri Paananen, Emma Eckhoff, and Ashley Colley. Balancing playability and historical accuracy in game design — developing the Struve geodetic arc mobile game. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):295:1– 295:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677060.

Maram:2024:TMA

[MPKEN24] Sai Siddartha Maram, Johannes Pfau, Mansi Rajendra Kasar, and Magy Seif El-Nasr. A topic modeling approach towards understanding the discourse between religion and videogames on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):289:1– 289:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677054.

Ma:2023:IRT

[MQX^{+}23] Yifan Ma, Batu Qi, Wenhua Xu, Mingjie Wang, Bowen Du, and Hongfei Fan. Integrating realtime and non-real-time collaborative programming: Workflow, techniques, and prototypes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):13:1–13:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567563.

Mendes:2020:CTB

[MRGN20] Daniel Mendes, Sofia Reis, João Guerreiro, and Hugo Nicolau. Collaborative tabletops for blind people: The effect of auditory design on workspace awareness. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 197:1–197:19, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427325.

Mosconi:2022:DDS

[MRK^{+}22] Gaia Mosconi, Dave Randall, Helena Karasti, Saja Aljuneidi, Tong Yu, Peter Tolmie, and Volkmar Pipek. Designing a data story: a storytelling approach to curation, sharing and data reuse in support of ethnographicallydriven research. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):289:1–289:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555180.

Mostaeen:2019:DRT

[MRRS19] Golam Mostaeen, Banani Roy, Chanchal Roy, and Kevin Schneider. Designing for real-time groupware systems to support complex scientific data analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):9:1–9:28, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3331151.

Muralidhar:2022:CIW

[MRS22] Srihari Hulikal Muralidhar, Sean Rintel, and Siddharth Suri. Collaboration, invisible work, and the costs of macrotask freelancing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):284:1–284:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555175.

Magrofuoco:2022:ARS

[MRV22] Nathan Magrofuoco, Paolo Roselli, and Jean Vanderdonckt. μV: an articulation, rotation, scaling, and translation invariant (ARST) multistroke gesture recognizer. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 150:1–150:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3532200.

Morrison:2023:EIH

[MSHP23] Katelyn Morrison, Donghoon Shin, Kenneth Holstein, and Adam Perer. Evaluating the impact of human explanation strategies on human-AI visual decision-making. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):48:1–48:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579481.

Meyer:2022:HIA

[MSK22] Johannes Meyer, Thomas Schlebusch, and Enkelejda Kasneci. A highly integrated ambient light robust eye-tracking sensor for retinal projection AR glasses based on laser feedback interferometry. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):140:1–140:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530881.

Miceli:2020:BSI

[MSY20] Milagros Miceli, Martin Schuessler, and Tianling Yang. Between subjectivity and imposition: Power dynamics in data annotation for computer vision. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):115:1–115:25, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415186.

Marathe:2021:SRE

[MT21] Megh Marathe and Kentaro Toyama. The situated, relational, and evolving nature of epilepsy diagnosis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):217:1–217:18, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432916.

Maquil:2017:CRI

[MTA^{+}17] Valerie Maquil, Eric Tobias, Dimitra Anastasiou, Hélène Mayer, and Thibaud Latour. COPSE: Rapidly instantiating problem solving activities based on tangible tabletop interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):6:1–6:16, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3095808.

Magrofuoco:2019:GCP

[MV19] Nathan Magrofuoco and Jean Vanderdonckt. Gelicit: a cloud platform for distributed gesture elicitation studies. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):6:1–6:41, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331148.

Maris:2023:TWP

[MWBB23] Elena Maris, Kelly B.

Wagman, Rachel Bergmann, and Danielle Bragg. Tech worker perspectives on considering the interpersonal implications of communication technologies. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):16:1–16:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567566.

MacArthur:2019:MQI

[MWH19] Cayley MacArthur, Caroline Wong, and Mark Hancock. Makers and quilters: Investigating opportunities for improving gender-imbalanced maker groups. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):29:1–29:24, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359131.

Moitra:2022:ADR

[MWK^{+}22] Aparna Moitra, Dennis Wagenaar, Manveer Kalirai, Syed Ishtiaque Ahmed, and Robert Soden. AI and disaster risk: a practitioner perspective. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):272:1–272:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555163.

Mao:2019:HDS

[MWM^{+}19] Yaoli Mao, Dakuo Wang, Michael Muller, Kush R. Varshney, Ioana Baldini, Casey Dugan, and Aleksandra Mojsilović. How data scientists work together with domain experts in scientific collaborations: To find the right answer or to ask the right question? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):237:1–237:23, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361118.

Min:2023:SWI

[MXYS23] Tian Min, Chengshuo Xia, Takumi Yamamoto, and Yuta Sugiura. Seeing the wind: an interactive mist interface for airflow input. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):444:1–444:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626480.

Marathe:2021:TVT

[MYS^{+}21] Megh Marathe, Yoonseon Yi, Chia-Hsuan Su, Ting-Wei Chang, and Gabriela Marcu. Tedious versus taxing: The nature of work in a behavioral health context. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):302:1–302:24, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476043.

Nwagu:2023:EBB

[NAO23] Chukwuemeka Nwagu, Alaa AlSlaity, and Rita Orji. EEG-based braincomputer interactions in immersive virtual and augmented reality: a systematic review. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 174:1–174:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593226.

Neumayr:2022:SAA

[NAS^{+}22] Thomas Neumayr, Mirjam Augstein, Johannes Schönböck, Sean Rintel, Helmut Leeb, and Thomas Teichmeister. Semi-automated analysis of collaborative interaction: Are we there yet? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):571:1–571:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567724.

Nejati:2020:WFG

[NB20] Javad Nejati and Aruna Balasubramanian. WProfX: a fine-grained visualization tool for Web page loads. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):73:1–73:22, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394975.

Nova:2022:CCI

[NCM^{+}22] Fayika Farhat Nova, Amanda Coupe, Elizabeth D. Mynatt, Shion Guha, and Jessica A. Pater. Cultivating the community: Inferring influence within eating disorder networks on Twitter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):7:1–7:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3492826.

Nunes:2024:EPE

[ND24] Caio Nunes and Ticianne Darin. Echoes of player experience: a literature review on audio assessment and player experience in games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):304:1– 304:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677069.

Nagai:2022:HGT

[NFTK22] Takahiro Nagai, Kazuyuki Fujita, Kazuki Takashima, and Yoshifumi Kitamura. HandyGaze: a gaze tracking technique for room-scale environments using a single smartphone. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):562:1–562:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567715.

Nolte:2020:HSN

[NHH20] Alexander Nolte, Linda Bailey Hayden, and James D. Herbsleb. How to support newcomers in scientific hackathons — an action research study on expert mentoring. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):025:1–025:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392830.

Nepal:2024:BCI

[NHL^{+}24] Subigya Nepal, Javier Hernandez, Robert Lewis, Ahad Chaudhry, Brian Houck, Eric Knudsen, Raul Rojas, Ben Tankus, Hemma Prafullchandra, and Mary Czerwinski. Burnout in cybersecurity incident responders: Exploring the factors that light the fire. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):27:1–27:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637304.

Ntouros:2021:DSE

[NKV21] Vasilis Ntouros, Hara Kouki, and Vasilis Vlachokyriakos. Designing sharing economy platforms through a ‘solidarity HCI’ lens. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):23:1–23:25, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449097.

Norambuena:2021:NMA

[NM21] Brian Felipe Keith Norambuena and Tanushree Mitra. Narrative maps: an algorithmic approach to represent and extract information narratives. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):228:1–228:33, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432927.

Ng:2022:EUD

[NM22a] Chloe Ng and Nicolai Marquardt. Eliciting user-defined touch and mid-air gestures for colocated mobile gaming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):569:1–569:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567722.

Nielsen:2022:DLU

[NM22b] Trine Rask Nielsen and Naja Holten Møller. Data as a lens for understanding what constitutes credibility in asylum decision-making. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):6:1–6:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3492825.

Niyazov:2021:DDP

[NMBS21] Aziz Niyazov, Nicolas Mellado, Loic Barthe, and Marcos Serrano. Dynamic decals: Pervasive freeform interfaces using constrained deformable graphical elements. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 493:1–493:27, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488538.

Narkar:2024:GAD

[NMPDJ24] Anish S. Narkar, Jan J.

Michalak, Candace E. Peacock, and Brendan David-John. GazeIntent: Adapting dwelltime selection in VR interaction with real-time intent modeling. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):226:1–226:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655600.

Norton:2019:IGS

[NPT19] Juliet Norton, Birgit Penzenstadler, and Bill Tomlinson. Implications of grassroots sustainable agriculture community values on the design of information systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 34:1–34:22, November 2019. CODEN ????

ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359136.

Nelimarkka:2021:ENV

[NSJ21] Matti Nelimarkka, Antti Salovaara, and Giulio Jacucci. Exploring norm violations and norm management in collocated synchronous communication: Dualchannel communication as hybrid space with spillover and asymmetries. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(GROUP): 221:1–221:24, July 2021. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3463933.

Nicholas:2022:FEM

[NSV22] Molly Jane Nicholas, Brian A. Smith, and Rajan Vaish. Friendscope: Exploring in-the-moment experience sharing on camera glasses via a shared camera. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):56:1–56:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512903.

Ozdel:2024:PPS

[ÖBK24] Süleyman Özdel, Efe Bozkir, and Enkelejda Kasneci. Privacypreserving scanpath comparison for pervasive eye tracking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):231:1–231:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655605.

Ouverson:2021:CFC

[OG21] Kaitlyn M. Ouverson and Stephen B. Gilbert. A composite framework of co-located asymmetric virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):5:1–5:20, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449079.

Ortegat:2022:ETI

[OGRV22] Guillaume Ortegat, Donatien Grolaux, Etienne Riviere, and Jean Vanderdonckt. Engineering the transition of interactive collaborative software from cloud computing to edge computing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 160:1–160:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3532210.

Oomori:2023:IAO

[OKM^{+}23] Kotaro Oomori, Wataru Kawabe, Fabrice Matulic, Takeo Igarashi, and Keita Higuchi. Interactive 3D annotation of objects in moving videos from sparse multiview frames. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):440:1–440:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626476.

Oprea:2020:ESV

[OM20] Silviu Vlad Oprea and Walid Magdy. The effect of sociocultural variables on sarcasm communication online. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):029:1–029:22, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392834.

Osinski:2019:TSK

[OR19] Meike Osinski and Nikol Rummel. Towards successful knowledge integration in online collaboration: an experiment on the role of metaknowledge. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 31:1–31:17, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359133.

Ong:2024:EDO

[OS24] Elyssia Barrie H. Ong and Briane Paul V. Samson. Exploring design opportunities for improved self-motivation in selftracking and health goal achievement. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 256:1–256:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676501.

Ousmer:2020:RTM

[OSM^{+}20] Mehdi Ousmer, Arthur Sluÿters, Nathan Magrofuoco, Paolo Roselli, and Jean Vanderdonckt. Recognizing 3D trajectories as 2D multi-stroke gestures. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 198:1–198:21, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427326.

Oppenlaender:2020:CSW

[OTH20]  Jonas  Oppenlaender,

Thanassis Tiropanis, and Simo Hosio. CrowdUI: Supporting Web design with the crowd. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):76:1–76:28, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394978.

Pandey:2022:DES

[PA22] Laxmi Pandey and Ahmed Sabbir Arif. Design and evaluation of a silent speechbased selection method for eye-gaze pointing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):570:1–570:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567723.

Perovich:2022:SQP

[PAC22] Laura J. Perovich, Meryl Alper, and Corey Cleveland. “Self-Quaranteens” process COVID-19: Understanding information visualization language in memes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):47:1–47:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512894.

Pucihar:2022:SSM

[PAK^{+}22] Klen Čopič Pucihar, Nuwan T. Attygalle, Matjaz Kljun, Christian Sandor, and Luis A. Leiva. Solids on Soli: Millimetre-wave radar sensing through materials. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 156:1–156:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3532212.

Prun:2022:FVG

[PB22] Daniel Prun and Pascal Béger. Formal verification of graphical properties of interactive systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 167:1–167:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3534521.

Panicker:2020:CRC

[PBC20] Aswati Panicker, Kavya Basu, and Chia-Fang Chung. Changing roles and contexts: Symbolic interactionism in the sharing of food and eating practices between remote, intergenerational family members. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):043:1–043:19, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392848.

Pruszko:2021:MHS

[PCL^{+}21] Laura Pruszko, Céline Coutrix, Yann Laurillau, Benoît Piranda, and Julien Bourgeois. Molecular HCI: Structuring the cross-disciplinary space of modular shapechanging user interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 211:1–211:33, June 2021.

CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461733.

Parvin:2018:RTA

[PCMP18] Parvaneh Parvin, Stefano Chessa, Marco Manca, and Fabio Paterno’. Real-time anomaly detection in elderly behavior with the support of task models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):15:1–15:18, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229097.

Perelman:2022:VTA

[PDPS22] Gary Perelman, Emmanuel Dubois, Alice Probst, and Marcos Serrano. Visual transitions around tabletops in mixed reality: Study on a visual acquisition task between vertical virtual displays and horizontal tabletops. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):585:1–585:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567738.

Panchanadikar:2024:ISD

[PF24] Ruchi Panchanadikar and Guo Freeman. “I’m a Solo Developer but AI is My New Ill-Informed Co-Worker”: Envisioning and designing generative AI to support indie game development. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):317:1– 317:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677082.

Pascher:2024:ATX

[PGK^{+}24] Max Pascher, Felix Ferdinand Goldau, Kirill Kronhardt, Udo Frese, and Jens Gerken. AdaptiX — a transitional XR framework for development and evaluation of shared control applications in assistive robotics. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 244:1–244:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3660243.

Polychronis:2023:WTU

[PK23] Cole Polychronis and Marina Kogan. Working together (to undermine democratic institutions): Challenging the social bot paradigm in SSIO research. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):242:1–242:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610033.

Park:2019:GMU

[PKB^{+}19] Chang Min Park, Taeyeon Ki, Ali J. Ben Ali, Nikhil Sunil Pawar, Karthik Dantu, Steven Y. Ko, and Lukasz Ziarek. Gesto: Mapping UI events to gestures and voice commands. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):5:1–5:22, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3300964.

Pena:2021:PSW

[PKC^{+}21] Belén Barros Pena, Bailey Kursar, Rachel E. Clarke, Katie Alpin, Merlyn Holkar, and John Vines. “Pick Someone Who Can Kick Your Ass”

— moneywork in financial third party access. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):218:1–218:28, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3432917.

Park:2022:END

[PKLK22] Solip Park, Annakaisa Kultima, Miikka J. Lehtonen, and Jeanine Krath. Everywhere but nowhere: Development experiences of the international game developers in Finland during the Covid-19 pandemic and remote work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):233:1– 233:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549496.

Poon:2023:CMS

[PLL^{+}23] Anthony Poon, Matthew Luebke, Julia Loughman, Ann Lee, Lourdes Guerrero, Madeline Sterling, and Nicola Dell. Computer-mediated sharing circles for intersectional peer support with home care workers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):39:1–39:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579472.

Pei:2023:TTC

[PLY23] Weiping Pei, Yanina Likhtenshteyn, and Chuan Yue. A tale of two communities: Privacy of third party app users in crowdsourcing — the case of receipt transcription. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):253:1–253:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610044.

Peacock:2022:GII

[PLZ^{+}22] Candace E. Peacock, Ben Lafreniere, Ting Zhang, Stephanie Santosa, Hrvoje Benko, and Tanya R. Jonker. Gaze as an indicator of input recognition errors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):142:1–142:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530883.

Phadke:2021:ESF

[PM21] Shruti Phadke and Tanushree Mitra. Educators, solicitors, flamers, motivators, sympathizers: Characterizing roles in online extremist movements. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):310:1–310:35, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476051.

Paudel:2024:AIR

[PNS24a] Shreyasha Paudel, Wendy Norris, and Robert Soden. Aftermath: Infrastructure, resources, and organizational adaptation in the wake of disaster. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):17:1–17:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637294.

Prakash:2024:AOP

[PNS^{+}24b] Yash Prakash, Akshay Kolgar Nayak, Mohan Sunkara, Sampath Jayarathna, Hae-Na Lee, and Vikas Ashok. All in one place: Ensuring usable access to online shopping items for blind users. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 257:1–257:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3664639.

Paananen:2021:IHS

[POG^{+}21] Ville Paananen, Jonas Oppenlaender, Jorge Goncalves, Danula Hettiachchi, and Simo Hosio. Investigating human scale spatial experience. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 496:1–496:18, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488541.

Pohl:2024:BBA

[Poh24] Henning Pohl. Bodybased augmented reality feedback during conversations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 246:1–246:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676491.

Pater:2023:EIS

[PP23] Jessica Pater and Raquel Oliveira Prates. Editorial introduction — supporting group work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):1:1–1:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3567551.

Pasumarthy:2022:GGT

[PPT^{+}22] Nandini Pasumarthy, Rakesh Patibanda, Yi Ling (Ellie) Tai, Elise van den Hoven, Jessica Danaher, and Rohit Ashok Khot. Gooey gut trail: Board game play to understand human-microbial interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):239:1– 239:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549502.

Prilla:2019:SWW

[Pri19] Michael Prilla. “I simply watched where she was looking at”: Coordination in short-term synchronous cooperative mixed reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):246:1–246:21, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361127.

Phadke:2021:WMP

[PSM21] Shruti Phadke, Mattia Samory, and Tanushree Mitra. What makes people join conspiracy communities?: Role of social factors in conspiracy engagement. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):223:1–223:30, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432922.

Prinster:2024:CAE

[PSTK24] Gale H. Prinster, C. Estelle Smith, Chenhao Tan, and Brian C. Keegan. Community archetypes: an empirical framework for guiding research methodologies to reflect user experiences of sense of virtual community on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):33:1–33:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637310.

Piorkowski:2023:AES

[PVC^{+}23] David Piorkowski, Inge Vejsbjerg, Owen Cornec, Elizabeth M. Daly, and Öznur Alkan. AIMEE: an exploratory study of how rules support AI developers to explain and edit models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):255:1–255:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3610046.

P:2022:CCA

[PVP22] Azhagu Meena S. P., Palashi Vaghela, and Joyojeet Pal. Counting to be counted: Anganwadi workers and digital infrastructures of ambivalent care. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):286:1–286:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555177.

Qiu:2020:ECS

[QGB20] Sihang Qiu, Ujwal Gadiraju, and Alessandro Bozzon. Estimating conversational styles in conversational microtask crowdsourcing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):032:1–032:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392837.

Ren:2021:SST

[RA21a] Yuan Ren and Ahmed Sabbir Arif. Stepper, swipe, tilt, force: Comparative evaluation of four number pickers for smartwatches. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 500:1–500:21, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488545.

Roemmich:2021:DSC

[RA21b] Kat Roemmich and Nazanin Andalibi. Data subjects’ conceptualizations of and attitudes toward automatic emotion recognition-enabled wellbeing interventions on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):308:1–308:34, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476049.

Rifat:2022:SPS

[RAA22] Mohammad Rashidujjaman Rifat, Mohammad Ruhul Amin, and Syed Ishtiaque Ahmed. Situating public speaking: The politics and poetics of the digital Islamic sermons in Bangladesh. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):64:1–64:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512911.

Roth:2020:ENM

[RAZ^{+}20] Rebecca Roth, Jaclyn Abraham, Heidi Zinzow, Pamela Wisniewski, Amro Khasawneh, and Kapil Chalil Madathil. Evaluating news media reports on the ’blue whale challenge’ for adherence to suicide prevention safe messaging guidelines. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):026:1–026:27, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392831.

Richards:2021:SUC

[RCM21] Olivia K. Richards, Adrian Choi, and Gabriela Marcu. Shared understanding in care coordination for children’s behavioral health. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):21:1–21:25, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449095.

Rogoda:2022:FQE

[RDF^{+}22] Kamil Rogoda, Piotr Daniszewski, Kamil Florowski, Rishab Mathur, Kourosh Amouzgar, James Mackenzie, Kim Sauvé, and Abhijit Karnik. Food-Choices(Q): Exploring the design of a serious game proxy for Likert-style survey questionnaires. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):236:1– 236:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549499.

Roels:2018:IUF

[RDS18] Reinout Roels, Arno De Witte, and Beat Signer. INFEX: a unifying framework for crossdevice information exploration and exchange. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):2:1–2:26, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3179427.

Roohi:2021:PGD

[RGR^{+}21] Shaghayegh Roohi, Christian Guckelsberger, Asko Relas, Henri Heiskanen, Jari Takatalo, and Perttu Hämäläinen. Predicting game difficulty and engagement using AI players. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):231:1– 231:17, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474658.

Register:2024:BIR

[RGW^{+}24] Yim Register, Izzi Grasso, Lauren N. Weingarten, Lilith Fury, Constanza Eliana Chinea, Tuck J. Malloy, and Emma S. Spiro. Beyond initial removal: Lasting impacts of discriminatory content moderation to marginalized creators on Instagram. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):23:1–23:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637300.

Raffaillac:2019:PPI

[RH19] Thibault Raffaillac and Stéphane Huot. Polyphony: Programming interfaces and interactions with the entity-component-system model. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS):8:1–8:22, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331150.

Raffaillac:2022:WDR

[RH22] Thibault Raffaillac and Stéphane Huot. What do researchers need when implementing novel interaction techniques? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (EICS):159:1–159:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3532209.

Rahman:2023:FUL

[RH23] Adil Rahman and Seongkook Heo. Frappé: an ultra lightweight mobile UI framework for rapid API-based prototyping and environmental deployment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI): 211:1–211:??, September 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3604258.

Rankin:2020:STB

[RI20] Yolanda A. Rankin and India Irish. A seat at the table: Black feminist thought as a critical framework for inclusive game design. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):117:1–117:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3415188.

Ribes:2019:HLW

[Rib19] David Ribes. How I learned what a domain was. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 38:1–38:12, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359140.

Richardson:2024:DVS

[RIC^{+}24] Dan Richardson, Md Adnanul Islam, Bronwyn J. Cumbo, Pranita Shrestha, Delvin Varghese, Tom Bartindale, and Patrick Olivier. A design vocabulary for scaffolding group interaction archetypes through synchronous telephony. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):12:1–12:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637289.

Ravi:2021:PSR

[RIK21] Prerna Ravi, Azra Ismail, and Neha Kumar. The pandemic shift to remote learning under resource constraints. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):314:1–314:28, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476055.

Rezaei:2024:IDI

[RJVS24] Parisa Pour Rezaei, Tero Jokela, Akos Vetek, and Marja Salmimaa. Informing the design of intervention solutions for body-focused repetitive behaviors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 268:1–268:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3676513.

Ribeiro:2021:DPM

[RJZ^{+}21] Manoel Horta Ribeiro, Shagun Jhaver, Savvas Zannettou, Jeremy Blackburn, Gianluca Stringhini, Emiliano De Cristofaro, and Robert West. Do platform migrations compromise content moderation? Evidence from r/The Donald and r/Incels. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):316:1–316:24, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476057.

Rong:2022:WWD

[RKFK22] Yao Rong, Naemi-Rebecca Kassautzki, Wolfgang Fuhl, and Enkelejda Kasneci. Where and what: Driver attention-based object detection. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):146:1–146:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530887.

Rashidi:2020:ETC

[RKNES20] Yasmeen Rashidi, Apu Kapadia, Christena Nippert-Eng, and Norman Makoto Su. “It’s easier than causing confrontation”: Sanctioning strategies to maintain social norms and privacy on social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):023:1–023:25, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392827.

Rekik:2021:MCT

[RLGG21] Yosra Rekik, Edward Lank, Adnane Guettaf, and Laurent Grisoni. Multi-channel tactile feedback based on user finger speed. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS):

504:1–504:17, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488549.

Rateau:2022:LSE

[RLL22] Hanae Rateau, Edward Lank, and Zhe Liu. Leveraging smartwatch and earbuds gesture capture to support wearable interaction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):557:1–557:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567710.

Reid:2022:FGC

[RMB^{+}22] Elizabeth Reid, Regan L. Mandryk, Nicole A. Beres, Madison Klarkowski, and Julian Frommel. Feeling good and in control: In-game tools to support targets of toxicity. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY): 235:1–235:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3549498.

Rakhmetulla:2023:GOH

[RRA23] Gulnar Rakhmetulla, Yuan Ren, and Ahmed Sabbir Arif. GeShort: Onehanded mobile text editing and formatting with gestural shortcuts and a floating clipboard. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):212:1–212:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604259.

Reibert:2020:MIP

[RRF20] Joshua Reibert, Patrick Riehmann, and Bernd Froehlich. Multitouch interaction with parallel coordinates on large vertical displays. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (ISS):199:1–199:22, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427327.

Roels:2019:CFC

[RS19] Reinout Roels and Beat Signer. A conceptual framework and content model for next generation presentation solutions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(EICS): 7:1–7:22, June 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3331149.

Rane:2024:AAE

[RSL24] Dharma Rane, Madhu Singh, and Uttama Lahiri. Automated assesment of eye-hand coordination skill using a vertical tracing task on a gaze-sensitive human computer interaction platform for children with autism. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):224:1–224:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655598.

Reetz:2021:NVS

[RVBC^{+}21] Adrian Reetz, Deltcho Valtchanov, Michael Barnett-Cowan, Mark Hancock, and James R. Wallace. Nature vs. stress: Investigating the use of biophilia in non-violent exploration games to reduce stress. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):247:1– 247:13, September 2021.

CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474674.

Rit:2024:DPI

[RVH24] Geertje Peters Rit, Anja Volk, and Hanna Hauptmann. Diverse perspectives: an inclusive serious game for (international) students. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):302:1– 302:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677067.

Rijnders:2022:LFT

[RWB22] Frans Rijnders, Günter Wallner, and Regina Bernhaupt. Live feedback for training through real-time data visualizations: a study with league of legends. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):243:1– 243:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549506.

Ruffin:2023:EWF

[RWL23] Margie Ruffin, Gang Wang, and Kirill Levchenko.

Explaining why fake photos are fake: Does it work? Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):8:1–8:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567558.

Rakova:2021:WRA

[RYCC21] Bogdana Rakova, Jingying Yang, Henriette Cramer, and Rumman Chowdhury. Where responsible AI meets reality: Practitioner perspectives on enablers for shifting organizational practices. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):7:1–7:23, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449081.

Rabaan:2021:DMS

[RYD21] Hawra Rabaan, Alyson L. Young, and Lynn Dombrowski. Daughters of men: Saudi Women’s sociotechnical agency practices in addressing domestic abuse. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):224:1–224:31, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432923.

Rey:2022:UAB

[RZP^{+}22] Bradley Rey, Kening Zhu, Simon Tangi Perrault, Sandra Bardot, Ali Neshati, and Pourang Irani. Understanding and adapting bezel-to-bezel interactions for circular smartwatches in mobile and encumbered scenarios. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 201:1–201:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3546736.

Soubutts:2021:APT

[SAE^{+}21] Ewan Soubutts, Amid Ayobi, Rachel Eardley, Kirsten Cater, and Aisling Ann O’Kane. Aging in place together: The journey towards adoption and acceptance of stairlifts in multiresident homes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):320:1–320:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476061.

Silva:2023:EOM

[SAHE23] Lucas M. Silva, Elizabeth A. Ankrah, Yuqi Huai, and Daniel A. Epstein. Exploring opportunities for multimodality and multiple devices in food journaling. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):209:1–209:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604256.

Salem:2017:UIO

[Sal17] Paulo Salem. User interface optimization using genetic programming with an application to landing pages. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):13:1–13:17, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3099583.

Sien:2023:CDM

[SAM23] Sang-Wha Sien, Jessica Y. Ahn, and Joanna McGrenere. Co-designing mental health technologies with international University students in Canada. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):258:1–258:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610049.

Sultana:2021:SCT

[SAR21] Sharifa Sultana, Syed Ishtiaque Ahmed, and Jeffrey M. Rzeszotarski. Seeing in context: Traditional visual communication practices in rural Bangladesh. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):214:1–214:31, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432913.

Seo:2023:HER

[SAR^{+}23] Jwawon Seo, Ignacio Avellino, Damaruka Priya Rajasagi, Anita Komlodi, and Helena M. Mentis. HoloMentor: Enabling remote instruction through augmented reality mobile views. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):11:1–11:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567561.

Schubhan:2024:AVB

[SAR^{+}24] Marc Schubhan, Maximilian Altmeyer, Katja Rogers, Donald Degraen, Pascal Lessel, and Antonio Krüger. Auditory, visual, or both? Comparing visual and auditory representations of game elements in a gamified image-tagging task. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):294:1– 294:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677059.

Sato:2024:EUD

[SAY^{+}24] Yukina Sato, Takashi Amesaka, Takumi Yamamoto, Hiroki Watanabe, and Yuta Sugiura. Exploring userdefined gestures as input for hearables and recognizing ear-level gestures with IMUs. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 258:1–258:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676503.

Segura:2017:HIV

[SB17] Vinícius Segura and Simone D. J. Barbosa. HistoryViewer: Instrumenting a visual analytics application to support revisiting a session of interactive data analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):11:1–11:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095813.

Samuelsson:2020:ESE

[SB20a] Sigurdur Gauti Samuelsson and Matthias Book. Eliciting sketched expressions of command intentions in an IDE. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (ISS):200:1–200:25, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427328.

Skeba:2020:IFL

[SB20b] Patrick Skeba and Eric P. S. Baumer. Informational friction as a lens for studying algorithmic aspects of privacy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):101:1–101:22, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415172.

Sandbhor:2021:LWR

[SBAK21] Prasad Sandbhor, Priti Bangal, Deepti Aggarwal, and Rohit Ashok Khot. Life on wings: Relating to a bird’s life in a city through a board game. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):232:1– 232:28, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474659.

Strohm:2024:LUE

[SBB24] Florian Strohm, Mihai Bâce, and Andreas Bulling. Learning user embeddings from human gaze for personalised saliency prediction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):229:1–229:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655603.

Shivakumar:2020:RTI

[SBPP20] Ashutosh Shivakumar, Aishwarya Bositty, Nia S. Peters, and Yong Pei. Real-time interruption management system for efficient distributed collaboration in multitasking environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):039:1–039:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392844.

Slavuljica:2024:NEO

[SBSM24] Aleksandar Slavuljica, Kenan Bektas, Jannis Strecker, and Simon Mayer. NeighboAR: Efficient object retrieval using proximity- and gazebased object grouping with an AR system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):225:1–225:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655599.

Stone:2023:UFD

[SC23] Scott A. Stone and Craig S. Chapman. Unconscious frustration: Dynamically assessing user experience using eye and mouse tracking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA):168:1–168:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3591137.

Sanei:2024:CUI

[SC24] Arghavan Sanei and Jinghui Cheng. Characterizing usability issue discussions in open source software projects. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):30:1–30:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637307.

Serpi:2018:WFF

[SCMS18] Matteo Serpi, Alessandro Carcangiu, Alessio Murru, and Lucio Davide Spano. Web5VR: a flexible framework for integrating virtual reality input and output devices on the Web. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):4:1–4:19, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3179429.

Saha:2021:UAS

[SCP^{+}21] Manaswi Saha, Devanshi Chauhan, Siddhant Patil, Rachel Kangas, Jeffrey Heer, and Jon E. Froehlich. Urban accessibility as a sociopolitical problem: a multi-stakeholder analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW3): 209:1–209:26, January 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3432908.

Stampf:2022:TII

[SCR22] Annika Stampf, Mark Colley, and Enrico Rukzio. Towards implicit interaction in highly automated vehicles — a systematic literature review. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):191:1–191:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546726.

Seberger:2024:BLT

[SCSD24] John S. Seberger, Hyesun Choung, Jaime Snyder, and Prabu David.

Better living through creepy technology? Exploring tensions between a novel class of well-being apps and affective discomfort in app culture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):22:1–22:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637299.

Speicher:2018:MAH

[SCY^{+}18]  Maximilian  Speicher,

Jingchen Cao, Ao Yu, Haihua Zhang, and Michael Nebeling. 360Anywhere: Mobile ad-hoc collaboration in any environment using 360 video and augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):9:1–9:20, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229091.

Santoro:2024:PEI

[SD24] Carmen Santoro and Anke Dittmar. PACMHCI — engineering interactive computing systems, June 2024: Editorial introduction. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 241:1–241:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3664470.

Schneider:2021:MPT

[SDF^{+}21] Nathan Schneider, Primavera De Filippi, Seth Frey, Joshua Z. Tan, and Amy X. Zhang. Modular politics: Toward a governance layer for online communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):16:1–16:26, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449090.

Shiripour:2021:GBG

[SDO21] Morteza Shiripour, Niraj Ramesh Dayama, and Antti Oulasvirta. Gridbased genetic operators for graphical layout generation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 208:1–208:30, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461730.

Satriadi:2020:MAM

[SEC^{+}20] Kadek Ananta Satriadi, Barrett Ens, Maxime Cordeil, Tobias Czauderna, and Bernhard Jenny. Maps around me: 3D multiview layouts in immersive spaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 201:1–201:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427329.

Seering:2020:RSM

[See20] Joseph Seering. Reconsidering self-moderation: the role of research in supporting communitybased models for online content moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):107:1–107:28, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415178.

Saleh:2024:GRN

[SELS24] Mennatallah Saleh, Passant Elagroudy, Paul Lukowicz, and Christian Sturm. Ghost readers of the Nile: Decrypting password sharing habits in chatting applications among Egyptian women. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):261:1–261:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676506.

Spektor:2023:CAH

[SFA^{+}23] Franchesca Spektor, Sarah E.

Fox, Ezra Awumey, Ben Begleiter, Chinmay Kulkarni, Betsy Stringam, Christine A. Riordan, Hye Jin Rho, Hunter Akridge, and Jodi Forlizzi. Charting the automation of hospitality: an interdisciplinary literature review examining the evolution of frontline service work in the face of algorithmic management. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):33:1–33:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579466.

Schulenberg:2023:CTM

[SFLB23] Kelsea Schulenberg, Guo Freeman, Lingyuan Li, and Catherine Barwulor. “Creepy Towards My Avatar Body, Creepy Towards My Body”: How women experience and manage harassment risks in social virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):236:1–236:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3610027.

Schelble:2022:LTT

[SFM^{+}22] Beau G. Schelble, Christopher Flathmann, Nathan J. McNeese, Guo Freeman, and Rohit Mallick. Let’s think together! Assessing shared mental models, performance, and trust in humanagent teams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):13:1–13:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3492832.

Simons:2020:HHU

[SGF20] Rachel N. Simons, Danna Gurari, and Kenneth R. Fleischmann. “I Hope This Is Helpful”: Understanding crowdworkers’ challenges and motivations for an image description task. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):105:1–105:26, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415176.

Stemasov:2023:BES

[SHC^{+}23] Evgeny Stemasov, Jessica Hohn, Maurice Cordts, Anja Schikorr, Enrico Rukzio, and Jan Gugenheimer. Brick-StARt: Enabling insitu design and tangible exploration for personal fabrication using mixed reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):429:1–429:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626465.

Scheuerman:2021:DDP

[SHD21] Morgan Klaus Scheuerman, Alex Hanna, and Emily Denton. Do datasets have politics? disciplinary values in computer vision dataset development. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):317:1–317:37, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476058.

Su:2022:WYE

[SHJ^{+}22] Zhaoyuan Su, Lu He, Sunit P. Jariwala, Kai Zheng, and Yunan Chen. “What is Your Envisioned Future?”: Toward Human-AI enrichment in data work of asthma care. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):267:1–267:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555157.

Smith:2023:TPH

[SHL23] C. Estelle Smith, Hannah Miller Hillberg, and Zachary Levonian. “Thoughts & Prayers” or “[heart emoji] & [unknown emoji]”: How the release of new reactions on Caring-Bridge reshapes supportive communication in health crises. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):244:1–244:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3610035.

Salim:2024:BMM

[SHM24] Shahreen Salim, Md Naimul Hoque, and Klaus Mueller. Belief Miner: a methodology for discovering causal beliefs and causal illusions from general populations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):21:1–21:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637298.

Simpson:2022:HTY

[SHS22] Ellen Simpson, Andrew Hamann, and Bryan Semaan. How to tame “Your” algorithm: LGBTQ+ users’ domestication of TikTok. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):22:1–22:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492841.

Speicher:2018:XAC

[SHY^{+}18]  Maximilian  Speicher,

Brian D. Hall, Ao Yu, Bowen Zhang, Haihua Zhang, Janet Nebeling, and Michael Nebeling. XD-AR: Challenges and opportunities in crossdevice augmented reality application development. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS):7:1–7:24, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3229089.

Singh:2021:SLI

[SJ21] Ranjit Singh and Steven Jackson. Seeing like an infrastructure: Lowresolution citizens and the Aadhaar Identification Project. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):315:1–315:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476056.

Sheng:2020:VSI

[SK20] Jeff T. Sheng and Sanjay R. Kairam. From virtual strangers to IRL friends: Relationship development in livestreaming communities on Twitch. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2): 94:1–94:34, October 2020. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3415165.

Seering:2023:WMT

[SK23] Joseph Seering and Sanjay R. Kairam. Who moderates on Twitch and what do they do?: Quantifying practices in community moderation on Twitch. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):18:1–18:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567568.

Saveski:2021:SCC

[SKV^{+}21] Martin Saveski, Farshad Kooti, Sylvia Morelli Vitousek, Carlos Diuk, Bryce Bartlett, and Lada A. Adamic. Social catalysts: Characterizing people who spark conversations among others. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):292:1–292:20, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476023.

Shaikh:2023:WMP

[SLB23] Riyaj Shaikh, Airi Lampinen, and Barry Brown. The work to make piecework work: an ethnographic study of food delivery work in India during the COVID-19 pandemic. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):243:1–243:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610034.

Soden:2021:PAN

[SLD21] Robert Soden, Nicolas James LaLone, and Dharma Dailey. A patent application for NEXTGEN flood early warning system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):216:1–216:16, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463928.

Sorries:2024:AVT

[SLMB24] Peter Sörries, David Leimstädtner, and Claudia Müller-Birn. Advocating values through meaningful participation: Introducing a method to elicit and analyze values for enriching data donation practices in healthcare. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):16:1–16:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637293.

Speicher:2021:DSY

[SLN21]  Maximilian  Speicher,

Katy Lewis, and Michael Nebeling. Designers, the stage is yours! mediumfidelity prototyping of augmented & virtual reality interfaces with 360theater. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 205:1–205:25, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461727.

Smiley:2021:MAM

[SLT^{+}21] Jim Smiley, Benjamin Lee, Siddhant Tandon, Maxime Cordeil, Lonni Besançon, Jarrod Knibbe, Bernhard Jenny, and Tim Dwyer. The MADE-Axis: a modular actuated device to embody the axis of a data dimension. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 501:1–501:23, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488546.

Steuerlein:2022:CFT

[SM22] Benedict Steuerlein and Sven Mayer. Conductive fiducial tangibles for everyone: a data simulation-based toolkit using deep learning. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):183:1–183:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3546718.

Sun:2023:DWF

[SMLH23] Yuling Sun, Xiaojuan Ma, Silvia Lindtner, and Liang He. Data work of frontline care workers: Practices, problems, and opportunities in the context of datadriven long-term care. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):42:1–42:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3579475.

Seidelin:2022:ARP

[SMSM22] Cathrine Seidelin, Therese Moreau, Irina Shklovski, and Naja Holten Møller.

Auditing risk prediction of long-term unemployment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):8:1–8:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3492827.

Stumpf:2018:WL

[SN18] Simone Stumpf and Jeffrey Nichols. Welcome letter. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2(EICS): 1:1, June 2018. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3179426.

Sluyters:2022:QFE

[SORV22] Arthur Sluÿters, Mehdi Ousmer, Paolo Roselli, and Jean Vanderdonckt. QuantumLeap, a framework for engineering gestural user interfaces based on the leap motion controller. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 161:1–161:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3532211.

Sunkara:2023:ECD

[SPL^{+}23] Mohan Sunkara, Yash Prakash, Hae-Na Lee, Sampath Jayarathna, and Vikas Ashok. Enabling customization of discussion forums for blind users. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 176:1–176:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593228.

Swofford:2020:ISA

[SPT^{+}20] Mason Swofford, John Peruzzi, Nathan Tsoi, Sydney Thompson, Roberto Martín-Martín, Silvio Savarese, and Marynel Vázquez. Improving social awareness through DANTE: Deep affinity network for clustering conversational interactants. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):020:1–020:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392824.

Sivek:2022:SMP

[SR22] Gary Sivek and Michael Riley. Spatial model personalization in Gboard.

Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):202:1–202:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546737.

Sakashita:2022:RRE

[SRAG22] Mose Sakashita, E. Andy Ricci, Jatin Arora, and François Guimbretière. RemoteCoDe: Robotic embodiment for enhancing peripheral awareness in remote collaboration tasks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):63:1–63:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512910.

Skovfoged:2022:CKH

[SRKK22] Milo M. Skovfoged, Alexander S. Rasmussen, David Kirsh, and Hendrik Knoche. The cost of knowing: How obstacle alerts reduce walking speeds of augmented white cane users. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):192:1–192:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3546727.

Scott:2024:DDI

[SRP^{+}24] Ava Elizabeth Scott, Leon Reicherts, Aditya Kumar Purohit, Elahi Hossain, Evropi Stefanidi, Nadine Wagener, Johannes Schöning, Yvonne Rogers, and Adrian Holzer. DIY digital interventions: Behaviour change with trigger-action programming. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 252:1–252:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3676497.

Sargolzaei:2024:AMR

[SRZ24] Parisa Sargolzaei, Mudit Rastogi, and Loutfouz Zaman. Advancing mixed reality game development: an evaluation of a visual game analytics tool in action-adventure and FPS genres. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):290:1– 290:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677055.

Sher:2019:SCH

[SS19] Stephen Tsung-Han Sher and Norman Makoto Su. Speedrunning for charity: How donations gather around a live streamed couch. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (CSCW):48:1–48:26, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359150.

Stewart:2021:NPR

[SS21] Leo G. Stewart and Emma S. Spiro. Nobody puts Redditor in a binary: Digital demography, collective identities, and gender in a Subreddit network. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):8:1–8:31, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449082.

Sabuncuoglu:2022:KEP

[SS22] Alpay Sabuncuoglu and T. Metin Sezgin. Kart-ON: an extensible paper programming strategy for affordable early programming education. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (EICS):170:1–170:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3534524.

Sabuncuoglu:2023:DMC

[SS23] Alpay Sabuncuoglu and T. Metin Sezgin. Developing a multimodal classroom engagement analysis dashboard for highereducation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 188:1–188:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593240.

Schmid:2022:DSP

[SSB^{+}22] Damaris Schmid, Dario Staehelin, Andreas Bucher, Mateusz Dolata, and Gerhard Schwabe. Does social presence increase perceived competence?: Evaluating conversational agents in advice giving through a videobased survey. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):26:1–26:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492845.

Shelby:2024:CMA

[SSB^{+}24a] Renee Shelby, Ramya Srinivasan, Katharina Burgdorf, Jennifer C. Lena, and Negar Rostamzadeh. Creative ML assemblages: The interactive politics of people, processes, and products. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):38:1–38:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637315.

Smith:2024:DTO

[SSB24b] Ian Smith, Erik Scheme, and Scott Bateman. Designing a techniqueoriented sport training game for motivating a change in running technique. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):299:1– 299:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677064.

Schoenebeck:2023:ESM

[SSD23] Sarita Schoenebeck, Yu Yin Shen, and Jill Davidson. Evaluating the social media profiles of online harassers: an experimental study of attention and attitudes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (GROUP):7:1–7:??, January 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567557.

Saxena:2020:MGT

[SSGB20] Devansh Saxena, Patrick Skeba, Shion Guha, and Eric P. S. Baumer. Methods for generating typologies of non/use. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):027:1–027:26, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392832.

Skeba:2021:WCS

[SSGB21] Patrick Skeba, Devansh Saxena, Shion Guha, and Eric P. S. Baumer. Who has a choice?: Surveybased predictors of volitionality in Facebook use and non-use. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (GROUP):223:1–223:25, July 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3463935.

Schoenebeck:2021:YTS

[SSH^{+}21] Sarita Schoenebeck, Carol F.

Scott, Emma Grace Hurley, Tammy Chang, and Ellen Selkie. Youth trust in social media companies and expectations of justice: Accountability and repair after online harassment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):2:1–2:18, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449076.

Sun:2020:RDB

[SSNC20] David Sun, Chengzheng Sun, Agustina Ng, and Weiwei Cai. Real differences between OT and CRDT in correctness and complexity for consistency maintenance in co-editors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):021:1–021:30, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392825.

Stanke:2022:TTR

[SSR22] Dennis Stanke, Peer Schroth, and Michael Rohs. TrackballWatch: Trackball and rotary knob as a non-occluding input method for smartwatches in map navigation scenarios. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):199:1–199:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3546734.

Sidji:2024:HAC

[SSR24] Matthew Sidji, Wally Smith, and Melissa J. Rogerson. Human-AI collaboration in cooperative games: a study of playing codenames with an LLM assistant. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):316:1– 316:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677081.

Swaminathan:2019:LUM

[STM^{+}19] Saiganesh Swaminathan, Indrani Medhi Thies, Devansh Mehta, Edward Cutrell, Amit Sharma, and William Thies. Learn2Earn: Using mobile airtime incentives to bolster public awareness campaigns. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 49:1–49:20, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359151.

Stefanidi:2024:ARM

[STM24] Helen Stefanidi, Markus Tatzgern, and Alexander Meschtscherjakov. Augmented reality on the move: a systematic literature review for vulnerable road users. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):245:1–245:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676490.

Stellmacher:2022:EPP

[STSS22] Carolin Stellmacher, Jette Ternieten, Daria Soroko, and Johannes Schöning. Escaping the privacy paradox: Evaluating the learning effects of privacy policies with serious games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):232:1– 232:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549495.

Schnitzer:2024:PZI

[SVS^{+}24] Benjamin Schnitzer, Umut Can Vural, Bastian Schnitzer, Muhammad Usman Sardar, Oren Fuerst, and Oliver Korn. Prototyping a zoomorphic interactive robot companion with emotion recognition and affective voice interaction for elderly people. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 242:1–242:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3660244.

Sluyters:2021:ESG

[SVV21] Arthur Sluÿters, Jean Vanderdonckt, and Radu-Daniel Vatavu. Engineering slidable graphical user interfaces with slime. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 200:1–200:29, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3457147.

Schipor:2019:STS

[SVW19] Ovidiu-Andrei Schipor, Radu-Daniel Vatavu, and Wenjun Wu. SAPI-ENS: Towards software architecture to support peripheral interaction in smart environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):11:1–11:24, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331153.

Suh:2020:PJP

[SWF^{+}20] Jina Suh, Spencer Williams, Jesse R. Fann, James Fogarty, Amy M. Bauer, and Gary Hsieh. Parallel journeys of patients with cancer and depression: Challenges and opportunities for technology-enabled collaborative care. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):038:1–038:36, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392843.

Shi:2023:EGA

[SWQ^{+}23] Rongkai Shi, Yushi Wei, Xueying Qin, Pan Hui, and Hai-Ning Liang. Exploring gaze-assisted and hand-based region selection in augmented reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(ETRA): 160:1–160:??, May 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3591129.

Silva:2020:ECB

[SWT20] Thiago Rocha Silva, Marco Winckler, and Hallvard Trætteberg. Ensuring the consistency between user requirements and task models: a behavior-based automated approach. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):77:1–77:32, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394979.

Schmidt:2024:DUE

[SY24] Leonard Schmidt and Enes Yigitbas. Development and usability evaluation of transitional cross-reality interfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS):

263:1–263:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3664637.

Tang:2021:UTE

[Tan21] John Tang. Understanding the telework experience of people with disabilities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):30:1–30:27, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449104.

Taugerbeck:2019:DPP

[TAS^{+}19] Sebastian Taugerbeck, Michael Ahmadi, Marén Schorch, David Unbehaun, Konstantin Aal, and Volker Wulf. Digital participation in prison — a public discourse analysis of the use of ICT by inmates. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):233:1–233:26, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361114.

Taylor:2021:MMD

[TBDB21] Lauren C. Taylor, Kelsie Belan, Munmun De Choudhury, and Eric P. S. Baumer. Misfires, missed data, misaligned treatment: Disconnects in collaborative treatment of eating disorders. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):31:1–31:28, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449105.

Terzimehic:2023:MME

[TBHM23] Nada Terzimehic, Florian Bemmann, Miriam Halsner, and Sven Mayer. A mixed-method exploration into the mobile phone rabbit hole. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):194:1–194:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604241.

Tseng:2024:DSC

[TBL^{+}24] Emily Tseng, Rosanna Bellini, Yeuk-Yu Lee, Alana Ramjit, Thomas Ristenpart, and Nicola Dell. Data stewardship in clinical computer security: Balancing benefit and burden in participatory systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):39:1–39:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637316.

Turner:2020:MBT

[TBR20] Jessica Turner, Judy Bowen, and Steve Reeves. Model-based testing of interactive systems using interaction sequences. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):85:1–85:37, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397873.

Toch:2020:CYT

[TCH20] Eran Toch, Hadas Chassidim, and Tali Hatuka. Can you turn it off?: The spatial and social context of mobile disturbance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(CSCW2):91:1–91:18, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415162.

Theis:2019:BPN

[TD19] Julian Theis and Houshang Darabi. Behavioral Petri net mining and automated analysis for human-computer interaction recommendations in multi-application environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):13:1–13:16, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331155.

Terrier:2017:CLB

[TDC17] Lénaïc Terrier, Alexandre Demeure, and Sybille Caffiau. CCBL: a language for better supporting context centered programming in the smart home. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):14:1–14:18, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3099584.

Tsuchida:2022:TMB

[TFI^{+}22] Taichi Tsuchida, Kazuyuki Fujita, Kaori Ikematsu, Sayan Sarcar, Kazuki Takashima, and Yoshifumi Kitamura. TetraForce: a magnetic-based interface enabling pressure force and shear force input applied to front and back of a smartphone. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):564:1–564:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567717.

Tomlein:2018:ARS

[TG18] Matús Tomlein and Kaj Grønbæk. Augmented reality supported modeling of industrial systems to infer software configuration. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 2 (EICS):5:1–5:17, June 2018. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3229087.

TeBlunthuis:2022:NCC

[TKB^{+}22]  Nathan  TeBlunthuis,

Charles Kiene, Isabella Brown, Laura (Alia) Levi, Nicole McGinnis, and Benjamin Mako Hill. No community can do everything: Why people participate in similar online communities. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):61:1–61:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512908.

Turner:2022:YJA

[TPNL22] Sarah Turner, Nandita Pattnaik, Jason R. C. Nurse, and Shujun Li. “You Just Assume It Is In There, I Guess”: Understanding UK families’ application and knowledge of smart home cyber security. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):269:1–269:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555159.

Tan:2022:CSI

[TRN^{+}22] Evelyn T. S. Tan, Katja Rogers, Lennart E. Nacke, Anders Drachen, and Alex Wade. Communication sequences indicate team cohesion: a mixed-methods study of ad hoc league of legends teams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):225:1– 225:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549488.

Thevin:2021:IAE

[TRO^{+}21] Lauren Thevin, Nicolas Rodier, Bernard Oriola, Martin Hachet, Christophe Jouffrais, and Anke M. Brock. Inclusive adaptation of existing board games for gamers with and without visual impairments using a spatial augmented reality framework for touch detection and audio feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 505:1–505:33, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3488550.

Tazi:2023:CSP

[TSD23] Faiza Tazi, Sunny Shrestha, and Sanchari Das. Cybersecurity, safety, & privacy concerns of student support structure for information and communication technologies in online education. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):264:1–264:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610055.

To:2020:TJD

[TSHK20] Alexandra To, Wenxia Sweeney, Jessica Hammer, and Geoff Kaufman. “They Just Don’t Get It”: Towards social technologies for coping with interpersonal racism. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):024:1–024:29, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392828.

Toxtli:2021:QIL

[TSS21] Carlos Toxtli, Siddharth Suri, and Saiph Savage. Quantifying the invisible labor in crowd work. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):319:1–319:26, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476060.

Trullemans:2017:CMT

[TVS17] Sandra Trullemans, Lars Van Holsbeeke, and Beat Signer. The context modelling toolkit: a unified multi-layered context modelling approach. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1(EICS):8:1–8:16, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3095810.

Uckun:2020:BAB

[UAAR20] Utku Uckun, Ali Selman Aydin, Vikas Ashok, and I. V. Ramakrishnan. Breaking the accessibility barrier in nonvisual interaction with PDF forms. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):80:1–80:16, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3397868.

Uhde:2023:BHR

[UHH23] Alarith Uhde, Tim Zum Hoff, and Marc Hassenzahl. Beyond hiding and revealing: Exploring effects of visibility and form of interaction on the witness experience. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):200:1–200:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604247.

Uttarapong:2022:TUM

[ULG^{+}22] Jirassaya Uttarapong, Nina LaMastra, Reesha Gandhi, Yu hao Lee, Chien Wen (Tina) Yuan, and Donghee Yvette Wohn. Twitch users’ motivations and practices during community mental health discussions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (GROUP):5:1–5:??, January 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3492824.

Usuba:2021:MMT

[UYM21] Hiroki Usuba, Shota Yamanaka, and Homei Miyashita. Modeling movement times and success rates for acquisition of one-dimensional targets with uncertain touchable sizes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (ISS):487:1–487:15, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3486953.

Usuba:2023:CEE

[UYS23] Hiroki Usuba, Shota Yamanaka, and Junichi Sato. Clarifying the effect of edge targets in touch pointing through crowdsourced experiments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):433:1–433:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626469.

Usuba:2022:PTA

[UYSM22] Hiroki Usuba, Shota Yamanaka, Junichi Sato, and Homei Miyashita. Predicting touch accuracy for rectangular targets by using onedimensional task results. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):579:1–579:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567732.

Volkmar:2022:EPC

[VAE^{+}22] Georg Volkmar, Dmitry Alexandrovsky, Asmus Eike Eilks, Dirk Queck, Marc Herrlich, and Rainer Malaka. Effects of PCG on creativity in playful city-building environments in VR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):230:1– 230:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549493.

vanBerkel:2019:CPF

[vBGH^{+}19] Niels van Berkel, Jorge Goncalves, Danula Hettiachchi, Senuri Wijenayake, Ryan M. Kelly, and Vassilis Kostakos. Crowdsourcing perceptions of fair predictors for machine learning: a recidivism case study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 28:1–28:21, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359130.

vanDeurzen:2022:CRF

[vDBL22] Bram van Deurzen, Herman Bruyninckx, and Kris Luyten. Choreobot: a reference framework and online visual dashboard for supporting the design of intelligible robotic systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 151:1–151:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3532201.

Vincent:2021:DII

[VH21] Nicholas Vincent and Brent Hecht. A deeper investigation of the importance of Wikipedia links to search engine results. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):4:1–4:15, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449078.

Villareale:2022:WSH

[VHZ22] Jennifer Villareale, Casper Harteveld, and Jichen Zhu. “I Want To See How Smart This AI Really Is”: Player mental model development of an adversarial AI player. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):219:1– 219:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3549482.

Vayanou:2021:PSC

[VKI21] Maria Vayanou, Akrivi Katifori, and Yannis Ioannidis. Perspective sharing in culture group games: Passing around the social mediator role. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CHI PLAY):249:1– 249:24, September 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3474676.

Vallett:2020:TAS

[VMD^{+}20] Richard Vallett, Denisa Qori McDonald, Genevieve Dion, Youngmoo Kim, and Ali Shokoufandeh. Toward accurate sensing with knitted fabric: Applications and technical considerations. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):79:1–79:26, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394981.

Villa:2022:EMA

[VMHO^{+}22] Steeven Villa, Sven Mayer, Jess Hartcher-O’Brien, Albrecht Schmidt, and Tonja-Katrin Machulla. Extended mid-air ultrasound haptics for virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):578:1–578:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567731.

Vaghela:2021:BCH

[VMP21] Palashi Vaghela, Ramaravind K. Mothilal, and Joyojeet Pal. Birds of a caste — how caste hierarchies manifest in retweet behavior of Indian politicians. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):212:1–212:24, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432911.

Vanderdonckt:2019:MDC

[VN19] Jean Vanderdonckt and Thanh-Diane Nguyen. MoCaDiX: Designing cross-device user interfaces of an information system based on its class diagram. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):17:1–17:40, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331159.

Villarreal-Narvaez:2022:TDV

[VNSVL22] Santiago Villarreal-Narvaez, Arthur Sluÿters, Jean Vanderdonckt, and Efrem Mbaki Luzayisu. Theoreticallydefined vs. user-defined squeeze gestures. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):559:1–559:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567805.

Vaiani:2024:EUD

[VP24a] Giacomo Vaiani and Fabio Paternò. End-user development for humanrobot interaction: Results and trends in an emerging field. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (EICS):252:1–252:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3661146.

Vella:2024:RDO

[VP24b] Daniel Vella and Chris Porter. Remapping the document object model using geometric and hierarchical data structures for efficient eye control. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):234:1–234:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655608.

Vyas:2023:JLD

[VQ23] Dhaval Vyas and Diogo Quental. “It’s Just Like doing Meditation”: Making at a community men’s shed. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):14:1–14:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567564.

VanDriel:2023:UAA

[VVBK23] Martine Van Driel, John Vines, Belén Barros Pena, and Nelya Koteyko. Understanding autistic adults’ use of social media. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):257:1–257:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610048.

Villa:2024:EUM

[VWHW24] Steeven Villa, Yannick Weiss, Niklas Hirsch, and Alexander Wiethoff. An examination of ultrasound mid-air haptics for enhanced material and temperature perception in virtual environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):243:1–243:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676488.

Vaccaro:2021:CCM

[VXHK21] Kristen Vaccaro, Ziang Xiao, Kevin Hamilton, and Karrie Karahalios. Contestability for content moderation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):318:1–318:28, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476059.

Vyas:2019:AWC

[Vya19] Dhaval Vyas. Altruism and wellbeing as care work in a craft-based maker culture. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):239:1–239:12, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361120.

Vanderdonckt:2019:ALB

[VZV19] Jean Vanderdonckt, Mathieu Zen, and Radu-Daniel Vatavu. AB4Web: an on-line A/B tester for comparing user interface design alternatives. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):18:1–18:28, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331160.

Welsford-Ackroyd:2021:SVE

[WACdA^{+}21] Finn Welsford-Ackroyd, Andrew Chalmers, Rafael Kuffner dos Anjos, Daniel Medeiros, Hyejin Kim, and Taehyun Rhee. Spectator View: Enabling asymmetric interaction between HMD wearers and spectators with a large display. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 485:1–485:17, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3486951.

Weibert:2023:LPM

[WAD^{+}23] Anne Weibert, Sandra Juliet Ahiataku, Godson Rashid Dawuni, Konstantin Aal, Kaoru Misaki, and Volker Wulf. Looking past the miracle box: an exploration of tools and practices along the e-waste value chain in Ghana. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(GROUP):15:1–15:??, January 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567565.

Wu:2021:BFN

[WB21] Y. Wayne Wu and Brian P. Bailey. Better feedback from nicer people: Narrative empathy and ingroup framing improve feedback exchange. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW3):236:1–236:20, January 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3432935.

Wagman:2023:WPC

[WBS^{+}23] Kelly B. Wagman, Elana B. Blinder, Kevin Song, Antoine Vignon, Solomon Dworkin, Tamara Clegg, Jessica Vitak, and Marshini Chetty. “We picked community over privacy”: Privacy and security concerns emerging from remote learning sociotechnical infrastructure during COVID-19. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):245:1–245:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610036.

Weber:2024:SPG

[WBSM24] Thomas Weber, Maximilian Brandmaier, Albrecht Schmidt, and Sven Mayer. Significant productivity gains through programming with large language models. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(EICS): 256:1–256:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3661145.

Wu:2022:AWU

[WDD^{+}22] Yiying Wu, Xianghua (Sharon)

Ding, Xuelan Dai, Peng Zhang, Tun Lu, and Ning Gu. Alignment work for urban accessibility: a study of how wheelchair users travel in urban spaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):274:1–274:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555165.

Wozniak:2020:SAP

[WDP^{+}20] Mikołaj P. Wożniak, Julia Dominiak, Michał Pieprzowski, Piotr Ładoński, Krzysztof Grudzień, Lars Lischke, Andrzej Romanowski, and Paweł W. Wożniak. Subtletee: Augmenting posture awareness for beginner golfers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (ISS):204:1–204:24, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427332.

Weber:2024:EJM

[WEM24] Thomas Weber, Janina Ehe, and Sven Mayer. Extending Jupyter with multi-paradigm editors. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (EICS):245:1–245:??, June 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3660247.

Wang:2022:PIC

[WFL^{+}22] Serena Wang, Cori Faklaris, Junchao Lin, Laura Dabbish, and Jason I. Hong. ’It’s problematic but I’m not concerned’: University perspectives on account sharing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):68:1–68:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512915.

Wyche:2020:EPR

[WGG20] Susan Wyche, April Greenwood, and Brian Samuel Geyer. Exploring photography in rural Kenyan households: Considering “relational objects” in CSCW and HCI. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):046:1–046:26, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3392852.

Weninger:2020:EIM

[WGGS20] Markus Weninger, Paul Grünbacher, Elias Gander, and Andreas Schörgenhumer. Evaluating an interactive memory analysis tool: Findings from a cognitive walkthrough and a user study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):75:1–75:37, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3394977.

Weninger:2021:GEM

[WGM21] Markus Weninger, Elias Gander, and Hanspeter Mössenböck. Guided exploration: a method for guiding novice users in interactive memory monitoring tools. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 209:1–209:34, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461731.

Willatt:2024:ILL

[WGM^{+}24] Alice Willatt, Stuart Iain Gray, Helen Manchester, Tot Foster, and Kirsten Cater. An intersectional lifecourse lens and participatory methods as the foundations for codesigning with and for minoritised older adults. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):18:1–18:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637295.

Wambecke:2021:MEG

[WGN^{+}21] Jérémy Wambecke, Alix Goguey, Laurence Nigay, Lauren Dargent, Daniel Hauret, Stéphanie Lafon, and Jean-Samuel Louis de Visme. M[eye]cro: Eye-gaze+Microgestures for multitasking and interruptions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(EICS): 210:1–210:22, June 2021. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3461732.

Weixelbraun:2024:DPR

[WGS^{+}24] Petra F. Weixelbraun, Barbara Göbl, Matthias Steinböck, Mirjam Duvivié, and Fares Kayali. Discussing the protagonist role of students in game-based learning. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):300:1– 300:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677065.

Windl:2022:SIP

[WHW^{+}22] Maximiliane Windl, Alexander Hiesinger, Robin Welsch, Albrecht Schmidt, and Sebastian S. Feger. SaferHome: Interactive physical and digital smart home dashboards for communicating privacy assessments to owners and bystanders. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):586:1–586:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567739.

Wilson:2020:CSC

[Wil20] Andrew D. Wilson. Combating the spread of coronavirus by modeling fomites with depth cameras. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 203:1–203:13, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3427331.

Wang:2024:VAP

[WJH^{+}24] Yao Wang, Yue Jiang, Zhiming Hu, Constantin Ruhdorfer, Mihai Bâce, and Andreas Bulling. VisRecall++: Analysing and predicting visualisation recallability from gaze behaviour. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(ETRA):239:1–239:??, May 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3655613.

Wu:2023:IAD

[WLZL23] Liwei Wu, Qing Liu, Jian Zhao, and Edward Lank. Interactions across displays and space: a study of virtual reality streaming practices on Twitch. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):437:1–437:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626473.

Windl:2022:SPC

[WM22] Maximiliane Windl and Sven Mayer. The skewed privacy concerns of bystanders in smart environments. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 184:1–184:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3546719.

Wang:2019:HDS

[WMBO19] April Yi Wang, Anant Mittal, Christopher Brooks, and Steve Oney. How data scientists use computational notebooks for real-time collaboration. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 39:1–39:30, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359141.

Williams:2019:PWL

[WMM^{+}19] Alex C. Williams, Gloria Mark, Kristy Milland, Edward Lank, and Edith Law. The perpetual work life of crowdworkers: How tooling practices increase fragmentation in crowdwork. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 24:1–24:28, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3359126.

Wang:2022:CLC

[WNT^{+}22] Ye Wang, Srichakradhar Reddy Nagireddy, Charan Tej Thota, Duy H. Ho, and Yugyung Lee. Community-in-the-loop: Creating artificial process intelligence for coproduction of city service. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW2): 285:1–285:??, November 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3555176.

Williams:2020:UGS

[WO20] Adam S. Williams and Francisco R. Ortega. Understanding gesture and speech multimodal interactions for manipulation tasks in augmented reality using unconstrained elicitation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 202:1–202:21, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427330.

Wells:2022:SEM

[WPH22a] Thomas Wells, Dominic Potts, and Steven Houben. A study into the effect of mobile device configurations on colocated collaboration using AR. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI): 200:1–200:??, September 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3546735.

Winckler:2022:EAG

[WPH^{+}22b] Marco Winckler, Philippe Palanque, Jean Luc Hak, Eric Barboni, Olivier Nicolas, and Laurent Goncalves. Engineering annotations: a generic framework for gluing design artefacts of interactive systems. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 174:1–174:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3535063.

Wang:2024:MBD

[WQCZ24] Ruotong Wang, Lin Qiu, Justin Cranshaw, and Amy X. Zhang. Meeting bridges: Designing information artifacts that bridge from synchronous meetings to asynchronous collaboration. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):35:1–35:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3637312.

Wu:2022:DGU

[WRF^{+}22] Hui-Yin Wu, Florent Robert, Théo Fafet, Brice Graulier, Barthelemy Passin-Cauneau, Lucile Sassatelli, and Marco Winckler. Designing guided user tasks in VR embodied experiences. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 158:1–158:??, June 2022. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3532208.

Wijkstra:2024:HTT

[WRM^{+}24] Michel Wijkstra, Katja Rogers, Regan L. Mandryk, Remco C. Veltkamp, and Julian Frommel. How to tame a toxic player? A systematic literature review on intervention systems for toxic behaviors in online video games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):315:1– 315:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677080.

Warren:2022:TFF

[WS22] Rachel B. Warren and Niloufar Salehi. Trial by file formats: Exploring public defenders’ challenges working with novel surveillance data. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):67:1–67:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512914.

Wu:2023:HDY

[WS23] Qunfang Wu and Bryan Semaan. “How Do You Quantify How Racist Something Is?”: Colorblind moderation in decentralized governance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):239:1–239:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3610030.

Windl:2024:EUM

[WSM24] Maximiliane Windl, Magdalena Schlegel, and Sven Mayer. Exploring users’ mental models and privacy concerns during interconnected interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI): 259:1–259:??, September 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3676504.

Woehler:2022:AGC

[WvECM22] Leslie Woehler, Moritz von Estorff, Susana Castillo, and Marcus Magnor. Automatic generation of customized areas of interest and evaluation of observers’ gaze in portrait videos. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ETRA):144:1–144:??, May 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3530885.

Weiss:2024:ERS

[WVG^{+}24] Yannick Weiss, Steeven Villa, Jesse W Grootjen, Matthias Hoppe, Yasin Kale, and Florian Müller. Exploring redirection and shifting techniques to mask hand movements from shoulder-surfing attacks during PIN authentication in virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):257:1–257:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3676502.

Wang:2023:MES

[WWH^{+}23] Yiyuan Wang, Senuri Wijenayake, Marius Hoggenmüller, Luke Hespanhol, Stewart Worrall, and Martin Tomitsch. My eyes speak: Improving perceived sociability of autonomous vehicles in shared spaces through emotional robotic eyes. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):214:1–214:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604261.

Wessel:2021:DDM

[WWSG21] Mairieli Wessel, Igor Wiese, Igor Steinmacher, and Marco Aurelio Gerosa. Don’t disturb me: Challenges of interacting with software bots on open source software projects. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):301:1–301:21, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476042.

Wu:2022:CAC

[WWSS22] Qunfang Wu, Louisa Kayah Williams, Ellen Simpson, and Bryan Semaan. Conversations about crime: Re-enforcing and fighting against platformed racism on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):54:1–54:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3512901.

Wang:2021:CTB

[WWT^{+}21] Liuping Wang, Dakuo Wang, Feng Tian, Zhenhui Peng, Xiangmin Fan, Zhan Zhang, Mo Yu, Xiaojuan Ma, and Hongan Wang. CASS: Towards building a socialsupport chatbot for online health community. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):9:1–9:31, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449083.

Wu:2022:PPO

[WXL^{+}22] Qin Wu, Rao Xu, Yuantong Liu, Danielle Lottridge, and Suranga Nanayakkara. Players and performance: Opportunities for social interaction with augmented tabletop games at centres for children with autism. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):563:1–563:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567716.

Wang:2021:TEQ

[WYY^{+}21] Liang Wang, Zhiwen Yu, Dingqi Yang, Tian Wang, En Wang, Bin Guo, and Daqing Zhang. Task execution quality maximization for mobile crowdsourcing in geosocial networks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):312:1–312:29, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3476053.

Wang:2022:VAV

[WZFA22] Cheng Yao Wang, Qian Zhou, George Fitzmaurice, and Fraser Anderson. VideoPoseVR: Authoring virtual reality character animations with online videos. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):575:1–575:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567728.

Wen:2024:FML

[WZT^{+}24] Jinhe Wen, Zhongyang Zhang, Tuan M. Tran, Lianrui Mu, Tauhidur Rahman, and Haojian Jin. Folk models of loot boxes in video games. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):307:1– 307:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677072.

Xiao:2024:CDC

[XDC^{+}24] Yunpeng Xiao, Bufan Deng, Siqi Chen, Kyrie Zhixuan Zhou, Ray L. C., Luyao Zhang, and Xin Tong. “Centralized or Decentralized?”: Concerns and value judgments of stakeholders in the non-fungible tokens (NFTs) market. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CSCW1):28:1–28:??, April 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3637305.

Xiao:2017:SRC

[XHH17] Robert Xiao, Scott Hudson, and Chris Harrison. Supporting responsive cohabitation between virtual interfaces and physical objects on everyday surfaces. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 1 (EICS):12:1–12:17, June 2017. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3095814.

Xiao:2022:EPE

[XJBH22] Ruowei Xiao, Sangwon Jung, Oguz ‘Oz’ Buruk, and Juho Hamari. Exploring the player experiences of wearable gaming interfaces: a user elicitation study. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY):234:1– 234:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3549497.

Xing:2023:KPG

[XKRW23] Yushan Xing, Ryan M.

Kelly, Melissa J. Rogerson, and Jenny Waycott. “Keeping the Program Going”: Technology use by community organizations to support the social connectedness of older adults during the COVID-19 pandemic. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):241:1–241:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610032.

Xia:2024:AAS

[XMS24] Chengshuo Xia, Tian Min, and Yuta Sugiura. AudioMove: Applying the spatial audio to multi-directional limb exercise guidance. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):244:1–244:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676489.

Xie:2022:SEI

[XOEK22] Benjamin Xie, Alannah Oleson, Jayne Everson, and Amy J. Ko. Surfacing equity issues in large computing courses with peerranked, demographicallylabeled student feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CSCW1):65:1–65:??, April 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3512912.

Xenopoulos:2022:GAL

[XRS22] Peter Xenopoulos, João Rulff, and Claudio Silva. ggViz: Accelerating large-scale esports game analysis. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(CHI PLAY): 238:1–238:??, October 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3549501.

Xu:2023:TMS

[XYDV23] Tian Xu, Junnan Yu, Dylan Thomas Doyle, and Stephen Voida. Technology-mediated strategies for coping with mental health challenges: Insights from people with bipolar disorder. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):240:1–240:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610031.

Xia:2020:EAC

[XZL^{+}20] Yan Xia, Haiyi Zhu, Tun Lu, Peng Zhang, and Ning Gu. Exploring antecedents and consequences of toxicity in online discussions: a case study on Reddit. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW2):108:1–108:23, October 2020. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3415179.

Yoo:2023:DSM

[YBE^{+}23] Dong Whi Yoo, Aditi Bhatnagar, Sindhu Kiranmai Ernala, Asra Ali, Michael L. Birnbaum, Gregory D. Abowd, and Munmun De Choudhury. Discussing social media during psychotherapy consultations: Patient narratives and privacy implications. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):46:1–46:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3579479.

Yang:2022:TIC

[YDW^{+}22] Ying Yang, Tim Dwyer, Michael Wybrow, Benjamin Lee, Maxime Cordeil, Mark Billinghurst, and Bruce H. Thomas. Towards immersive collaborative sensemaking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):588:1–588:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3567741.

Yu:2021:PFY

[YGHR21] Junnan Yu, Julisa Granados, Ronni Hayden, and Ricarose Roque. Parental facilitation of young children’s technologybased learning experiences from nondominant groups during the COVID-19 pandemic. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW2):307:1–307:27, October 2021. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3476048.

Yigitbas:2019:CDD

[YHR^{+}19] Enes Yigitbas, André Hottung, Sebastian Mansfield Rojas, Anthony Anjorin, Stefan Sauer, and Gregor Engels. Contextand data-driven satisfaction analysis of user interface adaptations based on instant user feedback. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):19:1–19:20, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331161.

Yang:2023:ISU

[YK23] Boyin Yang and Per Ola Kristensson. Imperfect surrogate users: Understanding performance implications of augmentative and alternative communication systems through bounded rationality, human error, and interruption modeling. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):213:1–213:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3604260.

Young:2020:DMA

[YMG^{+}20] Gareth Young, Hamish Milne, Daniel Griffiths, Elliot Padfield, Robert Blenkinsopp, and Orestis Georgiou. Designing mid-air haptic gesture controlled user interfaces for cars. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (EICS):81:1–81:23, June 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3397869.

Yu:2023:DEV

[YML^{+}23] Neng-Hao Yu, Shih-Yu Ma, Cong-Min Lin, Chi-Aan Fan, Luca E. Taglialatela, Tsai-Yuan Huang, Carolyn Yu, Yun-Ting Cheng, Ya-Chi Liao, and Mike Y. Chen. DrivingVibe: Enhancing VR driving experience using inertiabased vibrotactile feedback around the head. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI):206:1–206:??, September 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3604253.

Yin:2019:EMF

[YOFC19] Guofan Yin, Martin J.- D. Otis, Pascal E. Fortin, and Jeremy R. Cooperstock. Evaluating multimodal feedback for assembly tasks in a virtual environment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (EICS):21:1–21:11, June 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3331163.

Yang:2022:CES

[YPS^{+}22] Qi Yang, Michalis Papakostas, Jack M. Scott, Erin R. O’Neill, Kirill Sergeyevich Kondrashov, Victor A. Mateevitsi, Gregory Olsen, and Andrew Burke Dittberner. CheckMyFit: Ear selfie to assist user insertion of hearing aids. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(MHCI):186:1–186:??, September 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3546721.

Yeh:2020:EST

[YRIV20] Yen-Ting Yeh, Quentin Roy, Antony Albert Raj Irudayaraj, and Daniel Vogel. Expanding side touch input on mobile phones: Finger reachability and twodimensional taps and flicks using the index and thumb. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS):

206:1–206:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427334.

Yamanaka:2023:EAG

[YTF^{+}23] Shota Yamanaka, Takumi Takaku, Yukina Funazaki, Noboru Seto, and Satoshi Nakamura. Evaluating the applicability of GUI-based steering laws to VR car driving: a case of curved constrained paths. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):430:1–430:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626466.

Yamanaka:2020:RDG

[YU20] Shota Yamanaka and Hiroki Usuba. Rethinking the dual Gaussian distribution model for predicting touch accuracy in on-screen-start pointing tasks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 205:1–205:20, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427333.

Yamanaka:2022:EPS

[YUSM22] Shota Yamanaka, Hiroki Usuba, Wolfgang Stuerzlinger, and Homei Miyashita. The effectiveness of path-segmentation for modeling lasso times in width-varying paths. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):584:1–584:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567737.

Yan:2023:NEE

[YWL^{+}23] Zihan Yan, Yanhong Wu, Danli Luo, Chao Zhang, Qihang Jin, Wei Chen, Yingcai Wu, Xiang “Anthony” Chen, Guanyun Wang, and Haipeng Mi. NaCanva: Exploring and enabling the natureinspired creativity for children. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(MHCI): 215:1–215:??, September 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3604262.

Yamagishi:2024:AAM

[YYT^{+}24] Takeru Yamagishi, Yuki Yasunaka, Eri Takayama, Takashi Nomaru, Takanori Komatsu, and Keita Watanabe. Ah-aloud method to comprehend time-series emotion observation during gameplay: an initial investigation with Japanese speakers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):291:1– 291:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3677056.

Zhou:2021:LSE

[ZAS^{+}21] Ke Zhou, Luca Maria Aiello, Sanja Scepanovic, Daniele Quercia, and Sara Konrath. The language of situational empathy. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(CSCW1):13:1–13:19, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449087.

Zakaria:2019:SSD

[ZBL19] Camellia Zakaria, Rajesh Balan, and Youngki Lee. StressMon: Scalable detection of perceived stress and depression using passive sensing of changes in work routines and group interactions. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3(CSCW): 37:1–37:29, November 2019. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3359139.

Zen:2023:QMB

[ZBV23] Mathieu Zen, Nicolas Burny, and Jean Vanderdonckt. A quality model-based approach for measuring user interface aesthetics with grace. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 172:1–172:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

URL https://dl.acm.
org/doi/10.1145/3593224.

Zytko:2019:DSM

[ZD19] Douglas Zytko and Leanne DeVreugd. Designing a social matching system to connect academic researchers with local community collaborators. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 3 (GROUP):236:1–236:15, December 2019. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3361117.

Zhang:2023:EPI

[ZES^{+}23] Yidan Zhang, Barrett Ens, Kadek Ananta Satriadi, Ying Yang, and Sarah Goodwin. Embodied provenance for immersive sensemaking. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):435:1–435:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3626471.

Zhang:2022:BUO

[ZGL^{+}22] Peng Zhang, Zhengqing Guan, Baoxi Liu, Xianghua (Sharon) Ding, Tun Lu, Hansu Gu, and Ning Gu. Building user-oriented personalized machine translator based on user-generated textual content. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):280:1–280:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3555171.

Zhang:2023:UEM

[Zha23] Xinyong Zhang. Understanding the effects of movement direction on 2D touch pointing tasks. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (ISS):446:1–446:??, December 2023. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3626482.

Zhang:2022:CIB

[ZKL22] Futian Zhang, Keiko Katsuragawa, and Edward Lank. Conductor: Intersection-based bimanual pointing in augmented and virtual reality. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ISS): 560:1–560:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567713.

Zhang:2022:EDC

[ZLB^{+}22] Wenxiao Zhang, Sikun Lin, Farshid Hassani Bijarbooneh, Hao-Fei Cheng, Tristan Braud, Pengyuan Zhou, Lik-Hang Lee, and Pan Hui. EdgeXAR: a 6-DoF camera multi-target interaction framework for MAR with userfriendly latency compensation. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(EICS): 152:1–152:??, June 2022.

CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3532202.

Zhang:2024:HDU

[ZMGK24] Zinan Zhang, Sam Moradzadeh, Xinning Gui, and Yubo Kou. Harmful design in user-generated games and its ethical and governance challenges: an investigation of design co-ideation of game creators on Roblox. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(CHI PLAY):311:1– 311:??, 2024. CODEN ???? ISSN 2573-0142 (electronic). URL

                                        https://dl.acm.org/
                                        doi/10.1145/3677076.

Zhang:2021:JAU

[ZMR21] Xinlei Zhang, Takashi Miyaki, and Jun Rekimoto. JustSpeak: Automated, user-configurable, interactive agents for speech tutoring. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (EICS):202:1–202:24, June 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3459744.

Zhao:2022:RGB

[ZMS^{+}22] Kaixing Zhao, Julie Mulet, Clara Sorita, Bernard Oriola, Marcos Serrano, and Christophe Jouffrais. Remote graphicbased teaching for pupils with visual impairments: Understanding current practices and co-designing an accessible tool with special education teachers. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(ISS): 580:1–580:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3567733.

Zytko:2022:DAU

[ZMTH22] Douglas Zytko, Nicholas Mullins, Shelnesha Taylor, and Richard H. Holler. Dating apps are used for more than dating: How users disclose and detect (non-)Sexual interest in people-nearby applications. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6(GROUP):30:1–30:??, January 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3492849.

Zhang:2020:HDD

[ZMW20] Amy X. Zhang, Michael Muller, and Dakuo Wang.

How do data science workers collaborate? roles, workflows, and tools. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4 (CSCW1):022:1–022:23, May 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3392826.

Zhang:2021:AMO

[ZMZL21] Futian Zhang, Sachi Mizobuchi, Wei Zhou, and Edward Lank. Analyzing midair object pointing mappings for smart display input. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5(ISS): 490:1–490:18, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488535.

Zieher:2024:UOM

[ZP24] Tamara Zieher and Kathrin Probst. Usability optimization for mobile menu design: an empirical study of hand grips and user preferences. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8(MHCI):263:1–263:??, September 2024. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3676508.

Zhang:2024:HFY

[ZPMK24] Qianqia (Queenie) Zhang, Soya Park, Michael Muller, and David R. Karger. “How fancy you are to make us use your fancy tool”: Coordinating individuals’ tool preference over group boundaries. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 8 (GROUP):4:1–4:??, January 2024. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3633069.

Zand:2022:TOT

[ZRA22] Ghazal Zand, Yuan Ren, and Ahmed Sabbir Arif. TiltWalker: Operating a telepresence robot with one-hand by tilt controls on a smartphone. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (ISS):572:1–572:??, December 2022. CODEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3567725.

Zhou:2021:TTB

[ZRL21] Xingchen Zhou, Pei-Luen Patrick Rau, and Xueqian Liu. “Time to Take a Break”: How heavy adult gamers react to a built-in gaming gradual intervention system. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (CSCW1):3:1–3:30, April 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3449077.

Zhong:2022:BAH

[ZRP^{+}22] Sailin Zhong, Loïc Rosset, Michael Papinutto, Denis Lalanne, and Hamed Seiied Alavi. Binaural audio in hybrid meetings: Effects on speaker identification, comprehension, and user experience. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):279:1–279:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555170.

Zellner:2023:UDA

[ZSK^{+}23] Katherine A. Zellner, Aleksandra Sarcevic, Megan A. Krentsa, Travis M. Sullivan, and Randall S. Burd. Understanding delay awareness and mitigation mechanisms through an iterative design and evaluation of a prototype alert system for complex teamwork. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (CSCW2):260:1–260:??, October 2023. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3610051.

Zhao:2020:VHV

[ZSOJ20] Kaixing Zhao, Marcos Serrano, Bernard Oriola, and Christophe Jouffrais. VibHand: Onhand vibrotactile interface enhancing nonvisual exploration of digital graphics. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 4(ISS): 207:1–207:19, November 2020. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3427335.

Zhao:2023:CPW

[ZWG^{+}23] Yuheng Zhao, Xinyu Wang, Chen Guo, Min Lu, and Siming Chen. ContextWing: Pair-wise visual comparison for evolving sequential patterns of contexts in social media data streams. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(CSCW1):40:1–40:??, April 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3579473.

Zhao:2021:SBH

[ZWZ^{+}21] Run Zhao, Dong Wang, Qian Zhang, Xueyi Jin, and Ke Liu. Smartphonebased handwritten signature verification using acoustic signals. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 5 (ISS):499:1–499:26, November 2021. CODEN ???? ISSN 2573-0142 (electronic). URL https:

                                        //dl.acm.org/doi/10.
                                        1145/3488544.

Zhao:2023:LCN

[ZYY23] Guanhua Zhao, Yueli Yan, and Zhice Yang. Low-cost and non-visual labels using magnetic printing. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7(EICS): 180:1–180:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic).

                                        URL https://dl.acm.
                                        org/doi/10.1145/3593232.

Zander:2023:DOB

[ZZ23] Patrik Zander and Valentin Zieglmeier. Data owner benefit-driven design of people analytics. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 7 (EICS):173:1–173:??, June 2023. CODEN ???? ISSN 2573-0142 (electronic). URL https:

//dl.acm.org/doi/10.
1145/3593225.

Zhang:2022:CCG

[ZZK^{+}22] Yang Zhang, Ruohan Zong, Ziyi Kou, Lanyu Shang, and Dong Wang. CrowdNAS: a crowdguided neural architecture searching approach to disaster damage assessment. Proceedings of the ACM on Human-Computer Interaction (PACMHCI), 6 (CSCW2):288:1–288:??, November 2022. CO-DEN ???? ISSN 2573-0142 (electronic). URL

https://dl.acm.org/
doi/10.1145/3555179.